### v1

# Local analysis of iterative reconstruction from discrete generalized Radon transform data in the plane

Abstract.
Local reconstruction analysis (LRA) is a powerful and flexible technique to study images reconstructed from discrete generalized Radon transform (GRT) data, $g=\mathcal{R}f$. The main idea of LRA is to obtain a simple formula to accurately approximate an image, $f_{\epsilon}(x)$, reconstructed from discrete data $g(y_{j})$ in an $\epsilon$-neighborhood of a point, $x_{0}$. The points $y_{j}$ lie on a grid with step size of order $\epsilon$ in each direction. In this paper we study an iterative reconstruction algorithm, which consists of minimizing a quadratic cost functional. The cost functional is the sum of a data fidelity term and a Tikhonov regularization term. The function $f$ to be reconstructed has a jump discontinuity across a smooth surface $\mathcal{S}$. Fix a point $x_{0}\in\mathcal{S}$ and any $A>0$. The main result of the paper is the computation of the limit $\Delta F_{0}(\check{x};x_{0}):=\lim_{\epsilon\to 0}(f_{\epsilon}(x_{0}+\epsilon\check{x})-f_{\epsilon}(x_{0}))$, where $f_{\epsilon}$ is the solution to the minimization problem and $|\check{x}|\leq A$. A numerical experiment with a circular GRT demonstrates that $\Delta F_{0}(\check{x};x_{0})$ accurately approximates the actual reconstruction obtained by the cost functional minimization.
Key words and phrases:
Generalized Radon transform, discrete data, resolution analysis, Fourier integral operators, singularities
2020 Mathematics Subject Classification:
44A12, 65R10, 92C55
1Department of Mathematics, University of Central Florida, Orlando, FL 32816 (Alexander.Katsevich@ucf.edu).
1. Introduction
1.1. Local reconstruction analysis (LRA)
Let $\mathcal{R}$ be the generalized Radon transform (GRT); $f(x)$, $x\in\mathcal{U}\subset\mathbb{R}^{n}$, be the function to be reconstructed; and $g(y)$, $y\in\mathcal{V}\subset\mathbb{R}^{n}$, where $g=\mathcal{R}f$, be the GRT of $f$. The open sets $\mathcal{U}$ and $\mathcal{V}$ represent the image and data domains, respectively. The values of $g(y)$ are given on a rectangular grid $y_{j}$, $j\in\mathbb{Z}^{n}$. The step size of the grid along each coordinate is proportional to some small parameter $\epsilon>0$.
Consider a reconstruction formula of the form $f_{0}=\mathcal{R}^{*}\mathcal{B}g$, where $\mathcal{R}$ is the adjoint transform and $\mathcal{B}$ is a Pseudo-Differential Operator ($\Psi$DO). Depending on the choice of $\mathcal{B}$, the reconstruction can be theoretically exact ($f_{0}=f$), quasi-exact ($f_{0}-f$ is less singular than $f$), edge-enhancing ($f_{0}$ is more singular than $f$), and smoothing of a finite degree ($f_{0}$ is less singular than $f$).
Next, let $\varphi$ be a compactly supported, sufficiently smooth interpolation kernel, and let $g_{\epsilon}(y)$, $y\in\mathcal{V}$, denote the function obtained by interpolating the values $g(y_{j})$. Denote $f_{\epsilon}=\mathcal{R}^{*}\mathcal{B}g_{e}$. Thus, $f_{\epsilon}$ is reconstructed from discrete data.
LRA is a powerful and flexible technique proposed by the author to study images $f_{\epsilon}$ reconstructed from discrete GRT data. The main idea of LRA is to obtain a simple formula to accurately approximate $f_{\epsilon}$ in an $\epsilon$-neighborhood of a point, $x_{0}\in\mathcal{U}$. To illustrate the main idea of LRA consider an example. Let $f$ be a piece-wise smooth real-valued function with singularities along some hypersurface $\mathcal{S}$. More precisely, $f$ is conormal with respect to $\mathcal{S}$ [21, Definition 18.2.6]. Fix any $A>0$. We show that, in a variety of settings, the following limit exists
(1.1) $\lim_{\epsilon\to 0}\epsilon^{\nu}f_{\epsilon}(x_{0}+\epsilon\check{x}),\ |\check{x}|<A,$
the limit is uniform in $\check{x}$, and compute the limit explicitly. Here $\nu\geq 0$ is some constant, which depends on the strength of the singularity of $f_{0}$ at $x_{0}$.
To illustrate one such setting, suppose $f$ has a jump across $\mathcal{S}$, where $\mathcal{S}\subset\mathbb{R}^{2}$ is a smooth curve. Suppose also that the reconstruction is either exact or quasi-exact. In this case $\mathcal{R}^{*}\mathcal{B}\mathcal{R}$ is a $\Psi$DO with the principal symbol 1 and $\nu=0$. Pick $x_{0}\in\mathcal{S}$. Under some mild conditions on $\mathcal{S}$ and $x_{0}$, we show
(1.2) $\lim_{\epsilon\to 0}f_{\epsilon}(x_{0}+\epsilon\check{x})=f_{0}(x_{0}^{-})+\Delta f(x_{0})\text{DTB}(\check{x};x_{0}),\ |\check{x}|<A,$
uniformly in $\check{x}$. Here
  1. (1)
    $f_{0}(x_{0}^{-})$ is the value of the reconstruction from continuous data, $f_{0}$, on one side of $\mathcal{S}$,
  2. (2)
    $\Delta f(x_{0}):=f_{0}(x_{0}^{+})-f_{0}(x_{0}^{-})$ is the value of the jump of $f$ across $\mathcal{S}$ at $x_{0}$, and
  3. (3)
    DTB, which stands for the Discrete Transition Behavior, is an easily computable function independent of $f$.
The DTB function depends only on the curvature of $\mathcal{S}$ at $x_{0}$ and on the interpolation kernel, $\varphi$. When $\epsilon$ is sufficiently small, the right-hand side of (1.2) is an accurate approximation of $f_{\epsilon}$, and the DTB function accurately describes the smoothing of the jump of $f$ in $f_{\epsilon}$ due to the discrete nature of the tomographic data.
As is seen from (1.2), LRA provides a uniform approximation to $f_{\epsilon}(x)$ in domains of size $\sim\epsilon$, which is comparable to the data step size, i.e. at native resolution. Given the DTB function, one can study local properties of reconstruction from discrete data, such as spatial resolution.
In [24, 25, 27, 26, 30] we compute limits of the kind (1.1) when $f$ has jumps across smooth curves (and surfaces in higher dimensions). We consider GRT in any $\mathbb{R}^{n}$, $n\geq 2$, which integrates over submanifolds of any dimension $N$, $11\leq N\leq n-$ 1, $f$ may have a fairly general (conormal) singularity at $\mathcal{S}$, and the reconstruction operator is a fairly general Fourier Integral Operator (FIO).
In many applications, discontinuities of $f$ occur across rough surfaces. Examples include soil and rock imaging, where the surfaces of cracks and pores, and the boundaries between different regions are highly irregular and frequently simulated by fractals [15, 36, 39, 42]. Another example is cancer detection in CT. Cancerous lesions have rougher boundaries than benign ones [10, 9]. In [31, 29, 28], we extend LRA to functions on the plane with jumps across rough boundaries (i.e., sing $\text{sing\hskip 1.0ptsupp}f$ is no longer a smooth curve).
In [33], LRA is extended to view aliasing by obtaining a formula for a limit similar to (1.1) that accurately describes aliasing artifacts away from $\mathcal{S}$, i.e. $x_{0}\not\in\mathcal{S}$. These are rapidly oscillating artifacts (ripples) that are caused by aliasing from $\mathcal{S}$.
Besides the discrete nature of observed data, the second major factor affecting image quality is noise in the data. Very recently, Anuj Abhishek, James Webber and the author showed that the LRA approach allows one to obtain an accurate, complete, and simple description of the reconstructed noise in an $O(\epsilon)$-size domain [2, 32] if $g(y_{j})$ are known with random errors.
Some of the above questions have been investigated using other techniques, most notably within the framework of sampling theory. See, for example, [38, 40, 12] just to name a few papers, and more recently, [51, 37]. However, these approaches rely on certain assumptions, such as $f$ being bandlimited or essentially bandlimited. Assumptions of this sort usually hold only approximately and do not apply to functions with jumps. In contrast, LRA does not require $f$ to be bandlimited in any way, either exactly or approximately.
Even though LRA applies in a wide variety of settings, one of its major limitations is that so far it has only been applied when $f_{\epsilon}$ is computed by a linear reconstruction algorithm, i.e., by an application of a reconstruction formula to data: $f_{\epsilon}=\mathcal{R}^{*}\mathcal{B}g_{\epsilon}$. Very few GRTs admit an exact inversion formula. In such cases, at best, formula-based reconstruction can only guarantee an accurate recovery of the singularities of $f$ (which we called quasi-exact reconstruction above).
1.2. New result - LRA for iterative reconstruction
A popular flexible approach, which can numerically exactly invert a wide range of GRTs without relying on an inversion formula, consists of iterative minimization of a cost functional, see [17] and references therein. This approach is called iterative reconstruction (IR). A typical cost functional is the sum of a data fidelity and regularization terms.
In this paper we extend LRA to the study of 2D IR when $f$ has jumps across smooth boundaries, $\mathcal{S}$. We consider a GRT, $\mathcal{R}$, in the plane. The transform $\mathcal{R}$ maps a function $f(x)$, $x\in\mathcal{U}\subset\mathbb{R}^{2}$, into its weighted integrals, $g(y)=(\mathcal{R}f)(y)$, along a family of curves $\mathcal{S}_{y}$, $y\in\mathcal{V}\subset\mathbb{R}^{2}$. A more precise description of $\mathcal{V}$ is at the beginning of section 3.1. The discrete data, $g(y_{j})$, are given at the points $y_{j}$, $j\in\mathbb{Z}^{2}$. The reconstructed function, still denoted $f_{\epsilon}$, is computed as the solution to the following quadratic minimization problem:
(1.3) $f_{\epsilon}=\text{argmin}_{f\in H_{c}^{1}(\mathcal{U}_{b})}\Psi(f),\ \Psi(f):=\|\mathcal{R}f-g_{\epsilon}\|_{L^{2}(\mathcal{V})}^{2}+\kappa\epsilon^{3}\|\nabla f\|_{L^{2}(\mathcal{U})}^{2},$
where $\kappa>0$ is the regularization parameter (which is predetermined and independent of $\epsilon$). In the spirit of LRA, the factor $\epsilon^{3}$ in front of the Tikhonov regularization term ensures that the resolution of the reconstruction is at the native scale $\epsilon$.
To solve (1.3) numerically, one usually assumes that $f_{\epsilon}(x)=\sum_{l}c_{l}\psi_{l}(x)$, where $\psi_{l}$ are some basis functions, and their number is of the same order of magnitude as the number of data points. In this case, the reconstruction consists of computing the coefficients $c_{l}$. For conventional analysis of reconstruction at a scale $\gg\epsilon$, the effects of this finite-dimensional approximation of $f_{\epsilon}$ are largely invisible. However, LRA is designed to study reconstruction at the native scale, $\epsilon$. This is the scale of greatest practical importance. At this scale, the two types of effects, (a) due to the discrete nature of data and (b) due to the representation of $f_{\epsilon}$ as a linear combination of basis functions, become comparable and relevant.
It is clear that the two effects are fundamentally different. The first effect cannot be overcome, because the data are always discrete. The second effect can be largely mitigated by selecting a sufficiently fine reconstruction grid, making it negligible compared to the first one.
An important feature of the minimization problem (1.3) is that it looks for a solution, $f_{\epsilon}$, in an (infinite dimensional) Hilbert space. Of course, this is numerically impossible. In the end, we still have to find a solution in a finite-dimensional vector space. However, numerically, the effects of this replacement of the solution space can be made as small as one likes relative to the effects of discrete data. We consider (1.3) because our goal is to study the most fundamental limitation on image quality, namely the limitation due to discrete data. While the effects of finite-dimensional approximation of a solution are less fundamental (because they can be effectively mitigated), they are important from the point of view of computational complexity. Nevertheless, this line of research is beyond the scope of the paper and will be the subject of future work.
The main result of the paper, stated in Theorems 3.6 and 3.7, consists of formulas of the type (1.2) for the solution $f_{\epsilon}$ to the minimization problem (1.3).
To the best of the author’s knowledge, this is the first paper that investigates IR at native resolution in a mathematically rigorous fashion. Usually, analysis of IR consists of showing that the algorithm converges in some global norm [6]. Due to its practical significance, the local analysis of resolution of various tomographic modalities has been extensively studied in the applied literature [44, 13, 50, 49, 47, 48, 3, 8]. However, rigorous theoretical analysis has been absent.
1.3. Organization of the paper
The paper is organized as follows. In section 2, we introduce various function spaces used in the paper as well as remind the reader the definitions of a $\Psi$DO and FIO. These are not the most general definitions, but they will suffice for our purposes. In section 3 we describe the setting of the problem, including the GRT $\mathcal{R}$, the class of functions $f$ to be reconstructed, and the minimization problem, whose solution, $f_{\epsilon}$, approximates $f$. We also state our main results, Theorems 3.6 and 3.7.
The rest of the paper contains the proofs of the two theorems. In section 4 we obtain a formula for $\mathcal{R}f$ in a neighborhood of its singular support, (4.13). In section 5 we obtain an alternative formula for computing $f_{\epsilon}$, equation (5.6). The equation is of the form $\mathcal{T}f_{\epsilon}=F_{\epsilon}$, where $\mathcal{T}$ is some $\Psi$DO, and $F_{\epsilon}$ is computed from the data $g(y_{j})$. Also we find another function, $G_{\epsilon}$, such that (a) $G_{\epsilon}$ is easier to analyze than $f_{\epsilon}$ and (b) the local behavior of $f_{\epsilon}$ and $G_{\epsilon}$ near $x_{0}$ are the same from the perspective of LRA, see Lemma 5.2. In section 6 we prove that $\|F_{\epsilon}\|_{L^{\infty}(\mathcal{U})}$ is uniformly bounded when $0<\epsilon\ll 1$. In section 7 we compute the limit
(1.4) $\Delta F_{0}(\check{x}):=\lim_{\epsilon\to 0}\big{(}F_{\epsilon}^{(l)}(x_{0}+\epsilon\check{x})-F_{\epsilon}^{(l)}(x_{0})\big{)},$
where $F_{\epsilon}^{(l)}(x)$ is the leading singular term of $F_{\epsilon}(x)$ near $x_{0}$, see (7.18). In section 8 we compute the DTB function, $\Delta G_{0}$, as the limit
(1.5) $\Delta G_{0}(\check{x}):=\lim_{\epsilon\to 0}\big{(}G_{\epsilon}(x_{0}+\epsilon\check{x})-G_{\epsilon}(x_{0})\big{)},$
see equation (LABEL:rhs_lim_5). This calculation assumes that the singularity of $f$ at $x_{0}$ is visible from the data only once. This completes the proof of Theorem 3.6.
A more general formula for the DTB function when the data are redundant (i.e., the singularity is visible several times) is obtained in section 9, see equation (LABEL:upsilon_simple_mult). This completes the proof of Theorem 3.7. Results of a numerical experiment, which validate the developed theory, are presented in section 10. Finally, the proofs of all the technical lemmas are included in the appendices at the end of the paper.
2. Basic notation, function spaces, symbols, operators
Denote =$\mathbb{N}=\{1,2,\dots\}$ and $\mathbb{N}_{0}=\{0\}\cup\mathbb{N}$. Let $U\subset\mathbb{R}^{n}$, $n\in\mathbb{N}$, be a domain, which is defined as a non-empty, connected, open set. For clarity, the zero vector in $\mathbb{R}^{n}$, $n\geq 2$, is denoted $0_{n}$.
Definition 2.1.
Let $f$ and $g$ be two functions defined on a domain $U$. We say $f(x)\asymp g(x)$ for $x\in U$ if there exist $c_{1,2}>0$ such that
(2.1) $c_{1}\leq f(x)/g(x)\leq c_{2}\text{ if }g(x)\not=0\text{ and }f(x)=0\text{ if }g(x)=0$
for any $x\in U$.
We use several types of function spaces. First, $C^{k}(U)$, $k\in\mathbb{N}$, is the space of functions with bounded derivatives up to order $k$ with the norm
(2.2) $\|h\|_{C^{k}(U)}:=\max_{|m|\leq k}\|h^{(m)}\|_{L^{\infty}(U)}.$
The subscript ‘0’ in $C_{0}^{k}$ means that we consider the subspace of compactly supported functions, $C_{0}^{k}(U)\subset C^{k}(U)$. Further, $C^{\infty}(U)=\bigcap_{k\geq 1}C^{k}(U)$ and $C_{0}^{\infty}(U)=\bigcap_{k\geq 1}C_{0}^{k}(U)$.
Another type is the Hölder spaces $C^{s}(\mathbb{R}^{n})$, $s>0$, $s\not\in\mathbb{N}$. If $s=k+\gamma$, $k\in\mathbb{N}_{0}$, $10<\gamma<$ 1, then $C^{s}(\mathbb{R}^{n})$ is the space of $C^{k}(\mathbb{R}^{n})$ functions (or, $L^{\infty}(\mathbb{R}^{n})$ functions if $k=0$), which have Hölder continuous $k$-th order derivatives (see [4, Definition 5.28]), with the norm
(2.3) $\begin{split}\|f\|_{C^{s}}:=&\|f\|_{C^{k}}+\max_{|m|=k}\sup_{x\in\mathbb{R}^{n},|h|>0}\frac{|f^{(m)}(x+h)-f^{(m)}(x)|}{|h|^{\gamma}},\ k>0,\\ \|f\|_{C^{s}}:=&\|f\|_{L^{\infty}}+\sup_{x\in\mathbb{R}^{n},|h|>0}\frac{|f(x+h)-f(x)|}{|h|^{\gamma}},\ k=0.\end{split}$
As before, $C_{0}^{s}(U)$ denotes the subspace of $C^{s}(\mathbb{R}^{n})$ functions supported in $U$.
For $s>0$, $s\not\in\mathbb{N}$, Hölder spaces coincide with Zygmund spaces, $C_{*}^{s}(\mathbb{R}^{n})$. To define the latter, pick any $\mu_{0}\in C_{0}^{\infty}(\mathbb{R}^{n})$ such that $\mu_{0}(\eta)=1$ for $|\eta|\leq 1$, $\mu_{0}(\eta)=0$ for $|\eta|\geq 2$. Define $\mu_{j}(\eta):=\mu_{0}(2^{-j}\eta)-\mu_{0}(2^{-j+1}\eta)$, $j\in\mathbb{N}$ [1, Section 5.4]. Then
(2.4) $\begin{split}&C_{*}^{s}(\mathbb{R}^{n}):=\{f\in L^{\infty}(\mathbb{R}^{n}):\,\|f\|_{C_{*}^{s}(\mathbb{R}^{n})}<\infty\},\\ &\|f\|_{C_{*}^{s}(\mathbb{R}^{n})}:=\sup_{j\in\mathbb{N}_{0}}2^{js}\|\mathcal{F}^{-1}(\mu_{j}(\eta)\tilde{f}(\eta))\|_{L^{\infty}(\mathbb{R}^{n})},\end{split}$
where $\tilde{f}=\mathcal{F}_{n}f$. Here $\mathcal{F}_{n}$ is the Fourier transform in $\mathbb{R}^{n}$:
(2.5) $\tilde{f}(\xi)=\frac{1}{(2\pi)^{n}}\int_{\mathbb{R}^{n}}f(x)e^{i\xi\cdot x}\text{d}x,\ f\in L^{1}(\mathbb{R}^{n}),$
which extends to tempered distributions [20, section 7.1]. We have $C_{*}^{s}(\mathbb{R}^{n})=C^{s}(\mathbb{R}^{n})$ and the norms $\|\cdot\|_{C_{*}^{s}(\mathbb{R}^{n})}$, $\|\cdot\|_{C^{s}(\mathbb{R}^{n})}$ are equivalent, for any $s>0$, $s\not\in\mathbb{N}$ [4, Remark 5.31].
The Hölder-Zygmund spaces are a particular case of the Besov spaces: $C^{s}(\mathbb{R}^{n})=B^{s}_{p,q}(\mathbb{R}^{n})$, where $p,q=\infty$ [1, item 2 in Remark 6.4].
The Sobolev space $H^{s}(\mathbb{R}^{n})$, $s\in\mathbb{R}$, is the space of all tempered distributions $f$ for which
(2.6) $\|f\|_{H^{s}(\mathbb{R}^{n})}^{2}=(2\pi)^{-n}\int_{\mathbb{R}^{n}}|\tilde{f}(\xi)|^{2}(1+|\xi|^{2})^{s}\text{d}\xi<\infty.$
$H_{c}^{s}(U)$ is the collection of all $f\in H^{s}(\mathbb{R}^{n})$ such that $f\equiv 0$ on $\mathbb{R}^{n}\setminus K$ for some compact $K\subset U$. $H_{loc}^{s}(U)$ is the space of all distributions on $U$ such that $\chi f\in H^{s}(U)$ for any $\chi\in C_{0}^{\infty}(U)$. See [52, section 5, Notation and Background]
Definition 2.2.
Given a domain $U\subset\mathbb{R}^{n}$, $r\in\mathbb{R}$, and $N\in\mathbb{N}$, $S^{r}(U\times\mathbb{R}^{N})$ denotes the vector-space of $C^{\infty}(U\times(\mathbb{R}^{N}\setminus 0_{N}))$ functions, $\tilde{B}(x,\xi)$, having the following properties
(2.7) $\begin{split}&|\partial_{x}^{m}\tilde{B}(x,\xi)|\leq c_{m}\ \forall m\in\mathbb{N}_{0}^{n},x\in U,0<|\xi|\leq 1;\\ &|\partial_{x}^{m_{1}}\partial_{\xi}^{m_{2}}\tilde{B}(x,\xi)|\leq c_{m_{1},m_{2}}|\xi|^{r-|m_{2}|},\ \forall m_{1}\in\mathbb{N}_{0}^{n},m_{2}\in\mathbb{N}_{0}^{N},x\in U,|\xi|\geq 1;\end{split}$
for some constants $c_{m},c_{m_{1},m_{2}}>0$.
The elements of $S^{r}$ are called symbols of order $r$. When we are talking about symbols of $\Psi$DOs (in which case $N=n$), we use the notation $S^{r}(U)$. We modify the conventional definition slightly to allow for symbols to be non-smooth at the origin. This makes our analysis more streamlined, but otherwise has no effects. The space of the corresponding $\Psi$DOs, given by
(2.8) $(\text{Op}(\tilde{B}(x,\xi))f)(y):=\frac{1}{(2\pi)^{N}}\int_{\mathbb{R}^{N}}\tilde{B}(x,\xi)\tilde{f}(\xi)e^{-i\xi\cdot y}\text{d}\xi,\ f\in C_{0}^{\infty}(U),$
is denoted $L^{r}(U)$. As is known, $L^{r}(U):H_{c}^{s}(U)\to H_{loc}^{s-r}(U)$ is continuous [52, Theorem 2.1, Section I.2].
Let $U,V\subset\mathbb{R}^{n}$ be two domains, and let
(2.9) $\Phi(x,y,\Theta)\in C^{\infty}(U\times V\times(\mathbb{R}^{N}\setminus 0_{N})),\quad\Phi:U\times V\times\mathbb{R}^{N}\to\mathbb{R},$
be a nondegenerate phase function [16, Definition 11.1]. The latter property means that
  1. (1)
    $\Phi$ is positively homogeneous of degree one in $\Theta$: $\Phi(x,y,\lambda\Theta)=\lambda\Phi(x,y,\Theta)$, $\lambda>0$.
  2. (2)
    $\text{d}_{(x,y,\Theta)}\Phi$ does not vanish anywhere.
  3. (3)
    The differentials $\text{d}_{(x,y,\Theta)}(\partial\Phi/\partial\Theta_{l})$, $1\leq l\leq N$, are linearly independent on the set $\Sigma$:
    (2.10) $\Sigma:=\{(x,y,\Theta)\in U\times V\times(\mathbb{R}^{N}\setminus 0_{N}):\ \Phi_{\Theta}^{\prime}(x,y,\Theta)=0_{N}\}.$
The phase $\Phi$ determines a map $T^{*}U\to T^{*}V$ with the graph
(2.11) $\mathcal{C}:=\{(y,\text{d}_{y}\Phi(x,y,\Theta));(x,-\text{d}_{x}\Phi(x,y,\Theta)):(x,y,\Theta)\in\Sigma\}.$
$I^{r}(V\times U\times\mathcal{C})$ denotes the vector space of FIOs of order $r$ given by [16, section 11]
(2.12) $\begin{split}&C_{0}^{\infty}(U)\ni f\to\frac{1}{(2\pi)^{N}}\int_{\mathbb{R}^{N}}\int_{U}\tilde{B}(x,y,\Theta)f(x)e^{i\Phi(x,y,\Theta)}\text{d}x\text{d}\Theta\in\mathcal{D}^{\prime}(V),\\ &\tilde{B}\in S^{r+(n-N)/2}(U\times V\times\mathbb{R}^{N}).\end{split}$
For convenience, throughout the paper we use the following convention. If a constant $c$ is used in an equation, the qualifier ‘for some $c>0$’ is assumed. If several $c$ are used in a string of (in)equalities, then ‘for some’ applies to each of them, and the values of different $c$’s may all be different.
Additional notation and conventions are introduced as needed.
3. Setting of the problem, assumptions, and main results
3.1. GRT and its properties
Let $\Phi_{1}(x,y):\mathcal{U}\times\mathcal{V}\to\mathbb{R}^{2}$ be a defining function for a GRT $\mathcal{R}$. An open set $\mathcal{U}\subset\mathbb{R}^{2}$ is the image domain. We assume $y=(\alpha,p)$, and $\mathcal{V}=I_{\alpha}\times I_{p}$ is the data domain.
Assumption 3.1 (Properties of the GRT - I).
  1. (1)
    $I_{\alpha}\subset\mathbb{R}$ is a compact interval with the endpoints identified (so it is topologically equivalent to a circle),
  2. (2)
    $I_{p}\subset\mathbb{R}$ is an open interval (or all of $\mathbb{R}$), and
  3. (3)
    $\Phi_{1}(x,y)=p-\Phi(x,\alpha)$, $\Phi\in C^{\infty}(\mathcal{U}\times I_{\alpha})$.
Both $\mathcal{U}$ and $\mathcal{V}$ are endowed with the usual Euclidean metric. By a partition of unity we will always identify subsets of $\mathcal{V}$ with subsets of $\mathbb{R}^{2}$. The transform $\mathcal{R}$ integrates over smooth curves $\mathcal{S}_{(\alpha,p)}:=\{x\in\mathcal{U}:p=\Phi(x,\alpha)\}$:
(3.1) $\begin{split}\mathcal{R}f(y)=&\int_{\mathcal{U}}f(x)W(x,y)\delta(p-\Phi(x,\alpha))\text{d}x\\ =&\frac{1}{2\pi}\int_{\mathbb{R}}\int_{\mathcal{U}}f(x)W(x,y)e^{i\nu(p-\Phi(x,\alpha))}\text{d}x\text{d}\nu,\ y\in\mathcal{V},\end{split}$
where $W\in C^{\infty}(\mathcal{U}\times\mathcal{V})$.
Here and everywhere below we use the convention that whenever $y$, $\alpha$, and $p$ appear in the same equation or sentence, then $y=(\alpha,p)$. We assume $f$ is compactly supported, $\text{supp}f\subset\mathcal{U}$, and $f$ is sufficiently smooth, so $g(y):=\mathcal{R}f(y)$ is a continuous function. Assumption 3.1 implies that $\text{supp}\,(\mathcal{R}f)$ is compact if $\text{supp}\,f$ is compact.
The critical set of the phase $\nu\Phi_{1}$ is
(3.2) $\Sigma:=\{(y,x,\nu)\in\mathcal{V}\times\mathcal{U}\times\mathbb{R}:\ p-\Phi(x,\alpha)=0\}.$
Since $\partial_{\nu}(\nu\Phi_{1})=\Phi_{1}$ and $\text{d}_{(x,y,\nu)}\Phi_{1}\not=0$, the phase $\nu\Phi_{1}(x,\alpha)$ is clean and nondegenerate [21, Definition 21.2.15]. The phase $\nu\Phi_{1}(x,\alpha)$ determines a homogeneous canonical relation from $T^{*}\mathcal{U}$ to $T^{*}\mathcal{V}$ with the graph
(3.3) $\begin{split}\mathcal{C}:=&\{(y,\nu(-\Phi_{\alpha}^{\prime}(x,\alpha),1));(x,\nu\Phi_{x}^{\prime}(x,y)):(y,x,\nu)\in\Sigma\}.\end{split}$
Also, $\mathcal{R}\in I^{-1/2}(\mathcal{V}\times\mathcal{U},\mathcal{C})$ is an FIO of order $-1/2$ from $\mathcal{U}$ to $\mathcal{V}$ associated with $\mathcal{C}$ [53, Section 5.VIII]. Hence $\mathcal{R}:H_{c}^{s}(\mathcal{U})\to H_{loc}^{s+(1/2)}(\mathcal{V})$ is continuous for any $s\in\mathbb{R}$ [22, Corollary 25.3.2].
Next we compute the following determinant [53, equations (6.3) and (6.4), Chapter VIII]:
(3.4) $\begin{split}\Delta_{\Phi}(x,y,\nu):=&\text{det}\begin{pmatrix}[\nu\Phi_{1}(x,\alpha)]_{xy}^{\prime\prime}&[\nu\Phi_{1}(x,\alpha)]_{x\nu}^{\prime\prime}\\ [\nu\Phi_{1}(x,\alpha)]_{\nu y}^{\prime\prime}&[\nu\Phi_{1}(x,\alpha)]_{\nu\nu}^{\prime\prime}\end{pmatrix}\\ =&\text{det}\begin{pmatrix}-\nu\Phi_{x\alpha}^{\prime\prime}&(0,0)^{T}&-\Phi_{x}^{\prime}\\ -\Phi_{\alpha}^{\prime}&1&0\end{pmatrix}\\ =&\nu\text{det}\begin{pmatrix}\text{d}_{x}\Phi&\text{d}_{x}(\partial_{\alpha}\Phi)\end{pmatrix}.\end{split}$
Here $\Phi_{x}^{\prime}=(\partial_{x_{1}}\Phi,\partial_{x_{2}}\Phi)^{T}$, and all the other derivatives are defined similarly. The derivatives on the right are computed at $(x,\alpha,p=\Phi(x,\alpha))$. As is easily seen,
(3.5) $\Delta_{\Phi}(x,y,\nu)=\nu\Delta_{\Phi}(x,\alpha).$
Since the determinant on the last line in (3.4) does not depend on $p$ and $\nu$, we use the simpler notation $\Delta_{\Phi}(x,\alpha)$ instead of $\Delta_{\Phi}(x,y,1)$.
Assumption 3.2 (Properties of the GRT - II).
  1. (1)
    $\text{d}_{x}\Phi(x,\alpha)$ and $\text{d}_{x}\Phi_{\alpha}^{\prime}(x,\alpha)$ are linearly independent on $\mathcal{U}\times\ I_{\alpha}$.
  2. (2)
    The Bolker condition: if $\Phi(x_{1},\alpha)=\Phi(x_{2},\alpha)$ for some $x_{1},x_{2}\in\mathcal{U}$, $x_{1}\not=x_{2}$, and $\alpha\in I_{\alpha}$, then $\Phi_{\alpha}^{\prime}(x_{1},\alpha)\not=\Phi_{\alpha}^{\prime}(x_{2},\alpha)$.
  3. (3)
    For each $(x,\xi)\in T^{*}\mathcal{U}$, there exist $\alpha\in I_{\alpha}$ and $\nu\in\mathbb{R}$ such that $\xi=\nu\text{d}_{x}\Phi(x,\alpha)$.
  4. (4)
    $W(x,y)>c$ on $\mathcal{U}\times\mathcal{V}$.
  5. (5)
    For any bounded domain $\mathcal{U}_{b}\subset\mathcal{U}$ there exists $c=c(\mathcal{U}_{b})>0$ such that
    (3.6) $\|f\|_{H^{-1/2}(\mathcal{U})}\leq c\|\mathcal{R}f\|_{L^{2}(\mathcal{V})}\text{for any }f\in H_{c}^{-1/2}(\mathcal{U}_{b}).$
Assumption 3.2(1) implies that $\Delta_{\Phi}(x,y,\nu)\not=0$ on $\mathcal{U}\times\mathcal{V}\times(\mathbb{R}\setminus 0)$. In turn, this implies that $\mathcal{C}$ is the graph of a local diffeomorphism [53, Section VI.4]. Assumption 3.2(2) implies that the natural projection $\mathcal{C}\to T^{*}\mathcal{V}$ is an embedding (injective immersion). Assumption 3.2(3) asserts that any singularity of $f$ is visible from the data. Assumption 3.2(5) implies that $\mathcal{R}$ is boundedly invertible in the scale of Sobolev spaces.
3.2. The function to be reconstructed
Suppose a compactly supported distribution, $f\in\mathcal{E}^{\prime}(\mathcal{U})$, is given by
(3.7) $f(x)=\frac{1}{2\pi}\int_{\mathbb{R}}\tilde{f}(x,\lambda)e^{i\lambda H(x)}\text{d}\lambda,\ x\in\mathcal{U},$
where
(3.8) $\begin{split}&H\in C^{\infty}(\mathcal{U}),\quad\text{d}H\not=0\text{ on }\mathcal{U};\\ &\tilde{f}(x,\lambda)=-i\tilde{f}_{0}(x)\lambda^{-1}+\tilde{R}(x,\lambda),\forall x\in\mathcal{U},|\lambda|\geq 1;\\ &\tilde{f}(x,\lambda)\equiv 0\ \forall x\in\mathcal{U}\setminus\mathcal{U}_{b}\text{ for some bounded domain }\mathcal{U}_{b}\subset\mathcal{U};\\ &\tilde{f}\in S^{-1}(\mathcal{U}\times\mathbb{R}),\ \tilde{f}_{0}\in C_{0}^{\infty}(\mathcal{U}_{b}),\ \tilde{R}\in S^{-2}(\mathcal{U}\times\mathbb{R}),\end{split}$
for some $\tilde{f}_{0}$ and $\tilde{R}$. From (3.7) and (3.8) it follows that
(3.9) sing $\text{sing\hskip 1.0ptsupp}(f)\subset\mathcal{S}:=\{x\in\mathcal{U}:\ H(x)=0\}.$
Thus $\mathcal{S}$ is a smooth curve (submanifold) with possibly several connected components.
Equations (3.7) and (3.8) imply that $f\in I^{-1}(\mathcal{U},\mathcal{S})$ is a conormal distribution [21, Section 18.2] and
(3.10) $\begin{split}&f(x+\Delta x)\sim\tilde{f}_{0}(x)(\text{sgn}h/2),\ h:=\text{d}H(x)\Delta x,|\Delta x|\to 0,\ \forall x\in\mathcal{S}.\end{split}$
Thus, $\tilde{f}_{0}(x)$ is the value of the jump of $f$ across $\mathcal{S}$ at $x\in\mathcal{S}$.
Pick any $x\in\mathcal{S}$ and $y\in\mathcal{V}$ such that $\mathcal{S}_{y}$ is tangent to $\mathcal{S}$ at $x$. The curvatures of $\mathcal{S}$ and $\mathcal{S}_{y}$ at $x$ are given by
(3.11) $\varkappa_{\mathcal{S}}(x)=-\frac{(H_{xx}^{\prime\prime}(x)e,e)}{|H_{x}^{\prime}(x)|},\ \varkappa_{\mathcal{S}_{y}}(x)=-\frac{(\Phi_{xx}^{\prime\prime}(x,\alpha)e,e)}{|\Phi_{x}^{\prime}(x,\alpha)|},$
where $e$ is a unit vector tangent to $\mathcal{S}$ at $x$. If $H_{x}^{\prime}(x)\cdot\Phi_{x}^{\prime}(x,\alpha)<0$, the curvatures are not consistent with each other. In this case we flip the $p$-axis and replace $\Phi$ with $-\Phi$ to make them consistent. Thus, we can assume without loss of generality that $H_{x}^{\prime}(x)\cdot\Phi_{x}^{\prime}(x,\alpha)>0$.
Assumption 3.3 (Properties of $f$).
  1. (1)
    $f$ satisfies (3.7), (3.8).
  2. (2)
    For each pair $(x,y)\in\mathcal{S}\times\mathcal{V}$ such that $\mathcal{S}_{y}$ is tangent to $\mathcal{S}$ at $x$, one has $\varkappa_{\mathcal{S}}(x)-\varkappa_{\mathcal{S}_{y}}(x)>0$.
The last condition is not restrictive. Basically, it says that $\varkappa_{\mathcal{S}}(x)\not=\varkappa_{\mathcal{S}_{y}}(x)$ whenever $\mathcal{S}$ and $\mathcal{S}_{y}$ are tangent at some $x$. If $\varkappa_{\mathcal{S}}(x)-\varkappa_{\mathcal{S}_{y}}(x)<0$, we change both functions: replace $H$ with $-H$, $\Phi$ with $-\Phi$, and flip the $p$-axis.
It may happen that for some $y\in\mathcal{V}$, $\mathcal{S}_{y}$ is tangent to $\mathcal{S}$ at several distinct points $x_{l}$, and $\Phi$ cannot be made consistent with $H$ in the above sense at all $x_{l}$. Due to the linearity of the GRT, we can use a partition of unity and, for each $l$, adjust $\Phi$ and $H$ based solely on the pair $(x_{l},y)$ independently of the adjustments at all the others pairs. This is always tacitly assumed in what follows.
3.3. GRT data and the reconstruction algorithm
Discrete data $g(y_{j})$ are given at the points
(3.12) $y_{j}=(\Delta\alpha j_{1},\Delta pj_{2})\in\mathcal{V},\ \Delta p=\epsilon,\ \Delta\alpha=\mu\epsilon,\ j=(j_{1},j_{2})\in\mathbb{Z}^{2},$
The interpolated data, denoted $g_{\epsilon}(y)$, is computed by:
(3.13) $\begin{split}&g_{\epsilon}(y):=\sum_{j}\varphi_{\epsilon}(y-y_{j})g(y_{j}),\ y\in\mathcal{V},\\ &\varphi_{\epsilon}(y):=\varphi_{\alpha}(\alpha/\Delta\alpha)\varphi_{p}(p/\Delta p),\end{split}$
where $\varphi$ is an interpolation kernel.
Assumption 3.4 (Properties of the interpolation kernel, $\varphi$).
Let $\varphi_{*}$ denote any of the functions $\varphi_{\alpha}$ and $\varphi_{p}$. One has
  1. (1)
    $\varphi_{*}\in C_{0}^{2}(\mathbb{R})$;
  2. (2)
    $\varphi_{*}$ is even: $\varphi_{*}(u)=\varphi_{*}(-u)$, $u\in\mathbb{R}$.
  3. (3)
    $\varphi_{*}$ is exact up to order one, i.e.
    (3.14) $\sum_{j\in\mathbb{Z}}j^{m}\varphi_{*}(u-j)\equiv u^{m},\ m=0,1,\ u\in\mathbb{R}.$
Assumption 3.4(2) with $m=0$ implies that $\varphi_{*}$ is normalized:
(3.15) $\begin{split}1&=\int_{0}^{1}\sum_{j\in\mathbb{Z}}\varphi_{*}(u-j)\text{d}u=\int_{\mathbb{R}}\varphi_{*}(u)\text{d}u.\end{split}$
Let $f$ satisfy assumption 3.3 and $\mathcal{R}$ satisfy assumptions 3.1, 3.2. Lemma 4.2 below implies $\|g_{\epsilon}\|_{L^{2}(\mathcal{V})}\leq c$, $0<\epsilon\ll 1$. Reconstruction is achieved by minimizing the quadratic functional
(3.16) $f_{\epsilon}=\text{argmin}_{f\in H_{c}^{1}(\mathcal{U}_{b})}\Psi(f),\ \Psi(f):=\|\mathcal{R}f-g_{\epsilon}\|_{L^{2}(\mathcal{V})}^{2}+\kappa\epsilon^{3}\|\nabla f\|_{L^{2}(\mathcal{U})}^{2},$
with some fixed regularization parameter $\kappa>0$ and with $\mathcal{U}_{b}$ the same as in (3.8) (or any other bounded domain containing $\mathcal{U}_{b}$).
The following result is proven in appendix A.
Lemma 3.5.
Suppose assumptions 3.13.4 are satisfied. The solution to (3.16), $f_{\epsilon}$, exists and is unique for each $0<\epsilon\ll 1$.
Our first main result is the following theorem, which is proven in sections 48.
Theorem 3.6.
Pick any $x_{0}\in\mathcal{S}$ and $A>0$. Suppose there is only one $y_{0}=(\alpha_{0},p_{0})\in\mathcal{V}$ such that $\mathcal{S}_{y_{0}}$ is tangent to $\mathcal{S}$ at $x_{0}$. Suppose assumptions 3.13.4 are satisfied and $\mu\Phi_{\alpha}^{\prime}(x_{0},\alpha_{0})$ is irrational. One has
(3.17) $\lim_{\epsilon\to 0}\big{(}f_{\epsilon}(x_{0}+\epsilon\check{x})-f_{\epsilon}(x_{0})\big{)}=\tilde{f}_{0}(x_{0})\Upsilon\big{(}\vec{\Theta}_{0}\cdot\check{x}\big{)},\ |\check{x}|<A,$
where the limit is uniform in $\check{x}$ and
(3.18) $\begin{split}&\Upsilon(r)=\int_{\mathbb{R}}h_{0}(u)\int_{u_{1}}^{u_{1}+r}R(t)\text{d}t\text{d}u,\ \vec{\Theta}_{0}:=\frac{\Phi_{x}^{\prime}(x_{0},\alpha_{0})}{|\Phi_{x}^{\prime}(x_{0},\alpha_{0})|},\\ &h_{0}(u)=\int_{\mathbb{R}}\varphi_{\alpha}(s)\varphi_{p}(\mu s\Phi_{\alpha}^{\prime}(x_{0},\alpha_{0})+u)\text{d}s,\ u_{1}=\frac{u}{|\Phi_{x}^{\prime}(x_{0},\alpha_{0})|},\\ &R(t)=\mathcal{F}_{\lambda\to t}^{-1}\bigg{(}\bigg{[}1+\frac{\kappa}{2\pi\rho_{0}}|\lambda|^{3}\bigg{]}^{-1}\bigg{)},\ \rho_{0}=\frac{W^{2}(x_{0},y_{0})|\Phi_{x}^{\prime}(x_{0},\alpha_{0})|}{|\Delta_{\Phi}(x_{0},\alpha_{0})|}.\end{split}$
Moreover,
(3.19) $\Upsilon(0)=0,\quad\Upsilon(\pm\infty)=\pm 1/2.$
Thus $\Upsilon(+\infty)-\Upsilon(-\infty)=1$, and $\Upsilon$ is a true DTB function.
In the case of redundant data we obtain the following generalization, which is proven in section 9.
Theorem 3.7.
Pick any $x_{0}\in\mathcal{S}$ and $A>0$. Suppose there exist finitely many $y_{l}=(\alpha_{l},p_{l})\in\mathcal{V}$ such that $\mathcal{S}_{y_{l}}$ is tangent to $\mathcal{S}$ at $x_{0}$. Suppose assumptions 3.13.4 are satisfied and all $\mu\Phi_{\alpha}^{\prime}(x_{0},\alpha_{l})$ are irrational. One has
(3.20) $\lim_{\epsilon\to 0}\big{(}f_{\epsilon}(x_{0}+\epsilon\check{x})-f_{\epsilon}(x_{0})\big{)}=\tilde{f}_{0}(x_{0})\frac{\textstyle\sum_{l}\rho_{l}\Upsilon_{l}\big{(}\vec{\Theta}_{0}\cdot\check{x}\big{)}}{\textstyle\sum_{l}\rho_{l}},\ |\check{x}|<A,$
where the limit is uniform in $\check{x}$ and
(3.21) $\begin{split}&\Upsilon_{l}(r)=\int_{\mathbb{R}}h_{l}(u)\int_{u_{l}}^{u_{l}+r}R(t)\text{d}t\text{d}u,\ \vec{\Theta}_{0}:=\frac{H_{x}^{\prime}(x_{0})}{|H_{x}^{\prime}(x_{0})|},\\ &h_{l}(u)=\int_{\mathbb{R}}\varphi_{\alpha}(s)\varphi_{p}(\mu s\Phi_{\alpha}^{\prime}(x_{0},\alpha_{l})+u)\text{d}s,\ u_{l}=\frac{u}{|\Phi_{x}^{\prime}(x_{0},\alpha_{l})|},\\ &R(t)=\mathcal{F}_{\lambda\to t}^{-1}\bigg{(}\bigg{[}1+\frac{\kappa}{2\pi\textstyle\sum_{l}\rho_{l}}|\lambda|^{3}\bigg{]}^{-1}\bigg{)},\ \rho_{l}:=\frac{W^{2}(x_{0},y_{l})|\Phi_{x}^{\prime}(x_{0},\alpha_{l})|}{|\Delta_{\Phi}(x_{0},\alpha_{l})|}.\end{split}$
4. Behavior of $\mathcal{R}f$ near its singular support
Let $\Gamma$ be the set of all $y\in\mathcal{V}$ such that $\mathcal{S}_{y}$ is tangent to $\mathcal{S}$. Since $\mathcal{R}$ is an FIO with the canonical relation (3.3), we have
(4.1) $\Gamma=\text{sing\hskip 1.0ptsupp}(\mathcal{R}f),\ WF(\mathcal{R}f)\subset N^{*}\Gamma=\mathcal{C}\circ N^{*}\mathcal{S}.$
Here $N^{*}\mathcal{S}$ is the conormal bundle of $\mathcal{S}$, and similarly for $\Gamma$ [52, section I.6]. In particular, $\mathcal{R}f$ is smooth away from $\Gamma$. In this section we describe the leading singular behavior of $\mathcal{R}f$ near $\Gamma$.
Using a partition of unity and the linearity of $\mathcal{R}$, we can assume without loss of generality that $\mathcal{S}_{y}$, $y\in\Gamma$, is tangent to $\mathcal{S}$ only at one point.
Fix any $\tilde{y}\in\Gamma$, and let $\tilde{x}\in\mathcal{S}$ be the corresponding point of tangency. Let $V$ be a sufficiently small neighborhood of $\tilde{y}$. Substitute (3.7) into (3.1). The asymptotics as $\lambda\to\infty$ of the resulting integral with respect to $x$ is computed by the stationary phase method [53, Chapter VIII, eqs. (2.14)–(2.20)]
(4.2) $\begin{split}(\mathcal{R}f)(y)=&\frac{1}{2\pi}\int_{\mathbb{R}}\tilde{G}(y,\lambda)\text{d}\lambda,\ y\in V,\\ \tilde{G}(y,\lambda):=&\frac{1}{2\pi}\int_{\mathbb{R}}\int_{\mathcal{U}}\tilde{f}(x,\lambda)e^{i\lambda H(x)}W(x,y)e^{i\nu(p-\Phi(x,\alpha))}\text{d}x\text{d}\nu\\ =&\frac{|\lambda|}{2\pi}\int_{\mathbb{R}}\int_{\mathcal{U}}W(x,y)\tilde{f}(x,\lambda)e^{i\lambda[\nu_{1}(p-\Phi(x,\alpha))+H(x)]}\text{d}x\text{d}\nu_{1}\\ =&\left(\tilde{f}(x_{*},\lambda)\frac{W(x_{*},y)}{|\det M(y)|^{1/2}}\left(\frac{2\pi}{|\lambda|}\right)^{1/2}e(-\text{sgn}\lambda/2)+\tilde{R}(y,\lambda)\right)\\ &\quad\times e^{i\lambda H(x_{*})},\ |\lambda|\geq 1,\ \tilde{R}\in S^{-3/2}(\mathcal{V}\times\mathbb{R}),\ x_{*}=x_{*}(y),\end{split}$
for some $\tilde{R}$ and $0<\delta\ll 1$. Here $e(t):=e(i\pi t/2)$, and
(4.3) $M(y)=\begin{pmatrix}0&(-\Phi_{x}^{\prime}(x,\alpha))^{T}\\ -\Phi_{x}^{\prime}(x,\alpha)&H_{xx}^{\prime\prime}(x,\alpha)-\nu_{1}\Phi_{xx}^{\prime\prime}(x,\alpha)\end{pmatrix},\ \nu_{1}=\nu_{1*}(y),x=x_{*}(y),$
is the Hessian matrix of the phase at the stationary point $(\nu_{1*}(y),x_{*}(y))$, which is found by solving
(4.4) $p=\Phi(x,\alpha),\ \text{d}_{x}H(x)-\nu_{1}\text{d}_{x}\Phi(x,\alpha)=0,\ y\in V.$
By construction, $x_{*}(\tilde{y})=\tilde{x}$. As before, the existence and local uniqueness of the solution follows from assumptions 3.2(1,3).
The following result is proven in appendix B.1.
Lemma 4.1.
Suppose assumptions 3.13.3 are satisfied. Let $\mathcal{S}_{y}$ be tangent to $\mathcal{S}$ at $x\in\mathcal{S}$. Let $\varkappa_{\mathcal{S}}(x)$ and $\varkappa_{\mathcal{S}_{y}}(x)$ denote the curvatures of $\mathcal{S}$ and $\mathcal{S}_{y}$ at $x$, respectively (see (3.11)). One has
(4.5) $\begin{split}\det M(y)=&(\varkappa_{\mathcal{S}}(x)-\varkappa_{\mathcal{S}_{y}}(x)){|\Phi_{x}^{\prime}(x,\alpha)|^{2}}{|H_{x}^{\prime}(x)|}>0,\\ \text{sgn}M(y)=&-1.\end{split}$
It follows from the proof of Lemma 4.1 (see (B.1)) that we can locally solve $H(x_{*}(\alpha,p))=0$ for $p$ in terms of $\alpha$: $p=P(\alpha)$, where $P(\alpha)$ is locally smooth. Therefore
(4.6) $H(x_{*}(y))=\psi(y)(p-P(\alpha)),\ y\in V,$
where $\psi(y)\not=0$, $y\in V$, and $\psi$ is locally smooth. Moreover, by (B.1) and (4.6),
(4.7) $\begin{split}\psi(\alpha,P(\alpha))=&\partial_{p}H(x_{*}(\alpha,p))|_{p=P(\alpha)}=\frac{|H_{x}^{\prime}(x_{*}(y))|}{|\Phi_{x}^{\prime}(x_{*}(y),\alpha)|}>0,\\ (\alpha,P(\alpha))\in&V.\end{split}$
Therefore, $\psi(y)>0$, $y\in V$.
It is clear that instead of finding $P(\alpha)$ through the intermediary function $x_{*}(y)$, we can find $P(\alpha)$ by solving
(4.8) $H(x(\alpha))=0,\ H_{x}^{\prime}(x(\alpha))=\nu_{1}(\alpha)\Phi_{x}^{\prime}(x(\alpha),\alpha),\ (\alpha,P(\alpha))\in V,$
for $\nu_{1}(\alpha)$, $x(\alpha)$ and setting $P(\alpha):=\Phi(x(\alpha),\alpha)$. Then $\nu_{1}(\alpha)\equiv\nu_{1*}(\alpha,P(\alpha))$. Since $x(\alpha)\in\mathcal{S}$, we can say that $x(\alpha)$ and $P(\alpha)$ are associated with a segment of $\mathcal{S}$ (see Figure 1). The size of the segment depends on the domain of $x(\alpha)$. If we have in mind one specific $\tilde{x}\in\mathcal{S}$, then we informally say that $x(\alpha)$ and $P(\alpha)$ are associated with $\tilde{x}$ (i.e., with a segment of $\mathcal{S}$ containing $\tilde{x}$).
Figure 1. Illustration of $x(\alpha)$. In the figure, $y_{k}=(\alpha_{k},P(\alpha_{k}))$, $k=1,2$ and $\tilde{y}=(\tilde{\alpha},P(\tilde{\alpha}))$.
By (4.2)–(4.6), $\mathcal{R}f$ can be written as
(4.9) $(\mathcal{R}f)(y)=\frac{1}{2\pi}\int_{\mathbb{R}}\tilde{\upsilon}(y,\lambda)e^{i\lambda(p-P(\alpha))}\text{d}\lambda,\ y\in V,$
and, with some $c_{m}$, $\tilde{R}$,
(4.10) $\begin{split}&\tilde{\upsilon}(y,\lambda)=\tilde{\upsilon}_{0}(y)e(-3/2)(\lambda-i0)^{-3/2}+\tilde{R}(y,\lambda),\ |\lambda|\geq 1;\\ &\tilde{\upsilon}\in S^{-3/2}(V\times\mathbb{R});\ \tilde{R}\in S^{-5/2}(V\times\mathbb{R});\\ &\tilde{\upsilon}_{0}(y)=(2\pi)^{1/2}\frac{W(x_{*}(y),y)\tilde{f}_{0}(x_{*}(y))\psi(y)^{1/2}}{|\det M(y)|^{1/2}}\in C^{\infty}(V).\end{split}$
The function $\psi(y)$ is absorbed into the frequency variable $\lambda$ leading to the relation
(4.11) $\tilde{\upsilon}(y,\lambda)=\tilde{G}(y,\lambda/\psi(y))/\psi(y),$
where $\tilde{G}$ is defined in (4.2). To obtain the formula for $\tilde{\upsilon}_{0}$ we use that the leading order term in $\tilde{\upsilon}$ is homogeneous of degree $-3/2$ and the distribution $(\lambda-i0)^{a}$ [14, equations (28) and (30), p. 336].
By (4.9),
(4.12) sing $\text{sing\hskip 1.0ptsupp}(\mathcal{R}f)\cap V\subset\Gamma:=\{(\alpha,p)\in V:p=P(\alpha)\}.$
All of the above properties of $\mathcal{R}f$ follow from the FIO calculus [53, Section VIII.5], but for our purposes it is more convenient to provide their explicit derivation (4.2)–(4.11).
By appealing to a partition of unity and the linearity of the GRT, the following global result is proven in appendix B.2.
Lemma 4.2.
Suppose assumptions 3.13.4 are satisfied. Let $g=\mathcal{R}f$. Then $g,g_{\epsilon}\in C_{0}^{1/2}(\mathcal{V})$ and $\|g_{\epsilon}\|_{C^{1/2}(\mathcal{V})}<c$, $0<\epsilon\ll 1$. Also,
(4.13) $\begin{split}&g(y)=\textstyle\sum_{l}a_{l}(y)(p-P_{l}(\alpha))_{+}^{1/2}+\Delta g(y),\\ &a_{l}(y)=4\pi^{1/2}\tilde{\upsilon}_{0,l}(y),\ a_{l}\in C_{0}^{\infty}(\mathcal{V}),\ \Delta g\in C_{0}^{3/2}(\mathcal{V}),\end{split}$
and $\|\Delta g_{\epsilon}\|_{C^{3/2}(\mathcal{V})}<c$ for all $0<\epsilon\ll 1$. Here, for each $l$,
  1. (1)
    $\text{supp}\,a_{l}$ is contained in a small neighborhood of some $y_{l}\in\Gamma$;
  2. (2)
    $p=P_{l}(\alpha)$ is a local solution to $H(x_{*}(\alpha,p))=0$ for $p$ in terms of $\alpha$, which is defined for all $\alpha$ such that $(\alpha,P_{l}(\alpha))\in\text{supp}\,a_{l}$; and
  3. (3)
    $\tilde{\upsilon}_{0,l}(y)$ is computed as in (4.10).
5. Regularized solution
5.1. Equation for the solution to (3.16), $f_{\epsilon}$
Let $\mathcal{R}^{*}$ be the adjoint of $\mathcal{R}:L^{2}(\mathcal{U})\to L^{2}(\mathcal{V})$:
(5.1) $\begin{split}(\mathcal{R}^{*}g)(x)=&\int_{\mathcal{V}}g(y)W(x,y)\delta(p-\Phi(x,\alpha))\text{d}y,\ x\in\mathcal{U}.\end{split}$
As is well-known, assumption 3.2 ensures that $\mathcal{R}^{*}\mathcal{R}$ is an elliptic pseudo-differential operator $(\Psi$DO) [46, 19], [53, Section VIII.6.2] of order one in $\mathcal{U}$, i.e. $\mathcal{R}^{*}\mathcal{R}\in L^{1}(\mathcal{U})$.
Clearly, the functional $\Psi$ in (3.16) is Fréchet-differentiable. Indeed, one has
(5.2) $\begin{split}&\Psi(f+h)-\Psi(f)=\Psi^{\prime}(f;h)+O(\|h\|_{H^{1}(\mathcal{U})}^{2}),\ f,h\in H_{c}^{1}(\mathcal{U}_{b}),\\ &(1/2)\Psi^{\prime}(f;h)=(\mathcal{R}f-g_{\epsilon},\mathcal{R}h)_{L^{2}(\mathcal{V})}+\kappa\epsilon^{3}(\nabla f,\nabla h)_{L^{2}(\mathcal{U})},\end{split}$
and the functional $H_{c}^{1}(\mathcal{U}_{b})\ni h\to\Psi^{\prime}(f;h)$ is linear and continuous [41, p. 14]. Recall that $g_{\epsilon}$ is defined in (3.13). Then the subgradient of $\Psi$ coincides with the gradient [41, Proposition 3.20], and $f$ is the minimizer if and only if the gradient equals zero [41, Theorem 3.24]. The facts that $\Psi$ is proper and convex, which are required for the latter conclusion, are established in the proof of Lemma 3.5 in appendix A. Therefore the solution to (3.16) is the unique solution to the following equation, which is equivalent to the first order optimality condition:
(5.3) $(\mathcal{R}^{*}\mathcal{R}-\kappa\epsilon^{3}\Delta)f_{\epsilon}=\mathcal{R}^{*}g_{\epsilon},\ x\in\mathcal{U}_{b},\ f_{\epsilon}\in H_{c}^{1}(\mathcal{U}_{b}).$
This follows by setting $\Psi^{\prime}(f_{\epsilon};h)\equiv 0$, $h\in C_{0}^{\infty}(\mathcal{U}_{b})$, integrating by parts, and using that $C_{0}^{\infty}(\mathcal{U}_{b})$ is dense in $H_{c}^{1}(\mathcal{U}_{b})$.
Lemma 5.1.
Suppose assumptions 3.13.4 are satisfied, and $f_{\epsilon}$ is obtained by solving (3.16) or, equivalently, (5.3). One has
[…]
Proof.
Since $\Psi(f\equiv 0)=\|g_{\epsilon}\|_{L^{2}(\mathcal{V})}^{2}$ (cf. (3.16)), the solution $f_{\epsilon}$ satisfies $\|\mathcal{R}f_{\epsilon}\|_{L^{2}(\mathcal{V})}\leq 2\|g_{\epsilon}\|_{L^{2}(\mathcal{V})}<c$, $0<\epsilon\ll 1$ (cf. the paragraph above (3.16)). By construction, $f_{\epsilon}\in H_{c}^{1}(\mathcal{U}_{b})\subset H_{c}^{-1/2}(\mathcal{U}_{b})$, where $\mathcal{U}_{b}$ is bounded. Application of (3.6) completes the proof. ∎
Let $\tilde{D}$ be the complete symbol of $\mathcal{R}^{*}\mathcal{R}$, i.e. $\mathcal{R}^{*}\mathcal{R}=\text{Op}(\tilde{D}(x,\xi))$ (see [16, Theorem 3.4]). Rewrite (5.3) in an alternative form:
(5.5) $\begin{split}\big{(}\text{Op}(\tilde{D}(x,\xi)+\kappa\epsilon^{3}|\xi|^{2})f_{\epsilon}\big{)}(x)=(\mathcal{R}^{*}g_{\epsilon})(x),\ x\in\mathcal{U}_{b}.\end{split}$
Let $\mathcal{U}_{0}$ be a sufficiently small neighborhood of $x_{0}\in\mathcal{U}_{b}$. Fix any $\chi\in C_{0}^{\infty}(\mathcal{U}_{b})$ such that $\chi(x)\equiv 1$ in a neighborhood of the closure of $\mathcal{U}_{0}$ (see Figure 4 in appendix C). Apply $\text{Op}(\chi(x)|\xi|)\chi(x)$ on both sides of (5.5):
(5.6) $\begin{split}&\text{Op}(\chi(x)|\xi|)\chi(x)\text{Op}\big{(}\tilde{D}(x,\xi)+\kappa\epsilon^{3}|\xi|^{2}\big{)}f_{\epsilon}=\text{Op}(\chi(x)|\xi|)\chi(x)\mathcal{R}^{*}g_{\epsilon}=:F_{\epsilon}.\end{split}$
Let $\tilde{D}_{0}(x,\xi):=\lim_{\lambda\to\infty}|\lambda\xi|\tilde{D}(x,\lambda\xi)$. The limit exists and $\tilde{D}_{0}$ is homogeneous of degree 0 in $\xi$ (see (C.6) below). The following lemma is proven in appendix C.
Lemma 5.2.
Suppose assumptions 3.13.4 are satisfied. Define
(5.7) $G_{\epsilon}:=\text{Op}\big{(}\big{[}\tilde{D}_{0}(x,\xi)+\kappa|\epsilon\xi|^{3}\big{]}^{-1}\big{)}F_{\epsilon}.$
Fix any $A>0$. One has:
[…]
6. Proving the boundedness of $F_{\epsilon}$
In this section we prove the following key result, which is used in the proof of Lemma 5.2.
Lemma 6.1.
Suppose assumptions 3.13.4 are satisfied. Then $\|F_{\epsilon}\|_{L^{\infty}(\mathcal{U})}\leq c$ for all $0<\epsilon\ll 1$.
6.1. Preliminary calculations
Our goal is to show that the right-hand side of (5.6) is bounded. Rewrite (5.6) as follows.
(6.1) $\begin{split}F_{\epsilon}(x)&=(\mathcal{T}_{1}g_{\epsilon})(x),\ \mathcal{T}_{1}:=\text{Op}(\chi(x)|\xi|)\chi(z)\mathcal{R}^{*},\\ (\mathcal{T}_{1}g_{\epsilon})(x)&=\frac{\chi(x)}{(2\pi)^{3}}\int_{\mathbb{R}^{2}}\int_{\mathcal{U}}|\xi|e^{i\xi\cdot(z-x)}\chi(z)\\ &\hskip 28.45274pt\times\frac{1}{2\pi}\int_{\mathbb{R}}\int_{\mathcal{V}}W(z,y)e^{i\nu(p-\Phi(z,\alpha))}g_{\epsilon}(y)\text{d}y\text{d}\nu\text{d}z\text{d}\xi.\end{split}$
Clearly, $\mathcal{T}_{1}\in I^{1/2}(\mathcal{U}\times\mathcal{V}\times\mathcal{C}^{*})$. Here the canonical relation $\mathcal{C}^{*}$ is the image of $\mathcal{C}$ (see (3.3)) under the exchange of $T^{*}\mathcal{U}$ and $T^{*}\mathcal{V}$. Next, consider the expression
(6.2) $\begin{split}\frac{\chi(x)}{(2\pi)^{3}}\int_{\mathbb{R}^{2}}\int_{\mathcal{U}}\chi(z)W(z,y)|\xi|e^{i\xi\cdot(z-x)-i\nu(\Phi(z,\alpha)-\Phi(x,\alpha))}\text{d}z\text{d}\xi.\end{split}$
Upon changing variables $\xi\to\xi_{1}=\xi/\nu$, the stationary point of the phase satisfies $z=x$, $\xi_{1}=\Phi_{z}^{\prime}(z,\alpha)$, and the stationary point is nondegenerate. By the stationary phase method (see also [53, Ch. VI, eqs. (4.1)–(4.4)]),
(6.3) $\mathcal{T}_{1}=\mathcal{T}_{11}+\mathcal{T}_{12},\ \mathcal{T}_{11}\in I^{1/2}(\mathcal{U}\times\mathcal{V}\times\mathcal{C}^{*}),\ \mathcal{T}_{12}\in I^{-1/2}(\mathcal{U}\times\mathcal{V}\times\mathcal{C}^{*}),$
where
(6.4) $\begin{split}&(\mathcal{T}_{11}g_{\epsilon})(x):=\frac{1}{2\pi}\int_{\mathbb{R}}|\nu|\int_{\mathcal{V}}K(x,\alpha)e^{i\nu(p-\Phi(x,\alpha))}g_{\epsilon}(y)\text{d}y\text{d}\nu,\ x\in\mathcal{U}_{0},\\ &K(x,\alpha):=\chi^{2}(x)|\Phi_{x}^{\prime}(x,\alpha)|W(x,(\Phi(x,\alpha),\alpha)),\end{split}$
and $\mathcal{T}_{12}:=\mathcal{T}_{1}-\mathcal{T}_{11}$. The inclusion $\mathcal{T}_{12}\in I^{-1/2}(\mathcal{U}\times\mathcal{V}\times\mathcal{C}^{*})$ follows by considering lower order terms in the stationary phase expansion as $\nu\to\infty$ of the integral in (6.2) [53, Chapter VIII, eqs. (2.14), (2.16)].
By the continuity of FIOs in Hölder-Zygmund spaces (see [23, Proposition 5.5] and (C.12), (C.13)), $\|\mathcal{T}_{12}g_{\epsilon}\|_{C^{1/2}(\mathcal{U})}<c$ for all $0<\epsilon\ll 1$. By Lemma 4.2, $\Delta g,\Delta g_{\epsilon}\in C_{0}^{3/2}(\mathcal{V})$ and $\|\Delta g_{\epsilon}\|_{C^{3/2}(\mathcal{V})}<c$, $0<\epsilon\ll 1$. In this case $\|\mathcal{T}_{1}\Delta g_{\epsilon}\|_{C^{1/2}(\mathcal{U})}<c$, $0<\epsilon\ll 1$.
Therefore, all that remains is to bound the following expression:
(6.5) $\begin{split}F_{\epsilon}^{(l)}(x):=-\frac{1}{\pi}\int_{\mathcal{V}}&\frac{\chi_{1\text{D}}(p-\Phi(x,\alpha))K(x,\alpha)}{p-\Phi(x,\alpha)}\\ &\times\sum_{j}\partial_{p}\varphi_{\epsilon}(y-y_{j})a_{0}(y_{j})(p_{j_{2}}-P(\alpha_{j_{1}}))_{+}^{1/2}\text{d}y.\end{split}$
Here and in the rest of this section, the pair $a_{0}(y),P(\alpha)$ stands for any one of the pairs $a_{l},P_{l}$ described in Lemma 4.2. This means that the domain of integration $\mathcal{V}$ in (6.5) can be replaced by a small neighborhood of $\text{supp}\,a_{l}$. This increase in support is due to interpolation.
To get (6.5) from (6.4) we expressed $\text{Op}(|\nu|)=-1/(\pi t^{2})$, integrated by parts with respect to $p$, and inserted a 1D cutoff function $\chi_{1\text{D}}$. The latter identically equals one in a neighborhood of zero. Also, we omitted smooth terms with uniformly bounded $\|\cdot\|_{C^{1}(\mathcal{U})}$ norm, $0<\epsilon\ll 1$, which arise due to the cutoff.
Summarizing the above results yields
[…]
for any fixed $A>0$.
6.2. Simplification of $F_{\epsilon}^{(l)}$, the leading order term of $F_{\epsilon}$
In what follows we use the function
(6.7) $\begin{split}&Q(x,\alpha):=\Phi(x,\alpha)-P(\alpha).\end{split}$
Since we want to prove that $F_{\epsilon}^{(l)}(x)$ is uniformly bounded for all $x\in\mathcal{U}$, we drop $x$ from notation (e.g., $K(x,\alpha)$ of (6.4) becomes $K(\alpha)$) and make sure that all the estimates are uniform with respect to $x$. It suffices to assume that $x$ is confined to a sufficiently small open subset $U^{\prime}\subset\mathcal{U}$.
Change variables $p=w+\Phi(\alpha)$ in (6.5):
(6.8) $\begin{split}F_{\epsilon}^{(l)}=-\frac{1}{\pi}\sum_{j_{1}}\int_{\mathbb{R}}&\frac{\chi_{1\text{D}}(w)}{w}\sum_{j_{2}}\bigg{[}\partial_{w}\int_{\mathbb{R}}K(\alpha)\varphi_{\epsilon}(y-y_{j})\text{d}\alpha\bigg{]}\\ &\times a_{0}(y_{j})(p_{j_{2}}-P(\alpha_{j_{1}}))_{+}^{1/2}\text{d}w,\ y=(\alpha,w+\Phi(\alpha)).\end{split}$
The sum with respect to $j_{1}$ is over $\alpha_{j_{1}}$ in the domain of the locally defined function $P(\alpha)$. Consider the integral with respect to $\alpha$:
(6.9) $\begin{split}\int_{\mathbb{R}}&K(\alpha)\varphi_{\alpha}\bigg{(}\frac{\alpha-\alpha_{j_{1}}}{\mu\epsilon}\bigg{)}\partial_{w}\varphi_{p}\bigg{(}\frac{w+\Phi(\alpha)-p_{j_{2}}}{\epsilon}\bigg{)}\text{d}\alpha\\ =&\epsilon^{-1}\int_{\mathbb{R}}K(\alpha)\varphi_{\alpha}\bigg{(}\frac{\alpha-\alpha_{j_{1}}}{\mu\epsilon}\bigg{)}\varphi_{p}^{\prime}\bigg{(}\frac{\Phi(\alpha)-\Phi(\alpha_{j_{1}})}{\epsilon}+u\bigg{)}\text{d}\alpha,\\ u:=&(w+\Phi(\alpha_{j_{1}})-p_{j_{2}})/\epsilon.\end{split}$
We have used that $\varphi$ is the product of two kernels, see (3.13). Replace $\alpha_{j_{1}}$ with $\theta$, change variable $\alpha=\theta+\epsilon\mu s$, and define
(6.10) $\begin{split}h(u;\theta,\epsilon):=&\mu\int_{\mathbb{R}}K(\theta+\epsilon\mu s)\varphi_{\alpha}(s)\varphi_{p}\bigg{(}\frac{\Phi(\theta+\epsilon\mu s)-\Phi(\theta)}{\epsilon}+u\bigg{)}\text{d}s.\end{split}$
The sum with respect to $j_{2}$ in (6.8) simplifies by the introduction of the following function
(6.11) $H(\check{t},\check{q};\theta,\epsilon):=\sum_{j_{2}}h^{\prime}(\check{t}+\check{q}-j_{2};\theta,\epsilon)a_{0}(\theta,\epsilon j_{2})(j_{2}-\check{q})_{+}^{1/2}.$
The prime in $h^{\prime}$ denotes the derivative with respect to the first argument. To get the sum in (6.8) we observe that $\epsilon$ is factored out from $p_{j_{2}}-P(\alpha_{j_{1}})$ and use (6.11) with
(6.12) $\check{t}=(w+Q(\alpha_{j_{1}}))/\epsilon,\ \check{q}=P(\alpha_{j_{1}})/\epsilon,\ \theta=\alpha_{j_{1}}.$
Some properties of the function $H$ are obtained in Lemma D.1 in appendix D.1.
Next we consider the integral with respect to $w$ in (6.8):
(6.13) $\begin{split}&\int_{\mathbb{R}}\frac{\chi_{1\text{D}}(w)}{w}H\bigg{(}\frac{w+Q(\theta)}{\epsilon},\frac{P(\theta)}{\epsilon};\theta,\epsilon\bigg{)}\text{d}w\\ &=\int_{\mathbb{R}}\frac{\chi_{1\text{D}}(\epsilon\check{w})}{\check{w}}H\bigg{(}\check{w}+\frac{Q(\theta)}{\epsilon},\frac{P(\theta)}{\epsilon};\theta,\epsilon\bigg{)}\text{d}\check{w},\end{split}$
where $\theta=\alpha_{j_{1}}$. Define a key intermediate function
(6.14) $\begin{split}G(\check{t},\check{q};\theta,\epsilon):=&\int_{\mathbb{R}}\chi_{1\text{D}}(\epsilon\check{w})\frac{H(\check{w}+\check{t},\check{q};\theta,\epsilon)}{\check{w}}\text{d}\check{w}.\end{split}$
In terms of $G$, (6.8) becomes
(6.15) $\begin{split}F_{\epsilon}^{(l)}=-\frac{\epsilon^{1/2}}{\pi}\sum_{j_{1}}G\bigg{(}\frac{Q(\alpha_{j_{1}})}{\epsilon},\frac{P(\alpha_{j_{1}})}{\epsilon},\alpha_{j_{1}},\epsilon\bigg{)}.\end{split}$
The extra factor $\epsilon^{1/2}$ appears because it has been factored out from the term $(p_{j_{2}}-P(\alpha_{j_{1}}))_{+}^{1/2}$ in (6.8) (cf. (6.11)).
Let $\text{dom}\,P$ denote the domain of the selected $P(\alpha)$. The following lemma is proven in appendix D.1.
Lemma 6.2.
Suppose assumptions 3.13.4 are satisfied. Pick any $A>0$. One has
(6.16) $\begin{split}G(\check{t},\check{q};\theta,\epsilon)=&\begin{cases}O(\epsilon^{1/2})+O(\check{t}^{-1}),&\check{t}\to\infty,\\ O((-\check{t})^{-1/2}),&\check{t}\to-\infty.\end{cases}\end{split}$
In the above formula the big-$O$ term is uniform in $\check{q}\in\mathbb{R}$, $\theta\in\text{dom}\,P$, $|\check{t}|\leq A/\epsilon$, and $0<\epsilon\ll 1$.
We use this lemma in the proof of the following result, see appendix D.3.
Lemma 6.3.
Suppose assumptions 3.13.4 are satisfied. Then $|F_{\epsilon}^{(l)}(x)|\leq c$ for all $x\in\mathcal{U}_{b}$ and $0<\epsilon\ll 1$.
Combining Lemma 6.3 with (6.6) proves Lemma 6.1.
7. Computing the leading term $\Delta F_{0}$
In this section we consider $x=x_{0}+\epsilon\check{x}$ and assume for simplicity $x_{0}=0_{2}$. The reconstruction point is then $x=\epsilon\check{x}$. We partially return the dependence on $x$ in the notation by keeping only the dependence on $\check{x}$. If an argument of a function is $x_{0}=0_{2}$, then it is omitted from notation. Fix any $A_{0}>0$. Throughout this section we assume $|\check{x}|\leq A_{0}$. Define
(7.1) $\Delta F_{0}(\check{x}):=\lim_{\epsilon\to 0}\big{[}F_{\epsilon}(\epsilon\check{x})-F_{\epsilon}(0_{2})\big{]}.$
By the second statement in (6.6),
(7.2) $\Delta F_{0}(\check{x})=\lim_{\epsilon\to 0}\Delta F_{\epsilon}^{(l)}(\epsilon\check{x}),\ \Delta F_{\epsilon}^{(l)}(\epsilon\check{x}):=F_{\epsilon}^{(l)}(\epsilon\check{x})-F_{\epsilon}^{(l)}(0_{2}).$
Using (6.11), define another intermediate function
(7.3) $\begin{split}\Delta G(r,\check{t},\check{q};\theta,\epsilon):=&\int_{\mathbb{R}}\chi_{1\text{D}}(\epsilon\check{w})\frac{\Delta H(r,\check{w}+\check{t},\check{q};\theta,\epsilon)}{\check{w}}\text{d}\check{w},\\ \Delta H(r,\check{t},\check{q};\cdot):=&H(r+\check{t},\check{q};\cdot)-H(\check{t},\check{q};\cdot).\end{split}$
We define $\Delta G$ not via $G$ in (6.14) but in terms of $\Delta H$ because this allows for more accurate bounds. By (6.15), the expression for $\Delta F_{\epsilon}^{(l)}$ becomes:
(7.4) $\begin{split}&\Delta F_{\epsilon}^{(l)}(\epsilon\check{x})=-\frac{\epsilon^{1/2}}{\pi}\sum_{j_{1}}\Delta G\bigg{(}\frac{Q(\epsilon\check{x},\alpha_{j_{1}})-Q(\alpha_{j_{1}})}{\epsilon},\frac{Q(\alpha_{j_{1}})}{\epsilon},\frac{P(\alpha_{j_{1}})}{\epsilon};\alpha_{j_{1}},\epsilon\bigg{)}.\end{split}$
As before, the sum is over $\alpha_{j_{1}}\in\text{dom}\,P$.
The following result is proven in appendix E.1.
Lemma 7.1.
Suppose assumptions 3.13.4 are satisfied. Pick any $A>0$. One has
(7.5) $\begin{split}\Delta G(r,\check{t},\check{q};\theta,\epsilon),\partial_{r}\Delta G(r,\check{t},\check{q};\theta,\epsilon)=O(|\check{t}|^{-1}),\ \check{t}\to\infty.\end{split}$
Also,
[…]
In the above formulas the big-$O$ terms are uniform in $\check{q}\in\mathbb{R}$, $\theta\in\text{dom}\,P$, $|\check{t}|\leq A/\epsilon$, $|r|\leq A$, and $0<\epsilon\ll 1$.
Recall that $Q(\alpha)$ stands for $Q(x_{0}=0_{2},\alpha)$. Suppose first that the selected $P(\alpha)$ is associated with $x_{0}$, so $Q(0)=0$. By Lemma D.3, $Q_{\alpha}^{\prime}(0)=0$. From
(7.7) $\begin{split}Q(\epsilon\check{x},\alpha)=&Q_{x}^{\prime}(0)\epsilon\check{x}+Q_{\alpha\alpha}^{\prime\prime}(0)\alpha^{2}/2+O(\epsilon^{3/2}),\\ Q_{\alpha}^{\prime}(\epsilon\check{x},\alpha)=&Q_{\alpha\alpha}^{\prime\prime}(0)\alpha+O(\epsilon),\quad\alpha=O(\epsilon^{1/2}),\end{split}$
we conclude
(7.8) $\alpha_{1}(\epsilon\check{x})=O(\epsilon^{1/2}),\ \alpha_{2}(\epsilon\check{x})=O(\epsilon),$
where $\alpha_{1}$, $\alpha_{2}$ are found by solving $Q(\epsilon\check{x},\alpha_{1})=0$ and $Q_{\alpha}^{\prime}(\epsilon\check{x},\alpha_{2})=0$. By (6.7), $Q_{x}^{\prime}(x,\alpha)=\Phi_{x}^{\prime}(x,\alpha)$.
Pick any $A\gg 1$. By Lemma D.3, $Q_{\alpha\alpha}^{\prime\prime}(0)>0$. Hence, by continuity,
(7.9) $Q(\theta)\geq c\theta^{2},\quad|\theta|\leq\delta,$
for some $0<\delta\ll 1$. Also, $Q(\theta)\geq c$, $|\theta|\geq\delta$, $\theta\in\text{dom}\,P$. By (7.5), this implies
(7.10) $\begin{split}&\epsilon^{1/2}\bigg{(}\sum_{A\epsilon^{1/2}\leq|\alpha_{j_{1}}|\leq\delta}+\sum_{|\alpha_{j_{1}}|>\delta}\bigg{)}|\Delta G|\\ &\leq c\epsilon^{1/2}\bigg{(}\sum_{j_{1}>A\epsilon^{-1/2}}(\epsilon j_{1}^{2})^{-1}+\sum_{j_{1}=\delta/\epsilon}^{1/\epsilon}\epsilon\bigg{)}=O(A^{-1})+O(\epsilon^{1/2}),\end{split}$
where the arguments of $\Delta G$ are the same as in (7.4).
By (7.5) and (7.6), for $|\alpha_{j_{1}}|\leq A\epsilon^{1/2}$:
(7.11) $\begin{split}&\Delta G\bigg{(}\frac{Q(\epsilon\check{x},\alpha_{j_{1}})-Q(\alpha_{j_{1}})}{\epsilon},\frac{Q(\alpha_{j_{1}})}{\epsilon},\frac{P(\alpha_{j_{1}})}{\epsilon};\alpha_{j_{1}},\epsilon\bigg{)}\\ &=\Delta G\bigg{(}\Phi_{x}^{\prime}\check{x}+O(\epsilon^{1/2}),\frac{Q_{\alpha\alpha}^{\prime\prime}}{2}\epsilon(\mu j_{1})^{2}+O(\epsilon^{1/2}),\frac{P(\epsilon\mu j_{1})}{\epsilon};\alpha_{j_{1}},\epsilon\bigg{)}\\ &=\Delta G\bigg{(}\Phi_{x}^{\prime}\check{x},\frac{Q_{\alpha\alpha}^{\prime\prime}}{2}\epsilon(\mu j_{1})^{2},\frac{P(\epsilon\mu j_{1})}{\epsilon};\alpha_{j_{1}},\epsilon\bigg{)}+\frac{O(\epsilon^{1/2}\ln(1/\epsilon))}{(1+\epsilon j_{1}^{2})^{3/2}}.\end{split}$
Pick any $P(\alpha)$, which is not associated with a segment of $\mathcal{S}$ through $x_{0}$. This means there is a neighborhood $U^{\prime}$ of $x_{0}$ such that none of the curves $\mathcal{S}_{(\alpha,P(\alpha))}$, $\alpha\in\text{dom}\,P$, is tangent to $\mathcal{S}\cap U^{\prime}$. There are two possibilities: (a) Neither of these curves passes through $x_{0}$ (in this case $Q(\alpha)$ is bounded away from zero on $\text{dom}\,P$) and (b) one of the curves contains $x_{0}$ (in this case $Q(\alpha)=0$ for some $\alpha\in\text{dom}\,P$).
If $|Q(\alpha)|>c$, $\alpha\in\text{dom}\,P$, then, similarly to (7.10),
(7.12) $\epsilon^{1/2}\textstyle\sum_{j_{1}}|\Delta G|\leq c\epsilon^{1/2}\textstyle\sum_{j_{1}=1}^{1/\epsilon}\epsilon=O(\epsilon^{1/2}).$
If $Q(\tilde{\alpha})=0$, but $|Q^{\prime}(\alpha)|>c$, then $|Q(\tilde{\alpha})|\asymp|\alpha-\tilde{\alpha}|$, $\alpha\in\text{dom}\,P$, and
(7.13) $\epsilon^{1/2}\textstyle\sum_{j_{1}}|\Delta G|\leq c\epsilon^{1/2}\ln(1/\epsilon)\textstyle\sum_{j_{1}=1}^{1/\epsilon}(1/j_{1})=O(\epsilon^{1/2}\ln^{2}(1/\epsilon)).$
Combine (7.10)–(7.13) and use (7.4)
(7.14) $\begin{split}\Delta F_{\epsilon}^{(l)}(\epsilon\check{x})=&-\frac{\epsilon^{1/2}}{\pi}\sum_{|\alpha_{j_{1}}|\leq A\epsilon^{1/2}}\Delta G\bigg{(}\Phi_{x}^{\prime}\check{x},\frac{Q_{\alpha\alpha}^{\prime\prime}}{2}\epsilon(\mu j_{1})^{2},\frac{P(\epsilon\mu j_{1})}{\epsilon};\alpha_{j_{1}},\epsilon\bigg{)}\\ &+O(\epsilon^{1/2}\ln^{2}(1/\epsilon))+O(1/A),\end{split}$
where $O(\epsilon^{1/2}\ln^{2}(1/\epsilon))$ is independent of $A$, and $P$ is associated with $x_{0}$. Here we have used that $\epsilon^{1/2}\sum_{j_{1}}(1+\epsilon j_{1}^{2})^{-3/2}<c$, $10<\epsilon<$ 1.
The following result is proven in appendix E.2.
Lemma 7.2.
Under the assumptions of Theorem 3.6 one has
(7.15) $\begin{split}\Delta F_{0}(\check{x})=-\frac{1}{\pi}\int_{|s|\leq A}\int_{0}^{1}&\big{[}G_{0}\big{(}\Phi_{x}^{\prime}\check{x}+(Q_{\alpha\alpha}^{\prime\prime}/2)s^{2},\check{q}\big{)}\\ &-G_{0}\big{(}(Q_{\alpha\alpha}^{\prime\prime}/2)s^{2},\check{q}\big{)}\big{]}\text{d}\check{q}\text{d}s+O(1/A),\end{split}$
where
(7.16) $\begin{split}&G_{0}(\check{t},\check{q}):=\int_{\mathbb{R}}\frac{H_{0}(\check{w}+\check{t},\check{q})}{\check{w}}\text{d}\check{w},\\ &H_{0}(\check{t},\check{q}):=a_{0}(y_{0})\sum_{j_{2}}h_{0}^{\prime}(\check{t}+\check{q}-j_{2})(j_{2}-\check{q})_{+}^{1/2},\\ &h_{0}(u):=K(x_{0},\alpha_{0})\int_{\mathbb{R}}\varphi_{\alpha}(s)\varphi_{p}(\mu s\Phi_{\alpha}^{\prime}(0)+u)\text{d}s.\end{split}$
See (6.4) for the definition of $K(x,\alpha)$. Letting $A\to\infty$ in (7.15) gives
(7.17) $\begin{split}\Delta F_{0}(\check{x})=&-\frac{1}{\pi}\int_{\mathbb{R}}\int_{0}^{1}\big{[}G_{0}\big{(}\Phi_{x}^{\prime}\check{x}+(Q_{\alpha\alpha}^{\prime\prime}/2)s^{2},\check{q}\big{)}-G_{0}\big{(}(Q_{\alpha\alpha}^{\prime\prime}/2)s^{2},\check{q}\big{)}\big{]}\text{d}\check{q}\text{d}s\\ =&-\frac{a_{0}(y_{0})}{\pi}\int_{\mathbb{R}}\int_{\mathbb{R}}\int_{0}^{\infty}\frac{1}{\check{w}}\big{[}h_{0}^{\prime}\big{(}\check{w}+\Phi_{x}^{\prime}\check{x}+(Q_{\alpha\alpha}^{\prime\prime}/2)s^{2}-\check{q}\big{)}\\ &\hskip 99.58464pt-h_{0}^{\prime}\big{(}\check{w}+(Q_{\alpha\alpha}^{\prime\prime}/2)s^{2}-\check{q}\big{)}\big{]}\check{q}^{1/2}\text{d}\check{q}\text{d}s\text{d}\check{w}.\end{split}$
It is easy to see that $G_{0}(\check{t}+r,\check{q})-G_{0}(\check{t},\check{q})$ satisfies the same estimates as $\Delta G$ in (7.5), so the integral in (7.17) is absolutely convergent.
The following result is proven in appendix E.3.
Lemma 7.3.
Under the assumptions of Theorem 3.6 one has
(7.18) $\Delta F_{0}(\check{x})=C\int_{0}^{\Phi_{x}^{\prime}\check{x}}h_{0}(-t)\text{d}t,\ C:=\frac{a_{0}(y_{0})}{4(Q_{\alpha\alpha}^{\prime\prime}/2)^{1/2}}.$
8. Computing the DTB function. End of proof of Theorem 3.6
Pick any $A_{0}>0$. Throughout this section we assume $|\check{x}|\leq A_{0}$.
8.1. Reduction of the DTB function to an integral
Now we study $G_{\epsilon}$ introduced in (5.7). As usual, suppose $x_{0}=0_{2}$. Change variables $\hat{\xi}=\epsilon\xi$ and $\check{x}=x/\epsilon$, $\check{z}=z/\epsilon$ in (5.7) and express $G_{\epsilon}$ as follows:
(8.1) $\begin{split}&G_{\epsilon}(\check{x}):=\frac{1}{(2\pi)^{2}}\int_{\mathcal{U}/\epsilon}\int_{\mathbb{R}^{2}}\big{(}\tilde{D}_{0}(\epsilon\check{x},\hat{\xi})+\kappa|\hat{\xi}|^{3}\big{)}^{-1}e^{-i\hat{\xi}\cdot(\check{x}-\check{z})}\text{d}\hat{\xi}\,F_{\epsilon}(\epsilon\check{z})\text{d}\check{z}.\end{split}$
For clarity, $G_{\epsilon}$ in (LABEL:to_scl_PDO_2) is different from the functions $G$ and $G_{0}$ used in sections 6 and 7. Since
(8.2) $\tilde{D}_{0}(\epsilon\check{x},\hat{\xi})+\kappa|\hat{\xi}|^{3}\asymp|\hat{\xi}|^{3},\ |\hat{\xi}|\to\infty,$
the integral with respect to $\hat{\xi}$ in (LABEL:to_scl_PDO_2) is absolutely convergent.
Our aim is to reconstruct the jump of $f$ at $x_{0}=0_{2}$, so we compute $\Delta G_{\epsilon}(\check{x}):=G_{\epsilon}(\check{x})-G_{\epsilon}(0_{2})$. By (LABEL:to_scl_PDO_2), $\Delta G_{\epsilon}(\check{x})=\Delta G_{\epsilon}^{(1)}(\check{x})+\Delta G_{\epsilon}^{(2)}(\check{x})$, where
(8.3) $\begin{split}\Delta G_{\epsilon}^{(1)}(\check{x})=\frac{1}{(2\pi)^{2}}\int_{\mathcal{U}/\epsilon}\int_{\mathbb{R}^{2}}&\big{[}(\tilde{D}_{0}(\epsilon\check{x},\hat{\xi})+\kappa|\hat{\xi}|^{3})^{-1}-(\tilde{D}_{0}(0_{2},\hat{\xi})+\kappa|\hat{\xi}|^{3})^{-1}\big{]}\\ &\times e^{-i\hat{\xi}\cdot(\check{x}-\check{z})}\text{d}\hat{\xi}\,F_{\epsilon}(\epsilon\check{z})\text{d}\check{z},\end{split}$
(8.4) $\begin{split}&\Delta G_{\epsilon}^{(2)}(\check{x})=\frac{1}{(2\pi)^{2}}\int_{\mathcal{U}/\epsilon}\int_{\mathbb{R}^{2}}\frac{e^{-i\hat{\xi}\cdot\check{x}}-1}{\tilde{D}_{0}(0_{2},\hat{\xi})+\kappa|\hat{\xi}|^{3}}e^{i\hat{\xi}\cdot\check{z}}\text{d}\hat{\xi}\,F_{\epsilon}(\epsilon\check{z})\text{d}\check{z}.\end{split}$
The following result is proven in appendix F.1.
Lemma 8.1.
Let $I=(a,b)$ be an interval where $-\pi<a<0<b<\pi$. Let $D(\lambda,\theta,\epsilon)\in C^{\infty}\big{(}(0,\infty)\times I\times(0,\epsilon_{0})\big{)}$ be a function which satisfies
(8.5) $\big{|}\partial_{\theta}^{l}\partial_{\lambda}^{k}D(\lambda,\theta,\epsilon)\big{|}\leq c_{k,l}(1+\lambda)^{-3},\ (\lambda,\theta,\epsilon)\in(0,\infty)\times I\times(0,\epsilon_{0}),$
for any $k=0,1,2$, $l\in\mathbb{N}_{0}$, some $c_{k,l}$, which can be independent of $\epsilon$, and some $\epsilon_{0}>0$. Define
(8.6) $J(r):=\int_{0}^{\infty}\int_{I}D(\lambda,\theta,\epsilon)e^{ir\lambda\sin\theta}\text{d}\theta\lambda\text{d}\lambda.$
Then
(8.7) $J(r)=\begin{cases}O(r^{-2})&\\ O(r^{-3}),&\text{if }D(\lambda=0^{+},\theta,\epsilon)\equiv 0,\end{cases}\ r\to\infty,$
uniformly in $\epsilon\in(0,\epsilon_{0})$. Moreover, if $c_{k,l}=O(\epsilon)$, then
(8.8) $J(r)=\begin{cases}O(\epsilon)O(r^{-2})\\ O(\epsilon)O(r^{-3}),&\text{if }D(\lambda=0^{+},\theta,\epsilon)\equiv 0,\end{cases}\ \epsilon\to 0,r\to\infty.$
The expression in brackets in (8.3) satisfies (8.5) with $c_{k,l}=O(\epsilon)$ uniformly in $\check{x}$ within bounded sets (see (C.7)). Here $\hat{\xi}=\lambda(\cos\theta,\sin\theta)$, $\lambda=|\hat{\xi}|$. By the top case in (8.8), the integral with respect to $\hat{\xi}$ in (8.3) is $O(\epsilon)O(|\check{z}|^{-2})$, $|\check{z}|\to\infty$. Recall that $\check{x}$ is confined to a bounded set. Additionally, the support of $F_{\epsilon}(\epsilon\check{z})$ is of size $O(1/\epsilon)$. Hence
(8.9) $|\Delta G_{\epsilon}^{(1)}(\check{x})|\leq c\epsilon\int_{|\check{z}|\leq 1/\epsilon}(1+|\check{z}|)^{-2}\text{d}\check{z}=O(\epsilon\ln(1/\epsilon)).$
The fraction in (LABEL:del_bfe_2) satisfies (8.5) uniformly in $\check{x}$ within bounded sets. Moreover, its value at $\lambda=0^{+}$ equals zero. By the bottom case in (8.7), the integral with respect to $\hat{\xi}$ is $O(|\check{z}|^{-3})$, $|\check{z}|\to\infty$. This means that the integral with respect to $\check{z}$ in (LABEL:del_bfe_2) is absolutely convergent. In both integrals we have used Lemma 6.1.
In section 7 we computed the limit
(8.10) $\Delta F_{0}(\check{x})=\lim_{\epsilon\to 0}\Delta F_{\epsilon}(\epsilon\check{x}),\ |\check{x}|\leq A,$
for any $A>0$ (see (7.1) and (7.18)). Pick any $A\gg A_{0}$ and use (8.9) to obtain from (LABEL:to_scl_PDO_2), (8.3), (LABEL:del_bfe_2):
(8.11) $\begin{split}\Delta G_{\epsilon}(\check{x})=&\frac{1}{(2\pi)^{2}}\int_{|\check{z}|\leq A}\bigg{[}\int_{\mathbb{R}^{2}}\frac{e^{-i\hat{\xi}\cdot\check{x}}-1}{\tilde{D}_{0}(0_{2},\hat{\xi})+\kappa|\hat{\xi}|^{3}}e^{i\hat{\xi}\cdot\check{z}}\text{d}\hat{\xi}\bigg{]}\Delta F_{\epsilon}(\epsilon\check{z})\text{d}\check{z}\\ &+\frac{F_{\epsilon}(0)}{2\pi}\int_{\mathbb{R}^{2}}\frac{e^{-i\hat{\xi}\cdot\check{x}}-1}{\tilde{D}_{0}(0_{2},\hat{\xi})+\kappa|\hat{\xi}|^{3}}\frac{AJ_{1}(A|\hat{\xi}|)}{|\hat{\xi}|}\text{d}\hat{\xi}\\ &+O(1/A)+O(\epsilon\ln(1/\epsilon)),\end{split}$
where $J_{1}$ is the Bessel function of the first kind. By Lemma 8.1, the integral in brackets above is $O(|\check{z}|^{-3})$, $|\check{z}|\to\infty$. Use (8.10), and the Lebesgue dominated convergence theorem to take the limit as $\epsilon\to 0$ in the first integral in (8.11) to obtain:
(8.12) $\begin{split}\frac{1}{(2\pi)^{2}}\int_{|\check{z}|\leq A}\int_{\mathbb{R}^{2}}\frac{e^{-i\hat{\xi}\cdot\check{x}}-1}{\tilde{D}_{0}(0_{2},\hat{\xi})+\kappa|\hat{\xi}|^{3}}e^{i\hat{\xi}\cdot\check{z}}\text{d}\hat{\xi}\,\Delta F_{0}(\check{z})\text{d}\check{z}.\end{split}$
Since $A\gg A_{0}$ can be arbitrarily large and $AJ_{1}(A|\hat{\xi}|)/|\hat{\xi}|\to\delta(\hat{\xi})$ as $A\to\infty$, (8.11), (8.12), and Lemma 6.1 imply
(8.13) $\begin{split}\lim_{\epsilon\to 0}\Delta G_{\epsilon}(\check{x})=&\frac{1}{(2\pi)^{2}}\int_{\mathbb{R}^{2}}\int_{\mathbb{R}^{2}}\frac{e^{-i\hat{\xi}\cdot\check{x}}-1}{\tilde{D}_{0}(0_{2},\hat{\xi})+\kappa|\hat{\xi}|^{3}}e^{i\hat{\xi}\cdot\check{z}}\text{d}\hat{\xi}\,\Delta F_{0}(\check{z})\text{d}\check{z}.\end{split}$
8.2. Evaluation of the integral in (8.13)
The last integral simplifies using that $\Delta F_{0}(\check{x})=\varphi(\Phi_{x}^{\prime}\check{x})$ for some $\varphi$ (see (7.18)). Simple transformations give
(8.14) $\begin{split}\Delta G_{0}(\check{x}):=&\lim_{\epsilon\to 0}\Delta G_{\epsilon}(\check{x})\\ =&\frac{1}{2\pi}\int_{\mathbb{R}}\bigg{[}\int_{\mathbb{R}}\frac{e^{-i\lambda\vec{\Theta}\cdot\check{x}}-1}{\tilde{D}_{0}(0_{2},\vec{\Theta}_{0})+\kappa|\lambda|^{3}}e^{i\lambda t}\text{d}\lambda\bigg{]}\varphi(|\Phi_{x}^{\prime}|t)\text{d}t,\\ \varphi(t):=&C\int_{0}^{t}h_{0}(-s)\text{d}s,\ \vec{\Theta}_{0}:=\Phi_{x}^{\prime}/|\Phi_{x}^{\prime}|,\end{split}$
where $C$ is defined in (7.18). Denote
(8.15) $\tilde{R}(\lambda):=\big{[}\tilde{D}_{0}(0_{2},\vec{\Theta}_{0})+\kappa|\lambda|^{3}\big{]}^{-1},\ R:=\mathcal{F}^{-1}\tilde{R}.$
Since $\tilde{R}(\lambda)$ is $C^{2}$ near $\lambda=0$, $R(t)$ is absolutely integrable. Combine (7.18), (8.14), and (8.15) to obtain
(8.16) $\begin{split}&\Delta G_{0}(\check{x})=C\int_{\mathbb{R}}(R(r-t)-R(-t))\int_{-\infty}^{|\Phi_{x}^{\prime}|t}h_{0}(-s)\text{d}s\text{d}t,\ r:=\vec{\Theta}_{0}\cdot\check{x}.\end{split}$
We have extended the integral with respect to $s$ to $(-\infty,|\Phi_{x}^{\prime}|t]$ because $\int(R(r-t)-R(-t))\text{d}t\equiv 0$ for all $r$. Simple transformations give
(8.17) $\begin{split}&\Delta G_{0}(\check{x})=(C/|\Phi_{x}^{\prime}|)\int_{\mathbb{R}}h_{0}(s)\int_{s}^{s+\Phi_{x}^{\prime}\check{x}}R(t/|\Phi_{x}^{\prime}|)\text{d}t\text{d}s.\end{split}$
For convenience, we collect here all the constants that will be used in subsequent calculations.
(8.18) $\tilde{D}(0_{2},\Phi_{x}^{\prime})=2\pi\frac{W^{2}|\Phi_{x}^{\prime}|}{|\Delta_{\Phi}|}\quad$ (by (C.5), (C.6))
(8.19) $K(x_{0},\alpha_{0})=|\Phi_{x}^{\prime}|W\quad$ (by (6.4))
(8.20) $Q_{\alpha\alpha}^{\prime\prime}(x_{0},\alpha_{0})=\frac{|\Delta_{\Phi}||x_{\alpha}^{\prime}|}{|\Phi_{x}^{\prime}|}\quad$ (by (D.21), (D.22))
(8.21) $\psi(y_{0})=\frac{|H_{x}^{\prime}|}{|\Phi_{x}^{\prime}|}\quad$ (by (4.7))
[…]
For simplicity, we dropped all the arguments on the right.
Let $\bar{h}_{0}$ be defined the same way as $h_{0}$ in (7.16), but without the factor $K(x_{0},\alpha_{0})$. Then, by the above formulas and (8.15):
(8.23) $\begin{split}&\Delta G_{0}(\check{x})=\tilde{f}_{0}\Upsilon\big{(}\Phi_{x}^{\prime}\check{x}\big{)},\\ &\Upsilon(r)=2\pi\frac{W^{2}}{|\Delta_{\Phi}|}\int_{\mathbb{R}}\bar{h}_{0}(u)\int_{u}^{u+r}\mathcal{F}^{-1}\bigg{(}\frac{1}{2\pi\frac{W^{2}|\Phi_{x}^{\prime}|}{|\Delta_{\Phi}|}+\kappa|\lambda|^{3}}\bigg{)}(t/|\Phi_{x}^{\prime}|)\text{d}t\text{d}u.\end{split}$
A simple transformation yields:
(8.24) $\begin{split}\Upsilon(r)=&\frac{1}{|\Phi_{x}^{\prime}|}\int_{\mathbb{R}}\bar{h}_{0}(u)\int_{u}^{u+r}\mathcal{F}^{-1}\bigg{(}\frac{1}{1+\kappa C_{1}|\lambda|^{3}}\bigg{)}(t/|\Phi_{x}^{\prime}|)\text{d}t\text{d}u,\\ C_{1}=&\frac{|\Delta_{\Phi}|}{2\pi W^{2}|\Phi_{x}^{\prime}|}.\end{split}$
Further simple transformations prove (3.17), (3.18). The rest of the claims follow from the next lemma, which is proven in appendix F.2.
Lemma 8.2.
One has $\Upsilon(0)=0$ and $\Upsilon(\pm\infty)=\pm 1/2$.
9. Redundant data. Proof of Theorem 3.7
If there are multiple critical points $y_{l}$ from which the singularity at $(x_{0},\xi_{0})$ is visible, the analog of (5.7) becomes
(9.1) $\begin{split}&G_{\epsilon}=\text{Op}\big{(}\big{[}\textstyle\sum_{l}\tilde{D}_{l}(x,\xi/|\xi|)+\kappa|\epsilon\xi|^{3}\big{]}^{-1}\big{)}F_{\epsilon},\ x\in\mathcal{U}_{0},\\ &\tilde{D}_{l}(x_{0},\vec{\Theta}_{0}):=2\pi\frac{W^{2}(x_{0},y_{l})|\Phi_{x}^{\prime}(x_{0},\alpha_{l})|}{|\Delta_{\Phi}(x_{0},\alpha_{l})|}.\end{split}$
Here we have used that
(9.2) $\vec{\Theta}_{0}=\xi_{0}/|\xi_{0}|=H_{x}^{\prime}(x_{0})/|H_{x}^{\prime}(x_{0})|=\Phi_{x}^{\prime}(x_{0},\alpha_{l})/|\Phi_{x}^{\prime}(x_{0},\alpha_{l})|$
for all $l$, cf. the convention about the directions of $H_{x}^{\prime}$ and $\Phi_{x}^{\prime}$ stated above assumption 3.3. Therefore, we get from (8.15) that the analog of the function $R$ is
(9.3) $\tilde{R}(\lambda):=\big{[}\textstyle\sum_{l}\tilde{D}_{l}(x_{0},\vec{\Theta}_{0})+\kappa|\lambda|^{3}\big{]}^{-1},\ R:=\mathcal{F}^{-1}\tilde{R}.$
The analog (LABEL:rhs_lim_5) becomes
(9.4) $\begin{split}&\Delta G_{0}(\check{x})=\tilde{f}_{0}\textstyle\sum_{l}\Upsilon_{l}\big{(}\Phi_{x}^{\prime}(x_{0},\alpha_{l})\check{x}\big{)},\\ &\Upsilon_{l}(r)=2\pi\frac{W^{2}(x_{0},y_{l})}{|\Delta_{\Phi}(x_{0},\alpha_{l})|}\int_{\mathbb{R}}\bar{h}_{l}(u)\\ &\hskip 28.45274pt\times\int_{u}^{u+r}\mathcal{F}^{-1}\left(\big{[}\textstyle\sum_{l}\tilde{D}_{l}(x_{0},\vec{\Theta}_{0})+\kappa|\lambda|^{3}\big{]}^{-1}\right)\bigg{(}\frac{t}{|\Phi_{x}^{\prime}(x_{0},\alpha_{l})|}\bigg{)}\text{d}t\text{d}u,\\ &\bar{h}_{l}(u):=\int_{\mathbb{R}}\varphi_{\alpha}(s)\varphi_{p}(\mu s\Phi_{\alpha}^{\prime}(x_{0},\alpha_{l})+u)\text{d}s.\end{split}$
Simplifying we get similarly to (LABEL:rhs_lim_5), (8.24)
(9.5) $\begin{split}&\Delta G_{0}(\check{x})=\tilde{f}_{0}\frac{\textstyle\sum_{l}\rho_{l}\Upsilon_{l}\big{(}\Phi_{x}^{\prime}(x_{0},\alpha_{l})\check{x}/|\Phi_{x}^{\prime}(x_{0},\alpha_{l})|\big{)}}{\sum_{l}\rho_{l}},\\ &\Upsilon_{l}(r)=\int_{\mathbb{R}}\bar{h}_{0}(u)\int_{u_{l}}^{u_{l}+r}\mathcal{F}^{-1}\bigg{(}\frac{1}{1+\kappa C_{1}|\lambda|^{3}}\bigg{)}(t)\text{d}t\text{d}u,\\ &\rho_{l}:=\frac{W^{2}(x_{0},y_{l})|\Phi_{x}^{\prime}(x_{0},\alpha_{l})|}{|\Delta_{\Phi}(x_{0},\alpha_{l})|},\ u_{l}:=\frac{u}{|\Phi_{x}^{\prime}(x_{0},\alpha_{l})|},\ C_{1}=\frac{1}{2\pi\textstyle\sum_{l}\rho_{l}}.\end{split}$
Using (9.2) finishes the proof.
10. Numerical experiment
We choose $f$ to be the characteristic function of the disk centered at $x_{c}=(1,1)$ with radius $r=2$. Thus, $\mathcal{S}=\{x\in\mathbb{R}^{2}:\|x-x_{c}\|=r\}$. The center of a local region of interest (ROI) is $x_{0}=x_{c}+r\vec{\beta}_{0}$, where $\beta_{0}=0.17\pi$. The GRT integrates over circles $\mathcal{S}_{y}$, $y=(\alpha,\rho)$, with various radii $\rho>0$ and centers $R\vec{\alpha}$, $\alpha\in[0,2\pi)$, where $R=10$. Thus, $\rho=\Phi(x,\alpha)=\|x-R\vec{\alpha}\|$ and
(10.1) $\begin{split}&\Phi_{x}^{\prime}(x,\alpha)=\frac{x-R\vec{\alpha}}{\|x-R\vec{\alpha}\|}=:\vec{\Theta},\ \Phi_{\alpha}^{\prime}(x,\alpha)=R\vec{\Theta}\cdot\vec{\alpha}^{\perp},\\ &\Delta_{\Phi}(x,\alpha)=\det\begin{pmatrix}\vec{\Theta}&-\frac{R}{\|x-R\vec{\alpha}\|}\vec{\Theta}^{\perp}+c\vec{\Theta}\end{pmatrix}=-\frac{R}{\|x-R\vec{\alpha}\|}.\end{split}$
Here $\alpha^{\perp}=(-\sin\alpha,\cos\alpha)$, and $\Theta^{\perp}$ is defined similarly. The continuous data corresponds to $\alpha\in[0,2\pi)$ and $\rho\in(0,2R)$. The global reconstruction region is the square $\mathcal{U}_{b}:=\{(x_{1},x_{2})\in\mathbb{R}^{2}:|x_{1}|<R_{rec},\ |x_{2}|<R_{rec}\}$, $R_{rec}=3.7$. This corresponds to the set $\mathcal{U}_{b}$ used in (3.16). The discrete data are given at the points
[…]
Clearly, for each $x\in\mathcal{S}$ there are two points, $y_{l}=(\alpha_{l},\rho_{l})$, $l=1,2$, such that $\mathcal{S}_{y_{l}}$ is tangent to $\mathcal{S}$ at $x$. To find $y_{1,2}$ for the selected $x_{0}$ we first solve $\|x_{c}+t\vec{\beta}_{0}\|=R$. This gives two values $t_{1},t_{2}$, $t_{2}<0<t_{1}$. Then $\alpha_{l}$ are determined from $x_{c}+t_{l}\vec{\beta}_{0}=R\vec{\alpha}_{l}$, and $\rho_{l}=\|x_{0}-R\vec{\alpha}_{l}\|$, $l=1,2$. In view of (10.1), in numerical computations we use $\|x_{0}-R\vec{\alpha}_{1}\|=t_{1}-r$ and $\|x_{0}-R\vec{\alpha}_{2}\|=-t_{2}+r$.
The regularization parameter is set at $\kappa=0.1$. In order to approximate the continuous GRT $\mathcal{R}$ and its adjoint $\mathcal{R}^{*}$, the reconstruction is performed on a dense $801\times 801$ grid covering the same square region $\mathcal{U}_{b}$, and the interpolated data $g_{\epsilon}$ are assumed to be given on a dense grid of the kind (10.2) with $N_{\alpha}^{\prime}=800$, $N_{\rho}^{\prime}=1201$. Prior to the reconstruction, the GRT data $g(y_{j})$ are interpolated from the grid (10.2) to the more dense grid using (3.13). For this we use the Keys interpolation kernel [34, 5]
(10.3) $\varphi_{\alpha}(t)=\varphi_{\rho}(t)=\varphi(t)=3B_{3}(t+2)-(B_{2}(t+2)+B_{2}(t+1)),$
where $B_{n}$ is the cardinal $B$-spline of degree $n$ supported on $[0,n+1]$. Therefore $\text{supp}\,\varphi=[-2,2]$. The kernel is a piecewise-cubic polynomial with continuous $\varphi,\varphi^{\prime}$ and bounded $\varphi^{\prime\prime}$, so $\varphi\in C_{0}^{2}(\mathbb{R})$.
Minimization of the functional (3.16) is performed using gradient descent. Results of reconstructions are shown in Figures 2 and 3. Reconstruction of the entire region $\mathcal{U}_{b}$ is shown in Figure 2. A profile of the reconstruction through the center of the ball, $x_{c}$, is shown on the right. The location of the profile is shown in the left panel.
A small section of the reconstruction, which is located inside the square ROI shown in Figure 2, is extracted from the global reconstruction and shown in Figure 3. The right panel of Figure 3 shows the reconstructed (green) and predicted (blue) profiles. The latter is computed by using (10.1) in (3.20). The slight discrepancy in the interval $[1,2]$ is due to random numerical errors, which are clearly visible in the right panel of Figure 2. The plot in Figure 3 corresponds to the right jump in Figure 2. Overall, the match between the reconstruction and prediction is very accurate.
Refer to caption Refer to caption
Figure 2. Left: global reconstruction of the disk phantom on a $801\times 801$ grid. The region $\mathcal{U}_{b}=(-3.7,3.7)\times(-3.7,3.7)$ is shown. Right: profile of the reconstruction through the center of the ball, $x_{c}$. The $x$-axis is the distance along the profile with the origin at the center of the disk. The location of the profile is shown in the left panel.
Refer to caption Refer to caption
Figure 3. Left: local reconstruction of the disk phantom inside an ROI. The ROI is a $26\times 26$ square shown in the left panel of Figure 2. Right: profile of the reconstruction through the center of the local ROI, $x_{0}$. The $x$-axis is the (signed) distance in units of $\epsilon$ along the profile with the origin at $x_{0}$. The location of the profile is shown in the left panel. The numerically computed profile is in green, and the predicted profile is in blue.
Appendix A Proof of Lemma 3.5
It is easy to see that $\Psi:H_{c}^{1}(\mathcal{U}_{b})\to\mathbb{R}$ is strongly convex [7, Definition 1.1.48] and [41, Section 2.3] for each $\epsilon>0$. Strong convexity means that there exists $c>0$ such that
(A.1) $\Psi(\lambda f_{1}+(1-\lambda)f_{2})\leq\lambda\Psi(f_{1})+(1-\lambda)\Psi(f_{2})-c\lambda(1-\lambda)\|f_{2}-f_{1}\|_{H^{1}(\mathcal{U}_{b})}^{2}$
for any $f_{1,2}\in H_{c}^{1}(\mathcal{U}_{b})$ and $\lambda\in(0,1)$. By an easy calculation,
(A.2) $\begin{split}\|\lambda\nabla f_{1}+(1-\lambda)\nabla f_{2}\|_{L^{2}(\mathcal{U})}^{2}\leq&\lambda\|\nabla f_{1}\|_{L^{2}(\mathcal{U})}^{2}+(1-\lambda)\|\nabla f_{2}\|_{L^{2}(\mathcal{U})}^{2}\\ &-2\lambda(1-\lambda)\|\nabla(f_{1}-f_{2})\|_{L^{2}(\mathcal{U})}^{2}\end{split}$
and
(A.3) $\begin{split}\|\mathcal{R}(\lambda f_{1}&+(1-\lambda)f_{2})-g_{\epsilon}\|_{L^{2}(\mathcal{V})}^{2}\\ \leq&\lambda\|\mathcal{R}f_{1}-g_{\epsilon}\|_{L^{2}(\mathcal{V})}^{2}+(1-\lambda)\|\mathcal{R}f_{2}-g_{\epsilon}\|_{L^{2}(\mathcal{V})}^{2}\\ &-2\lambda(1-\lambda)\|\mathcal{R}(f_{1}-f_{2})\|_{L^{2}(\mathcal{V})}^{2}.\end{split}$
Therefore
(A.4) $\begin{split}\Psi(\lambda f_{1}&+(1-\lambda)f_{2})\leq\lambda\Psi(f_{1})+(1-\lambda)\Psi(f_{2})\\ &-c\lambda(1-\lambda)\big{(}\|\mathcal{R}(f_{1}-f_{2})\|_{L^{2}(\mathcal{V})}^{2}+\|\nabla(f_{2}-f_{1})\|_{L^{2}(\mathcal{U})}\big{)}.\end{split}$
By (3.6),
(A.5) $\|f\|_{H^{-1/2}(\mathcal{U})}\leq c\|\mathcal{R}f\|_{L^{2}(\mathcal{V})},\ f\in C_{0}^{\infty}(\mathcal{U}_{b}).$
Also
(A.6) $\|f\|_{H^{1}(\mathcal{U})}\leq c\big{(}\|f\|_{H^{-1/2}(\mathcal{U})}+\|\nabla f\|_{L^{2}(\mathcal{U})}\big{)},\ f\in C_{0}^{\infty}(\mathcal{U}_{b}).$
Applying (A.5) and (A.6) to (A.4) proves (A.1). Moreover, the functional is proper (its domain is not empty) and continuous $H_{c}^{1}(\mathcal{U}_{b})\to\mathbb{R}$. Thus the solution to (3.16) exists and is unique for each $\epsilon>0$ [7, Theorem 1.3.1], [41, Corollary 2.20].
Appendix B Proof of Lemmas 4.1 and 4.2
B.1. Proof of Lemma 4.1
Consider the function $H(x_{*}(y))$. Differentiating the first equation in (4.4) with respect to $p$ and then using the second equation gives
(B.1) $\begin{split}\partial_{p}H(x_{*}(y))=&\nu_{1*}(y)=\frac{|H_{x}^{\prime}(x_{*}(y))|}{|\Phi_{x}^{\prime}(x_{*}(y),\alpha)|}>0.\end{split}$
Another easy calculation, combined with (B.1) and (3.11), shows
(B.2) $\begin{split}\det M(y)&=-\big{(}[H_{xx}^{\prime\prime}(x_{*}(y))-\nu_{1*}(y)\Phi_{xx}^{\prime\prime}(x_{*}(y),\alpha)]e,e\big{)}|\Phi_{x}^{\prime}(x,\alpha)|^{2}\\ &=(\varkappa_{\mathcal{S}}(x)-\varkappa_{\mathcal{S}_{y}}(x)){|\Phi_{x}^{\prime}(x,\alpha)|^{2}}{|H_{x}^{\prime}(x)|},\end{split}$
where $e$ is a unit vector tangent to $\mathcal{S}$ at $x_{*}(y)$. Using assumption 3.3(2) proves the first statement.
Now we prove the second assertion in (4.5). Denote
(B.3) $e_{0}:=-\Phi_{x}^{\prime}(x_{*}(y),\alpha)/|\Phi_{x}^{\prime}(x_{*}(y),\alpha)|,\ V:=\begin{pmatrix}e_{0}&e\end{pmatrix}.$
Clearly, $V$ is a $2\times 2$ orthogonal matrix. By construction,
(B.4) $M_{1}:=\begin{pmatrix}1&0_{2}^{T}\\ 0_{2}&V^{T}\end{pmatrix}M\begin{pmatrix}1&0_{2}^{T}\\ 0_{2}&V\end{pmatrix}=\begin{pmatrix}0&|\Phi_{x}^{\prime}|&0\\ |\Phi_{x}^{\prime}|&*&*\\ 0&*&([H_{xx}^{\prime\prime}-\nu_{1*}\Phi_{xx}^{\prime\prime}]e,e)\end{pmatrix},$
$\text{sgn}M=\text{sgn}M_{1}$, and * denote various quantities, whose values are irrelevant. Next we use Haynsworth’s inertia additivity formula [18],
(B.5) $\begin{split}\text{sgn}M_{1}=&\text{sgn}\bigg{(}\big{(}\big{[}H_{xx}^{\prime\prime}-\nu_{1*}\Phi_{xx}^{\prime\prime}\big{]}e,e\big{)}-\begin{pmatrix}|\Phi_{x}^{\prime}|&0\end{pmatrix}\tilde{M}_{1}^{-1}\begin{pmatrix}|\Phi_{x}^{\prime}|\\ 0\end{pmatrix}\bigg{)}\\ &+\text{sgn}\tilde{M}_{1},\quad\tilde{M}_{1}:=\begin{pmatrix}0&|\Phi_{x}^{\prime}|\\ |\Phi_{x}^{\prime}|&*\end{pmatrix}.\end{split}$
An easy calculation shows that $\text{sgn}\tilde{M}_{1}=0$ and
(B.6) $\begin{split}\text{sgn}M_{1}=&\text{sgn}\big{(}\big{[}H_{xx}^{\prime\prime}-\nu_{1*}\Phi_{xx}^{\prime\prime}\big{]}e,e\big{)}=-1.\end{split}$
B.2. Proof of Lemma 4.2
Eq. (4.9) implies that $\mathcal{R}f\in I^{-3/2}(\mathcal{V},\Gamma)$ is a conormal (or, more generally, Lagrangian) distribution, see [22, Section 25.1]. By [22, Proposition 25.1.5], $\tilde{g}(\eta)=O(|\eta|^{-3/2})$, $|\eta|\to\infty$. Substituting this into (2.4) and using the equivalence of norms (2.3) and (2.4) gives $g\in C^{1/2}(\mathcal{V})$. As mentioned above, assumptions 3.1(1,3) imply that $\text{supp}\,g$ is compact.
The leading order term in (4.13) is obtained using [14, eq. 26, p. 360]:
(B.7) $\mathcal{F}((t-i0)^{a})=\frac{2\pi e(-a)}{\Gamma(-a)}\mu_{+}^{-(a+1)}.$
Using the decay of the lower order terms in $\tilde{\upsilon}$ as $\lambda\to\infty$ (see $\tilde{R}$ in (4.9) and [22, eq. (25.1.4)]), gives $\Delta g\in C_{0}^{3/2}(\mathcal{V})$.
It remains to prove the properties of the interpolated functions $g_{\epsilon}$ and $\Delta g_{\epsilon}$. Let $V\subset\mathbb{R}^{n}$ be any domain. Pick any $g\in C_{0}^{s}(V)$ and suppose first $10<s<$ 1. Let $\varphi\in C_{0}^{\lceil s\rceil}(\mathbb{R}^{n})$ be an interpolation kernel exact to degree $1$, and $g_{\epsilon}$ be the corresponding interpolated function. Denote $y_{h}:=y+h$, $\Delta g(y,h):=g(y_{h})-g(y)$ and similarly for $\Delta g_{\epsilon}(y,h)$. We have two equivalent representations
(B.8) $\begin{split}&\Delta g_{\epsilon}(y,h)-\Delta g(y,h)\\ &=\sum_{j}\varphi\bigg{(}\frac{y_{h}}{\epsilon}-j\bigg{)}\big{(}g(\epsilon j)-g(y_{h})\big{)}-\sum_{j}\varphi\bigg{(}\frac{y}{\epsilon}-j\bigg{)}\big{(}g(\epsilon j)-g(y)\big{)}\\ &=\sum_{j}\bigg{[}\varphi\bigg{(}\frac{y_{h}}{\epsilon}-j\bigg{)}-\varphi\bigg{(}\frac{y}{\epsilon}-j\bigg{)}\bigg{]}\big{(}g(\epsilon j)-g(y_{h})\big{)}\\ &\quad+\sum_{j}\varphi\bigg{(}\frac{y}{\epsilon}-j\bigg{)}\big{(}g(y_{h})-g(y)\big{)}.\end{split}$
We assume here for simplicity that $y_{j}=\epsilon j$, $j\in\mathbb{Z}^{2}$. When $h\geq\epsilon$, we use the second line above. Since $\varphi$ is compactly supported, the number of terms in each sum is finite. Moreover, $|\epsilon j-y_{h}|\leq c\epsilon$ in the first sum and $|\epsilon j-y|\leq c\epsilon$ in the second sum. Using that $g\in C_{0}^{s}(V)$, implies
(B.9) $|\Delta g_{\epsilon}(y,h)-\Delta g(y,h)|\leq ch^{s},\ h\geq\epsilon.$
When $0<h<\epsilon$, we use the second representation in (B.8). The kernel $\varphi$ is at least as smooth as $g$, so
(B.10) $\begin{split}\bigg{|}\frac{\varphi\big{(}(y_{h}/\epsilon)-j\big{)}-\varphi\big{(}(y/\epsilon)-j\big{)}}{(h/\epsilon)^{s}}\frac{g(\epsilon j)-g(y_{h})}{\epsilon^{s}}\bigg{|}<c.\end{split}$
The term on the last line in (B.8) is bounded by $ch^{s}$, and the assertion follows. The fact that $\|g_{\epsilon}\|_{L^{\infty}(V)}<c$ is obvious.
Suppose now $g\in C_{0}^{s}(V)$ and $21<s<$ 2. As above, $\|g_{\epsilon}\|_{L^{\infty}(V)}<c$. By the exactness of $\varphi$ up to order one, differentiating (3.13) gives
(B.11) $\begin{split}&g_{\epsilon}^{\prime}(y)=g^{\prime}(y)+\sum_{j}\varphi^{\prime}\bigg{(}\frac{y}{\epsilon}-j\bigg{)}\frac{g(\epsilon j)-(g(y)+(\epsilon j-y)\nabla g(y))}{\epsilon},\end{split}$
where all the primes stand for the same derivative $\partial_{y_{l}}$, $1\leq l\leq n$, and $\nabla g=(\partial_{y_{1}}g,\dots,\partial_{y_{n}}g)^{T}$. The fraction on the right is $O(\epsilon^{s-1})$, therefore $\|g_{\epsilon}\|_{C^{1}(V)}<c$, $0<\epsilon\ll 1$.
To prove $\|g_{\epsilon}^{\prime}\|_{C^{s-1}(V)}<c$, we argue similarly to (B.8). The case $h\geq\epsilon$ is completely analogous. If $0<h<\epsilon$, from (B.11) and the exactness of $\varphi$:
(B.12) $\begin{split}&\Delta g_{\epsilon}^{\prime}(y,h)-\Delta g^{\prime}(y,h)\\ &=\sum_{j}\bigg{[}\varphi^{\prime}\bigg{(}\frac{y_{h}}{\epsilon}-j\bigg{)}-\varphi^{\prime}\bigg{(}\frac{y}{\epsilon}-j\bigg{)}\bigg{]}\frac{g(\epsilon j)-[g(y_{h})+(\epsilon j-y_{h})\nabla g(y_{h})]}{\epsilon}\\ &\quad+\sum_{j}\varphi^{\prime}\bigg{(}\frac{y}{\epsilon}-j\bigg{)}\frac{\epsilon j-y}{\epsilon}[\nabla g(y_{h})-\nabla g(y)].\end{split}$
Dividing the first sum by $h^{s-1}=(h/\epsilon)^{s-1}\epsilon^{s-1}$ we get similarly to (B.10) that the ratio is bounded. The second sum divided by $h^{s-1}$ is bounded, because $g^{\prime}\in C^{s-1}(V)$. The proof is complete.
Appendix C Proof of Lemma 5.2
The idea of the proof is to reduce $f_{\epsilon}$ to $G_{\epsilon}$ by first neglecting smooth terms in $f_{\epsilon}$ and then by neglecting lower order terms of finite smoothness, which still do not contribute to the DTB. This is done in sections C.1 and C.2, respectively.
C.1. Simplification of $f_{\epsilon}$ by subtracting smooth terms
Let $\mathcal{T}_{\epsilon}$ denote the operator on the left in (5.6). By assumption 3.2, $\mathcal{R}^{*}\mathcal{R}\in L^{-1}(\mathcal{U})$ is elliptic. Clearly,
(C.1) $\mathcal{T}_{\epsilon}=\text{Op}(\chi(x)|\xi|)\chi(x)\text{Op}\big{(}\tilde{D}(x,\xi)+\kappa\epsilon^{3}|\xi|^{2}\big{)}\in L^{3}(\mathcal{U}).$
Let $\mathcal{B}_{\epsilon}\in L^{-3}(\mathcal{U}_{b})$ be a local parametrix for $\mathcal{T}_{\epsilon}$ on the open set $\{x\in\mathcal{U}_{b}:\chi(x)>1/2\}\supset\mathcal{U}_{0}$, see Figure 4. Below, we omit the part “$x\in\mathcal{U}_{b}$” in similar sets. Then
(C.2) $\begin{split}&f_{\epsilon}(x)=(\mathcal{B}_{\epsilon}F_{\epsilon})(x)+(\mathcal{H}_{\epsilon}f_{\epsilon})(x),\\ &\tilde{B}_{\epsilon}(x,\xi)-\big{(}|\xi|\tilde{D}(x,\xi)+\kappa|\epsilon\xi|^{3}\big{)}^{-1}\in S^{-1}(\mathcal{U}_{0}),\end{split}$
where $\tilde{B}_{\epsilon}\in S^{-3}(\mathcal{U}_{b})$ is the complete symbol of $\mathcal{B}_{\epsilon}$, and $\mathcal{H}_{\epsilon}:\mathcal{E}^{\prime}(\mathcal{U}_{b})\to C^{\infty}(\mathcal{U}_{0})$ is an operator with a smooth kernel, $H_{\epsilon}(x,z)$.
Figure 4. Illustration of the sets $\mathcal{U}_{0}$, $\mathcal{U}_{b}$, $\mathcal{U}$, $\{\chi(x)=1\}$, and the level set $\{\chi(x)=c\}$, $10<c<$ 1.
To prove the latter claim, use a partition of unity to write $f=f_{1}+f_{2}$, where
$\text{supp}\,f_{1}\subset\{\chi(x)>1/2\},\quad\text{supp}\,f_{2}\subset\{\chi(x)<3/4\}.$
Then $\mathcal{B}_{\epsilon}\mathcal{T}_{\epsilon}(f_{1}+f_{2})=f_{1}+\mathcal{H}_{\epsilon}^{\prime}f_{1}+\mathcal{B}_{\epsilon}\mathcal{T}_{\epsilon}f_{2}$, where $\mathcal{H}_{\epsilon}^{\prime}\in L^{-\infty}(\{\chi(x)>1/2\})$. The first two terms on the right follow by the properties of a parametrix. The fact that
$\mathcal{B}_{\epsilon}\mathcal{T}_{\epsilon}:\mathcal{E}^{\prime}(\{\chi(x)<3/4\})\to C^{\infty}(\mathcal{U}_{0})$
follows by the pseudolocality of $\Psi$DOs.
Derivatives of $H_{\epsilon}(x,z)$ with respect to $x$ and $z$ are uniformly bounded with respect to $x\in\mathcal{U}_{0}$, $z\in\mathcal{U}_{b}$, and $0<\epsilon\ll 1$. This follows from the way a symbol of $\mathcal{B}_{\epsilon}$ is constructed (see [52, Ch. I, eqs. (4.39), (4.40)]) and observing that all the seminorms of $\big{(}|\xi|\tilde{D}(x,\xi)+\kappa|\epsilon\xi|^{3}\big{)}^{-1}$ as member of $S^{0}(\mathcal{U}_{b})$ (i.e., the minimal constants $c_{m}$ and $c_{m_{1},m_{2}}$ in (2.7), where $r=0$) are uniformly bounded for $0<\epsilon\ll 1$.
Therefore, by Lemma 5.1,
[…]
C.2. Extracting the leading local singularity of $f_{\epsilon}$
In the preceding section we converted $f_{\epsilon}$ to $\mathcal{B}_{\epsilon}F_{\epsilon}$ modulo a smooth function. We now represent $\mathcal{B}_{\epsilon}=\mathcal{B}_{\epsilon}^{(0)}+\Delta\mathcal{B}_{\epsilon}$ with some convenient $\mathcal{B}_{\epsilon}^{(0)}\in L^{0}(\mathcal{U}_{b})$, $\Delta\mathcal{B}_{\epsilon}\in L^{-1}(\mathcal{U}_{b})$. Then we use the mapping properties of $\Delta\mathcal{B}_{\epsilon}$ to show that $\Delta\mathcal{B}_{\epsilon}F_{\epsilon}$ does not contribute to the DTB.
Using the results in [53, Section 6.2, Chapter VIII] and [11, Proposition 4.2.4, Section 4.2] and (3.5), the principal symbol of $\mathcal{R}^{*}\mathcal{R}$ is given by
(C.4) $\begin{split}\frac{2\pi}{|\nu|}\frac{W^{2}(x,y)}{|\Delta_{\Phi}(x,\alpha)|},\ y=y(x,\xi),\nu=\nu(x,\xi),\end{split}$
where the functions $y(x,\xi)\in\mathcal{V}$ and $\nu(x,\xi)\in\mathbb{R}$ are obtained by solving the equations:
(C.5) $p=\Phi(x,\alpha),\ \xi=-\nu\text{d}_{x}\Phi(x,\alpha).$
The existence of a solution follows from assumption 3.2(3), and its local uniqueness follows from assumption 3.2(1).
Define
(C.6) $\begin{split}\tilde{D}_{0}(x,\xi):=&2\pi\frac{|\xi|}{|\nu|}\frac{W^{2}(x,y)}{|\Delta_{\Phi}(x,\alpha)|},\ y=y(x,\xi),\nu=\nu(x,\xi),\\ (x,\xi)\in&\,\mathcal{U}\times(\mathbb{R}^{2}\setminus 0_{2}).\end{split}$
It is clear from (C.5) and (C.6) that
(C.7) $\begin{split}&\tilde{D}_{0}(x,\lambda\xi)=\tilde{D}_{0}(x,\xi),\ \lambda\not=0,x\in\mathcal{U},\xi\in\mathbb{R}^{2}\setminus 0_{2}.\end{split}$
Recall that $\tilde{D}$ is the complete symbol of $\mathcal{R}^{*}\mathcal{R}$. From (C.4) and (C.6),
(C.8) $|\xi|\tilde{D}(x,\xi)-\tilde{D}_{0}(x,\xi/|\xi|)\in S^{-1}(\mathcal{U}_{0}).$
Thus we can find $\Delta\tilde{B}_{\epsilon}\in S^{-1}(\mathcal{U}_{b})$ (with all seminorms uniformly bounded when $0<\epsilon\ll 1$) such that
(C.9) $\begin{split}&\tilde{B}_{\epsilon}(x,\xi)=\big{(}\tilde{D}_{0}(x,\xi/|\xi|)+\kappa|\epsilon\xi|^{3}\big{)}^{-1}+\Delta\tilde{B}_{\epsilon}(x,\xi),\ x\in\mathcal{U}_{0}.\end{split}$
Note that (LABEL:tilde_B) is required to hold only for $x\in\mathcal{U}_{0}$. Since $\mathcal{B}_{\epsilon}$ is a parametrix for $T_{\epsilon}$ on $\{\chi(x)>1/2\}$, we can select $\tilde{B}_{\epsilon}$ in (C.2) and $\Delta\tilde{B}_{\epsilon}$ so that
(C.10) $\begin{split}\tilde{B}_{\epsilon}(x,\xi)\equiv 0,\ \Delta\tilde{B}_{\epsilon}(x,\xi)\equiv 0,\ x\in\{\chi(x)<1/4\}.\end{split}$
From (5.6),
(C.11) $\text{Op}\big{(}\Delta\tilde{B}_{\epsilon}(x,\xi)\big{)}F_{\epsilon}=\mathcal{W}_{\epsilon}g_{\epsilon},\ \mathcal{W}_{\epsilon}:=\text{Op}\big{(}\Delta\tilde{B}_{\epsilon}(x,\xi)\big{)}\text{Op}(\chi(x)|\xi|)\chi(x)\mathcal{R}^{*}.$
By construction, $\mathcal{W}_{\epsilon}$ is an FIO of order $-1/2$, $\mathcal{W}_{\epsilon}\in I^{-1/2}(\mathcal{U}\times\mathcal{V}\times\mathcal{C}^{*})$. The canonical relation $\mathcal{C}^{*}$ is the image of $\mathcal{C}$ (see (3.3)) under the exchange of $T^{*}\mathcal{U}$ and $T^{*}\mathcal{V}$. By (LABEL:tilde_B), all the seminorms of a symbol of $\mathcal{W}_{\epsilon}$, $W_{\epsilon}\in S^{0}(\mathcal{U}\times\mathcal{V}\times\mathbb{R})$, $0<\epsilon\ll 1$, are bounded.
Recall that $B_{\infty,\infty}^{s}(\mathbb{R}^{n})=C^{s}(\mathbb{R}^{n})$, $s>0$, $s\not\in\mathbb{N}$ [1, Remark 6.4(2)]. We use [23, Proposition 5.5] with $p,q=\infty$, which asserts that
(C.12) $\mathcal{W}_{\epsilon}:C^{s-(1/2)-m_{c}(\infty)}(\mathbb{R}^{n})\to C^{s}(\mathbb{R}^{n}),\ m_{c}(p)=-(n-1)\big{|}(1/p)-(1/2)\big{|},$
is continuous. Here $s-(1/2)-m_{c}(\infty)>0$ and $s>0$. See also [45, Theorem, p. 173]. For the validity of (C.12) it is required that the phase function of $\mathcal{W}_{\epsilon}$ be non-degenerate (see the paragraph following (3.2)) and the amplitude be compactly supported in $x$ (see (C.10)).
Clearly, $m_{c}(\infty)=-1/2$. Using (C.12) with $s=1/2$ and that $g_{\epsilon}\in C_{0}^{1/2}(\mathcal{V})$, $\|g_{\epsilon}\|_{C^{1/2}(\mathcal{U}_{b})}<c$, $0<\epsilon\ll 1$ (see Lemma 4.2), we conclude that $\mathcal{W}_{\epsilon}g_{\epsilon}\in C_{0}^{1/2}(\mathcal{U}_{b})$ and $\|\mathcal{W}_{\epsilon}g_{\epsilon}\|_{C^{1/2}(\mathcal{U}_{b})}<c$, $0<\epsilon\ll 1$. Therefore, by (2.3),
(C.13) $\|\mathcal{W}_{\epsilon}g_{\epsilon}\|_{L^{\infty}(\mathcal{U}_{b})}<c,0<\epsilon\ll 1;\ |\mathcal{W}_{\epsilon}g_{\epsilon}(x_{0}+\epsilon\check{x})-\mathcal{W}_{\epsilon}g_{\epsilon}(x_{0})|=O(\epsilon^{1/2})$
when $\check{x}$ is confined to any bounded set.
Combining (C.3) and (C.13) finishes the proof.
Appendix D Proof of Lemmas 6.2, 6.3
D.1. Proof of Lemma 6.2
We begin by stating a lemma. Its proof is in appendix D.2.
Lemma D.1.
Suppose assumptions 3.13.4 are satisfied. Pick any $A>0$. One has
(D.1) $H(\check{t},\check{q};\cdot)\equiv 0,\ \check{t}<-c.$
Also, for some smooth and bounded $\varkappa(t,q;\cdot)$,
(D.2) $\begin{split}&\partial_{\check{t}}^{k}\big{[}H(\check{t},\check{q};\cdot)-\varkappa(\epsilon(\check{t}+\check{q});\cdot)\check{t}^{-1/2}\big{]}=O(\check{t}^{-3/2}),\\ &k=0,1,\ \check{t}\to+\infty,\ |\check{t}|<A/\epsilon,\end{split}$
and
(D.3) $\begin{split}&\partial_{\check{t}}\big{[}\varkappa(\epsilon(\check{t}+\check{q});\cdot)\check{t}^{-(1/2)}\big{]}=O(\check{t}^{-3/2}),\ \check{t}\to+\infty,\ |\check{t}|<A/\epsilon.\end{split}$
The big-$O$ terms are uniform with respect to $\check{q}$ and the second group of variables collectively denoted $\cdot$’, namely $\theta,\epsilon$. These variables are restricted to the sets indicated in Lemma 6.2.
For simplicity, we drop the arguments $\check{q}$, $\theta$ and $\epsilon$ from all the functions. The limit as $\check{t}\to-\infty$ in (6.16) follows immediately from (D.1). Next, suppose $\check{t}\to\infty$. Then
(D.4) $\begin{split}&G(\check{t})=\int_{\mathbb{R}}\chi_{1\text{D}}(\epsilon(s-\check{t}))\frac{H_{1}(s)}{s-\check{t}}\text{d}s+\int_{0}^{\infty}\chi_{1\text{D}}(\epsilon(s-\check{t}))\frac{\varkappa(\epsilon(s+\check{q}))}{(s-\check{t})s^{1/2}}\text{d}s,\\ &H_{1}(s):=H(s)-\varkappa(\epsilon(s+\check{q}))s_{+}^{-1/2},\end{split}$
where $\varkappa(\cdot)$ is the same as in (LABEL:H_props_2_a), (LABEL:H_props_2_b). In view of (D.1), the first of the integrals is split into five
(D.5) $\begin{split}&\bigg{(}\int_{-c}^{0}+\int_{0}^{1}+\int_{1}^{\check{t}-1}+\int_{\check{t}-1}^{\check{t}+1}+\int_{\check{t}+1}^{\infty}\bigg{)}\chi_{1\text{D}}(\epsilon(s-\check{t}))\frac{H_{1}(s)}{s-\check{t}}\text{d}s\\ &=J_{1}+\dots+J_{5}.\end{split}$
By (LABEL:H_props_2_a) with $k=0$, $H_{1}(s)=O(s^{-3/2})$, $s\to\infty$. It is trivial to see that $J_{1},J_{2}=O(1/\check{t})$. Breaking up the integral over $[1,\check{t}-1]$ into the integrals over $[1,\check{t}/2]$ and $[\check{t}/2,\check{t}-1]$ we get $J_{3}=O(1/\check{t})$. Breaking up the integral over $[\check{t}+1,\infty)$ into the integrals over $[\check{t}+1,2\check{t}]$ and $[2\check{t},\infty)$ gives $J_{5}=O(1/\check{t})$. Finally, using (LABEL:H_props_2_a) with $k=1$ and that $|s-\check{t}|\leq 1$, $\epsilon\check{t}\leq c$, we get
(D.6) $\partial_{s}\big{[}\chi_{1\text{D}}(\epsilon(s-\check{t}))H_{1}(s)\big{]}=O(\check{t}^{-3/2}),$
which implies $J_{4}=O(\check{t}^{-3/2})$. Combining the five estimates proves that the first integral in (D.4) is $O(1/\check{t})$.
Consider the second integral in (D.4). Recall that $\check{t}\to\infty$. Write the integral in the form
(D.7) $\begin{split}J(\check{t},\epsilon):=&\int_{0}^{\infty}\frac{\chi_{1\text{D}}(\epsilon(\check{s}-\check{t}))\varkappa(\epsilon(\check{s}+\check{q}))}{(\check{s}-\check{t})\check{s}^{1/2}}\text{d}\check{s}=\epsilon^{1/2}\int_{0}^{\infty}\frac{\chi_{1\text{D}}(s-t)\varkappa(s+q)}{(s-t)s^{1/2}}\text{d}s\\ =&\epsilon^{1/2}\int_{0}^{\infty}\frac{(\chi_{1\text{D}}(s-t)-\chi_{1\text{D}}(0))\varkappa(s+q)}{(s-t)s^{1/2}}\text{d}s\\ &+\epsilon^{1/2}\chi_{1\text{D}}(0)\int_{0}^{\infty}\frac{\varkappa(s+q)}{(s-t)s^{1/2}}\text{d}s,\ q=\epsilon\check{q}.\end{split}$
The integral on the second line is clearly uniformly bounded. The identity [43, Eq. 2.2.5.26] with $\alpha=1/2$:
(D.8) $\int_{0}^{\infty}\frac{\text{d}s}{(s-t)s^{1/2}}=0,\ t>0,$
implies
(D.9) $\int_{0}^{\infty}\frac{\varkappa(s+q)}{(s-t)s^{1/2}}\text{d}s=\int_{0}^{\infty}\frac{\varkappa(s+q)-\varkappa(t+q)}{(s-t)s^{1/2}}\text{d}s=O(1),$
Here the term $O(1)$ is uniform with respect to $t,q\in\mathbb{R}$. Therefore $J(\check{t},\epsilon)=O(\epsilon^{1/2})$. Combining the results proves (6.16).
D.2. Proof of Lemma D.1
Assertion (D.1) is obvious from (6.11). To prove (LABEL:H_props_2_a), (LABEL:H_props_2_b), begin by proving the following auxiliary lemma.
Lemma D.2.
Pick any $A>0$ and a function $f\in C^{2}(\mathbb{R})$. Let $\varphi\in C_{0}^{2}(\mathbb{R})$ be an interpolating kernel, which is exact up to degree one. One has
(D.10) $\begin{split}\partial_{t}^{k}\big{[}&\textstyle\sum_{j}\varphi(t+b-j)f(\epsilon j)(j-b)_{+}^{1/2}-f(\epsilon(t+b))t^{1/2}\big{]}=O\big{(}t^{-3/2}\big{)},\\ &t\to+\infty,\ \epsilon\to 0,\ \epsilon|t|<A,\ k=0,1,2.\end{split}$
The big-$O$ term is uniform in $b\in\mathbb{R}$.
Proof.
Denote $g(t):=f(\epsilon t)(t-b)^{1/2}$ and $u:=t+b$. Taylor expanding $g$ gives
(D.11) $\begin{split}&g(j)=g(u)+g^{\prime}(u)(j-u)+R_{1}(j,u),\\ &R_{1}(j,u)=\int_{u}^{j}g^{\prime\prime}(s)(j-s)\text{d}s.\end{split}$
The exactness of $\varphi$ up to degree one implies:
(D.12) $\sum_{j}\varphi(u-j)g(j)=g(u)+\textstyle\sum_{j}\varphi(u-j)R_{1}(j,u).$
Clearly
(D.13) $|g^{\prime\prime}(u)|=O\big{(}t^{-3/2}\big{)},\ \epsilon|t|<A,\ t\to\infty.$
Hence, by (D.11)
(D.14) $\begin{split}&R_{1}(j,u)=O(t^{-3/2}),\ |j-u|<A,\ \epsilon|t|<A,\ t\to\infty.\end{split}$
Furthermore, using that $\varphi$ is exact up to order one,
(D.15) $\begin{split}\partial_{u}R_{1}(j,u)=&-g^{\prime\prime}(u)(j-u),\\ \partial_{u}\textstyle\sum_{j}\varphi(u-j)R_{1}(j,u)=&\textstyle\sum_{j}\varphi^{\prime}(u-j)R_{1}(j,u)-g^{\prime\prime}(u)\textstyle\sum_{j}\varphi(u-j)(j-u)\\ \hskip 56.9055pt=&\textstyle\sum_{j}\varphi^{\prime}(u-j)R_{1}(j,u),\end{split}$
and
(D.16) $\begin{split}\partial_{u}^{2}\textstyle\sum_{j}\varphi(u-j)&R_{1}(j,u)\\ =&\textstyle\sum_{j}\varphi^{\prime\prime}(u-j)R_{1}(j,u)-g^{\prime\prime}(u)\textstyle\sum_{j}\varphi^{\prime}(u-j)(j-u)\\ \hskip 56.9055pt=&\textstyle\sum_{j}\varphi^{\prime\prime}(u-j)R_{1}(j,u)-g^{\prime\prime}(u).\end{split}$
Combining (D.12)–(D.16) and using that $\varphi^{\prime\prime}\in L^{\infty}(\mathbb{R})$ finishes the proof. ∎
To prove (LABEL:H_props_2_a), by (6.10) and (6.11) we consider
(D.17) $H_{aux}(\check{t}):=\textstyle\sum_{j_{2}}\varphi_{p}(\check{t}+\check{q}-j_{2})a_{0}(\theta,\epsilon j_{2})(j_{2}-\check{q})_{+}^{1/2}.$
By Lemma D.2,
(D.18) $\partial_{\check{t}}^{k-1}\left(\partial_{\check{t}}H_{aux}(\check{t})-\partial_{\check{t}}\big{[}a_{0}(\theta,\epsilon(\check{t}+\check{q}))\check{t}^{1/2}]\right)=O(\check{t}^{-3/2}),\ k=1,2.$
We moved one derivative inside the parentheses because $H$ is defined in terms of $h^{\prime}$ in (6.11). Replacing $\check{t}$ with $[\Phi(\theta+\epsilon\mu s)-\Phi(\theta)]/\epsilon+\check{t}+\check{q}$, multiplying both sides by $K(\theta+\epsilon\mu s)\varphi_{\alpha}(s)$, and integrating with respect to $s$ (cf. (6.10), (6.11)) finishes the proof. This also gives a formula for $\varkappa$. Our derivation shows that $\varkappa$ is smooth and bounded with all derivatives, but the formula for $\varkappa$ is not needed.
The claim (LABEL:H_props_2_b) immediately follows because $\varkappa$ is smooth and $\epsilon|\check{t}|<A$.
D.3. Proof of Lemma 6.3
For simplicity, in what follows we omit the dependence of $G(\check{t},\check{q};\theta,\epsilon)$ on $\check{q},\theta$, and $\epsilon$, and denote $Q^{\prime}:=Q_{\alpha}^{\prime}$ and $Q^{\prime\prime}:=Q_{\alpha\alpha}^{\prime\prime}$. By (6.15) we study the convergence of the sum
(D.19) $J_{\epsilon}:=\epsilon^{1/2}\sum_{\alpha_{j_{1}}\in\text{dom}\,P}G(Q(\alpha_{j_{1}})/\epsilon).$
Recall that $Q$ depends on $x$, which is confined to a sufficiently small open set $U^{\prime}$ (see the paragraph following (6.7)), and on the selected $P(\alpha)$.
If $|Q(\alpha)|\geq c$ on $\text{dom}\,P$, using (6.16) in (D.19) gives for each of the partial sums, over $Q(\alpha_{j_{1}})>0$ and $Q(\alpha_{j_{1}})<0$, $|J_{\epsilon}|\leq c\epsilon^{1/2}(1/\epsilon)\epsilon^{1/2}=O(1)$. Thus, in what follows we assume $Q(a_{0})=0$ for some $a_{0}\in\text{dom}\,P$ and replace $\text{dom}\,P$ in (D.19) with $I_{\alpha}^{\prime}$, which is a sufficiently small neighborhood of $a_{0}$. If $Q^{\prime}(\alpha)\not=0$ on $I_{\alpha}^{\prime}$, then $|Q^{\prime}(\alpha)|\asymp|\alpha-a_{0}|$ on $I_{\alpha}^{\prime}$ and
(D.20) $|J_{\epsilon}|\leq c\epsilon^{1/2}\sum_{j_{1}=1}^{1/\epsilon}\big{(}j_{1}^{-1/2}+\epsilon^{1/2}+j_{1}^{-1}\big{)}<c,\ \epsilon\to 0.$
Hence, in what follows, we assume that $Q^{\prime}(\tilde{\alpha})=0$ for some $\tilde{\alpha}\in I_{\alpha}^{\prime}$ and $\tilde{x}\in U^{\prime}$.
As is easily seen, $Q(\tilde{\alpha})=0$ and $Q^{\prime}(\tilde{\alpha})=0$ for some $\tilde{\alpha}\in\text{dom}\,P$ and $\tilde{x}\in U^{\prime}$ if and only if $\tilde{x}\in\mathcal{S}$ and $P(\alpha)$ is associated with $\tilde{x}$. The necessity is established in Lemma D.3 below, and the sufficiency follows from the Bolker condition, assumption 3.2(2). Recall that due to possible data redundancy, there can be several $P(\alpha)$ associated with $\tilde{x}$. In the rest of this subsection we select one of them. We need the following lemma, which is proven in appendix D.4.
Lemma D.3.
Suppose assumptions 3.1, 3.2, and 3.3 are satisfied. Pick any $\tilde{y}\in\Gamma$. Let $\mathcal{S}_{\tilde{y}}$ be tangent to $\mathcal{S}$ at $\tilde{x}$. Let $x(\alpha)$ be the solution to (4.8) associated with $\tilde{x}$ which satisfies $\tilde{x}=x(\tilde{\alpha})$, see Figure 1. One has
(D.21) $Q_{\alpha}^{\prime}(\tilde{x},\tilde{\alpha})=0,\ Q_{\alpha\alpha}^{\prime\prime}(\tilde{x},\tilde{\alpha})=\det M(\tilde{y})\frac{|x_{\alpha}^{\prime}(\tilde{\alpha})|^{2}}{|\Phi_{x}^{\prime}(\tilde{x},\tilde{\alpha})||H_{x}^{\prime}(\tilde{x})|}>0$
and
(D.22) $|\Delta_{\Phi}(\tilde{x},\tilde{\alpha})|=|\det M(\tilde{y})|\frac{|x_{\alpha}^{\prime}(\tilde{\alpha})|}{|H_{x}^{\prime}(\tilde{x})|}.$
By continuity, $Q_{\alpha\alpha}^{\prime\prime}(x,\alpha)>0$ for all $(x,\alpha)\in U^{\prime}\times I_{\alpha}^{\prime}$.
D.3.1. Critical point in $I_{\alpha}^{\prime}$
Fix any $x\in U^{\prime}$. Suppose $Q^{\prime}(a_{1})=0$ for some $a_{1}\in I_{\alpha}^{\prime}$. By shifting $\alpha$, we can assume without loss of generality that $a_{1}=0$ and
(D.23) $Q(\theta)=-h+\int_{0}^{\theta}(\theta-s)Q^{\prime\prime}(s)\text{d}s$
for some $h$. The value of $h$ is determined from $Q(a_{0})=0$. Since $Q^{\prime\prime}(\alpha)\asymp 1$ on $I_{\alpha}^{\prime}$ (see the sentence following Lemma D.3) and $Q(a)=0$, we have $h\geq 0$.
For any $c>0$, the number of $\alpha_{j_{1}}$ inside the set $\{\alpha\in I_{\alpha}^{\prime}:|Q(\alpha)|\leq c\epsilon\}$ is $O(\epsilon^{-1/2})$ and is uniformly bounded with respect to $a_{0}\in I_{\alpha}^{\prime}$. Since $G$ is bounded, the total contribution of such $\alpha_{j_{1}}$ to $J_{\epsilon}$ is bounded. Hence it remains to estimate the following two sums
(D.24) $\begin{split}J_{\epsilon}^{+}:=&\epsilon^{1/2}\sum_{\alpha_{j_{1}}\in I_{\alpha}^{\prime},Q(\alpha_{j_{1}})>\epsilon}\bigg{[}\epsilon^{1/2}+\frac{\epsilon}{Q(\alpha_{j_{1}})}\bigg{]},\\ J_{\epsilon}^{-}:=&\epsilon^{1/2}\sum_{\alpha_{j_{1}}\in I_{\alpha}^{\prime},Q(\alpha_{j_{1}})<-\epsilon}\bigg{(}-\frac{\epsilon}{Q(\alpha_{j_{1}})}\bigg{)}^{1/2}.\end{split}$
Begin with $J_{\epsilon}^{+}$. Clearly the sums over $\alpha_{j_{1}}>0$ and $\alpha_{j_{1}}<0$ lead to the same upper bound, so we consider only $\alpha_{j_{1}}>0$. Also, for estimation purposes, we can assume $Q^{\prime\prime}(s)\geq 1$, $s\in I_{\alpha}^{\prime}$ (rather than the more cumbersome $Q^{\prime\prime}(s)\geq c$). Let $\theta_{1}>0$ satisfy
(D.25) $Q(\theta_{1})=-h+\int_{0}^{\theta_{1}}(\theta_{1}-s)Q^{\prime\prime}(s)\text{d}s=\epsilon.$
For any $\theta\geq\theta_{1}$ one has
(D.26) $Q(\theta)\geq-h+\int_{0}^{\theta_{1}}(\theta_{1}-s)Q^{\prime\prime}(s)\text{d}s+\int_{\theta_{1}}^{\theta}(\theta-s)\text{d}s=\epsilon+(\theta-\theta_{1})^{2}/2.$
Since the shift $\theta\to\theta-\theta_{1}$ does not affect the upper bound we get
(D.27) $\begin{split}|J_{\epsilon}^{+}|\leq&c\epsilon^{1/2}\sum_{j_{1}=0}^{1/\epsilon}\bigg{[}\epsilon^{1/2}+\frac{\epsilon}{\epsilon+(\epsilon j_{1})^{2}}\bigg{]}\leq c+c\epsilon^{1/2}\int_{0}^{1}\frac{\text{d}x}{\epsilon+x^{2}}=O(1).\end{split}$
Next, consider $J_{\epsilon}^{-}$. If $h<\epsilon$, then $Q(\theta)\geq-\epsilon$ for all $\alpha\in I_{\alpha}^{\prime}$ and $J_{\epsilon}^{-}=0$. Hence we can assume $h>\epsilon$. In this case we initially assume for simplicity that $Q(0)=Q(a)=-\epsilon$ for some $a>0$. Therefore
(D.28) $\begin{split}-(Q(\theta)+\epsilon)=&-\int_{0}^{\theta}(\theta-s)Q^{\prime\prime}(s)\text{d}s+\theta\int_{0}^{a}(1-(s/a))Q^{\prime\prime}(s)\text{d}s\\ =&\int_{0}^{\theta}s(1-(\theta/a))Q^{\prime\prime}(s)\text{d}s+\theta\int_{\theta}^{a}(1-(s/a))Q^{\prime\prime}(s)\text{d}s.\end{split}$
Both integrands on the second line are positive, so the minimum on the right is achieved when $Q^{\prime\prime}(s)\equiv 1$. Hence, shifting the argument of $Q$ again we can assume $Q(\theta)=-h+\theta^{2}/2$. Similarly to (D.27),
(D.29) $\begin{split}|J_{\epsilon}^{-}|\leq&c\epsilon^{1/2}\sum_{h-(\epsilon j_{1})^{2}>\epsilon}\bigg{[}\frac{\epsilon}{h-(\epsilon j_{1})^{2}}\bigg{]}^{1/2}\leq c\int_{0}^{h^{1/2}}\frac{\text{d}x}{(h-x^{2})^{1/2}}<\infty.\end{split}$
Note that the value of the last integral is independent of $h>0$.
It is clear that in the above arguments all the bounds can be made uniform with respect to $x\in U^{\prime}$.
D.4. Proof of Lemma D.3
Recall that $P(\alpha)=\Phi(x(\alpha),\alpha)$ (see the paragraph below (4.8)). Therefore
(D.30) $Q(\alpha)=\Phi(\tilde{x},\alpha)-P(\alpha)=\Phi(\tilde{x},\alpha)-\Phi(x(\alpha),\alpha).$
The first claim in (D.21) follows by differentiating (4.8) and (D.30) with respect to $\alpha$ and using that $x^{\prime}(\alpha)$ is tangent to $\mathcal{S}$, $x(\tilde{\alpha})=\tilde{x}$, and $\text{d}_{x}\Phi(\tilde{x},\tilde{\alpha})$ is conormal to $\mathcal{S}$ at $\tilde{x}$.
Differentiating (D.30) with respect to $\alpha$ twice and using (6.7) gives:
(D.31) $Q_{\alpha\alpha}^{\prime\prime}=-\big{[}(\Phi_{xx}^{\prime\prime}x_{\alpha}^{\prime},x_{\alpha}^{\prime})+2\Phi_{x\alpha}^{\prime\prime}x_{\alpha}^{\prime}+\Phi_{x}^{\prime}x_{\alpha\alpha}^{\prime\prime}\big{]}.$
Differentiate the second equation in (4.8) with respect to $\alpha$:
(D.32) $H_{xx}^{\prime\prime}x_{\alpha}^{\prime}=\nu_{1}^{\prime}\Phi_{x}^{\prime}+\nu_{1}\big{[}\Phi_{xx}^{\prime\prime}x_{\alpha}^{\prime}+\Phi_{x\alpha}^{\prime\prime}\big{]}.$
Compute the dot product of (D.32) with $x_{\alpha}^{\prime}$ and use that $\Phi_{x}^{\prime}x_{\alpha}^{\prime}\equiv H_{x}^{\prime}x_{\alpha}^{\prime}\equiv 0$ (which follows from (4.8) to get
(D.33) $(H_{xx}^{\prime\prime}x_{\alpha}^{\prime},x_{\alpha}^{\prime})=\nu_{1}\big{[}(\Phi_{xx}^{\prime\prime}x_{\alpha}^{\prime},x_{\alpha}^{\prime})+\Phi_{x\alpha}^{\prime\prime}x_{\alpha}^{\prime}\big{]}.$
Differentiate the identity $\Phi_{x}^{\prime}x_{\alpha}^{\prime}\equiv 0$ to get
(D.34) $(\Phi_{xx}^{\prime\prime}x_{\alpha}^{\prime},x_{\alpha}^{\prime})+\Phi_{x\alpha}^{\prime\prime}x_{\alpha}^{\prime}+\Phi_{x}^{\prime}x_{\alpha\alpha}^{\prime\prime}\equiv 0.$
Using (D.33) and (D.34) in (D.31) and applying Lemma 4.1 finally gives:
(D.35) $\begin{split}Q_{\alpha\alpha}^{\prime\prime}=&-\big{[}(1/\nu_{1})(H_{xx}^{\prime\prime}x_{\alpha}^{\prime},x_{\alpha}^{\prime})-(\Phi_{xx}^{\prime\prime}x_{\alpha}^{\prime},x_{\alpha}^{\prime})\big{]}\\ =&-\big{(}\big{[}H_{xx}^{\prime\prime}-\nu_{1}\Phi_{xx}^{\prime\prime}\big{]}e,e\big{)}\frac{|x_{\alpha}^{\prime}|^{2}}{\nu_{1}},\ e:=x_{\alpha}^{\prime}/|x_{\alpha}^{\prime}|.\end{split}$
Combining this with (B.1) and (B.2) proves the second claim in (D.21).
Next we prove (D.22). Similarly to (B.2),
(D.36) $|\Delta_{\Phi}(x,\alpha)|=|\Phi_{x}^{\prime}||\Phi_{x\alpha}^{\prime\prime}e|.$
Using (D.33) and then (B.2) gives
(D.37) $\begin{split}|\Phi_{x}^{\prime}||\Phi_{x\alpha}^{\prime\prime}e|&=\big{|}\big{(}[H_{xx}^{\prime\prime}-\nu_{1}\Phi_{xx}^{\prime\prime}]e,e\big{)}\big{|}\frac{|x_{\alpha}^{\prime}||\Phi_{x}^{\prime}|}{|\nu_{1}|}=\frac{|\text{det}M||x_{\alpha}^{\prime}|}{|\nu_{1}||\Phi_{x}^{\prime}|},\end{split}$
and (D.22) follows from (B.1).
Appendix E Proof of Lemmas 7.1, 7.2, and 7.3
E.1. Proof of Lemma 7.1
For convenience, we drop the arguments $\check{q}$, $\theta$ and $\epsilon$ from all the functions. By (LABEL:H_props_2_a) with $k=1$ and (LABEL:H_props_2_b), $\Delta H(r,\check{t})=O(\check{t}^{-3/2})$, $\check{t}\to+\infty$. The statement $\Delta G=O(1/|\check{t}|)$, $\check{t}\to-\infty$, follows immediately from (D.1). Next, suppose $\check{t}\to+\infty$. Then
(E.1) $\begin{split}&\Delta G(r,\check{t})=\int_{\mathbb{R}}\chi_{1\text{D}}(\epsilon(s-\check{t}))\frac{\Delta H(r,s)}{s-\check{t}}\text{d}s.\end{split}$
In view of (D.1), the integral in (E.1) is split into three
(E.2) $\begin{split}&\bigg{(}\int_{-c}^{\check{t}-1}+\int_{\check{t}-1}^{\check{t}+1}+\int_{\check{t}+1}^{\infty}\bigg{)}\chi_{1\text{D}}(\epsilon(s-\check{t}))\frac{\Delta H(r,s)}{s-\check{t}}\text{d}s\\ &=J_{1}+J_{2}+J_{3}.\end{split}$
Breaking up the integral over $[-c,\check{t}-1]$ into the integrals over $[-c,\check{t}/2]$ and $[\check{t}/2,\check{t}-1]$ we get $J_{1}=O(1/\check{t})$. Breaking up the integral over $[\check{t}+1,\infty)$ into the integrals over $[\check{t}+1,2\check{t}]$ and $[2\check{t},\infty)$ gives $J_{3}=O(1/\check{t})$.
Also, from (LABEL:H_props_2_a) with $k=1$ and (LABEL:H_props_2_b) we get
(E.3) $|H_{\check{t}}^{\prime}(\check{t}+r)-H_{\check{t}}^{\prime}(\check{t})|\leq O(\check{t}^{-3/2})+c\big{|}\partial_{\check{t}}\big{[}\chi(\epsilon(\check{t}+\check{q}))\check{t}^{-1/2}\big{]}\big{|}=O(\check{t}^{-3/2}).$
Therefore, using $|s-\check{t}|\leq 1$ gives
(E.4) $\partial_{s}\big{[}\chi_{1\text{D}}(\epsilon(s-\check{t}))\Delta H(r,s)\big{]}=\epsilon O(\check{t}^{-3/2})+O(\check{t}^{-3/2}),\ \check{t}\to+\infty.$
Combining the three estimates proves that the integral in (E.1) is $O(1/\check{t})$. Combining the results finishes the proof of the estimate of $\Delta G$.
The derivative $\partial_{r}\Delta G(r,\check{t})$ in (6.16) is estimated in exactly the same fashion. In particular, we use $\partial_{r}\Delta H(r,\check{t})=\partial_{r}H(r+\check{t})=O(\check{t}^{-3/2})$ as $\check{t}\to+\infty$.
To prove (7.6), use (E.1) to write
(E.5) $\begin{split}&\Delta G(r,\check{t}+h)-\Delta G(r,\check{t})\\ &=\bigg{(}\int_{-1}^{1}+\int_{|s|>1}\bigg{)}\bigg{[}\frac{\chi_{1\text{D}}(\epsilon(s-h))}{s-h}-\frac{\chi_{1\text{D}}(\epsilon s)}{s}\bigg{]}\Delta H(r,s+\check{t})\text{d}s\\ &=:J_{1}(\check{t},h)+J_{2}(\check{t},h).\end{split}$
When $0<\epsilon\ll 1$ and $|s|<1$, we have $\chi_{1\text{D}}(\epsilon(s-h))\equiv\chi_{1\text{D}}(\epsilon s)\equiv 1$. Then
(E.6) $\begin{split}J_{1}(\check{t},h)&=\bigg{[}\ln\bigg{|}\frac{s-h}{s}\bigg{|}\Delta H(r,s+\check{t})\bigg{]}\bigg{|}_{s=-1}^{1}-\int_{-1}^{1}\Delta H_{\check{t}}^{\prime}(r,s+\check{t})\ln\bigg{|}\frac{s-h}{s}\bigg{|}\text{d}s\\ &=:J_{11}(\check{t},h)+J_{12}(\check{t},h).\end{split}$
By (LABEL:H_props_2_a) with $k=0$ and (LABEL:H_props_2_b),
(E.7) $\begin{split}\bigg{|}\ln\bigg{|}\frac{s-h}{s}\bigg{|}\Delta H(r,s+\check{t})\bigg{|}\leq c\frac{h\ln(1/h)}{(1+|\check{t}|)^{3/2}},\ s=\pm 1.\end{split}$
To bound all the remaining terms, we can look at them individually instead of estimating differences at $\check{t}+r$ and $\check{t}$ as above. Indeed,
(E.8) $\begin{split}|J_{12}(\check{t},h)|&\leq\int_{-1}^{1}\big{|}H^{\prime}(s+\check{t})\ln|1-(h/s)|\big{|}\text{d}s\\ &\leq c\int_{-1}^{1}(1+|s+\check{t}|)^{-3/2}\big{|}\ln|1-(h/s)|\big{|}\text{d}s\\ &\leq c\frac{h}{(1+|\check{t}|)^{3/2}}\int_{-1/h}^{1/h}\big{|}\ln|1-(1/u)|\big{|}\text{d}u\leq c\frac{h\ln(1/h)}{(1+|\check{t}|)^{3/2}}\end{split}$
and
(E.9) $\begin{split}&|J_{2}(\check{t},h)|\leq ch\int_{|s|>1}\bigg{(}\frac{1}{s^{2}}+\frac{\epsilon}{|s|}\bigg{)}\frac{1}{(1+|s+\check{t}|)^{3/2}}\text{d}s\\ &\leq ch(|\check{t}|^{-3/2}+\epsilon|\check{t}|^{-1}).\end{split}$
Combining the results finishes the proof.
E.2. Proof of Lemma 7.2
First we investigate the dependence of $G$ (which is implicitly used in (7.14) as part of $\Delta G$) on $\epsilon$ in its last argument. We need the following technical result, which is proven at the end of this section. Recall that $G_{0}$ is defined in (7.16), and $G$ – in (6.14).
Lemma E.1.
Pick any $A_{1}>0$. Suppose $|\check{t}|<A_{1}$, $|\epsilon\check{q}-P(0)|<\epsilon^{1/2}A_{1}$, and $|\theta|<\epsilon^{1/2}A_{1}$. Under the assumptions of Theorem 3.6 one has
(E.10) $\begin{split}G(\check{t},\check{q};\theta,\epsilon)=\mu G_{0}(\check{t},\check{q})+O(\epsilon^{1/2}).\end{split}$
Partition the interval $[-A,A]$ into the union of $K\gg 1$ non-overlapping intervals of length $2A/K$. Let these intervals be denoted $B_{k}$, and $s_{k}$ be their centers, $k=1,2,\dots,K$. Recall that $P(\alpha)$ is associated with $x_{0}\in\mathcal{S}$. Clearly,
(E.11) $\begin{split}\frac{P(\epsilon^{1/2}s)}{\epsilon}=&\frac{P(\epsilon^{1/2}s_{k})}{\epsilon}+P^{\prime}(\epsilon^{1/2}s_{k})\frac{s-s_{k}}{\epsilon^{1/2}}+O(1/K^{2})\\ =&\frac{P(\epsilon^{1/2}s_{k})}{\epsilon}+\big{[}P^{\prime}(0)+O(\epsilon^{1/2}s_{k})\big{]}\frac{s-s_{k}}{\epsilon^{1/2}}+O(1/K^{2})\\ =&\frac{P(\epsilon^{1/2}s_{k})}{\epsilon}+P^{\prime}(0)\frac{s-s_{k}}{\epsilon^{1/2}}+O(1/K),\ s\in B_{k}.\end{split}$
Here we have used that $A>0$ is fixed and $|s_{k}|<A$. Otherwise, the term $O(1/K)$ may grow with $A$. The same applies to many other bounds and big-$O$ terms in this section. Using that $H_{0}$ is compactly supported and $|\check{t}|<A$, we get from (7.16)
(E.12) $\begin{split}&G_{0}(\check{t}+a,\check{q})-G_{0}(\check{t},\check{q})=-\int_{-c}^{c}\ln\bigg{|}\frac{w-a}{w}\bigg{|}\partial_{\check{t}}H_{0}(w+\check{t},\check{q})\text{d}w,\ |\check{t}|<A.\end{split}$
Hence, by assumption 3.4(1) and (7.16), $|\partial_{\check{t}}H_{0}(w+\check{t},\check{q})|<c$ and
(E.13) $\begin{split}&|G_{0}(\check{t}+a,\check{q})-G_{0}(\check{t},\check{q})|<c\int_{-c}^{c}\bigg{|}\ln\bigg{|}\frac{w-a}{w}\bigg{|}\bigg{|}\text{d}w<ca\ln(1/a),\\ &0<a<1,\ |\check{t}|<A.\end{split}$
Furthermore, assumption 3.4(1) implies
[…]
Hence
[…]
Combining (E.11) and (E.15) with (E.10) implies
(E.16) $\begin{split}G\bigg{(}\check{t},\frac{P(\epsilon^{1/2}s)}{\epsilon};\theta,\epsilon\bigg{)}=&\mu G_{0}\bigg{(}\check{t},\frac{P(\epsilon^{1/2}s)}{\epsilon}\bigg{)}+O(\epsilon^{1/2})\\ =&\mu G_{0}\bigg{(}\check{t},\frac{P(\epsilon^{1/2}s_{k})}{\epsilon}+P^{\prime}(0)\frac{s-s_{k}}{\epsilon^{1/2}}\bigg{)}\\ &+O(1/K^{1/2})+O(\epsilon^{1/2}),\end{split}$
where $s=\epsilon^{1/2}\mu j_{1}\in B_{k}$ and $|\theta|<\epsilon^{1/2}A$. The equation (E.10) applies because by the assumptions of the lemma,
(E.17) $\begin{split}&\check{t}=\frac{Q_{\alpha\alpha}^{\prime\prime}}{2}\epsilon(\mu j_{1})^{2}=O(1),\\ &\theta=\alpha_{j_{1}}=\epsilon\mu j_{1}=O(\epsilon^{1/2}),\\ &\epsilon\check{q}=P(\epsilon\mu j_{1})=P(0)+O(\epsilon^{1/2}),\ \epsilon^{1/2}\mu|j_{1}|\leq A,\end{split}$
see (7.14). Further, using that $s^{2}=s_{k}^{2}+O(1/K)$ and (E.13), gives
(E.18) $\begin{split}G\bigg{(}\frac{Q_{\alpha\alpha}^{\prime\prime}}{2}s^{2},\frac{P(\epsilon^{1/2}s)}{\epsilon};\theta,\epsilon\bigg{)}=&\mu G_{0}\bigg{(}\frac{Q_{\alpha\alpha}^{\prime\prime}}{2}s_{k}^{2},\frac{P(\epsilon^{1/2}s_{k})}{\epsilon}+P^{\prime}(0)\frac{s-s_{k}}{\epsilon^{1/2}}\bigg{)}\\ &+O(1/K^{1/2})+O(\epsilon^{1/2}),\ s\in B_{k},|\theta|<\epsilon^{1/2}A.\end{split}$
There are $|B_{k}|/(\epsilon^{1/2}\mu)$ values of $j_{1}$ such that $\epsilon^{1/2}\mu j_{1}\in B_{k}$. Sum (E.18) with $s=\epsilon^{1/2}\mu j_{1}$ to obtain
(E.19) $\begin{split}\epsilon^{1/2}&\sum_{\epsilon^{1/2}\mu j_{1}\in B_{k}}G\bigg{(}\frac{Q_{\alpha\alpha}^{\prime\prime}}{2}\epsilon(\mu j_{1})^{2},\frac{P(\epsilon\mu j_{1})}{\epsilon};\theta,\epsilon\bigg{)}\\ =&\epsilon^{1/2}\sum_{\epsilon^{1/2}\mu j_{1}\in B_{k}}G_{0}\bigg{(}\frac{Q_{\alpha\alpha}^{\prime\prime}}{2}s_{k}^{2},u_{\epsilon}+P^{\prime}(0)\mu j_{1}\bigg{)}\\ &+|B_{k}|\big{[}O(1/K^{1/2})+O(\epsilon^{1/2})\big{]},\ u_{\epsilon}:=\frac{P(\epsilon^{1/2}s_{k})-P^{\prime}(0)\epsilon^{1/2}s_{k}}{\epsilon}.\end{split}$
It follows from (7.16) and (E.15) that $G_{0}$ is 1-periodic in its last argument and Hölder continuous.
From (4.8) and (D.30),
$P^{\prime}(\alpha)=[\Phi(x(\alpha),\alpha)]_{\alpha}^{\prime}=\Phi_{x}^{\prime}(x(\alpha),\alpha)x_{\alpha}^{\prime}(\alpha)+\Phi_{\alpha}^{\prime}(x(\alpha),\alpha)=\Phi_{\alpha}^{\prime}(x(\alpha),\alpha).$
By the assumption of the theorem, $P^{\prime}(0)\mu$ is irrational, so the points $P^{\prime}(0)\mu j_{1}$ are uniformly distributed mod 1 [35, Exercise 5.6]. Take the limit $\epsilon\to 0$ in (E.19) and use [35, Corollary 1.2] to find
(E.20) $\begin{split}\lim_{\epsilon\to 0}\epsilon^{1/2}&\sum_{\epsilon^{1/2}\mu j_{1}\in B_{k}}G\bigg{(}\frac{Q_{\alpha\alpha}^{\prime\prime}}{2}\epsilon(\mu j_{1})^{2},\frac{P(\epsilon\mu j_{1})}{\epsilon};\epsilon\mu j_{1},\epsilon\bigg{)}\\ =&\frac{|B_{k}|}{\mu}\bigg{[}\mu\int_{0}^{1}G_{0}\bigg{(}\frac{Q_{\alpha\alpha}^{\prime\prime}}{2}s_{k}^{2},\check{q}\bigg{)}\text{d}\check{q}+O(1/K^{1/2})\bigg{]}.\end{split}$
Sum (E.20) over all $B_{k}$ and use that $K\gg 1$ can be arbitrarily large:
(E.21) $\begin{split}\lim_{\epsilon\to 0}\epsilon^{1/2}&\sum_{\epsilon^{1/2}\mu|j_{1}|\leq A}G\bigg{(}\frac{Q_{\alpha\alpha}^{\prime\prime}}{2}\epsilon(\mu j_{1})^{2},\frac{P(\epsilon\mu j_{1})}{\epsilon};\epsilon\mu j_{1},\epsilon\bigg{)}\\ =&\int_{|s|\leq A}\int_{0}^{1}G_{0}\bigg{(}\frac{Q_{\alpha\alpha}^{\prime\prime}}{2}s^{2},\check{q}\bigg{)}\text{d}\check{q}\text{d}s.\end{split}$
Clearly the same arguments apply when we add $\Phi_{x}^{\prime}\check{x}$ to the first argument of $G$. Therefore, combining (E.21), (7.14) and (7.2), proves (7.15).
E.2.1. Proof of Lemma E.1
Fix any $A_{1}$. Throughout the proof we assume $|\check{t}|<A_{1}$. Fix any $\bar{h}\in C_{0}^{2}(\mathbb{R})$ and denote
(E.22) $\begin{split}\mathcal{Q}(u):=\int_{\mathbb{R}}\frac{\chi_{1\text{D}}(\epsilon\check{w})}{\check{w}}\bar{h}^{\prime}(\check{w}+u)\text{d}\check{w}.\end{split}$
We make use of the following simple result, which is stated without proof:
(E.23) $\begin{split}&|\mathcal{Q}(u)|\leq c(1+u^{2})^{-1},\ |u|\leq c_{1}/\epsilon,\quad\mathcal{Q}(u)\equiv 0,\ |u|\geq c_{1}/\epsilon,\end{split}$
for some $c_{1}>0$, which is the same in both places.
In view of (6.11), (6.14), and (E.22), we estimate
(E.24) $\begin{split}&\bigg{|}\sum_{j_{2}}\mathcal{Q}(\check{t}+\check{q}-j_{2};\theta,\epsilon)\big{[}a_{0}(\theta,\epsilon j_{2})-a_{0}(\theta,\epsilon\check{q})\big{]}(j_{2}-\check{q})_{+}^{1/2}\bigg{|}\\ &\leq c\epsilon\sum_{j_{2}=1}^{1/\epsilon}\frac{1}{j_{2}^{2}}j_{2}^{3/2}=O(\epsilon^{1/2}).\end{split}$
This implies
(E.25) $\begin{split}G(\check{t},\check{q};\theta,\epsilon)=&a_{0}(\theta,\epsilon\check{q})\sum_{j_{2}}\mathcal{Q}(\check{t}+\check{q}-j_{2};\theta,\epsilon)(j_{2}-\check{q})_{+}^{1/2}+O(\epsilon^{1/2}).\end{split}$
Using the assumptions of Lemma E.1, we find
(E.26) $\begin{split}G(\check{t},\check{q};\theta,\epsilon)=&\big{[}a_{0}(0,P(0))+O(\epsilon^{1/2})\big{]}\\ &\times\sum_{j_{2}}\mathcal{Q}(\check{t}+\check{q}-j_{2};\theta,\epsilon)(j_{2}-\check{q})_{+}^{1/2}+O(\epsilon^{1/2})\\ =&a_{0}(0,P(0))\sum_{j_{2}}\mathcal{Q}(\check{t}+\check{q}-j_{2};\theta,\epsilon)(j_{2}-\check{q})_{+}^{1/2}+O(\epsilon^{1/2}).\end{split}$
To show that $\chi$ can be dropped from the integral in (E.26) (which is inside the definition of $\mathcal{Q}$), we consider similarly to (E.22):
(E.27) $\begin{split}\mathcal{Q}_{c}(u):=\int_{\mathbb{R}}\frac{1-\chi_{1\text{D}}(\epsilon\check{w})}{\check{w}}\bar{h}^{\prime}(\check{w}+u)\text{d}\check{w}=\int_{|\check{w}|\geq c/\epsilon}\frac{1-\chi_{1\text{D}}(\epsilon\check{w})}{\check{w}}\bar{h}^{\prime}(\check{w}+u)\text{d}\check{w}.\end{split}$
As is easily seen,
(E.28) $\begin{split}&|\mathcal{Q}_{c}(u)|\leq c\big{(}u^{-2}+(\epsilon/|u|)\big{)},\ |u|>c_{1}/\epsilon;\quad\mathcal{Q}_{c}(u)=0,\ |u|\leq c_{1}/\epsilon,\end{split}$
for some $c_{1}>0$, which is the same in both places. Given that $|\check{t}|<A_{1}$, this implies
(E.29) $\begin{split}\sum_{j_{2}}\mathcal{Q}_{c}(\check{t}+\check{q}-j_{2};\theta,\epsilon)(j_{2}-\check{q})_{+}^{1/2}=O(\epsilon^{1/2}).\end{split}$
Therefore, by (E.26) and (E.29),
(E.30) $\begin{split}G(\check{t},\check{q};\theta,\epsilon)=&a_{0}(0,P(0))\int_{\mathbb{R}}\frac{1}{\check{w}}\sum_{j_{2}}h^{\prime}(\check{w}+\check{t}+\check{q}-j_{2};\theta,\epsilon)(j_{2}-\check{q})_{+}^{1/2}\text{d}\check{w}\\ &+O(\epsilon^{1/2}).\end{split}$
Further,
(E.31) $\begin{split}K(\theta+\epsilon\mu s)=&K(\theta)+O(\epsilon),\ \frac{\Phi(\theta+\epsilon\mu s)-\Phi(\theta)}{\epsilon}=\mu s\Phi^{\prime}(\theta)+O(\epsilon),\end{split}$
where $s\in\text{supp}\,\varphi_{\alpha}$, which implies (see (6.10))
(E.32) $\begin{split}&\partial_{u}^{k}\big{(}h(u;\theta,\epsilon)-\mu h_{0}(u)\big{)}=O(\epsilon),\ k=0,1,\\ &|\partial_{u}^{2}\big{(}h(u;\theta,\epsilon)-\mu h_{0}(u)\big{)}|<c,\ u\in\mathbb{R},\theta=O(\epsilon^{1/2}),\end{split}$
Similarly to (E.22) and (E.23), pick any $\bar{h}\in C_{0}^{2}(\mathbb{R})$ and define
(E.33) $\begin{split}\mathcal{Q}_{\epsilon}(u):=\int_{\mathbb{R}}\frac{1}{\check{w}}\Delta\bar{h}_{\epsilon}^{\prime}(\check{w}+u)\text{d}\check{w},\ \Delta\bar{h}_{\epsilon}(u):=\bar{h}(u+\epsilon)-\bar{h}(u).\end{split}$
Then
(E.34) $\begin{split}&|\mathcal{Q}_{\epsilon}(u)|\leq c\epsilon\ln(1/\epsilon)(1+u^{2})^{-1},\ u\in\mathbb{R}.\end{split}$
For large $|u|$, (E.34) is proven integrating by parts in (E.33). For $|u|$ in a bounded set, we argue similarly to (E.12), (E.13).
Using (E.32) and (E.33), (E.34) with $h_{\epsilon}(u)=h(u;\theta,\epsilon)-\mu h_{0}(u)$ we get from (E.30)
(E.35) $\begin{split}&|G(\check{t},\check{q};\theta,\epsilon)-\mu G_{0}(\check{t},\check{q})|\leq c|J_{\epsilon}|+O(\epsilon^{1/2}),\\ &J_{\epsilon}:=\sum_{j_{2}}\mathcal{Q}_{\epsilon}(\check{t}+\check{q}-j_{2})(j_{2}-\check{q})_{+}^{1/2}=O(\epsilon\ln(1/\epsilon)).\end{split}$
This proves the lemma.
E.3. Proof of Lemma 7.3
We evaluate the integrals in (7.17) similarly to [25, eqs. (3.16)–(3.20)]. Begin by considering
(E.36) $\begin{split}J(r):=&\int_{\mathbb{R}}\int_{\mathbb{R}}\int_{0}^{\infty}\frac{1}{\check{w}}\big{[}h_{0}^{\prime}\big{(}\check{w}+r+(Q_{\alpha\alpha}^{\prime\prime}/2)s^{2}-\check{q}\big{)}\\ &\hskip 85.35826pt-h_{0}^{\prime}\big{(}\check{w}+(Q_{\alpha\alpha}^{\prime\prime}/2)s^{2}-\check{q}\big{)}\big{]}\check{q}^{1/2}\text{d}\check{q}\text{d}s\text{d}\check{w}\\ =&(Q_{\alpha\alpha}^{\prime\prime}/2)^{-1/2}\int_{\mathbb{R}}\int_{\mathbb{R}}\int_{\mathbb{R}}\frac{1}{\check{w}}\big{[}h_{0}^{\prime}\big{(}\check{w}+r+u-\check{q}\big{)}\\ &\hskip 85.35826pt-h_{0}^{\prime}\big{(}\check{w}+u-\check{q}\big{)}\big{]}u_{+}^{-1/2}\check{q}_{+}^{1/2}\text{d}\check{q}\text{d}u\text{d}\check{w}.\\ \end{split}$
Substitute $h_{0}$ expressed in terms of its Fourier transform, $\tilde{h}_{0}=\mathcal{F}h$, and use the identity [14, p. 360, eq. 20]:
(E.37) $\mathcal{F}(x_{+}^{a-1})=e^{ai\pi/2}\Gamma(a)(\lambda+i0)^{-a},$
to obtain
(E.38) $\begin{split}J(r)=\frac{1}{2\pi}\int_{\mathbb{R}}&\big{(}\mathcal{F}(1/\check{w})\big{)}(-\lambda)\big{(}\mathcal{F}u_{+}^{-1/2}\big{)}(-\lambda)\big{(}\mathcal{F}\check{q}_{+}^{1/2}\big{)}(\lambda)\\ &\times\big{(}e^{-i\lambda r}-1\big{)}(-i\lambda)\tilde{h}_{0}(\lambda)\text{d}\lambda\\ =\frac{1}{2\pi}\int_{\mathbb{R}}&\big{(}-i\pi\text{sgn}(\lambda)\big{)}\big{(}e^{3i\pi/4}\Gamma(3/2)(\lambda+i0)^{-3/2}\big{)}\\ &\times\big{(}e^{-i\pi/4}\Gamma(1/2)(\lambda-i0)^{-1/2}\big{)}\big{(}e^{-i\lambda r}-1\big{)}(-i\lambda)\tilde{h}_{0}(\lambda)\text{d}\lambda.\end{split}$
This simplifies to
(E.39) $\begin{split}J(r)=&-\frac{i\Gamma(1/2)\Gamma(3/2)}{2}\int_{\mathbb{R}}\frac{e^{-i\lambda r}-1}{\lambda}\tilde{h}_{0}(\lambda)\text{d}\lambda\\ =&\frac{\pi}{8}\int_{\mathbb{R}}\big{[}\text{sgn}(-r-t)-\text{sgn}(-t)\big{]}h_{0}(t)\text{d}t=-\frac{\pi}{4}\int_{0}^{r}h_{0}(-t)\text{d}t.\end{split}$
Substitution of (E.39) back into (7.17) gives (7.18).
Appendix F Proof of Lemmas 8.1 and 8.2
F.1. Proof of Lemma 8.1
Integrating by parts in (8.6) with respect to $\lambda$ twice and using a partition of unity we can assume that $I=(-\delta,\delta)$ for some small $0<\delta\ll 1$, and $D(\lambda,\theta,\epsilon)\in C_{0}^{\infty}(I)$ in $\theta$. On $I$ we can change variable $\theta\to t$ to get
(F.1) $J(r)=\int_{0}^{\infty}\int_{I_{1}}D_{1}(\lambda,t,\epsilon)e^{ir\lambda t}\text{d}t\lambda\text{d}\lambda=\int_{0}^{\infty}\tilde{D}_{1}(\lambda,r\lambda,\epsilon)\lambda\text{d}\lambda.$
Here
(F.2) $I_{1}=(-\sin\delta,\sin\delta),\ D_{1}(\lambda,t,\epsilon)=D(\lambda,\sin^{-1}t,\epsilon)/(1-t^{2})^{1/2},$
and $\tilde{D}_{1}(\lambda,\mu,\epsilon)$ is the Fourier transform of $D(\lambda,\theta,\epsilon)$ with respect to $\theta$. Thus $\tilde{D}_{1}(\lambda,\mu,\epsilon)$ decays rapidly as $\mu\to\infty$, i.e. for any $N\in\mathbb{N}_{0}$ there exists $c_{N}$ such that
(F.3) $|\tilde{D}_{1}(\lambda,\mu,\epsilon)|\leq c_{N}(1+|\mu|)^{-N},\ (\lambda,\mu,\epsilon)\in(0,\infty)\times\mathbb{R}\times(0,\epsilon_{0}).$
Selecting $N=3$ gives
(F.4) $|J(r)|\leq c\int_{0}^{\infty}\frac{\lambda}{1+|r\lambda|^{3}}\text{d}\lambda.$
Changing variable $\lambda\to r\lambda$ proves the top cases in (8.7).
All the other assertions of the lemma can be proven in a similar fashion. For example, if $c_{k,l}=O(\epsilon)$ in (8.5), then $c_{N}$ in (F.3) and $c$ in (F.4) are $O(\epsilon)$.
If $D(\lambda=0,\theta,\epsilon)\equiv 0$, then we integrate by parts in (8.6) three times instead of two. The analog of (F.1) becomes
(F.5) $J(r)=\int_{0}^{\infty}(\tilde{D}_{1}(\lambda,r\lambda,\epsilon)/\lambda)\lambda^{2}\text{d}\lambda.$
Selecting $N=4$ in (F.3) then gives
(F.6) $|J(r)|\leq c\int_{0}^{\infty}\frac{\lambda^{2}}{1+|r\lambda|^{4}}\text{d}\lambda=O(r^{-3}),\ r\to\infty.$
F.2. Proof of Lemma 8.2
The statement $\Upsilon(0)=0$ is immediate. To prove the second claim, compute
(F.7) $\begin{split}\Upsilon(+\infty)-\Upsilon(-\infty)&=\int_{\mathbb{R}}\bar{h}_{0}(u)\text{d}u=1.\end{split}$
The last equality follows from (3.15) and (7.16) (recall that $\bar{h}_{0}$ is defined without the prefactor in (7.16)). Similarly,
(F.8) $\begin{split}&\Upsilon(+\infty)+\Upsilon(-\infty)\\ &=\frac{1}{|\Phi_{x}^{\prime}|}\int_{\mathbb{R}}\bar{h}_{0}(u)\int_{\mathbb{R}}\text{sgn}(t-u)\mathcal{F}^{-1}\bigg{(}\frac{1}{1+\kappa C_{1}|\lambda|^{3}}\bigg{)}(t/|\Phi_{x}^{\prime}|)\text{d}t\text{d}u=0.\end{split}$
The last equality follows because the signum function is odd, while the functions $\bar{h}_{0}$ and $\mathcal{F}^{-1}\big{(}(1+\kappa C_{1}|\lambda|^{3})^{-1}\big{)}$ are even. The fact that $\bar{h}_{0}$ is even follows from assumption 3.4(2) and (7.16).

### v2

@@
 1.1. Local reconstruction analysis (LRA)
-Let $\mathcal{R}$ be the generalized Radon transform (GRT); $f(x)$, $x\in\mathcal{U}\subset\mathbb{R}^{n}$, be the function to be reconstructed; and $g(y)$, $y\in\mathcal{V}\subset\mathbb{R}^{n}$, where $g=\mathcal{R}f$, be the GRT of $f$.
+Let $\mathcal{R}$ be the generalized Radon transform (GRT), which integrates a function $f$ on $\mathcal{U}\subset\mathbb{R}^{n}$ over a family of surfaces parameterized by a set $\mathcal{V}\subset\mathbb{R}^{n}$.
+Thus, $f$ is the function to be reconstructed.
+Let $\hat{f}(y)=(\mathcal{R}f)(y)$, $y\in\mathcal{V}$, denote the GRT of $f$.
 The open sets $\mathcal{U}$ and $\mathcal{V}$ represent the image and data domains, respectively.
-The values of $g(y)$ are given on a rectangular grid $y_{j}$, $j\in\mathbb{Z}^{n}$.
-The step size of the grid along each coordinate is proportional to some small parameter $\epsilon>0$.
-Consider a reconstruction formula of the form $f_{0}=\mathcal{R}^{*}\mathcal{B}g$, where $\mathcal{R}$ is the adjoint transform and $\mathcal{B}$ is a Pseudo-Differential Operator ($\Psi$DO).
-Depending on the choice of $\mathcal{B}$, the reconstruction can be theoretically exact ($f_{0}=f$), quasi-exact ($f_{0}-f$ is less singular than $f$), edge-enhancing ($f_{0}$ is more singular than $f$), and smoothing of a finite degree ($f_{0}$ is less singular than $f$).
-Next, let $\varphi$ be a compactly supported, sufficiently smooth interpolation kernel, and let $g_{\epsilon}(y)$, $y\in\mathcal{V}$, denote the function obtained by interpolating the values $g(y_{j})$.
-Denote $f_{\epsilon}=\mathcal{R}^{*}\mathcal{B}g_{e}$.
+The values of $\hat{f}(y)$ are given on a rectangular grid $y_{j}$, $j\in\mathbb{Z}^{n}$.
+The step size of the grid along each coordinate is proportional to a small parameter $\epsilon>0$.
+Consider a reconstruction formula of the form $f_{0}=\mathcal{R}^{*}\mathcal{B}\hat{f}$, where $\mathcal{R}$ is the adjoint transform and $\mathcal{B}$ is a Pseudo-Differential Operator ($\Psi$DO).
+Depending on the choice of $\mathcal{B}$, the reconstruction can be theoretically exact ($f_{0}=f$), quasi-exact ($f_{0}-f$ is less singular than $f$), edge-enhancing ($f_{0}$ is more singular than $f$), or smoothing of a finite degree ($f_{0}$ is less singular than $f$).
+The common property of each of these kinds of reconstructions is that, microlocally, $\text{WF}(f)=\text{WF}(f_{0})$.
+Next, let $\varphi$ be a compactly supported, sufficiently smooth interpolation kernel, and let $\hat{f}_{\epsilon}(y)$, $y\in\mathcal{V}$, denote the function obtained by interpolating the values $\hat{f}(y_{j})$, see eq.
+(3.14).
+Denote $f_{\epsilon}=\mathcal{R}^{*}\mathcal{B}\hat{f}_{\epsilon}$.
 Thus, $f_{\epsilon}$ is reconstructed from discrete data.
 LRA is a powerful and flexible technique proposed by the author to study images $f_{\epsilon}$ reconstructed from discrete GRT data.
+The reconstruction can be of any kind, e.g., exact, quasi-exact, edge-enhancing, or smoothing as described above.
 The main idea of LRA is to obtain a simple formula to accurately approximate $f_{\epsilon}$ in an $\epsilon$-neighborhood of a point, $x_{0}\in\mathcal{U}$.
-To illustrate the main idea of LRA consider an example.
-Let $f$ be a piece-wise smooth real-valued function with singularities along some hypersurface $\mathcal{S}$.
-More precisely, $f$ is conormal with respect to $\mathcal{S}$ [21, Definition 18.2.6].
-Fix any $A>0$.
+We assume that $f$ is conormal with respect to some hypersurface $\mathcal{S}$ [24, Definition 18.2.6].
+Fix any $x_{0}\in\mathcal{S}$ and $A>0$.
 We show that, in a variety of settings, the following limit exists
@@
 the limit is uniform in $\check{x}$, and compute the limit explicitly. Here $\nu\geq 0$ is some constant, which depends on the strength of the singularity of $f_{0}$ at $x_{0}$.
-To illustrate one such setting, suppose $f$ has a jump across $\mathcal{S}$, where $\mathcal{S}\subset\mathbb{R}^{2}$ is a smooth curve.
+To help understand (1.1), we consider an example of a GRT in the plane.
+Suppose $f$ has a jump across $\mathcal{S}$, where $\mathcal{S}\subset\mathbb{R}^{2}$ is a smooth curve.
 Suppose also that the reconstruction is either exact or quasi-exact.
 In this case $\mathcal{R}^{*}\mathcal{B}\mathcal{R}$ is a $\Psi$DO with the principal symbol 1 and $\nu=0$.
-Pick $x_{0}\in\mathcal{S}$.
+Let $\vec{\Theta}_{0}$ be a unit vector perpendicular to $\mathcal{S}$ at $x_{0}$.
 Under some mild conditions on $\mathcal{S}$ and $x_{0}$, we show
-(1.2) $\lim_{\epsilon\to 0}f_{\epsilon}(x_{0}+\epsilon\check{x})=f_{0}(x_{0}^{-})+\Delta f(x_{0})\text{DTB}(\check{x};x_{0}),\ |\check{x}|<A,$
+(1.2) $\lim_{\epsilon\to 0}f_{\epsilon}(x_{0}+\epsilon\check{x})=f_{0}(x_{0}^{-})+\Delta f(x_{0})\text{DTB}(\vec{\Theta}_{0}\cdot\check{x};x_{0}),\ |\check{x}|<A,$
 uniformly in $\check{x}$. Here
   1. (1)
-    $f_{0}(x_{0}^{-})$ is the value of the reconstruction from continuous data, $f_{0}$, on one side of $\mathcal{S}$,
+    $f_{0}(x_{0}^{\pm}):=\lim_{t\to 0^{\pm}}f_{0}(x_{0}+t\vec{\Theta}_{0})$;
   2. (2)
-    $\Delta f(x_{0}):=f_{0}(x_{0}^{+})-f_{0}(x_{0}^{-})$ is the value of the jump of $f$ across $\mathcal{S}$ at $x_{0}$, and
+    $\Delta f(x_{0}):=f_{0}(x_{0}^{+})-f_{0}(x_{0}^{-})$ is the value of the jump of $f$ across $\mathcal{S}$ at $x_{0}$; and
   3. (3)
     DTB, which stands for the Discrete Transition Behavior, is an easily computable function independent of $f$.
-The DTB function depends only on the curvature of $\mathcal{S}$ at $x_{0}$ and on the interpolation kernel, $\varphi$.
+The DTB function depends only on the interpolation kernel, $\varphi$.
 When $\epsilon$ is sufficiently small, the right-hand side of (1.2) is an accurate approximation of $f_{\epsilon}$, and the DTB function accurately describes the smoothing of the jump of $f$ in $f_{\epsilon}$ due to the discrete nature of the tomographic data.
@@
 In contrast, LRA does not require $f$ to be bandlimited in any way, either exactly or approximately.
-Even though LRA applies in a wide variety of settings, one of its major limitations is that so far it has only been applied when $f_{\epsilon}$ is computed by a linear reconstruction algorithm, i.e., by an application of a reconstruction formula to data: $f_{\epsilon}=\mathcal{R}^{*}\mathcal{B}g_{\epsilon}$.
+Even though LRA applies in a wide variety of settings, one of its major limitations is that so far it has only been applied when $f_{\epsilon}$ is computed by a linear reconstruction algorithm, i.e., by an application of a reconstruction formula to data: $f_{\epsilon}=\mathcal{R}^{*}\mathcal{B}\hat{f}_{\epsilon}$.
 Very few GRTs admit an exact inversion formula.
@@
 The reconstructed function, still denoted $f_{\epsilon}$, is computed as the solution to the following quadratic minimization problem:
-(1.3) $f_{\epsilon}=\text{argmin}_{f\in H_{c}^{1}(\mathcal{U}_{b})}\Psi(f),\ \Psi(f):=\|\mathcal{R}f-g_{\epsilon}\|_{L^{2}(\mathcal{V})}^{2}+\kappa\epsilon^{3}\|\nabla f\|_{L^{2}(\mathcal{U})}^{2},$
-where $\kappa>0$ is the regularization parameter (which is predetermined and independent of $\epsilon$).
+(1.3) $f_{\epsilon}=\text{argmin}_{f\in H_{0}^{1}(\mathcal{U}_{b})}\Psi(f),\ \Psi(f):=\|\mathcal{R}f-\hat{f}_{\epsilon}\|_{L^{2}(\mathcal{V})}^{2}+\kappa\epsilon^{3}\|\partial_{x}f\|_{L^{2}(\mathcal{U})}^{2},$
+where $\kappa>0$ is the regularization parameter (which is predetermined and independent of $\epsilon$), $U_{b}\Subset U$ is a domain, and $H_{0}^{1}(\mathcal{U}_{b})$ is the closure of $C_{0}^{\infty}(U_{b})$ in the $H^{1}$ norm.
 In the spirit of LRA, the factor $\epsilon^{3}$ in front of the Tikhonov regularization term ensures that the resolution of the reconstruction is at the native scale $\epsilon$.
 To solve (1.3) numerically, one usually assumes that $f_{\epsilon}(x)=\sum_{l}c_{l}\psi_{l}(x)$, where $\psi_{l}$ are some basis functions, and their number is of the same order of magnitude as the number of data points.
-In this case, the reconstruction consists of computing the coefficients $c_{l}$.
+The reconstruction consists of computing the coefficients $c_{l}$.
 For conventional analysis of reconstruction at a scale $\gg\epsilon$, the effects of this finite-dimensional approximation of $f_{\epsilon}$ are largely invisible.
@@
 While the effects of finite-dimensional approximation of a solution are less fundamental (because they can be effectively mitigated), they are important from the point of view of computational complexity.
-Nevertheless, this line of research is beyond the scope of the paper and will be the subject of future work.
+Nevertheless, this line of research is beyond the scope of the paper.
 The main result of the paper, stated in Theorems 3.6 and 3.7, consists of formulas of the type (1.2) for the solution $f_{\epsilon}$ to the minimization problem (1.3).
@@
 In Section 5 we obtain an alternative formula for computing $f_{\epsilon}$, equation (5.6).
-The equation is of the form $\mathcal{T}f_{\epsilon}=F_{\epsilon}$, where $\mathcal{T}$ is some $\Psi$DO, and $F_{\epsilon}$ is computed from the data $g(y_{j})$.
-Also we find another function, $G_{\epsilon}$, such that (a) $G_{\epsilon}$ is easier to analyze than $f_{\epsilon}$ and (b) the local behavior of $f_{\epsilon}$ and $G_{\epsilon}$ near $x_{0}$ are the same from the perspective of LRA, see Lemma 5.2.
+The equation is of the form $\mathcal{T}f_{\epsilon}=F_{\epsilon}$, where $\mathcal{T}$ is some $\Psi$DO, and $F_{\epsilon}$ is computed from the data $\hat{f}(y_{j})$.
+Also, we find another function, $\mathcal{G}_{\epsilon}$, such that (a) $\mathcal{G}_{\epsilon}$ is easier to analyze than $f_{\epsilon}$ and (b) the local behavior of $f_{\epsilon}$ and $\mathcal{G}_{\epsilon}$ near $x_{0}$ are the same from the perspective of LRA, see Lemma 5.2.
 In Section 6 we prove that $\|F_{\epsilon}\|_{L^{\infty}(\mathcal{U})}$ is uniformly bounded when $0<\epsilon\ll 1$.
@@
 (1.4) $\Delta F_{0}(\check{x}):=\lim_{\epsilon\to 0}\big{(}F_{\epsilon}^{(l)}(x_{0}+\epsilon\check{x})-F_{\epsilon}^{(l)}(x_{0})\big{)},$
-where $F_{\epsilon}^{(l)}(x)$ is the leading singular term of $F_{\epsilon}(x)$ near $x_{0}$, see (7.18). In section 8 we compute the DTB function, $\Delta G_{0}$, as the limit
-(1.5) $\Delta G_{0}(\check{x}):=\lim_{\epsilon\to 0}\big{(}G_{\epsilon}(x_{0}+\epsilon\check{x})-G_{\epsilon}(x_{0})\big{)},$
-see equation (LABEL:rhs_lim_5). This calculation assumes that the singularity of $f$ at $x_{0}$ is visible from the data only once. This completes the proof of Theorem 3.6.
+where $F_{\epsilon}^{(l)}(x)$ is the leading singular term of $F_{\epsilon}(x)$ near $x_{0}$, see (7.20). In Section 8 we compute the DTB function, $\Delta\mathcal{G}_{0}$, as the limit
+(1.5) $\Delta\mathcal{G}_{0}(\check{x}):=\lim_{\epsilon\to 0}\big{(}\mathcal{G}_{\epsilon}(x_{0}+\epsilon\check{x})-\mathcal{G}_{\epsilon}(x_{0})\big{)},$
+see equation (LABEL:rhs_lim_5).
+The limits in (1.4) and (1.5) are uniform in $\check{x}$ confined to any bounded set.
+The calculation assumes that the singularity of $f$ at $x_{0}$ is visible from the data only once.
+This means that there is only one $y\in\mathcal{V}$ such that $\mathcal{S}_{y}$ is tangent to $\mathcal{S}$ at $x_{0}$.
+This completes the proof of Theorem 3.6.
 A more general formula for the DTB function when the data are redundant (i.e., the singularity is visible several times) is obtained in Section 9, see equation (LABEL:upsilon_simple_mult).
@@
 Definition 2.1.
-Let $f$ and $g$ be two functions defined on a domain $U$. We say $f(x)\asymp g(x)$ for $x\in U$ if there exist $c_{1,2}>0$ such that
-(2.1) $c_{1}\leq f(x)/g(x)\leq c_{2}\text{ if }g(x)\not=0\text{ and }f(x)=0\text{ if }g(x)=0$
-for any $x\in U$.
+Let $f$ and $g$ be two functions defined on a domain $U$. We say $f(x)\asymp g(x)$ for $x\in U$ if there exist $c_{1,2}>0$ such that $f(x)\leq c_{1}g(x)$ and $g(x)\leq c_{2}f(x)$ for any $x\in U$.
+We identify tangent and cotangent spaces on Euclidean spaces with the underlying Euclidean spaces. Thus, for a function $f(x)$ defined on $\mathbb{R}^{n}$ we have
+(2.1) $\begin{split}&\partial_{x}f(x):=f_{x}^{\prime}(x):=(\partial_{x_{1}}f(x),\dots,\partial_{x_{n}}f(x))^{T},\\ &\text{d}_{x}f=(\partial_{x_{1}}f(x),\dots,\partial_{x_{n}}f(x)),\\ &|\partial_{x}f(x)|=|\text{d}_{x}f(x)|=\big{[}(\partial_{x_{1}}f(x))^{2}+\dots+(\partial_{x_{n}}f(x))^{2}\big{]}^{1/2}.\end{split}$
+We use $f_{x}^{\prime}$ or $\partial_{x}f$ interchangeably depending on what is more convenient from the notational perspective.
 We use several types of function spaces. First, $C^{k}(U)$, $k\in\mathbb{N}$, is the space of functions with bounded derivatives up to order $k$ with the norm
-(2.2) $\|h\|_{C^{k}(U)}:=\max_{|m|\leq k}\|h^{(m)}\|_{L^{\infty}(U)}.$
+(2.2) $\|h\|_{C^{k}(U)}:=\max_{|m|\leq k}\|\partial_{x}^{m}h\|_{L^{\infty}(U)},\ m\in\mathbb{N}_{0}^{n}.$
 The subscript ‘0’ in $C_{0}^{k}$ means that we consider the subspace of compactly supported functions, $C_{0}^{k}(U)\subset C^{k}(U)$.
@@
 Another type is the Hölder spaces $C^{s}(\mathbb{R}^{n})$, $s>0$, $s\not\in\mathbb{N}$.
-If $s=k+\gamma$, $k\in\mathbb{N}_{0}$, $10<\gamma<$ 1, then $C^{s}(\mathbb{R}^{n})$ is the space of $C^{k}(\mathbb{R}^{n})$ functions (or, $L^{\infty}(\mathbb{R}^{n})$ functions if $k=0$), which have Hölder continuous $k$-th order derivatives (see [4, Definition 5.28]), with the norm
-(2.3) $\begin{split}\|f\|_{C^{s}}:=&\|f\|_{C^{k}}+\max_{|m|=k}\sup_{x\in\mathbb{R}^{n},|h|>0}\frac{|f^{(m)}(x+h)-f^{(m)}(x)|}{|h|^{\gamma}},\ k>0,\\ \|f\|_{C^{s}}:=&\|f\|_{L^{\infty}}+\sup_{x\in\mathbb{R}^{n},|h|>0}\frac{|f(x+h)-f(x)|}{|h|^{\gamma}},\ k=0.\end{split}$
+If $s=k+\gamma$, $k\in\mathbb{N}_{0}$, $10<\gamma<$ 1, then $C^{s}(\mathbb{R}^{n})$ is the space of $C^{k}(\mathbb{R}^{n})$ functions (or, $L^{\infty}(\mathbb{R}^{n})$ functions if $k=0$), which have Hölder continuous $k$-th order derivatives (see [5, Definition 5.28]), with the norm
+(2.3) $\begin{split}\|f\|_{C^{s}}:=&\|f\|_{C^{k}}+\max_{|m|=k}\sup_{x\in\mathbb{R}^{n},|h|>0}\frac{|\partial_{x}^{m}(f(x+h)-f(x))|}{|h|^{\gamma}},\ k>0,\\ \|f\|_{C^{s}}:=&\|f\|_{L^{\infty}}+\sup_{x\in\mathbb{R}^{n},|h|>0}\frac{|f(x+h)-f(x)|}{|h|^{\gamma}},\ k=0.\end{split}$
 As before, $C_{0}^{s}(U)$ denotes the subspace of $C^{s}(\mathbb{R}^{n})$ functions supported in $U$.
@@
 To define the latter, pick any $\mu_{0}\in C_{0}^{\infty}(\mathbb{R}^{n})$ such that $\mu_{0}(\eta)=1$ for $|\eta|\leq 1$, $\mu_{0}(\eta)=0$ for $|\eta|\geq 2$.
-Define $\mu_{j}(\eta):=\mu_{0}(2^{-j}\eta)-\mu_{0}(2^{-j+1}\eta)$, $j\in\mathbb{N}$ [1, Section 5.4].
+Define $\mu_{j}(\eta):=\mu_{0}(2^{-j}\eta)-\mu_{0}(2^{-j+1}\eta)$, $j\in\mathbb{N}$ [5, Definition 5.30] and [1, Section 6.2].
 Then
@@
 (2.5) $\tilde{f}(\xi)=\frac{1}{(2\pi)^{n}}\int_{\mathbb{R}^{n}}f(x)e^{i\xi\cdot x}\text{d}x,\ f\in L^{1}(\mathbb{R}^{n}),$
-which extends to tempered distributions [20, section 7.1].
-We have $C_{*}^{s}(\mathbb{R}^{n})=C^{s}(\mathbb{R}^{n})$ and the norms $\|\cdot\|_{C_{*}^{s}(\mathbb{R}^{n})}$, $\|\cdot\|_{C^{s}(\mathbb{R}^{n})}$ are equivalent, for any $s>0$, $s\not\in\mathbb{N}$ [4, Remark 5.31].
+which extends to tempered distributions [25, Section 7.1].
+We have $C_{*}^{s}(\mathbb{R}^{n})=C^{s}(\mathbb{R}^{n})$ and the norms $\|\cdot\|_{C_{*}^{s}(\mathbb{R}^{n})}$, $\|\cdot\|_{C^{s}(\mathbb{R}^{n})}$ are equivalent, for any $s>0$, $s\not\in\mathbb{N}$ [5, Remark 5.31].
 The Hölder-Zygmund spaces are a particular case of the Besov spaces: $C^{s}(\mathbb{R}^{n})=B^{s}_{p,q}(\mathbb{R}^{n})$, where $p,q=\infty$ [1, item 2 in Remark 6.4].
@@
 (2.6) $\|f\|_{H^{s}(\mathbb{R}^{n})}^{2}=(2\pi)^{-n}\int_{\mathbb{R}^{n}}|\tilde{f}(\xi)|^{2}(1+|\xi|^{2})^{s}\text{d}\xi<\infty.$
-$H_{c}^{s}(U)$ is the collection of all $f\in H^{s}(\mathbb{R}^{n})$ such that $f\equiv 0$ on $\mathbb{R}^{n}\setminus K$ for some compact $K\subset U$.
-$H_{loc}^{s}(U)$ is the space of all distributions on $U$ such that $\chi f\in H^{s}(U)$ for any $\chi\in C_{0}^{\infty}(U)$.
-See [52, section 5, Notation and Background]
+$H_{0}^{s}(U)$ is the closure of $C_{0}^{\infty}(U)$ in $H^{s}(\mathbb{R}^{n})$.
+$H_{loc}^{s}(U)$ is the space of all distributions on $U$ such that $\chi f\in H^{s}(\mathbb{R}^{n})$ for any $\chi\in C_{0}^{\infty}(U)$.
+See [55, Section 5, Notation and Background]
 Definition 2.2.
 Given a domain $U\subset\mathbb{R}^{n}$, $r\in\mathbb{R}$, and $N\in\mathbb{N}$, $S^{r}(U\times\mathbb{R}^{N})$ denotes the vector-space of $C^{\infty}(U\times(\mathbb{R}^{N}\setminus 0_{N}))$ functions, $\tilde{B}(x,\xi)$, having the following properties
-(2.7) $\begin{split}&|\partial_{x}^{m}\tilde{B}(x,\xi)|\leq c_{m}\ \forall m\in\mathbb{N}_{0}^{n},x\in U,0<|\xi|\leq 1;\\ &|\partial_{x}^{m_{1}}\partial_{\xi}^{m_{2}}\tilde{B}(x,\xi)|\leq c_{m_{1},m_{2}}|\xi|^{r-|m_{2}|},\ \forall m_{1}\in\mathbb{N}_{0}^{n},m_{2}\in\mathbb{N}_{0}^{N},x\in U,|\xi|\geq 1;\end{split}$
-for some constants $c_{m},c_{m_{1},m_{2}}>0$.
+(2.7) $\begin{split}&|\partial_{x}^{m}\tilde{B}(x,\xi)|\leq c_{m}|\xi|^{-n+\delta}\ \forall m\in\mathbb{N}_{0}^{n},x\in U,0<|\xi|\leq 1;\\ &|\partial_{x}^{m_{1}}\partial_{\xi}^{m_{2}}\tilde{B}(x,\xi)|\leq c_{m_{1},m_{2}}|\xi|^{r-|m_{2}|},\ \forall m_{1}\in\mathbb{N}_{0}^{n},m_{2}\in\mathbb{N}_{0}^{N},x\in U,|\xi|\geq 1;\end{split}$
+for some constants $\delta>0$ and $c_{m},c_{m_{1},m_{2}}>0$.
 The elements of $S^{r}$ are called symbols of order $r$.
@@
 (2.8) $(\text{Op}(\tilde{B}(x,\xi))f)(y):=\frac{1}{(2\pi)^{n}}\int_{\mathbb{R}^{n}}\tilde{B}(x,\xi)\tilde{f}(\xi)e^{-i\xi\cdot y}\text{d}\xi,\ f\in C_{0}^{\infty}(U),$
-is denoted $L^{r}(U)$. As is known, $L^{r}(U):H_{c}^{s}(U)\to H_{loc}^{s-r}(U)$ is continuous [52, Theorem 2.1, Section I.2].
+is denoted $L^{r}(U)$.
+By [55, Theorem 2.1, Section I.2], any $\Psi$DO in $L^{r}(U)$ is $H_{0}^{s}(U_{b})\to H_{loc}^{s-r}(U)$ continuous for any domain $U_{b}\Subset U$.
+The notation $\mathcal{U}_{b}\Subset\mathcal{U}$ means that the closure $\overline{\mathcal{U}_{b}}$ is compact and $\overline{\mathcal{U}_{b}}\subset\mathcal{U}$.
 Let $U,V\subset\mathbb{R}^{n}$ be two domains, and let
 (2.9) $\Phi(x,y,\Theta)\in C^{\infty}(U\times V\times(\mathbb{R}^{N}\setminus 0_{N})),\quad\Phi:U\times V\times\mathbb{R}^{N}\to\mathbb{R},$
-be a nondegenerate phase function [16, Definition 11.1]. The latter property means that
+be a nondegenerate phase function [19, Definition 11.1]. The latter property means that
   1. (1)
@@
   3. (3)
-    The differentials $\text{d}_{(x,y,\Theta)}(\partial\Phi/\partial\Theta_{l})$, $1\leq l\leq N$, are linearly independent on the set $\Sigma$:
-    (2.10) $\Sigma:=\{(x,y,\Theta)\in U\times V\times(\mathbb{R}^{N}\setminus 0_{N}):\ \Phi_{\Theta}^{\prime}(x,y,\Theta)=0_{N}\}.$
-The phase $\Phi$ determines a map $T^{*}U\to T^{*}V$ with the graph
-(2.11) $\mathcal{C}:=\{(y,\text{d}_{y}\Phi(x,y,\Theta));(x,-\text{d}_{x}\Phi(x,y,\Theta)):(x,y,\Theta)\in\Sigma\}.$
-$I^{r}(V\times U\times\mathcal{C})$ denotes the vector space of FIOs of order $r$ given by [16, section 11]
+    The differentials $\text{d}_{(x,y,\Theta)}\Phi_{\Theta_{l}}^{\prime}$, $1\leq l\leq N$, are linearly independent on the set $\Sigma$:
+    (2.10) $\Sigma:=\{(y,x,\Theta)\in V\times U\times(\mathbb{R}^{N}\setminus 0_{N}):\ \Phi_{\Theta}^{\prime}(x,y,\Theta)=0_{N}\}.$
+The phase $\Phi$ determines the homogeneous canonical relation [56, Chapter VIII, Definition 5.1]
+(2.11) $\begin{split}\mathcal{C}:=&\{(y,\text{d}_{y}\Phi(x,y,\Theta));(x,-\text{d}_{x}\Phi(x,y,\Theta)):(x,y,\Theta)\in\Sigma\}\\ \subset&(T^{*}V\setminus 0_{2n})\times(T^{*}U\setminus 0_{2n}).\end{split}$
+$I^{r}(V\times U\times\mathcal{C})$ denotes the vector space of FIOs of order $r$ given by [19, Section 11]
 (2.12) $\begin{split}&C_{0}^{\infty}(U)\ni f\to\frac{1}{(2\pi)^{N}}\int_{\mathbb{R}^{N}}\int_{U}\tilde{B}(x,y,\Theta)f(x)e^{i\Phi(x,y,\Theta)}\text{d}x\text{d}\Theta\in\mathcal{D}^{\prime}(V),\\ &\tilde{B}\in S^{r+(n-N)/2}(U\times V\times\mathbb{R}^{N}).\end{split}$
@@
 Let $\Phi_{1}(x,y):\mathcal{U}\times\mathcal{V}\to\mathbb{R}^{2}$ be a defining function for a GRT $\mathcal{R}$.
+This is a function such that $\mathcal{S}_{y}:=\{x\in\mathcal{U}:\Phi_{1}(x,y)=0\}$ for any $y\in\mathcal{V}$, and $\mathcal{R}$ integrates over $\mathcal{S}_{y}$.
 An open set $\mathcal{U}\subset\mathbb{R}^{2}$ is the image domain.
-We assume $y=(\alpha,p)$, and $\mathcal{V}=I_{\alpha}\times I_{p}$ is the data domain.
+We assume $y=(\alpha,p)$, and $\mathcal{V}=\mathcal{I}_{\alpha}\times\mathcal{I}_{p}$ is the data domain.
 Assumption 3.1 (Properties of the GRT - I).
   1. (1)
-    $I_{\alpha}\subset\mathbb{R}$ is a compact interval with the endpoints identified (so it is topologically equivalent to a circle),
+    $\mathcal{I}_{\alpha}\subset\mathbb{R}$ is a closed interval with the endpoints identified, so it can be viewed as a circle,
   2. (2)
-    $I_{p}\subset\mathbb{R}$ is an open interval (or all of $\mathbb{R}$), and
+    $\mathcal{I}_{p}\subset\mathbb{R}$ is an open interval (or all of $\mathbb{R}$), and
   3. (3)
-    $\Phi_{1}(x,y)=p-\Phi(x,\alpha)$, $\Phi\in C^{\infty}(\mathcal{U}\times I_{\alpha})$.
+    $\Phi_{1}(x,y)=p-\Phi(x,\alpha)$, $\Phi\in C^{\infty}(\mathcal{U}\times\mathcal{I}_{\alpha})$.
 Both $\mathcal{U}$ and $\mathcal{V}$ are endowed with the usual Euclidean metric.
-By a partition of unity we will always identify subsets of $\mathcal{V}$ with subsets of $\mathbb{R}^{2}$.
+By a partition of unity we will always identify small subsets of $\mathcal{V}$ with subsets of $\mathbb{R}^{2}$.
 The transform $\mathcal{R}$ integrates over smooth curves $\mathcal{S}_{(\alpha,p)}:=\{x\in\mathcal{U}:p=\Phi(x,\alpha)\}$:
@@
 Here and everywhere below we use the convention that whenever $y$, $\alpha$, and $p$ appear in the same equation or sentence, then $y=(\alpha,p)$.
-We assume $f$ is compactly supported, $\text{supp}f\subset\mathcal{U}$, and $f$ is sufficiently smooth, so $g(y):=\mathcal{R}f(y)$ is a continuous function.
-Assumption 3.1 implies that $\text{supp}\,(\mathcal{R}f)$ is compact if $\text{supp}\,f$ is compact.
+We assume $f$ is compactly supported, $\text{supp}f\subset\mathcal{U}$, and $f$ is sufficiently smooth, so $\hat{f}(y):=\mathcal{R}f(y)$ is a continuous function.
+Assumption 3.1 implies that $\text{supp}\,\hat{f}$ is compact if $\text{supp}\,f$ is compact.
+The smoothness of $\mathcal{S}_{y}$ follows from Assumption 3.2(1) below.
 The critical set of the phase $\nu\Phi_{1}$ is
-(3.2) $\Sigma:=\{(y,x,\nu)\in\mathcal{V}\times\mathcal{U}\times\mathbb{R}:\ p-\Phi(x,\alpha)=0\}.$
-Since $\partial_{\nu}(\nu\Phi_{1})=\Phi_{1}$ and $\text{d}_{(x,y,\nu)}\Phi_{1}\not=0$, the phase $\nu\Phi_{1}(x,\alpha)$ is clean and nondegenerate [21, Definition 21.2.15].
-The phase $\nu\Phi_{1}(x,\alpha)$ determines a homogeneous canonical relation from $T^{*}\mathcal{U}$ to $T^{*}\mathcal{V}$ with the graph
-(3.3) $\begin{split}\mathcal{C}:=&\{(y,\nu(-\Phi_{\alpha}^{\prime}(x,\alpha),1));(x,\nu\Phi_{x}^{\prime}(x,y)):(y,x,\nu)\in\Sigma\}.\end{split}$
-Also, $\mathcal{R}\in I^{-1/2}(\mathcal{V}\times\mathcal{U},\mathcal{C})$ is an FIO of order $-1/2$ from $\mathcal{U}$ to $\mathcal{V}$ associated with $\mathcal{C}$ [53, Section 5.VIII].
-Hence $\mathcal{R}:H_{c}^{s}(\mathcal{U})\to H_{loc}^{s+(1/2)}(\mathcal{V})$ is continuous for any $s\in\mathbb{R}$ [22, Corollary 25.3.2].
-Next we compute the following determinant [53, equations (6.3) and (6.4), Chapter VIII]:
-(3.4) $\begin{split}\Delta_{\Phi}(x,y,\nu):=&\text{det}\begin{pmatrix}[\nu\Phi_{1}(x,\alpha)]_{xy}^{\prime\prime}&[\nu\Phi_{1}(x,\alpha)]_{x\nu}^{\prime\prime}\\ [\nu\Phi_{1}(x,\alpha)]_{\nu y}^{\prime\prime}&[\nu\Phi_{1}(x,\alpha)]_{\nu\nu}^{\prime\prime}\end{pmatrix}\\ =&\text{det}\begin{pmatrix}-\nu\Phi_{x\alpha}^{\prime\prime}&(0,0)^{T}&-\Phi_{x}^{\prime}\\ -\Phi_{\alpha}^{\prime}&1&0\end{pmatrix}\\ =&\nu\text{det}\begin{pmatrix}\text{d}_{x}\Phi&\text{d}_{x}(\partial_{\alpha}\Phi)\end{pmatrix}.\end{split}$
-Here $\Phi_{x}^{\prime}=(\partial_{x_{1}}\Phi,\partial_{x_{2}}\Phi)^{T}$, and all the other derivatives are defined similarly.
+(3.2) $\Sigma:=\{(y,x,\nu)\in\mathcal{V}\times\mathcal{U}\times(\mathbb{R}\setminus 0):\ p-\Phi(x,\alpha)=0\}.$
+Since $\partial_{\nu}(\nu\Phi_{1})=\Phi_{1}$ and $\text{d}_{(x,y,\nu)}\Phi_{1}\not=0$, the phase $\nu\Phi_{1}(x,\alpha)$ is clean and nondegenerate [24, Definition 21.2.15].
+The phase $\nu\Phi_{1}(x,\alpha)$ determines the homogeneous canonical relation from $T^{*}\mathcal{U}$ to $T^{*}\mathcal{V}$
+(3.3) $\begin{split}\mathcal{C}:=&\{(y,\nu(-\text{d}_{\alpha}\Phi(x,\alpha),1));(x,\nu\text{d}_{x}\Phi(x,\alpha)):(y,x,\nu)\in\Sigma\}.\end{split}$
+Also, $\mathcal{R}\in I^{-1/2}(\mathcal{V}\times\mathcal{U},\mathcal{C})$ is an FIO of order $-1/2$ from $\mathcal{U}$ to $\mathcal{V}$ associated with $\mathcal{C}$ [56, Section VIII.5].
+Hence $\mathcal{R}:H_{0}^{s}(\mathcal{U}_{b})\to H_{loc}^{s+(1/2)}(\mathcal{V})$ is continuous for any $s\in\mathbb{R}$ and any domain $U_{b}\Subset U$ [26, Corollary 25.3.2].
+Next we compute the following determinant [56, equations (6.3) and (6.4), Chapter VIII]:
+(3.4) $\begin{split}\Delta_{\Phi}(x,y,\nu):=&\text{det}\begin{pmatrix}[\nu\Phi_{1}(x,\alpha)]_{xy}^{\prime\prime}&[\nu\Phi_{1}(x,\alpha)]_{x\nu}^{\prime\prime}\\ [\nu\Phi_{1}(x,\alpha)]_{\nu y}^{\prime\prime}&[\nu\Phi_{1}(x,\alpha)]_{\nu\nu}^{\prime\prime}\end{pmatrix}\\ =&\text{det}\begin{pmatrix}-\nu\text{d}_{x}\Phi_{\alpha}^{\prime}&-\Phi_{\alpha}^{\prime}\\ 0\quad 0&1\\ -\text{d}_{x}\Phi&0\end{pmatrix}=\nu\text{det}\begin{pmatrix}\text{d}_{x}\Phi\\ \text{d}_{x}\Phi_{\alpha}^{\prime}\end{pmatrix}.\end{split}$
+We will impose conditions to ensure that the last determinant is not zero.
+In this case, $\mathcal{C}$ is a local canonical graph [56, Definition 6.1, Chapter VIII].
 The derivatives on the right are computed at $(x,\alpha,p=\Phi(x,\alpha))$.
 As is easily seen,
-(3.5) $\Delta_{\Phi}(x,y,\nu)=\nu\Delta_{\Phi}(x,\alpha).$
+(3.5) $\Delta_{\Phi}(x,y,\nu){\color[rgb]{0,0,0}=\nu\Delta_{\Phi}(x,y,1)=:}\nu\Delta_{\Phi}(x,\alpha).$
 Since the determinant on the last line in (3.4) does not depend on $p$ and $\nu$, we use the simpler notation $\Delta_{\Phi}(x,\alpha)$ instead of $\Delta_{\Phi}(x,y,1)$.
@@
   3. (3)
-    For each $(x,\xi)\in T^{*}\mathcal{U}$, there exist $\alpha\in I_{\alpha}$ and $\nu\in\mathbb{R}$ such that $\xi=\nu\text{d}_{x}\Phi(x,\alpha)$.
+    For each $(x,\xi)\in T^{*}\mathcal{U}$, there exist $\alpha\in\mathcal{I}_{\alpha}$ and $\nu\in\mathbb{R}$ such that $\xi=\nu\text{d}_{x}\Phi(x,\alpha)$.
   4. (4)
-    $W(x,y)>c$ on $\mathcal{U}\times\mathcal{V}$.
+    $W\in C^{\infty}(\mathcal{U}\times\mathcal{V})$ and $|W(x,y)|>c$ on $\mathcal{U}\times\mathcal{V}$.
   5. (5)
-    For any bounded domain $\mathcal{U}_{b}\subset\mathcal{U}$ there exists $c=c(\mathcal{U}_{b})>0$ such that
-    (3.6) $\|f\|_{H^{-1/2}(\mathcal{U})}\leq c\|\mathcal{R}f\|_{L^{2}(\mathcal{V})}\text{for any }f\in H_{c}^{-1/2}(\mathcal{U}_{b}).$
+    For any domain $\mathcal{U}_{b}\Subset\mathcal{U}$ there exists $c=c(\mathcal{U}_{b})>0$ such that
+    (3.6) $\|f\|_{H^{-1/2}(\mathcal{U})}\leq c\|\mathcal{R}f\|_{L^{2}(\mathcal{V})}\text{for any }f\in H_{0}^{-1/2}(\mathcal{U}_{b}).$
 Assumption 3.2(1) implies that $\Delta_{\Phi}(x,y,\nu)\not=0$ on $\mathcal{U}\times\mathcal{V}\times(\mathbb{R}\setminus 0)$.
-In turn, this implies that $\mathcal{C}$ is the graph of a local diffeomorphism [53, Section VI.4].
+In turn, this implies that locally $\mathcal{C}$ is the graph of a diffeomorphism [56, Section VI.4].
 Assumption 3.2(2) implies that the natural projection $\mathcal{C}\to T^{*}\mathcal{V}$ is an embedding (injective immersion).
@@
 Assumption 3.2(5) implies that $\mathcal{R}$ is boundedly invertible in the scale of Sobolev spaces.
+Let us discuss the last assumption in more detail.
+The most difficult part of the assumption is to check the injectivity of the GRT.
+There is no general theory that can ascertain whether a given GRT is injective or not.
+The question of injectivity is studied on a case by case basis.
+See, e.g., [15, 23] where the injectivity (and stability) are shown for some families of GRTs.
+If the injectivity is established, stability estimates of the kind (3.6) are obtained using the calculus of FIO, see [15, proof of Theorem 2.3] and [23, proof of Theorem 2].
+When the GRT integrates over simple curves, the injectivity is easier to establish.
+The injectivity of the classical Radon transform is well known [42].
+The case when the GRT integrates over circles has been studied as well, see [3, 6] and references therein.
+Consider a circular GRT, $\mathcal{R}$.
+Suppose $\mathcal{R}$ integrates $C_{0}^{\infty}(\mathbb{R}^{2})$ functions over circles of arbitrary radius, and their centers are confined to a set $\Xi$.
+One calls $\Xi$ a set of injectivity of $\mathcal{R}$ if $\mathcal{R}$ is injective.
+A complete description of sets of injectivity was first obtained in [3].
+For example, the boundary of any bounded set in the plane is an injectivity set for $\mathcal{R}$ [6, Corollary 5].
 3.2. The function to be reconstructed
@@
 where
-(3.8) $\begin{split}&H\in C^{\infty}(\mathcal{U}),\quad\text{d}H\not=0\text{ on }\mathcal{U};\\ &\tilde{f}(x,\lambda)=-i\tilde{f}_{0}(x)\lambda^{-1}+\tilde{R}(x,\lambda),\forall x\in\mathcal{U},|\lambda|\geq 1;\\ &\tilde{f}(x,\lambda)\equiv 0\ \forall x\in\mathcal{U}\setminus\mathcal{U}_{b}\text{ for some bounded domain }\mathcal{U}_{b}\subset\mathcal{U};\\ &\tilde{f}\in S^{-1}(\mathcal{U}\times\mathbb{R}),\ \tilde{f}_{0}\in C_{0}^{\infty}(\mathcal{U}_{b}),\ \tilde{R}\in S^{-2}(\mathcal{U}\times\mathbb{R}),\end{split}$
+(3.8) $\begin{split}&H\in C^{\infty}(\mathcal{U}),\quad\text{d}H\not=0\text{ on }\mathcal{U};\\ &\tilde{f}(x,\lambda)=-i\tilde{f}_{0}(x)\lambda^{-1}+\tilde{R}(x,\lambda),\forall x\in\mathcal{U},|\lambda|\geq 1;\\ &\tilde{f}(x,\lambda)\equiv 0\ \forall x\in\mathcal{U}\setminus\mathcal{U}_{b}\text{ for some domain }\mathcal{U}_{b}\Subset\mathcal{U};\\ &\tilde{f}\in S^{-1}(\mathcal{U}\times\mathbb{R}),\ \tilde{f}_{0}\in C_{0}^{\infty}(\mathcal{U}_{b}),\ \tilde{R}\in S^{-2}(\mathcal{U}\times\mathbb{R}),\end{split}$
 for some $\tilde{f}_{0}$ and $\tilde{R}$. From (3.7) and (3.8) it follows that
 (3.9) sing $\text{sing\hskip 1.0ptsupp}(f)\subset\mathcal{S}:=\{x\in\mathcal{U}:\ H(x)=0\}.$
-Thus $\mathcal{S}$ is a smooth curve (submanifold) with possibly several connected components.
-Equations (3.7) and (3.8) imply that $f\in I^{-1}(\mathcal{U},\mathcal{S})$ is a conormal distribution [21, Section 18.2] and
-(3.10) $\begin{split}&f(x+\Delta x)\sim\tilde{f}_{0}(x)(\text{sgn}h/2),\ h:=\text{d}H(x)\Delta x,|\Delta x|\to 0,\ \forall x\in\mathcal{S}.\end{split}$
+Thus $\mathcal{S}$ is a smooth curve (submanifold).
+Equations (3.7) and (3.8) imply that $f\in I^{-1}(\mathcal{U},\mathcal{S})$ is a conormal distribution [24, Section 18.2] and
+(3.10) $\begin{split}&f(x+\Delta x)\sim\tilde{f}_{0}(x)(\text{sgn}h/2),\ h:=\text{d}H(x)\cdot\Delta x,|\Delta x|\to 0,\ \forall x\in\mathcal{S}.\end{split}$
 Thus, $\tilde{f}_{0}(x)$ is the value of the jump of $f$ across $\mathcal{S}$ at $x\in\mathcal{S}$.
+For this reason, in what follows we use the notation $\Delta f(x)$ instead of $\tilde{f}_{0}(x)$.
+The direction for computing the jump value across $\mathcal{S}$ is encoded in (3.7), (3.8).
 Pick any $x\in\mathcal{S}$ and $y\in\mathcal{V}$ such that $\mathcal{S}_{y}$ is tangent to $\mathcal{S}$ at $x$. The curvatures of $\mathcal{S}$ and $\mathcal{S}_{y}$ at $x$ are given by
-(3.11) $\varkappa_{\mathcal{S}}(x)=-\frac{(H_{xx}^{\prime\prime}(x)e,e)}{|H_{x}^{\prime}(x)|},\ \varkappa_{\mathcal{S}_{y}}(x)=-\frac{(\Phi_{xx}^{\prime\prime}(x,\alpha)e,e)}{|\Phi_{x}^{\prime}(x,\alpha)|},$
+(3.11) $\varkappa_{\mathcal{S}}(x)=-\frac{\text{d}_{x}^{2}H(x)(e,e)}{|\text{d}_{x}H(x)|},\quad\varkappa_{\mathcal{S}_{y}}(x)=-\frac{\text{d}_{x}^{2}\Phi(x,\alpha)(e,e)}{|\text{d}_{x}\Phi(x,\alpha)|},$
 where $e$ is a unit vector tangent to $\mathcal{S}$ at $x$.
-If $H_{x}^{\prime}(x)\cdot\Phi_{x}^{\prime}(x,\alpha)<0$, the curvatures are not consistent with each other.
+If $\text{d}_{x}H(x)\cdot\text{d}_{x}\Phi(x,\alpha)<0$, the curvatures are not consistent with each other.
 In this case we flip the $p$-axis and replace $\Phi$ with $-\Phi$ to make them consistent.
-Thus, we can assume without loss of generality that $H_{x}^{\prime}(x)\cdot\Phi_{x}^{\prime}(x,\alpha)>0$.
+Thus, we can assume without loss of generality that
+(3.12) $\text{d}_{x}H(x)\cdot\text{d}_{x}\Phi(x,\alpha)>0.$
 Assumption 3.3 (Properties of $f$).
@@
     For each pair $(x,y)\in\mathcal{S}\times\mathcal{V}$ such that $\mathcal{S}_{y}$ is tangent to $\mathcal{S}$ at $x$, one has $\varkappa_{\mathcal{S}}(x)-\varkappa_{\mathcal{S}_{y}}(x)>0$.
-The last condition is not restrictive.
+The requirement that the difference of the curvatures in Assumption 3.3(2) be positive is not restrictive.
 Basically, it says that $\varkappa_{\mathcal{S}}(x)\not=\varkappa_{\mathcal{S}_{y}}(x)$ whenever $\mathcal{S}$ and $\mathcal{S}_{y}$ are tangent at some $x$.
@@
 This is always tacitly assumed in what follows.
+Assumption 3.3(2) imposes a nontrivial restriction on the types of functions $f$ we consider.
+The assumption prevents any $\mathcal{S}_{y}$ from being tangent to $\mathcal{S}$ of order higher than one, i.e., only simple tangency between $\mathcal{S}_{y}$ and $\mathcal{S}$ is allowed.
+In particular, $\mathcal{S}_{y}$ cannot be tangent to $\mathcal{S}$ along an entire curve segment.
+While the latter is generally not a significant issue in medical CT, such situations are more common in industrial applications of X-ray CT when scanning artificial (man-made) objects.
+In these cases, $\mathcal{S}_{y}$ are straight lines, and $\mathcal{S}$ may contain a flat segment.
+Mathematical analysis of these cases presents additional challenges.
+This is because $\hat{f}(y)$ is generally more singular in a neighborhood of the exceptional $y$ described above than in the case of simple tangency.
+An analysis of these more singular cases is beyond the scope of the paper.
 3.3. GRT data and the reconstruction algorithm
@@
 (3.13) $y_{j}=(\Delta\alpha j_{1},\Delta pj_{2})\in\mathcal{V},\ \Delta p=\epsilon,\ \Delta\alpha=\mu\epsilon,\ j=(j_{1},j_{2})\in\mathbb{Z}^{2},$
-The interpolated data, denoted $g_{\epsilon}(y)$, is computed by:
-(3.13) $\begin{split}&g_{\epsilon}(y):=\sum_{j}\varphi_{\epsilon}(y-y_{j})g(y_{j}),\ y\in\mathcal{V},\\ &\varphi_{\epsilon}(y):=\varphi_{\alpha}(\alpha/\Delta\alpha)\varphi_{p}(p/\Delta p),\end{split}$
+The interpolated data, denoted $\hat{f}_{\epsilon}(y)$, is computed by:
+(3.14) $\begin{split}&\hat{f}_{\epsilon}(y):=\sum_{y_{j}\in\mathcal{V}}\varphi_{\epsilon}(y-y_{j})\hat{f}(y_{j}),\ y\in\mathcal{V},\\ &\varphi_{\epsilon}(y):=\varphi(y/\epsilon),\ \varphi(y):=\varphi_{\alpha}(\alpha/\mu)\varphi_{p}(p),\end{split}$
 where $\varphi$ is an interpolation kernel.
@@
   1. (1)
-    $\varphi_{*}\in C_{0}^{2}(\mathbb{R})$;
+    $\varphi_{*}\in C_{0}^{2}(\mathbb{R})$.
   2. (2)
@@
 Let $f$ satisfy Assumption 3.3 and $\mathcal{R}$ satisfy Assumptions 3.1, 3.2.
-Lemma 4.2 below implies $\|g_{\epsilon}\|_{L^{2}(\mathcal{V})}\leq c$, $0<\epsilon\ll 1$.
+By Lemma 4.2 below, $\|\hat{f}_{\epsilon}\|_{L^{2}(\mathcal{V})}\leq c$, $0<\epsilon\ll 1$.
 Reconstruction is achieved by minimizing the quadratic functional
-(3.16) $f_{\epsilon}=\text{argmin}_{f\in H_{c}^{1}(\mathcal{U}_{b})}\Psi(f),\ \Psi(f):=\|\mathcal{R}f-g_{\epsilon}\|_{L^{2}(\mathcal{V})}^{2}+\kappa\epsilon^{3}\|\nabla f\|_{L^{2}(\mathcal{U})}^{2},$
-with some fixed regularization parameter $\kappa>0$ and with $\mathcal{U}_{b}$ the same as in (3.8) (or any other bounded domain containing $\mathcal{U}_{b}$).
+(3.17) $f_{\epsilon}=\text{argmin}_{f\in H_{0}^{1}(\mathcal{U}_{b})}\Psi(f),\ \Psi(f):=\|\mathcal{R}f-\hat{f}_{\epsilon}\|_{L^{2}(\mathcal{V})}^{2}+\kappa\epsilon^{3}\|\partial_{x}f\|_{L^{2}(\mathcal{U})}^{2},$
+with some fixed regularization parameter $\kappa>0$ and with $\mathcal{U}_{b}$ the same as in (3.8) (or with any other $\mathcal{U}_{b}^{\prime}$ that satisfies $\mathcal{U}_{b}\subset\mathcal{U}_{b}^{\prime}\Subset U$).
 The following result is proven in Appendix A.
@@
 Theorem 3.6.
-Pick any $x_{0}\in\mathcal{S}$ and $A>0$.
+Pick any $x_{0}\in\mathcal{S}$ and $A_{0}>0$.
 Suppose there is only one $y_{0}=(\alpha_{0},p_{0})\in\mathcal{V}$ such that $\mathcal{S}_{y_{0}}$ is tangent to $\mathcal{S}$ at $x_{0}$.
@@
 One has
-(3.17) $\lim_{\epsilon\to 0}\big{(}f_{\epsilon}(x_{0}+\epsilon\check{x})-f_{\epsilon}(x_{0})\big{)}=\tilde{f}_{0}(x_{0})\Upsilon\big{(}\vec{\Theta}_{0}\cdot\check{x}\big{)},\ |\check{x}|<A,$
+(3.18) $\lim_{\epsilon\to 0^{+}}\big{(}f_{\epsilon}(x_{0}+\epsilon\check{x})-f_{\epsilon}(x_{0})\big{)}=\Delta f(x_{0})\Upsilon\big{(}\vec{\Theta}_{0}\cdot\check{x}\big{)},\ |\check{x}|<A_{0},$
 where the limit is uniform in $\check{x}$ and
-(3.18) $\begin{split}&\Upsilon(r)=\int_{\mathbb{R}}h_{0}(u)\int_{u_{1}}^{u_{1}+r}R(t)\text{d}t\text{d}u,\ \vec{\Theta}_{0}:=\frac{\Phi_{x}^{\prime}(x_{0},\alpha_{0})}{|\Phi_{x}^{\prime}(x_{0},\alpha_{0})|},\\ &h_{0}(u)=\int_{\mathbb{R}}\varphi_{\alpha}(s)\varphi_{p}(\mu s\Phi_{\alpha}^{\prime}(x_{0},\alpha_{0})+u)\text{d}s,\ u_{1}=\frac{u}{|\Phi_{x}^{\prime}(x_{0},\alpha_{0})|},\\ &R(t)=\mathcal{F}_{\lambda\to t}^{-1}\bigg{(}\bigg{[}1+\frac{\kappa}{2\pi\rho_{0}}|\lambda|^{3}\bigg{]}^{-1}\bigg{)},\ \rho_{0}=\frac{W^{2}(x_{0},y_{0})|\Phi_{x}^{\prime}(x_{0},\alpha_{0})|}{|\Delta_{\Phi}(x_{0},\alpha_{0})|}.\end{split}$
+(3.19) $\begin{split}&\Upsilon(r)=\int_{\mathbb{R}}h_{0}(u)\int_{u_{1}}^{u_{1}+r}R(t)\text{d}t\text{d}u,\ \vec{\Theta}_{0}:=\frac{\text{d}_{x}\Phi(x_{0},\alpha_{0})}{|\text{d}_{x}\Phi(x_{0},\alpha_{0})|},\\ &h_{0}(u)=\int_{\mathbb{R}}\varphi_{\alpha}(s)\varphi_{p}(\mu s\Phi_{\alpha}^{\prime}(x_{0},\alpha_{0})+u)\text{d}s,\ u_{1}=\frac{u}{|\text{d}_{x}\Phi(x_{0},\alpha_{0})|},\\ &R(t)=\mathcal{F}_{\lambda\to t}^{-1}\left(\left[1+\frac{\kappa}{2\pi\rho_{0}}|\lambda|^{3}\right]^{-1}\right),\ \rho_{0}=\frac{W^{2}(x_{0},y_{0})|\text{d}_{x}\Phi(x_{0},\alpha_{0})|}{|\Delta_{\Phi}(x_{0},\alpha_{0})|}.\end{split}$
 Moreover,
 (3.20) $\Upsilon(0)=0,\quad\Upsilon(\pm\infty)=\pm 1/2.$
-Thus $\Upsilon(+\infty)-\Upsilon(-\infty)=1$, and $\Upsilon$ is a true DTB function.
-In the case of redundant data we obtain the following generalization, which is proven in section 9.
+Thus $\Upsilon(+\infty)-\Upsilon(-\infty)=1$, and $\Upsilon$ is indeed a DTB function.
+In the proof of Theorem 3.6 we assume without loss of generality that $\alpha_{0}=0$. In the case of redundant data we obtain the following generalization, which is proven in Section 9.
 Theorem 3.7.
-Pick any $x_{0}\in\mathcal{S}$ and $A>0$.
+Pick any $x_{0}\in\mathcal{S}$ and $A_{0}>0$.
 Suppose there exist finitely many $y_{l}=(\alpha_{l},p_{l})\in\mathcal{V}$ such that $\mathcal{S}_{y_{l}}$ is tangent to $\mathcal{S}$ at $x_{0}$.
@@
 One has
-(3.20) $\lim_{\epsilon\to 0}\big{(}f_{\epsilon}(x_{0}+\epsilon\check{x})-f_{\epsilon}(x_{0})\big{)}=\tilde{f}_{0}(x_{0})\frac{\textstyle\sum_{l}\rho_{l}\Upsilon_{l}\big{(}\vec{\Theta}_{0}\cdot\check{x}\big{)}}{\textstyle\sum_{l}\rho_{l}},\ |\check{x}|<A,$
+(3.21) $\lim_{\epsilon\to 0^{+}}\big{(}f_{\epsilon}(x_{0}+\epsilon\check{x})-f_{\epsilon}(x_{0})\big{)}=\Delta f(x_{0})\frac{\textstyle\sum_{l}\nu_{l}\Upsilon_{l}\big{(}\vec{\Theta}_{0}\cdot\check{x}\big{)}}{\textstyle\sum_{l}\nu_{l}},\ |\check{x}|<A_{0},$
 where the limit is uniform in $\check{x}$ and
-(3.21) $\begin{split}&\Upsilon_{l}(r)=\int_{\mathbb{R}}h_{l}(u)\int_{u_{l}}^{u_{l}+r}R(t)\text{d}t\text{d}u,\ \vec{\Theta}_{0}:=\frac{H_{x}^{\prime}(x_{0})}{|H_{x}^{\prime}(x_{0})|},\\ &h_{l}(u)=\int_{\mathbb{R}}\varphi_{\alpha}(s)\varphi_{p}(\mu s\Phi_{\alpha}^{\prime}(x_{0},\alpha_{l})+u)\text{d}s,\ u_{l}=\frac{u}{|\Phi_{x}^{\prime}(x_{0},\alpha_{l})|},\\ &R(t)=\mathcal{F}_{\lambda\to t}^{-1}\bigg{(}\bigg{[}1+\frac{\kappa}{2\pi\textstyle\sum_{l}\rho_{l}}|\lambda|^{3}\bigg{]}^{-1}\bigg{)},\ \rho_{l}:=\frac{W^{2}(x_{0},y_{l})|\Phi_{x}^{\prime}(x_{0},\alpha_{l})|}{|\Delta_{\Phi}(x_{0},\alpha_{l})|}.\end{split}$
+(3.22) $\begin{split}&\Upsilon_{l}(r)=\int_{\mathbb{R}}h_{l}(u)\int_{u_{l}}^{u_{l}+r}R(t)\text{d}t\text{d}u,\ \vec{\Theta}_{0}:=\frac{\text{d}_{x}H(x_{0})}{|\text{d}_{x}H(x_{0})|},\\ &h_{l}(u)=\int_{\mathbb{R}}\varphi_{\alpha}(s)\varphi_{p}(\mu s\Phi_{\alpha}^{\prime}(x_{0},\alpha_{l})+u)\text{d}s,\ u_{l}=\frac{u}{|\text{d}_{x}\Phi(x_{0},\alpha_{l})|},\\ &R(t)=\mathcal{F}_{\lambda\to t}^{-1}\bigg{(}\bigg{[}1+\frac{\kappa}{2\pi\textstyle\sum_{l}\nu_{l}}|\lambda|^{3}\bigg{]}^{-1}\bigg{)},\ \nu_{l}:=\frac{W^{2}(x_{0},y_{l})|\text{d}_{x}\Phi(x_{0},\alpha_{l})|}{|\Delta_{\Phi}(x_{0},\alpha_{l})|}.\end{split}$
+Taken together, Theorems 3.6 and 3.7 provide an easy to use characterization of the reconstructed image in an $O(\epsilon)$-size neighborhood of an edge.
+This characterization is in terms of universal functions $\Upsilon_{0}$ and $\Upsilon_{l}$, which are independent of $f$ and can be easily precomputed.
+The only contribution from $f$ is $\Delta f(x_{0})$, the value of the jump of $f$ at $x_{0}$, which enters the formulas (3.18) and (3.21) as a multiplicative factor.
+Although the proofs are quite extensive, the final result is satisfying due to its simple form and straightforward applicability.
+3.4. Example
+To briefly illustrate the above theorems, we consider the classical Radon transform in the plane that integrates over the lines $\mathcal{S}_{(\alpha,p)}=\{x\in\mathbb{R}^{2}:\vec{\alpha}\cdot x=p\}$, $\vec{\alpha}:=(\cos\alpha,\sin\alpha)$.
+Then $\mathcal{I}_{\alpha}=[0,2\pi)\sim S^{1}$, $\mathcal{I}_{p}=\mathbb{R}$, $\Phi(x,\alpha)=\vec{\alpha}\cdot x$, $W(x,y)\equiv 1$, and
+(3.23) $\begin{split}&\text{d}_{x}\Phi(x,\alpha)=\vec{\alpha},\ |\text{d}_{x}\Phi(x,\alpha)|=1,\ \Phi_{\alpha}^{\prime}(x,\alpha)=\vec{\alpha}^{\perp}\cdot x,\\ &\Delta_{\Phi}=\det\begin{pmatrix}\vec{\alpha}\\ \vec{\alpha}^{\perp}\end{pmatrix}=1,\ \vec{\alpha}^{\perp}:=(-\sin\alpha,\cos\alpha).\end{split}$
+Pick any $x_{0}\in\mathcal{S}$.
+Let $\vec{\alpha}_{0}$ be perpendicular to $\mathcal{S}$ at $x_{0}$.
+Strictly speaking, there are two data points $y_{1}=(\alpha_{0},\vec{\alpha}_{0}\cdot x_{0})$ and $y_{2}=(\alpha_{0}+\pi,-\vec{\alpha}_{0}\cdot x_{0})$ such that $\mathcal{S}_{y_{l}}$ is tangent to $\mathcal{S}$ at $x_{0}$.
+Hence we should use Theorem 3.7.
+From (3.22) and (3.23), $u_{l}=u$, $\nu_{l}=1$, and
+(3.24) $\begin{split}&\Upsilon_{l}(r)=\int_{\mathbb{R}}h_{l}(u)\int_{u}^{u+r}R(t)\text{d}t\text{d}u,\\ &h_{l}(u)=\int_{\mathbb{R}}\varphi_{\alpha}(s)\varphi_{p}(\mu s\vec{\alpha}_{l}^{\perp}\cdot x_{0}+u)\text{d}s,\ l=1,2,\\ &R(t)=\mathcal{F}_{\lambda\to t}^{-1}\bigg{(}\big{[}1+\frac{\kappa/2}{2\pi}|\lambda|^{3}\big{]}^{-1}\bigg{)}.\end{split}$
+By Assumption 3.4(2), $\varphi_{\alpha}$ is even and $\vec{\alpha}_{1}^{\perp}\cdot x_{0}=-\vec{\alpha}_{2}^{\perp}\cdot x_{0}$. Therefore, $h_{1}(u)\equiv h_{2}(u)$ and (cf. (3.21))
+(3.25) $\frac{\textstyle\sum_{l}\nu_{l}\Upsilon_{l}(r)}{\textstyle\sum_{l}\nu_{l}}=\int_{\mathbb{R}}h_{1}(u)\int_{u}^{u+r}R(t)\text{d}t\text{d}u.$
+This is the formula we would have obtained by formally using Theorem 3.6 and replacing $\kappa$ with $\kappa/2$.
+The result is not surprising, as $\mathcal{S}_{y_{1}}$ and $\mathcal{S}_{y_{2}}$ define the same tangent line.
+See Section 10 for an example involving the GRT that integrates over circles.
+The functions $\Upsilon_{l}$ have been used to compute the predicted behavior of the reconstruction (the DTB function), which is shown as a blue curve in Figure 3.
 4. Behavior of $\mathcal{R}f$ near its singular support
 Let $\Gamma$ be the set of all $y\in\mathcal{V}$ such that $\mathcal{S}_{y}$ is tangent to $\mathcal{S}$. Since $\mathcal{R}$ is an FIO with the canonical relation (3.3), we have
-(4.1) $\Gamma=\text{sing\hskip 1.0ptsupp}(\mathcal{R}f),\ WF(\mathcal{R}f)\subset N^{*}\Gamma=\mathcal{C}\circ N^{*}\mathcal{S}.$
-Here $N^{*}\mathcal{S}$ is the conormal bundle of $\mathcal{S}$, and similarly for $\Gamma$ [52, section I.6].
-In particular, $\mathcal{R}f$ is smooth away from $\Gamma$.
-In this section we describe the leading singular behavior of $\mathcal{R}f$ near $\Gamma$.
-Using a partition of unity and the linearity of $\mathcal{R}$, we can assume without loss of generality that $\mathcal{S}_{y}$, $y\in\Gamma$, is tangent to $\mathcal{S}$ only at one point.
-Fix any $\tilde{y}\in\Gamma$, and let $\tilde{x}\in\mathcal{S}$ be the corresponding point of tangency.
+(4.1) $\Gamma=\text{sing\hskip 1.0ptsupp}\hat{f},\ \text{WF}(\hat{f})\subset N^{*}\Gamma=\mathcal{C}\circ N^{*}\mathcal{S}.$
+Here $N^{*}\mathcal{S}$ is the conormal bundle of $\mathcal{S}$, and similarly for $\Gamma$ [55, section I.6].
+In particular, $\hat{f}$ is smooth away from $\Gamma$.
+In this section we describe the leading singular behavior of $\hat{f}$ near $\Gamma$.
+Fix any $\tilde{y}\in\Gamma$, and let $\tilde{x}\in\mathcal{S}$ be the corresponding point of tangency, see Figure 1.
 Let $V$ be a sufficiently small neighborhood of $\tilde{y}$.
+Using a partition of unity, the linearity of $\mathcal{R}$, and Assumption 3.3(2), we can assume without loss of generality that $\mathcal{S}_{y}$, $y\in\Gamma\cap V$, is tangent to $\mathcal{S}$ only at one point.
+We underscore that the main goal of this section is to study the singularities of $\mathcal{R}f$.
+The way these singularities are mapped into the reconstructed image is studied in Sections 7, 8, and 9.
 Substitute (3.7) into (3.1).
-The asymptotics as $\lambda\to\infty$ of the resulting integral with respect to $x$ is computed by the stationary phase method [53, Chapter VIII, eqs.
-(2.14)–(2.20)]
-(4.2) $\begin{split}(\mathcal{R}f)(y)=&\frac{1}{2\pi}\int_{\mathbb{R}}\tilde{G}(y,\lambda)\text{d}\lambda,\ y\in V,\\ \tilde{G}(y,\lambda):=&\frac{1}{2\pi}\int_{\mathbb{R}}\int_{\mathcal{U}}\tilde{f}(x,\lambda)e^{i\lambda H(x)}W(x,y)e^{i\nu(p-\Phi(x,\alpha))}\text{d}x\text{d}\nu\\ =&\frac{|\lambda|}{2\pi}\int_{\mathbb{R}}\int_{\mathcal{U}}W(x,y)\tilde{f}(x,\lambda)e^{i\lambda[\nu_{1}(p-\Phi(x,\alpha))+H(x)]}\text{d}x\text{d}\nu_{1}\\ =&\left(\tilde{f}(x_{*},\lambda)\frac{W(x_{*},y)}{|\det M(y)|^{1/2}}\left(\frac{2\pi}{|\lambda|}\right)^{1/2}e(-\text{sgn}\lambda/2)+\tilde{R}(y,\lambda)\right)\\ &\quad\times e^{i\lambda H(x_{*})},\ |\lambda|\geq 1,\ \tilde{R}\in S^{-3/2}(\mathcal{V}\times\mathbb{R}),\ x_{*}=x_{*}(y),\end{split}$
-for some $\tilde{R}$ and $0<\delta\ll 1$. Here $e(t):=e(i\pi t/2)$, and
-(4.3) $M(y)=\begin{pmatrix}0&(-\Phi_{x}^{\prime}(x,\alpha))^{T}\\ -\Phi_{x}^{\prime}(x,\alpha)&H_{xx}^{\prime\prime}(x,\alpha)-\nu_{1}\Phi_{xx}^{\prime\prime}(x,\alpha)\end{pmatrix},\ \nu_{1}=\nu_{1*}(y),x=x_{*}(y),$
+The asymptotics as $\lambda\to\infty$ of the resulting integral with respect to $x$ is computed by the stationary phase method [56, Chapter VIII, eqs.
+(2.14)–(2.20)] using the properties of the Hessian at the stationary point established in Lemma 4.1 below:
+(4.2) $\begin{split}\hat{f}(y)=&\frac{1}{2\pi}\int_{\mathbb{R}}\tilde{G}(y,\lambda)\text{d}\lambda,\ y\in V,\\ \tilde{G}(y,\lambda):=&\frac{1}{2\pi}\int_{\mathbb{R}}\int_{\mathcal{U}}\tilde{f}(x,\lambda)e^{i\lambda H(x)}W(x,y)e^{i\nu(p-\Phi(x,\alpha))}\text{d}x\text{d}\nu\\ =&\frac{|\lambda|}{2\pi}\int_{\mathbb{R}}\int_{\mathcal{U}}W(x,y)\tilde{f}(x,\lambda)e^{i\lambda[\nu_{1}(p-\Phi(x,\alpha))+H(x)]}\text{d}x\text{d}\nu_{1}\\ =&\left(\tilde{f}(x_{*},\lambda)\frac{W(x_{*},y)}{|\det M(y)|^{1/2}}\left(\frac{2\pi}{|\lambda|}\right)^{1/2}e(-\text{sgn}\lambda/2)+\tilde{R}(y,\lambda)\right)\\ &\quad\times e^{i\lambda H(x_{*})},\ |\lambda|\geq 1,\ \tilde{R}\in S^{-3/2}(\mathcal{V}\times\mathbb{R}),\ x_{*}=x_{*}(y),\end{split}$
+for some $\tilde{R}$. Here $e(t):=\exp(i\pi t/2)$, and
+(4.3) $\begin{split}&M(y)=\begin{pmatrix}0&-\text{d}_{x}\Phi(x,\alpha)\\ (-\text{d}_{x}\Phi(x,\alpha))^{T}&\text{d}_{x}^{2}\big{[}H(x,\alpha)-\nu_{1}\Phi(x,\alpha)\big{]}\end{pmatrix},\\ &\nu_{1}=\nu_{1*}(y),\ x=x_{*}(y),\end{split}$
 is the Hessian matrix of the phase at the stationary point $(\nu_{1*}(y),x_{*}(y))$, which is found by solving
 (4.4) $p=\Phi(x,\alpha),\ \text{d}_{x}H(x)-\nu_{1}\text{d}_{x}\Phi(x,\alpha)=0,\ y\in V.$
-By construction, $x_{*}(\tilde{y})=\tilde{x}$. As before, the existence and local uniqueness of the solution follows from assumptions 3.2(1,3).
+By construction, $x_{*}(\tilde{y})=\tilde{x}$. As before, the existence, local uniqueness, and smoothness of the solution follows from Assumptions 3.2(1,3) and 3.3(2).
+Given a nondegenerate symmetric matrix $M$, let $n_{+}$ and $n_{-}$ denote the number of positive and negative eigenvalues of $M$, respectively.
+Then the signature of $M$, denoted $\text{sgn}\,M$, is the difference $n_{+}-n_{-}$.
 The following result is proven in Appendix B.1.
@@
 Suppose Assumptions 3.13.3 are satisfied.
-Let $\mathcal{S}_{y}$ be tangent to $\mathcal{S}$ at $x\in\mathcal{S}$.
+Let $\mathcal{S}_{y}$ be tangent to $\mathcal{S}$ at $x$.
 Let $\varkappa_{\mathcal{S}}(x)$ and $\varkappa_{\mathcal{S}_{y}}(x)$ denote the curvatures of $\mathcal{S}$ and $\mathcal{S}_{y}$ at $x$, respectively (see (3.11)).
 One has
-(4.5) $\begin{split}\det M(y)=&(\varkappa_{\mathcal{S}}(x)-\varkappa_{\mathcal{S}_{y}}(x)){|\Phi_{x}^{\prime}(x,\alpha)|^{2}}{|H_{x}^{\prime}(x)|}>0,\\ \text{sgn}M(y)=&-1.\end{split}$
+(4.5) $\begin{split}\det M(y)=&(\varkappa_{\mathcal{S}}(x)-\varkappa_{\mathcal{S}_{y}}(x)){|\text{d}_{x}\Phi(x,\alpha)|^{2}}{|\text{d}_{x}H(x)|}>0,\\ \text{sgn}\,M(y)=&-1.\end{split}$
 It follows from the proof of Lemma 4.1 (see (B.1)) that we can locally solve $H(x_{*}(\alpha,p))=0$ for $p$ in terms of $\alpha$: $p=P(\alpha)$, where $P(\alpha)$ is locally smooth. Therefore
@@
 where $\psi(y)\not=0$, $y\in V$, and $\psi$ is locally smooth. Moreover, by (B.1) and (4.6),
-(4.7) $\begin{split}\psi(\alpha,P(\alpha))=&\partial_{p}H(x_{*}(\alpha,p))|_{p=P(\alpha)}=\frac{|H_{x}^{\prime}(x_{*}(y))|}{|\Phi_{x}^{\prime}(x_{*}(y),\alpha)|}>0,\\ (\alpha,P(\alpha))\in&V.\end{split}$
+(4.7) $\begin{split}\psi(\alpha,P(\alpha))=&\partial_{p}H(x_{*}(\alpha,p))|_{p=P(\alpha)}=\frac{|\text{d}_{x}H(x_{*}(y))|}{|\text{d}_{x}\Phi(x_{*}(y),\alpha)|}>0,\\ (\alpha,P(\alpha))\in&V.\end{split}$
 Therefore, $\psi(y)>0$, $y\in V$.
-It is clear that instead of finding $P(\alpha)$ through the intermediary function $x_{*}(y)$, we can find $P(\alpha)$ by solving
-(4.8) $H(x(\alpha))=0,\ H_{x}^{\prime}(x(\alpha))=\nu_{1}(\alpha)\Phi_{x}^{\prime}(x(\alpha),\alpha),\ (\alpha,P(\alpha))\in V,$
+Instead of finding $P(\alpha)$ via the intermediary function $x_{*}(y)$, we can find $P(\alpha)$ by locally solving
+(4.8) $H(x(\alpha))=0,\ \text{d}_{x}H(x(\alpha))=\nu_{1}(\alpha)\text{d}_{x}\Phi(x(\alpha),\alpha),$
 for $\nu_{1}(\alpha)$, $x(\alpha)$ and setting $P(\alpha):=\Phi(x(\alpha),\alpha)$.
 Then $\nu_{1}(\alpha)\equiv\nu_{1*}(\alpha,P(\alpha))$.
-Since $x(\alpha)\in\mathcal{S}$, we can say that $x(\alpha)$ and $P(\alpha)$ are associated with a segment of $\mathcal{S}$ (see Figure 1).
-The size of the segment depends on the domain of $x(\alpha)$.
-If we have in mind one specific $\tilde{x}\in\mathcal{S}$, then we informally say that $x(\alpha)$ and $P(\alpha)$ are associated with $\tilde{x}$ (i.e., with a segment of $\mathcal{S}$ containing $\tilde{x}$).
+Since $x(\alpha)\in\mathcal{S}$, we can say that $P(\alpha)$ and the segment of $\Gamma$ it defines are associated with a segment of $\mathcal{S}$ (see Figure 1).
+If we have in mind one specific $\tilde{x}\in\mathcal{S}$, then we informally say that $P(\alpha)$ and the corresponding segment of $\Gamma$ are associated with $\tilde{x}$ (i.e., with a segment of $\mathcal{S}$ containing $\tilde{x}$).
 Figure 1. Illustration of $x(\alpha)$. In the figure, $y_{k}=(\alpha_{k},P(\alpha_{k}))$, $k=1,2$ and $\tilde{y}=(\tilde{\alpha},P(\tilde{\alpha}))$.
-By (4.2)–(4.6), $\mathcal{R}f$ can be written as
-(4.9) $(\mathcal{R}f)(y)=\frac{1}{2\pi}\int_{\mathbb{R}}\tilde{\upsilon}(y,\lambda)e^{i\lambda(p-P(\alpha))}\text{d}\lambda,\ y\in V,$
+By (4.2)–(4.6), $\hat{f}$ can be written as
+(4.9) $\hat{f}(y)=\frac{1}{2\pi}\int_{\mathbb{R}}\tilde{\upsilon}(y,\lambda)e^{i\lambda(p-P(\alpha))}\text{d}\lambda,\ y\in V,$
 and, with some $c_{m}$, $\tilde{R}$,
-(4.10) $\begin{split}&\tilde{\upsilon}(y,\lambda)=\tilde{\upsilon}_{0}(y)e(-3/2)(\lambda-i0)^{-3/2}+\tilde{R}(y,\lambda),\ |\lambda|\geq 1;\\ &\tilde{\upsilon}\in S^{-3/2}(V\times\mathbb{R});\ \tilde{R}\in S^{-5/2}(V\times\mathbb{R});\\ &\tilde{\upsilon}_{0}(y)=(2\pi)^{1/2}\frac{W(x_{*}(y),y)\tilde{f}_{0}(x_{*}(y))\psi(y)^{1/2}}{|\det M(y)|^{1/2}}\in C^{\infty}(V).\end{split}$
+(4.10) $\begin{split}&\tilde{\upsilon}(y,\lambda)=\tilde{\upsilon}_{0}(y)e(-3/2)(\lambda-i0)^{-3/2}+\tilde{R}(y,\lambda),\ |\lambda|\geq 1;\\ &\tilde{\upsilon}\in S^{-3/2}(V\times\mathbb{R});\ \tilde{R}\in S^{-5/2}(V\times\mathbb{R});\\ &\tilde{\upsilon}_{0}(y)=(2\pi)^{1/2}\frac{W(x_{*}(y),y)\Delta f(x_{*}(y))\psi(y)^{1/2}}{|\det M(y)|^{1/2}}\in C^{\infty}(V).\end{split}$
 The function $\psi(y)$ is absorbed into the frequency variable $\lambda$ leading to the relation
@@
 where $\tilde{G}$ is defined in (4.2).
-To obtain the formula for $\tilde{\upsilon}_{0}$ we use that the leading order term in $\tilde{\upsilon}$ is homogeneous of degree $-3/2$ and the distribution $(\lambda-i0)^{a}$ [14, equations (28) and (30), p.
+To obtain the formula for $\tilde{\upsilon}_{0}$ we use that the leading order term in $\tilde{\upsilon}$ is homogeneous of degree $-3/2$ and the definition of the distribution $(\lambda-i0)^{a}$ [17, equations (28) and (30), p.
 336].
 By (4.9),
-(4.12) sing $\text{sing\hskip 1.0ptsupp}(\mathcal{R}f)\cap V\subset\Gamma:=\{(\alpha,p)\in V:p=P(\alpha)\}.$
-All of the above properties of $\mathcal{R}f$ follow from the FIO calculus [53, Section VIII.5], but for our purposes it is more convenient to provide their explicit derivation (4.2)–(4.11).
-By appealing to a partition of unity and the linearity of the GRT, the following global result is proven in appendix B.2.
+(4.12) sing $\text{sing\hskip 1.0ptsupp}\hat{f}\cap V\subset\Gamma:=\{(\alpha,p)\in V:p=P(\alpha)\}.$
+All of the above properties of $\hat{f}$ follow from the FIO calculus [56, Section VIII.5], but for our purposes it is more convenient to provide their explicit derivation (4.2)–(4.11).
+As mentioned above, Assumptions 3.1(1,3) and 3.3(1) imply $\text{supp}\,\hat{f}$ is compact.
+Consider an open cover of $\text{supp}\,\hat{f}$ by sufficiently small open sets $V_{l}:=I_{\alpha}^{l}\times I_{p}^{l}$, $1\leq l\leq L$, for some large $L$.
+Here $I_{\alpha}^{l}$ and $I_{p}^{l}$ are intervals.
+Using a partition of unity subordinate to this cover, we can consider one such set, $V_{l}$.
+Generally, the intersection $\Gamma\cap V_{l}$ is a single smooth curve segment.
+If $\Gamma$ self-intersects in $V_{l}$, we formally duplicate $V_{l}$ as many times as needed to make sure each intersection $\Gamma\cap V_{l}$ is identified with no more than one smooth curve segment, denoted $\Gamma_{l}$.
+Let $p=P_{l}(\alpha)$, $\alpha\in I_{\alpha}^{l}$, be the local parameterization for $\Gamma_{l}$.
+Unless stated otherwise, we consider only one set from the partition, denoted $V=I_{\alpha}\times I_{p}$.
+The corresponding $P_{l}(\alpha)$ is denoted $P(\alpha)$.
+Also, we denote $J_{\alpha}:=\{j_{1}\in\mathbb{Z}:\alpha_{j_{1}}\in I_{\alpha}\}$, and similarly for $J_{p}$.
+There are $O(1/\epsilon)$ points in $J_{\alpha}$ and $J_{p}$.
+The following result is proven in Appendix B.2.
 Lemma 4.2.
-Suppose assumptions 3.13.4 are satisfied. Let $g=\mathcal{R}f$. Then $g,g_{\epsilon}\in C_{0}^{1/2}(\mathcal{V})$ and $\|g_{\epsilon}\|_{C^{1/2}(\mathcal{V})}<c$, $0<\epsilon\ll 1$. Also,
-(4.13) $\begin{split}&g(y)=\textstyle\sum_{l}a_{l}(y)(p-P_{l}(\alpha))_{+}^{1/2}+\Delta g(y),\\ &a_{l}(y)=4\pi^{1/2}\tilde{\upsilon}_{0,l}(y),\ a_{l}\in C_{0}^{\infty}(\mathcal{V}),\ \Delta g\in C_{0}^{3/2}(\mathcal{V}),\end{split}$
-and $\|\Delta g_{\epsilon}\|_{C^{3/2}(\mathcal{V})}<c$ for all $0<\epsilon\ll 1$. Here, for each $l$,
-  1. (1)
-    $\text{supp}\,a_{l}$ is contained in a small neighborhood of some $y_{l}\in\Gamma$;
-  2. (2)
-    $p=P_{l}(\alpha)$ is a local solution to $H(x_{*}(\alpha,p))=0$ for $p$ in terms of $\alpha$, which is defined for all $\alpha$ such that $(\alpha,P_{l}(\alpha))\in\text{supp}\,a_{l}$; and
-  3. (3)
-    $\tilde{\upsilon}_{0,l}(y)$ is computed as in (4.10).
+Suppose Assumptions 3.13.4 are satisfied.
+Pick any $\chi_{V}\in C_{0}^{\infty}(V)$.
+Let $g:=\chi_{V}\hat{f}$.
+Then $g,g_{\epsilon}\in C_{0}^{1/2}(V)$ and $\|g_{\epsilon}\|_{C^{1/2}(V)}<c$, $0<\epsilon\ll 1$.
+Also,
+(4.13) $\begin{split}&g(y)=a_{0}(y)(p-P(\alpha))_{+}^{1/2}+\Delta g(y),\\ &a(y)=2\pi^{-1/2}\tilde{\upsilon}_{0}(y),\ a\in C_{0}^{\infty}(V),\ \Delta g\in C_{0}^{3/2}(\mathcal{V}),\end{split}$
+and $\|\Delta g_{\epsilon}\|_{C^{3/2}(V)}<c$ for all $0<\epsilon\ll 1$, where $\tilde{\upsilon}_{0}(y)$ is given in (4.10). Here
+(4.14) $g_{\epsilon}(y):=\textstyle\sum_{j\in V}\varphi_{\epsilon}(y-y_{j})g(y_{j}),$
+and $\Delta g_{\epsilon}$ is defined analogously, see (3.14).
+If $\Gamma\cap V=\varnothing$, we assume in (4.13) that $a_{0}\equiv 0$.
 5. Regularized solution
@@
 Let $\mathcal{R}^{*}$ be the adjoint of $\mathcal{R}:L^{2}(\mathcal{U})\to L^{2}(\mathcal{V})$:
-(5.1) $\begin{split}(\mathcal{R}^{*}g)(x)=&\int_{\mathcal{V}}g(y)W(x,y)\delta(p-\Phi(x,\alpha))\text{d}y,\ x\in\mathcal{U}.\end{split}$
-As is well-known, assumption 3.2 ensures that $\mathcal{R}^{*}\mathcal{R}$ is an elliptic pseudo-differential operator $(\Psi$DO) [46, 19], [53, Section VIII.6.2] of order one in $\mathcal{U}$, i.e.
-$\mathcal{R}^{*}\mathcal{R}\in L^{1}(\mathcal{U})$.
+(5.1) $\begin{split}(\mathcal{R}^{*}g)(x)=&\int_{\mathcal{V}}g(y)W(x,y)\delta(p-\Phi(x,\alpha))\text{d}y\\ =&\int_{\mathcal{I}_{\alpha}}g(\alpha,\Phi(x,\alpha))W(x,(\alpha,\Phi(x,\alpha)))\text{d}\alpha,\ x\in\mathcal{U}.\end{split}$
+As is well-known, Assumption 3.2 ensures that $\mathcal{R}^{*}\mathcal{R}$ is an elliptic pseudo-differential operator $(\Psi$DO) [49, 22], [56, Section VIII.6.2] of order -1 in $\mathcal{U}$, i.e.
+$\mathcal{R}^{*}\mathcal{R}\in L^{-1}(\mathcal{U})$.
 Clearly, the functional $\Psi$ in (3.17) is Fréchet-differentiable. Indeed, one has
-(5.2) $\begin{split}&\Psi(f+h)-\Psi(f)=\Psi^{\prime}(f;h)+O(\|h\|_{H^{1}(\mathcal{U})}^{2}),\ f,h\in H_{c}^{1}(\mathcal{U}_{b}),\\ &(1/2)\Psi^{\prime}(f;h)=(\mathcal{R}f-g_{\epsilon},\mathcal{R}h)_{L^{2}(\mathcal{V})}+\kappa\epsilon^{3}(\nabla f,\nabla h)_{L^{2}(\mathcal{U})},\end{split}$
-and the functional $H_{c}^{1}(\mathcal{U}_{b})\ni h\to\Psi^{\prime}(f;h)$ is linear and continuous [41, p.
+(5.2) $\begin{split}&\Psi(f+h)-\Psi(f)=\Psi^{\prime}(f;h)+O(\|h\|_{H^{1}(\mathcal{U})}^{2}),\ f,h\in H_{0}^{1}(\mathcal{U}_{b}),\\ &(1/2)\Psi^{\prime}(f;h)=(\mathcal{R}f-\hat{f}_{\epsilon},\mathcal{R}h)_{L^{2}(\mathcal{V})}+\kappa\epsilon^{3}(\partial_{x}f,\partial_{x}h)_{L^{2}(\mathcal{U})},\end{split}$
+and the functional $H_{0}^{1}(\mathcal{U}_{b})\ni h\to\Psi^{\prime}(f;h)$ is linear and continuous [45, p.
 14].
-Recall that $g_{\epsilon}$ is defined in (3.13).
-Then the subgradient of $\Psi$ coincides with the gradient [41, Proposition 3.20], and $f$ is the minimizer if and only if the gradient equals zero [41, Theorem 3.24].
-The facts that $\Psi$ is proper and convex, which are required for the latter conclusion, are established in the proof of Lemma 3.5 in appendix A.
+Recall that $\hat{f}_{\epsilon}$ is defined in (3.14).
+Then the subgradient of $\Psi$ coincides with the gradient [45, Proposition 3.20], and $f$ is the minimizer if and only if the gradient equals zero [45, Theorem 3.24].
+The fact that $\Psi$ is proper and convex, which is required for the latter conclusion, is established in the proof of Lemma 3.5 in Appendix A.
 Therefore the solution to (3.17) is the unique solution to the following equation, which is equivalent to the first order optimality condition:
-(5.3) $(\mathcal{R}^{*}\mathcal{R}-\kappa\epsilon^{3}\Delta)f_{\epsilon}=\mathcal{R}^{*}g_{\epsilon},\ x\in\mathcal{U}_{b},\ f_{\epsilon}\in H_{c}^{1}(\mathcal{U}_{b}).$
-This follows by setting $\Psi^{\prime}(f_{\epsilon};h)\equiv 0$, $h\in C_{0}^{\infty}(\mathcal{U}_{b})$, integrating by parts, and using that $C_{0}^{\infty}(\mathcal{U}_{b})$ is dense in $H_{c}^{1}(\mathcal{U}_{b})$.
+(5.3) $(\mathcal{R}^{*}\mathcal{R}-\kappa\epsilon^{3}\Delta)f_{\epsilon}=\mathcal{R}^{*}\hat{f}_{\epsilon},\ x\in\mathcal{U}_{b},\ f_{\epsilon}\in H_{0}^{1}(\mathcal{U}_{b}).$
+This follows by setting $\Psi^{\prime}(f_{\epsilon};h)\equiv 0$, $h\in C_{0}^{\infty}(\mathcal{U}_{b})$, integrating by parts, and using that $C_{0}^{\infty}(\mathcal{U}_{b})$ is dense in $H_{0}^{1}(\mathcal{U}_{b})$.
 Lemma 5.1.
@@
 Proof.
-Since $\Psi(f\equiv 0)=\|g_{\epsilon}\|_{L^{2}(\mathcal{V})}^{2}$ (cf.
-(3.16)), the solution $f_{\epsilon}$ satisfies $\|\mathcal{R}f_{\epsilon}\|_{L^{2}(\mathcal{V})}\leq 2\|g_{\epsilon}\|_{L^{2}(\mathcal{V})}<c$, $0<\epsilon\ll 1$ (cf.
-the paragraph above (3.16)).
-By construction, $f_{\epsilon}\in H_{c}^{1}(\mathcal{U}_{b})\subset H_{c}^{-1/2}(\mathcal{U}_{b})$, where $\mathcal{U}_{b}$ is bounded.
+Since $\Psi(f\equiv 0)=\|\hat{f}_{\epsilon}\|_{L^{2}(\mathcal{V})}^{2}$ (cf.
+(3.17)), the solution $f_{\epsilon}$ satisfies $\|\mathcal{R}f_{\epsilon}\|_{L^{2}(\mathcal{V})}\leq 2\|\hat{f}_{\epsilon}\|_{L^{2}(\mathcal{V})}<c$, $0<\epsilon\ll 1$ (cf.
+the paragraph preceding (3.17)).
+For clarity, $\mathcal{R}f_{\epsilon}$ is the GRT of the solution $f_{\epsilon}$, while $\hat{f}_{\epsilon}$ is the interpolated version of the exact discrete data, i.e.
+$\hat{f}_{\epsilon}=(\mathcal{R}f)_{\epsilon}$.
+By construction, $f_{\epsilon}\in H_{0}^{1}(\mathcal{U}_{b})\subset H_{0}^{-1/2}(\mathcal{U}_{b})$, where $\mathcal{U}_{b}$ is bounded.
 Application of (3.6) completes the proof.
 ∎
-Let $\tilde{D}$ be the complete symbol of $\mathcal{R}^{*}\mathcal{R}$, i.e. $\mathcal{R}^{*}\mathcal{R}=\text{Op}(\tilde{D}(x,\xi))$ (see [16, Theorem 3.4]). Rewrite (5.3) in an alternative form:
-(5.5) $\begin{split}\big{(}\text{Op}(\tilde{D}(x,\xi)+\kappa\epsilon^{3}|\xi|^{2})f_{\epsilon}\big{)}(x)=(\mathcal{R}^{*}g_{\epsilon})(x),\ x\in\mathcal{U}_{b}.\end{split}$
+Let $\tilde{D}$ be the complete symbol of $\mathcal{R}^{*}\mathcal{R}$, i.e. $\mathcal{R}^{*}\mathcal{R}=\text{Op}(\tilde{D}(x,\xi))$ (see [19, Theorem 3.4]). Rewrite (5.3) in an alternative form:
+(5.5) $\begin{split}\big{(}\text{Op}(\tilde{D}(x,\xi)+\kappa\epsilon^{3}|\xi|^{2})f_{\epsilon}\big{)}(x)=(\mathcal{R}^{*}\hat{f}_{\epsilon})(x),\ x\in\mathcal{U}_{b}.\end{split}$
 Let $\mathcal{U}_{0}$ be a sufficiently small neighborhood of $x_{0}\in\mathcal{U}_{b}$.
@@
 Apply $\text{Op}(\chi(x)|\xi|)\chi(x)$ on both sides of (5.5):
-(5.6) $\begin{split}&\text{Op}(\chi(x)|\xi|)\chi(x)\text{Op}\big{(}\tilde{D}(x,\xi)+\kappa\epsilon^{3}|\xi|^{2}\big{)}f_{\epsilon}=\text{Op}(\chi(x)|\xi|)\chi(x)\mathcal{R}^{*}g_{\epsilon}=:F_{\epsilon}.\end{split}$
+(5.6) $\begin{split}&\text{Op}(\chi(x)|\xi|)\chi(x)\text{Op}\big{(}\tilde{D}(x,\xi)+\kappa\epsilon^{3}|\xi|^{2}\big{)}f_{\epsilon}=\text{Op}(\chi(x)|\xi|)\chi(x)\mathcal{R}^{*}\hat{f}_{\epsilon}=:F_{\epsilon}.\end{split}$
 Let $\tilde{D}_{0}(x,\xi):=\lim_{\lambda\to\infty}|\lambda\xi|\tilde{D}(x,\lambda\xi)$.
@@
 Suppose Assumptions 3.13.4 are satisfied. Define
-(5.7) $G_{\epsilon}:=\text{Op}\big{(}\big{[}\tilde{D}_{0}(x,\xi)+\kappa|\epsilon\xi|^{3}\big{]}^{-1}\big{)}F_{\epsilon}.$
+(5.7) $\mathcal{G}_{\epsilon}:=\text{Op}\big{(}\big{[}\tilde{D}_{0}(x,\xi)+\kappa|\epsilon\xi|^{3}\big{]}^{-1}\big{)}F_{\epsilon}.$
 Fix any $A>0$. One has:
@@
 Suppose Assumptions 3.13.4 are satisfied. Then $\|F_{\epsilon}\|_{L^{\infty}(\mathcal{U})}\leq c$ for all $0<\epsilon\ll 1$.
+By the linearity of (5.6), to prove the lemma it suffices to consider $F_{\epsilon}$ as defined on the right in (5.6), but with $\hat{f}_{\epsilon}$ replaced by $g_{\epsilon}$ using a local piece of the data $g=\chi_{V}\mathcal{R}f$ introduced in Lemma 4.2.
 6.1. Preliminary calculations
 Our goal is to show that the right-hand side of (5.6) is bounded. Rewrite (5.6) as follows.
-(6.1) $\begin{split}F_{\epsilon}(x)&=(\mathcal{T}_{1}g_{\epsilon})(x),\ \mathcal{T}_{1}:=\text{Op}(\chi(x)|\xi|)\chi(z)\mathcal{R}^{*},\\ (\mathcal{T}_{1}g_{\epsilon})(x)&=\frac{\chi(x)}{(2\pi)^{3}}\int_{\mathbb{R}^{2}}\int_{\mathcal{U}}|\xi|e^{i\xi\cdot(z-x)}\chi(z)\\ &\hskip 28.45274pt\times\frac{1}{2\pi}\int_{\mathbb{R}}\int_{\mathcal{V}}W(z,y)e^{i\nu(p-\Phi(z,\alpha))}g_{\epsilon}(y)\text{d}y\text{d}\nu\text{d}z\text{d}\xi.\end{split}$
-Clearly, $\mathcal{T}_{1}\in I^{1/2}(\mathcal{U}\times\mathcal{V}\times\mathcal{C}^{*})$.
-Here the canonical relation $\mathcal{C}^{*}$ is the image of $\mathcal{C}$ (see (3.3)) under the exchange of $T^{*}\mathcal{U}$ and $T^{*}\mathcal{V}$.
-Next, consider the expression
-(6.2) $\begin{split}\frac{\chi(x)}{(2\pi)^{3}}\int_{\mathbb{R}^{2}}\int_{\mathcal{U}}\chi(z)W(z,y)|\xi|e^{i\xi\cdot(z-x)-i\nu(\Phi(z,\alpha)-\Phi(x,\alpha))}\text{d}z\text{d}\xi.\end{split}$
-Upon changing variables $\xi\to\xi_{1}=\xi/\nu$, the stationary point of the phase satisfies $z=x$, $\xi_{1}=\Phi_{z}^{\prime}(z,\alpha)$, and the stationary point is nondegenerate.
-By the stationary phase method (see also [53, Ch.
-VI, eqs.
-(4.1)–(4.4)]),
-(6.3) $\mathcal{T}_{1}=\mathcal{T}_{11}+\mathcal{T}_{12},\ \mathcal{T}_{11}\in I^{1/2}(\mathcal{U}\times\mathcal{V}\times\mathcal{C}^{*}),\ \mathcal{T}_{12}\in I^{-1/2}(\mathcal{U}\times\mathcal{V}\times\mathcal{C}^{*}),$
-where
-(6.4) $\begin{split}&(\mathcal{T}_{11}g_{\epsilon})(x):=\frac{1}{2\pi}\int_{\mathbb{R}}|\nu|\int_{\mathcal{V}}K(x,\alpha)e^{i\nu(p-\Phi(x,\alpha))}g_{\epsilon}(y)\text{d}y\text{d}\nu,\ x\in\mathcal{U}_{0},\\ &K(x,\alpha):=\chi^{2}(x)|\Phi_{x}^{\prime}(x,\alpha)|W(x,(\Phi(x,\alpha),\alpha)),\end{split}$
-and $\mathcal{T}_{12}:=\mathcal{T}_{1}-\mathcal{T}_{11}$.
-The inclusion $\mathcal{T}_{12}\in I^{-1/2}(\mathcal{U}\times\mathcal{V}\times\mathcal{C}^{*})$ follows by considering lower order terms in the stationary phase expansion as $\nu\to\infty$ of the integral in (6.2) [53, Chapter VIII, eqs.
-(2.14), (2.16)].
-By the continuity of FIOs in Hölder-Zygmund spaces (see [23, Proposition 5.5] and (C.12), (C.13)), $\|\mathcal{T}_{12}g_{\epsilon}\|_{C^{1/2}(\mathcal{U})}<c$ for all $0<\epsilon\ll 1$.
-By Lemma 4.2, $\Delta g,\Delta g_{\epsilon}\in C_{0}^{3/2}(\mathcal{V})$ and $\|\Delta g_{\epsilon}\|_{C^{3/2}(\mathcal{V})}<c$, $0<\epsilon\ll 1$.
-In this case $\|\mathcal{T}_{1}\Delta g_{\epsilon}\|_{C^{1/2}(\mathcal{U})}<c$, $0<\epsilon\ll 1$.
+(6.1) $\begin{split}F_{\epsilon}(x)&=(\mathcal{T}g_{\epsilon})(x),\ \mathcal{T}:=\text{Op}(\chi(x)|\xi|)\chi(z)\mathcal{R}^{*},\\ (\mathcal{T}g_{\epsilon})(x)&=\frac{\chi(x)}{(2\pi)^{3}}\int_{\mathbb{R}^{2}}\int_{\mathcal{U}}|\xi|e^{i\xi\cdot(z-x)}\chi(z)\\ &\hskip 28.45274pt\times\frac{1}{2\pi}\int_{\mathbb{R}}\int_{V}W(z,y)e^{i\nu(p-\Phi(z,\alpha))}g_{\epsilon}(y)\text{d}y\text{d}\nu\text{d}z\text{d}\xi.\end{split}$
+Next, consider the integral
+(6.2) $\begin{split}K_{1}(x,\alpha,\nu):=\frac{\chi(x)}{(2\pi)^{3}}\int_{\mathbb{R}^{2}}\int_{\mathcal{U}}\chi(z)W(z,y)|\xi|e^{i\xi\cdot(z-x)-i\nu(\Phi(z,\alpha)-\Phi(x,\alpha))}\text{d}z\text{d}\xi.\end{split}$
+Upon changing variables $\xi\to\xi_{1}=\xi/\nu$, the stationary point of the phase satisfies $z=x$, $\xi_{1}=\text{d}_{x}\Phi(x,\alpha)$, and the stationary point is nondegenerate.
+By the stationary phase method,
+(6.3) $\begin{split}&K_{1}(x,\alpha,\nu)=|\nu|K(x,\alpha)+\Delta K(x,\alpha;\nu),\\ &K(x,\alpha):=\chi^{2}(x)|\text{d}_{x}\Phi(x,\alpha)|W(x,(\alpha,\Phi(x,\alpha))),\ x\in\mathcal{U},\\ &K\in C_{0}^{\infty}(\mathcal{U}_{b}\times\mathcal{I}_{\alpha}),\ \Delta K\in C_{0}^{\infty}(\mathcal{U}_{b}\times\mathcal{I}_{\alpha}\times(\mathbb{R}\setminus 0)),\ \Delta K\in S^{0}(\mathbb{R}).\end{split}$
+By (6.1)–(6.3),
+(6.4) $\begin{split}(\mathcal{T}g_{\epsilon})(x)=&\int_{I_{\alpha}}K(x,\alpha)(\Lambda g_{\epsilon})(\alpha,\Phi(x,\alpha))\text{d}\alpha\\ &+\int_{I_{\alpha}}(\Delta\mathcal{K}_{x,\alpha}g_{\epsilon})(\alpha,\Phi(x,\alpha))\text{d}\alpha,\ x\in\mathcal{U}.\end{split}$
+Here $\Lambda\in L^{1}(\mathcal{V})$, and $\mathcal{K}_{x,\alpha}\in L^{1}(\mathcal{V})$ is a $\Psi$DO, which depends smoothly on $(x,\alpha)\in\mathcal{U}_{b}\times\mathcal{I}_{\alpha}$,
+(6.5) $\begin{split}(\Lambda g)(\alpha,t)=&\frac{1}{2\pi}\int_{\mathcal{I}_{p}}|\nu|e^{i\nu(p-t)}g(\alpha,p)\text{d}p\text{d}\nu,\\ (\mathcal{K}_{x,\alpha^{\prime}}g)(\alpha,t)=&\frac{1}{2\pi}\int_{\mathcal{I}_{p}}\Delta K(x,\alpha^{\prime};\nu)e^{i\nu(p-t)}g(\alpha,p)\text{d}p\text{d}\nu,\ (\alpha,t)\in\mathcal{V}.\end{split}$
+By Lemma 4.2,
+[…]
+Recall that $B_{\infty,\infty}^{s}(\mathbb{R}^{n})=C^{s}(\mathbb{R}^{n})$, $s>0$, $s\not\in\mathbb{N}$ [1, Remark 6.4(2)].
+By the continuity of $\Psi$DOs in Hölder-Zygmund spaces (see [1, Theorem 6.19]), $\|\Lambda\Delta g_{\epsilon}\|_{C^{1/2}(V_{1})}<c$, $\|\Delta\mathcal{K}_{x,\alpha^{\prime}}g_{\epsilon}\|_{C^{1/2}(V_{1})}<c$, $0<\epsilon\ll 1$, for any compact $V_{1}\subset\mathcal{V}$.
+It is now obvious that
+(6.7) $\begin{split}\int_{I_{\alpha}}(\Lambda\Delta g_{\epsilon})(\alpha,\Phi(x,\alpha))\text{d}\alpha+\int_{I_{\alpha}}(\Delta\mathcal{K}_{x,\alpha}g_{\epsilon})(\alpha,\Phi(x,\alpha))\text{d}\alpha\in C_{0}^{1/2}(\mathcal{U}),\end{split}$
+with its $C^{1/2}(\mathcal{U})$ norm uniformly bounded for $0<\epsilon\ll 1$.
 Therefore, all that remains is to bound the following expression:
-(6.5) $\begin{split}F_{\epsilon}^{(l)}(x):=-\frac{1}{\pi}\int_{\mathcal{V}}&\frac{\chi_{1\text{D}}(p-\Phi(x,\alpha))K(x,\alpha)}{p-\Phi(x,\alpha)}\\ &\times\sum_{j}\partial_{p}\varphi_{\epsilon}(y-y_{j})a_{0}(y_{j})(p_{j_{2}}-P(\alpha_{j_{1}}))_{+}^{1/2}\text{d}y.\end{split}$
-Here and in the rest of this section, the pair $a_{0}(y),P(\alpha)$ stands for any one of the pairs $a_{l},P_{l}$ described in Lemma 4.2.
-This means that the domain of integration $\mathcal{V}$ in (6.5) can be replaced by a small neighborhood of $\text{supp}\,a_{l}$.
-This increase in support is due to interpolation.
-To get (6.5) from (6.4) we expressed $\text{Op}(|\nu|)=-1/(\pi t^{2})$, integrated by parts with respect to $p$, and inserted a 1D cutoff function $\chi_{1\text{D}}$.
+(6.8) $\begin{split}F_{\epsilon}^{(l)}(x):=&-\frac{1}{\pi}\int_{V}\frac{\chi_{1\text{D}}(p-\Phi(x,\alpha))K(x,\alpha)}{p-\Phi(x,\alpha)}\\ &\times\sum_{y_{j}\in V}\partial_{p}\varphi_{\epsilon}(y-y_{j})a_{0}(y_{j})(p_{j_{2}}-P(\alpha_{j_{1}}))_{+}^{1/2}\text{d}y,\ x\in\mathcal{U}_{b}.\end{split}$
+Recall that $K(x,\alpha)\equiv 0$ for any $x\in\mathcal{U}\setminus\mathcal{U}_{b}$, cf.
+(6.3).
+To get (6.8) from (6.4), (6.5), we used the identity $\text{Op}(|\nu|)=-1/(\pi t^{2})$, integrated by parts with respect to $p$, and inserted a 1D cutoff function $\chi_{1\text{D}}$.
 The latter identically equals one in a neighborhood of zero.
@@
 […]
-for any fixed $A>0$.
+for any fixed $A>0$. The $O(\epsilon^{1/2})$ term is uniform in $\check{x}$.
 6.2. Simplification of $F_{\epsilon}^{(l)}$, the leading order term of $F_{\epsilon}$
 In what follows we use the function
-(6.7) $\begin{split}&Q(x,\alpha):=\Phi(x,\alpha)-P(\alpha).\end{split}$
-Since we want to prove that $F_{\epsilon}^{(l)}(x)$ is uniformly bounded for all $x\in\mathcal{U}$, we drop $x$ from notation (e.g., $K(x,\alpha)$ of (6.4) becomes $K(\alpha)$) and make sure that all the estimates are uniform with respect to $x$.
-It suffices to assume that $x$ is confined to a sufficiently small open subset $U^{\prime}\subset\mathcal{U}$.
-Change variables $p=w+\Phi(\alpha)$ in (6.5):
-(6.8) $\begin{split}F_{\epsilon}^{(l)}=-\frac{1}{\pi}\sum_{j_{1}}\int_{\mathbb{R}}&\frac{\chi_{1\text{D}}(w)}{w}\sum_{j_{2}}\bigg{[}\partial_{w}\int_{\mathbb{R}}K(\alpha)\varphi_{\epsilon}(y-y_{j})\text{d}\alpha\bigg{]}\\ &\times a_{0}(y_{j})(p_{j_{2}}-P(\alpha_{j_{1}}))_{+}^{1/2}\text{d}w,\ y=(\alpha,w+\Phi(\alpha)).\end{split}$
-The sum with respect to $j_{1}$ is over $\alpha_{j_{1}}$ in the domain of the locally defined function $P(\alpha)$. Consider the integral with respect to $\alpha$:
-(6.9) $\begin{split}\int_{\mathbb{R}}&K(\alpha)\varphi_{\alpha}\bigg{(}\frac{\alpha-\alpha_{j_{1}}}{\mu\epsilon}\bigg{)}\partial_{w}\varphi_{p}\bigg{(}\frac{w+\Phi(\alpha)-p_{j_{2}}}{\epsilon}\bigg{)}\text{d}\alpha\\ =&\epsilon^{-1}\int_{\mathbb{R}}K(\alpha)\varphi_{\alpha}\bigg{(}\frac{\alpha-\alpha_{j_{1}}}{\mu\epsilon}\bigg{)}\varphi_{p}^{\prime}\bigg{(}\frac{\Phi(\alpha)-\Phi(\alpha_{j_{1}})}{\epsilon}+u\bigg{)}\text{d}\alpha,\\ u:=&(w+\Phi(\alpha_{j_{1}})-p_{j_{2}})/\epsilon.\end{split}$
+(6.10) $\begin{split}&Q(x,\alpha):=\Phi(x,\alpha)-P(\alpha)=\Phi(x,\alpha)-\Phi(x(\alpha),\alpha).\end{split}$
+To prove that $F_{\epsilon}^{(l)}(x)$ is uniformly bounded for all $x\in\mathcal{U}_{b}$, it suffices to assume that $x$ is confined to a sufficiently small open subset $U^{\prime}\subset\mathcal{U}_{b}$.
+Change variable $p=w+\Phi(\alpha)$ in (6.8):
+(6.11) $\begin{split}F_{\epsilon}^{(l)}(x)=-\frac{1}{\pi}\sum_{j_{1}\in J_{\alpha}}\int_{\mathbb{R}}&\frac{\chi_{1\text{D}}(w)}{w}\sum_{j_{2}\in J_{p}}\bigg{[}\partial_{w}\int_{\mathcal{I}_{\alpha}}K(x,\alpha)\varphi_{\epsilon}(y-y_{j})\text{d}\alpha\bigg{]}\\ &\times a_{0}(y_{j})(p_{j_{2}}-P(\alpha_{j_{1}}))_{+}^{1/2}\text{d}w,\ y=(\alpha,w+\Phi(\alpha)).\end{split}$
+Consider the integral with respect to $\alpha$:
+(6.12) $\begin{split}\int_{\mathcal{I}_{\alpha}}&K(x,\alpha)\varphi_{\alpha}\bigg{(}\frac{\alpha-\alpha_{j_{1}}}{\mu\epsilon}\bigg{)}\partial_{w}\varphi_{p}\bigg{(}\frac{w+\Phi(\alpha)-p_{j_{2}}}{\epsilon}\bigg{)}\text{d}\alpha\\ =&\epsilon^{-1}\int_{\mathcal{I}_{\alpha}}K(x,\alpha)\varphi_{\alpha}\bigg{(}\frac{\alpha-\alpha_{j_{1}}}{\mu\epsilon}\bigg{)}\varphi_{p}^{\prime}\bigg{(}\frac{\Phi(\alpha)-\Phi(\alpha_{j_{1}})}{\epsilon}+u\bigg{)}\text{d}\alpha,\\ u:=&(w+\Phi(\alpha_{j_{1}})-p_{j_{2}})/\epsilon.\end{split}$
 We have used that $\varphi$ is the product of two kernels, see (3.14). Replace $\alpha_{j_{1}}$ with $\theta$, change variable $\alpha=\theta+\epsilon\mu s$, and define
-(6.10) $\begin{split}h(u;\theta,\epsilon):=&\mu\int_{\mathbb{R}}K(\theta+\epsilon\mu s)\varphi_{\alpha}(s)\varphi_{p}\bigg{(}\frac{\Phi(\theta+\epsilon\mu s)-\Phi(\theta)}{\epsilon}+u\bigg{)}\text{d}s.\end{split}$
+(6.13) $\begin{split}h(u;*):=&\mu\int_{\mathbb{R}}K(x,\theta+\epsilon\mu s)\varphi_{\alpha}(s)\varphi_{p}\bigg{(}\frac{\Phi(\theta+\epsilon\mu s)-\Phi(\theta)}{\epsilon}+u\bigg{)}\text{d}s,\\ *=&(x,\theta,\epsilon),\ u\in\mathbb{R},\ x\in U^{\prime},\ \theta\in I_{\alpha},\ 0<\epsilon\ll 1.\end{split}$
 The sum with respect to $j_{2}$ in (6.11) simplifies by the introduction of the following function
-(6.11) $H(\check{t},\check{q};\theta,\epsilon):=\sum_{j_{2}}h^{\prime}(\check{t}+\check{q}-j_{2};\theta,\epsilon)a_{0}(\theta,\epsilon j_{2})(j_{2}-\check{q})_{+}^{1/2}.$
+(6.14) $H(\check{t},\check{q};*):=\sum_{j_{2}\in J_{p}}h^{\prime}(\check{t}+\check{q}-j_{2};*)a_{0}(\theta,\epsilon j_{2})(j_{2}-\check{q})_{+}^{1/2}.$
 The prime in $h^{\prime}$ denotes the derivative with respect to the first argument.
 To get the sum in (6.11) we observe that $\epsilon$ is factored out from $p_{j_{2}}-P(\alpha_{j_{1}})$ and use (6.14) with
-(6.12) $\check{t}=(w+Q(\alpha_{j_{1}}))/\epsilon,\ \check{q}=P(\alpha_{j_{1}})/\epsilon,\ \theta=\alpha_{j_{1}}.$
+(6.15) $\check{t}=(w+Q(x,\alpha_{j_{1}}))/\epsilon,\ \check{q}=P(\alpha_{j_{1}})/\epsilon,\ \theta=\alpha_{j_{1}},\ j_{1}\in J_{\alpha}.$
 Some properties of the function $H$ are obtained in Lemma D.1 in Appendix D.1.
 Next we consider the integral with respect to $w$ in (6.11):
-(6.13) $\begin{split}&\int_{\mathbb{R}}\frac{\chi_{1\text{D}}(w)}{w}H\bigg{(}\frac{w+Q(\theta)}{\epsilon},\frac{P(\theta)}{\epsilon};\theta,\epsilon\bigg{)}\text{d}w\\ &=\int_{\mathbb{R}}\frac{\chi_{1\text{D}}(\epsilon\check{w})}{\check{w}}H\bigg{(}\check{w}+\frac{Q(\theta)}{\epsilon},\frac{P(\theta)}{\epsilon};\theta,\epsilon\bigg{)}\text{d}\check{w},\end{split}$
+(6.16) $\begin{split}&\int_{\mathbb{R}}\frac{\chi_{1\text{D}}(w)}{w}H\bigg{(}\frac{w+Q(x,\theta)}{\epsilon},\frac{P(\theta)}{\epsilon};*\bigg{)}\text{d}w\\ &=\int_{\mathbb{R}}\frac{\chi_{1\text{D}}(\epsilon\check{w})}{\check{w}}H\bigg{(}\check{w}+\frac{Q(x,\theta)}{\epsilon},\frac{P(\theta)}{\epsilon};*\bigg{)}\text{d}\check{w},\end{split}$
 where $\theta=\alpha_{j_{1}}$. Define a key intermediate function
-(6.14) $\begin{split}G(\check{t},\check{q};\theta,\epsilon):=&\int_{\mathbb{R}}\chi_{1\text{D}}(\epsilon\check{w})\frac{H(\check{w}+\check{t},\check{q};\theta,\epsilon)}{\check{w}}\text{d}\check{w}.\end{split}$
+(6.17) $\begin{split}G(\check{t},\check{q};*):=&\int_{\mathbb{R}}\chi_{1\text{D}}(\epsilon\check{w})\frac{H(\check{w}+\check{t},\check{q};*)}{\check{w}}\text{d}\check{w}.\end{split}$
 In terms of $G$, (6.11) becomes
-(6.15) $\begin{split}F_{\epsilon}^{(l)}=-\frac{\epsilon^{1/2}}{\pi}\sum_{j_{1}}G\bigg{(}\frac{Q(\alpha_{j_{1}})}{\epsilon},\frac{P(\alpha_{j_{1}})}{\epsilon},\alpha_{j_{1}},\epsilon\bigg{)}.\end{split}$
+(6.18) $\begin{split}F_{\epsilon}^{(l)}(x)=-\frac{\epsilon^{1/2}}{\pi}\sum_{j_{1}\in J_{\alpha}}G\bigg{(}\frac{Q(x,\alpha_{j_{1}})}{\epsilon},\frac{P(\alpha_{j_{1}})}{\epsilon};x,\alpha_{j_{1}},\epsilon\bigg{)}.\end{split}$
 The extra factor $\epsilon^{1/2}$ appears because it has been factored out from the term $(p_{j_{2}}-P(\alpha_{j_{1}}))_{+}^{1/2}$ in (6.11) (cf. (6.14)).
-Let $\text{dom}\,P$ denote the domain of the selected $P(\alpha)$. The following lemma is proven in appendix D.1.
+The following lemma is proven in Appendix D.1.
 Lemma 6.2.
 Suppose Assumptions 3.13.4 are satisfied. Pick any $A>0$. One has
-(6.16) $\begin{split}G(\check{t},\check{q};\theta,\epsilon)=&\begin{cases}O(\epsilon^{1/2})+O(\check{t}^{-1}),&\check{t}\to\infty,\\ O((-\check{t})^{-1/2}),&\check{t}\to-\infty.\end{cases}\end{split}$
-In the above formula the big-$O$ term is uniform in $\check{q}\in\mathbb{R}$, $\theta\in\text{dom}\,P$, $|\check{t}|\leq A/\epsilon$, and $0<\epsilon\ll 1$.
+(6.19) $\begin{split}|G(\check{t},\check{q};*)|\leq&c\begin{cases}\epsilon^{1/2}+(1+\check{t})^{-1},&\check{t}>0,\\ (1-\check{t})^{-1/2},&\check{t}<0,\end{cases}\end{split}$
+for some $c$ independent of $|\check{t}|\leq A/\epsilon$, $|\check{q}|\leq A/\epsilon$, $x\in U^{\prime}$, $\theta\in I_{\alpha}$, and $0<\epsilon\ll 1$.
 We use this lemma in the proof of the following result, see Appendix D.3.
@@
 Suppose Assumptions 3.13.4 are satisfied. Then $|F_{\epsilon}^{(l)}(x)|\leq c$ for all $x\in\mathcal{U}_{b}$ and $0<\epsilon\ll 1$.
-Combining Lemma 6.3 with (6.6) proves Lemma 6.1.
+The number of segments $\Gamma_{l}$ described above Lemma 4.2 is uniformly bounded, so combining Lemma 6.3 with (6.9) proves Lemma 6.1.
 7. Computing the leading term $\Delta F_{0}$
@@
 The reconstruction point is then $x=\epsilon\check{x}$.
-We partially return the dependence on $x$ in the notation by keeping only the dependence on $\check{x}$.
-If an argument of a function is $x_{0}=0_{2}$, then it is omitted from notation.
-Fix any $A_{0}>0$.
-Throughout this section we assume $|\check{x}|\leq A_{0}$.
+Fix any $A_{1}>0$.
+Throughout this section we assume $|\check{x}|<A_{1}$, including in Lemmas 7.2, 7.3.
+From the proof of Lemma 6.2 and (6.13), (6.14), (6.17) it follows that
+(7.1) $G(\check{t},\check{q};\epsilon\check{x},\theta,\epsilon)=G(\check{t},\check{q};0_{2},\theta,\epsilon)+O(\epsilon).$
+By (6.18),
+(7.2) $\begin{split}F_{\epsilon}^{(l)}(\epsilon\check{x})=-\frac{\epsilon^{1/2}}{\pi}\sum_{j_{1}\in J_{\alpha}}G\bigg{(}\frac{Q(x,\alpha_{j_{1}})}{\epsilon},\frac{P(\alpha_{j_{1}})}{\epsilon};0_{2},\alpha_{j_{1}},\epsilon\bigg{)}+O(\epsilon^{1/2}).\end{split}$
 Define
 (7.3) $\Delta F_{0}(\check{x}):=\lim_{\epsilon\to 0}\big{[}F_{\epsilon}(\epsilon\check{x})-F_{\epsilon}(0_{2})\big{]}.$
-By the second statement in (6.6),
+By the second line in (6.9),
 (7.4) $\Delta F_{0}(\check{x})=\lim_{\epsilon\to 0}\Delta F_{\epsilon}^{(l)}(\epsilon\check{x}),\ \Delta F_{\epsilon}^{(l)}(\epsilon\check{x}):=F_{\epsilon}^{(l)}(\epsilon\check{x})-F_{\epsilon}^{(l)}(0_{2}).$
+From (7.2) it follows that when computing $\Delta F_{0}(\check{x})$, we can replace $\epsilon\check{x}$ with $0_{2}$ in the list of arguments of $G$ when representing $F_{\epsilon}^{(l)}(\epsilon\check{x})$ in terms of $G$ in (7.4).
+Therefore, we can drop $x$ from the list of arguments of $h$, $H$, and $G$ and change the meaning of $*=(x,\theta,\epsilon)$ to $*=(\theta,\epsilon)$.
+Also, if the first argument of $Q$ is $x=0_{2}$, then it is omitted from notation.
 Using (6.14), define another intermediate function
-(7.3) $\begin{split}\Delta G(r,\check{t},\check{q};\theta,\epsilon):=&\int_{\mathbb{R}}\chi_{1\text{D}}(\epsilon\check{w})\frac{\Delta H(r,\check{w}+\check{t},\check{q};\theta,\epsilon)}{\check{w}}\text{d}\check{w},\\ \Delta H(r,\check{t},\check{q};\cdot):=&H(r+\check{t},\check{q};\cdot)-H(\check{t},\check{q};\cdot).\end{split}$
-We define $\Delta G$ not via $G$ in (6.14) but in terms of $\Delta H$ because this allows for more accurate bounds. By (6.15), the expression for $\Delta F_{\epsilon}^{(l)}$ becomes:
-(7.4) $\begin{split}&\Delta F_{\epsilon}^{(l)}(\epsilon\check{x})=-\frac{\epsilon^{1/2}}{\pi}\sum_{j_{1}}\Delta G\bigg{(}\frac{Q(\epsilon\check{x},\alpha_{j_{1}})-Q(\alpha_{j_{1}})}{\epsilon},\frac{Q(\alpha_{j_{1}})}{\epsilon},\frac{P(\alpha_{j_{1}})}{\epsilon};\alpha_{j_{1}},\epsilon\bigg{)}.\end{split}$
-As before, the sum is over $\alpha_{j_{1}}\in\text{dom}\,P$.
+(7.5) $\begin{split}\Delta G(r,\check{t},\check{q};*):=&G(r+\check{t},\check{q};*)-G(\check{t},\check{q};*)\\ =&\int_{\mathbb{R}}\chi_{1\text{D}}(\epsilon\check{w})\frac{\Delta H(r,\check{w}+\check{t},\check{q};*)}{\check{w}}\text{d}\check{w},\\ \Delta H(r,\check{t},\check{q};*):=&H(r+\check{t},\check{q};*)-H(\check{t},\check{q};*).\end{split}$
+We underscore that all the functions introduced in (7.3)–(7.5) are computed using a local patch of data $g=\chi_{V}\hat{f}$ as described in Lemma 4.2.
+By (6.18), the expression for $\Delta F_{\epsilon}^{(l)}$ becomes:
+(7.6) $\begin{split}&\Delta F_{\epsilon}^{(l)}(\epsilon\check{x})\\ &=-\frac{\epsilon^{1/2}}{\pi}\sum_{j_{1}\in J_{\alpha}}\Delta G\bigg{(}\frac{Q(\epsilon\check{x},\alpha_{j_{1}})-Q(\alpha_{j_{1}})}{\epsilon},\frac{Q(\alpha_{j_{1}})}{\epsilon},\frac{P(\alpha_{j_{1}})}{\epsilon};\alpha_{j_{1}},\epsilon\bigg{)}.\end{split}$
 The following result is proven in Appendix E.1.
 Lemma 7.1.
-Suppose assumptions 3.13.4 are satisfied. Pick any $A>0$. One has
-(7.5) $\begin{split}\Delta G(r,\check{t},\check{q};\theta,\epsilon),\partial_{r}\Delta G(r,\check{t},\check{q};\theta,\epsilon)=O(|\check{t}|^{-1}),\ \check{t}\to\infty.\end{split}$
-Also,
-[…]
-In the above formulas the big-$O$ terms are uniform in $\check{q}\in\mathbb{R}$, $\theta\in\text{dom}\,P$, $|\check{t}|\leq A/\epsilon$, $|r|\leq A$, and $0<\epsilon\ll 1$.
-Recall that $Q(\alpha)$ stands for $Q(x_{0}=0_{2},\alpha)$. Suppose first that the selected $P(\alpha)$ is associated with $x_{0}$, so $Q(0)=0$. By Lemma D.3, $Q_{\alpha}^{\prime}(0)=0$. From
-(7.7) $\begin{split}Q(\epsilon\check{x},\alpha)=&Q_{x}^{\prime}(0)\epsilon\check{x}+Q_{\alpha\alpha}^{\prime\prime}(0)\alpha^{2}/2+O(\epsilon^{3/2}),\\ Q_{\alpha}^{\prime}(\epsilon\check{x},\alpha)=&Q_{\alpha\alpha}^{\prime\prime}(0)\alpha+O(\epsilon),\quad\alpha=O(\epsilon^{1/2}),\end{split}$
-we conclude
-(7.8) $\alpha_{1}(\epsilon\check{x})=O(\epsilon^{1/2}),\ \alpha_{2}(\epsilon\check{x})=O(\epsilon),$
-where $\alpha_{1}$, $\alpha_{2}$ are found by solving $Q(\epsilon\check{x},\alpha_{1})=0$ and $Q_{\alpha}^{\prime}(\epsilon\check{x},\alpha_{2})=0$.
-By (6.7), $Q_{x}^{\prime}(x,\alpha)=\Phi_{x}^{\prime}(x,\alpha)$.
+Suppose Assumptions 3.13.4 are satisfied. Pick any $A_{2}>0$. One has:
+(7.7) $\begin{split}|\Delta G(r,\check{t},\check{q};*)|\leq c\frac{|r|\ln(1/|r|)}{1+|\check{t}|},\end{split}$
+(7.8) $\begin{split}|\Delta G(r+h,\check{t},\check{q};*)-\Delta G(r,\check{t},\check{q};*)|\leq c\frac{|h|\ln(1/|h|)}{1+|\check{t}|},\end{split}$
+and
+(7.9) $|\Delta G(r,\check{t}+h,\check{q};*)-\Delta G(r,\check{t},\check{q};*)|\leq c\frac{|h|\ln(1/|h|)}{1+|\check{t}|}.$
+In the above formulas, the $c$’s are independent of $|r|\leq A_{2}$, $|\check{t}|\leq A_{2}/\epsilon$, $|\check{q}|\leq A_{2}/\epsilon$, $\theta\in I_{\alpha}$, $0<\epsilon\ll 1$, and $0<h\ll 1$.
+We will apply Lemma 7.1 to (7.6), so we select $A_{2}$ as follows:
+(7.10) $A_{2}=\max_{|\check{x}|\leq A_{1},\alpha_{j_{1}}\in I_{\alpha},0<\epsilon\ll 1}\max\bigg{(}\frac{|Q(\epsilon\check{x},\alpha_{j_{1}})-Q(\alpha_{j_{1}})|}{\epsilon},|Q(\alpha_{j_{1}})|,|P(\alpha_{j_{1}})|\bigg{)}.$
+Recall that $Q(\alpha)$ stands for $Q(x_{0}=0_{2},\alpha)$. Since $\check{x}$ is confined to a bounded set and $Q$ is smooth, the right-hand side of (7.10) is bounded.
+Suppose first that $P(\alpha)$ is not associated with $x_{0}$.
+This means there is a neighborhood $U^{\prime}$ of $x_{0}$ such that no curve $\mathcal{S}_{(\alpha,P(\alpha))}$, $\alpha\in I_{\alpha}$, is tangent to $\mathcal{S}\cap U^{\prime}$.
+There are two possibilities: (a) Neither of these curves passes through $x_{0}$ (in this case $Q(\alpha)$ is bounded away from zero on $I_{\alpha}$) and (b) one of the curves contains $x_{0}$ (in this case $Q(\tilde{\alpha})=0$ for some $\tilde{\alpha}\in I_{\alpha}$).
+(a) If $|Q(\alpha)|>c$, $\alpha\in I_{\alpha}$, the inequality (7.7) implies
+(7.11) $\epsilon^{1/2}\textstyle\sum_{j_{1}\in J_{\alpha}}|\Delta G|\leq c\epsilon^{1/2}\textstyle\sum_{j_{1}=1}^{1/\epsilon}\epsilon=O(\epsilon^{1/2}).$
+Here and below, the arguments of $\Delta G$ are the same as in (7.6).
+(b) If $Q(\tilde{\alpha})=0$, but $|Q^{\prime}(\alpha)|>c$ on $I_{\alpha}$, then $|Q(\alpha)|\asymp|\alpha-\tilde{\alpha}|$, $\alpha\in I_{\alpha}$, and (7.7) leads to
+(7.12) ${\color[rgb]{0,0,0}\epsilon^{1/2}\textstyle\sum_{j_{1}\in J_{\alpha}}|\Delta G|\leq c\epsilon^{1/2}\textstyle\sum_{j_{1}=1}^{1/\epsilon}\frac{\epsilon}{\epsilon j_{1}}=O(\epsilon^{1/2}\ln(1/\epsilon)).}$
+Suppose now $P(\alpha)$ is associated with $x_{0}$.
+Therefore, $I_{\alpha}$ is a small neighborhood of $\alpha_{0}$.
+Recall that we assumed $\alpha_{0}=0$.
+By the definition of the function $Q$ (see (6.10)), $Q(0)=0$.
+Applying Lemma D.3 with $\tilde{\alpha}=\alpha_{0}=0$ and $\tilde{x}=x_{0}=0_{2}$ gives $Q_{\alpha}^{\prime}(0)=0$.
 Pick any $A\gg 1$. By Lemma D.3, $Q_{\alpha\alpha}^{\prime\prime}(0)>0$. Hence, by continuity,
-(7.9) $Q(\theta)\geq c\theta^{2},\quad|\theta|\leq\delta,$
-for some $0<\delta\ll 1$. Also, $Q(\theta)\geq c$, $|\theta|\geq\delta$, $\theta\in\text{dom}\,P$. By (7.5), this implies
-(7.10) $\begin{split}&\epsilon^{1/2}\bigg{(}\sum_{A\epsilon^{1/2}\leq|\alpha_{j_{1}}|\leq\delta}+\sum_{|\alpha_{j_{1}}|>\delta}\bigg{)}|\Delta G|\\ &\leq c\epsilon^{1/2}\bigg{(}\sum_{j_{1}>A\epsilon^{-1/2}}(\epsilon j_{1}^{2})^{-1}+\sum_{j_{1}=\delta/\epsilon}^{1/\epsilon}\epsilon\bigg{)}=O(A^{-1})+O(\epsilon^{1/2}),\end{split}$
-where the arguments of $\Delta G$ are the same as in (7.4).
-By (7.5) and (7.6), for $|\alpha_{j_{1}}|\leq A\epsilon^{1/2}$:
-(7.11) $\begin{split}&\Delta G\bigg{(}\frac{Q(\epsilon\check{x},\alpha_{j_{1}})-Q(\alpha_{j_{1}})}{\epsilon},\frac{Q(\alpha_{j_{1}})}{\epsilon},\frac{P(\alpha_{j_{1}})}{\epsilon};\alpha_{j_{1}},\epsilon\bigg{)}\\ &=\Delta G\bigg{(}\Phi_{x}^{\prime}\check{x}+O(\epsilon^{1/2}),\frac{Q_{\alpha\alpha}^{\prime\prime}}{2}\epsilon(\mu j_{1})^{2}+O(\epsilon^{1/2}),\frac{P(\epsilon\mu j_{1})}{\epsilon};\alpha_{j_{1}},\epsilon\bigg{)}\\ &=\Delta G\bigg{(}\Phi_{x}^{\prime}\check{x},\frac{Q_{\alpha\alpha}^{\prime\prime}}{2}\epsilon(\mu j_{1})^{2},\frac{P(\epsilon\mu j_{1})}{\epsilon};\alpha_{j_{1}},\epsilon\bigg{)}+\frac{O(\epsilon^{1/2}\ln(1/\epsilon))}{(1+\epsilon j_{1}^{2})^{3/2}}.\end{split}$
-Pick any $P(\alpha)$, which is not associated with a segment of $\mathcal{S}$ through $x_{0}$.
-This means there is a neighborhood $U^{\prime}$ of $x_{0}$ such that none of the curves $\mathcal{S}_{(\alpha,P(\alpha))}$, $\alpha\in\text{dom}\,P$, is tangent to $\mathcal{S}\cap U^{\prime}$.
-There are two possibilities: (a) Neither of these curves passes through $x_{0}$ (in this case $Q(\alpha)$ is bounded away from zero on $\text{dom}\,P$) and (b) one of the curves contains $x_{0}$ (in this case $Q(\alpha)=0$ for some $\alpha\in\text{dom}\,P$).
-If $|Q(\alpha)|>c$, $\alpha\in\text{dom}\,P$, then, similarly to (7.10),
-(7.12) $\epsilon^{1/2}\textstyle\sum_{j_{1}}|\Delta G|\leq c\epsilon^{1/2}\textstyle\sum_{j_{1}=1}^{1/\epsilon}\epsilon=O(\epsilon^{1/2}).$
-If $Q(\tilde{\alpha})=0$, but $|Q^{\prime}(\alpha)|>c$, then $|Q(\tilde{\alpha})|\asymp|\alpha-\tilde{\alpha}|$, $\alpha\in\text{dom}\,P$, and
-(7.13) $\epsilon^{1/2}\textstyle\sum_{j_{1}}|\Delta G|\leq c\epsilon^{1/2}\ln(1/\epsilon)\textstyle\sum_{j_{1}=1}^{1/\epsilon}(1/j_{1})=O(\epsilon^{1/2}\ln^{2}(1/\epsilon)).$
-Combine (7.10)–(7.13) and use (7.4)
-(7.14) $\begin{split}\Delta F_{\epsilon}^{(l)}(\epsilon\check{x})=&-\frac{\epsilon^{1/2}}{\pi}\sum_{|\alpha_{j_{1}}|\leq A\epsilon^{1/2}}\Delta G\bigg{(}\Phi_{x}^{\prime}\check{x},\frac{Q_{\alpha\alpha}^{\prime\prime}}{2}\epsilon(\mu j_{1})^{2},\frac{P(\epsilon\mu j_{1})}{\epsilon};\alpha_{j_{1}},\epsilon\bigg{)}\\ &+O(\epsilon^{1/2}\ln^{2}(1/\epsilon))+O(1/A),\end{split}$
-where $O(\epsilon^{1/2}\ln^{2}(1/\epsilon))$ is independent of $A$, and $P$ is associated with $x_{0}$.
-Here we have used that $\epsilon^{1/2}\sum_{j_{1}}(1+\epsilon j_{1}^{2})^{-3/2}<c$, $10<\epsilon<$ 1.
+(7.13) $Q(\alpha)\geq c\alpha^{2},\quad|\alpha|\leq\delta,$
+for some $0<\delta\ll 1$. Also, $Q(\alpha)\geq c$, $|\alpha|\geq\delta$, $\alpha\in I_{\alpha}$. By (7.7), this implies
+(7.14) $\begin{split}&\epsilon^{1/2}\bigg{(}\sum_{A\epsilon^{1/2}\leq|\alpha_{j_{1}}|\leq\delta}+\sum_{j_{1}\in J_{\alpha},|\alpha_{j_{1}}|>\delta}\bigg{)}|\Delta G|\\ &\leq c\epsilon^{1/2}\bigg{(}\sum_{j_{1}>A\epsilon^{-1/2}}\frac{\epsilon}{(\epsilon j_{1})^{2}}+\sum_{j_{1}=\delta/\epsilon}^{1/\epsilon}\epsilon\bigg{)}=O(A^{-1})+O(\epsilon^{1/2}),\end{split}$
+By (6.10), (7.8) and (7.9), for $|\alpha_{j_{1}}|\leq A\epsilon^{1/2}$:
+(7.15) $\begin{split}&\Delta G\bigg{(}\frac{Q(\epsilon\check{x},\alpha_{j_{1}})-Q(\alpha_{j_{1}})}{\epsilon},\frac{Q(\alpha_{j_{1}})}{\epsilon},\frac{P(\alpha_{j_{1}})}{\epsilon};\alpha_{j_{1}},\epsilon\bigg{)}\\ &=\Delta G\bigg{(}\text{d}_{x}\Phi\cdot\check{x}+O(\epsilon^{1/2}),\frac{Q_{\alpha\alpha}^{\prime\prime}}{2}\epsilon(\mu j_{1})^{2}+O(\epsilon^{1/2}),\frac{P(\epsilon\mu j_{1})}{\epsilon};\alpha_{j_{1}},\epsilon\bigg{)}\\ &=\Delta G\bigg{(}\text{d}_{x}\Phi\cdot\check{x},\frac{Q_{\alpha\alpha}^{\prime\prime}}{2}\epsilon(\mu j_{1})^{2},\frac{P(\epsilon\mu j_{1})}{\epsilon};\alpha_{j_{1}},\epsilon\bigg{)}+O(\epsilon^{1/2}\ln(1/\epsilon)).\end{split}$
+Combine (7.11), (7.12), (7.14), (7.15), and use (7.6) to obtain
+(7.16) $\begin{split}\Delta F_{\epsilon}^{(l)}(\epsilon\check{x})=&-\frac{\epsilon^{1/2}}{\pi}\sum_{|\alpha_{j_{1}}|\leq A\epsilon^{1/2}}\Delta G\bigg{(}\text{d}_{x}\Phi\cdot\check{x},\frac{Q_{\alpha\alpha}^{\prime\prime}}{2}\epsilon(\mu j_{1})^{2},\frac{P(\epsilon\mu j_{1})}{\epsilon};\alpha_{j_{1}},\epsilon\bigg{)}\\ &+O(\epsilon^{1/2}\ln(1/\epsilon))+O(1/A).\end{split}$
+Here we have used that there are $O(\epsilon^{1/2})$ values of $j_{1}$ such that $|\alpha_{j_{1}}|\leq A\epsilon^{1/2}$.
+Our analysis shows (see (7.11), (7.12)) that if $\Gamma_{l}$ is not associated with $x_{0}$, then the data in $V_{l}=I_{\alpha}^{l}\times I_{p}^{l}$ does not contribute to the DTB at $x_{0}$.
 The following result is proven in Appendix E.2.
@@
 Under the assumptions of Theorem 3.6 one has
-(7.15) $\begin{split}\Delta F_{0}(\check{x})=-\frac{1}{\pi}\int_{|s|\leq A}\int_{0}^{1}&\big{[}G_{0}\big{(}\Phi_{x}^{\prime}\check{x}+(Q_{\alpha\alpha}^{\prime\prime}/2)s^{2},\check{q}\big{)}\\ &-G_{0}\big{(}(Q_{\alpha\alpha}^{\prime\prime}/2)s^{2},\check{q}\big{)}\big{]}\text{d}\check{q}\text{d}s+O(1/A),\end{split}$
+(7.17) $\begin{split}\Delta F_{0}(\check{x})=-\frac{1}{\pi}\int_{|s|\leq A}\int_{0}^{1}&\big{[}G_{0}\big{(}\text{d}_{x}\Phi\cdot\check{x}+(Q_{\alpha\alpha}^{\prime\prime}/2)s^{2},\check{q}\big{)}\\ &-G_{0}\big{(}(Q_{\alpha\alpha}^{\prime\prime}/2)s^{2},\check{q}\big{)}\big{]}\text{d}\check{q}\text{d}s+O(1/A),\end{split}$
 where
-(7.16) $\begin{split}&G_{0}(\check{t},\check{q}):=\int_{\mathbb{R}}\frac{H_{0}(\check{w}+\check{t},\check{q})}{\check{w}}\text{d}\check{w},\\ &H_{0}(\check{t},\check{q}):=a_{0}(y_{0})\sum_{j_{2}}h_{0}^{\prime}(\check{t}+\check{q}-j_{2})(j_{2}-\check{q})_{+}^{1/2},\\ &h_{0}(u):=K(x_{0},\alpha_{0})\int_{\mathbb{R}}\varphi_{\alpha}(s)\varphi_{p}(\mu s\Phi_{\alpha}^{\prime}(0)+u)\text{d}s.\end{split}$
+(7.18) $\begin{split}&G_{0}(\check{t},\check{q}):=\int_{\mathbb{R}}\frac{H_{0}(\check{w}+\check{t},\check{q})}{\check{w}}\text{d}\check{w},\\ &H_{0}(\check{t},\check{q}):=a_{0}(y_{0})\sum_{j_{2}\in\mathbb{Z}}h_{0}^{\prime}(\check{t}+\check{q}-j_{2})(j_{2}-\check{q})_{+}^{1/2},\\ &h_{0}(u):=K(x_{0},\alpha_{0})\int_{\mathbb{R}}\varphi_{\alpha}(s)\varphi_{p}(\mu s\Phi_{\alpha}^{\prime}(0)+u)\text{d}s.\end{split}$
 See (6.3) for the definition of $K(x,\alpha)$. Letting $A\to\infty$ in (7.17) gives
-(7.17) $\begin{split}\Delta F_{0}(\check{x})=&-\frac{1}{\pi}\int_{\mathbb{R}}\int_{0}^{1}\big{[}G_{0}\big{(}\Phi_{x}^{\prime}\check{x}+(Q_{\alpha\alpha}^{\prime\prime}/2)s^{2},\check{q}\big{)}-G_{0}\big{(}(Q_{\alpha\alpha}^{\prime\prime}/2)s^{2},\check{q}\big{)}\big{]}\text{d}\check{q}\text{d}s\\ =&-\frac{a_{0}(y_{0})}{\pi}\int_{\mathbb{R}}\int_{\mathbb{R}}\int_{0}^{\infty}\frac{1}{\check{w}}\big{[}h_{0}^{\prime}\big{(}\check{w}+\Phi_{x}^{\prime}\check{x}+(Q_{\alpha\alpha}^{\prime\prime}/2)s^{2}-\check{q}\big{)}\\ &\hskip 99.58464pt-h_{0}^{\prime}\big{(}\check{w}+(Q_{\alpha\alpha}^{\prime\prime}/2)s^{2}-\check{q}\big{)}\big{]}\check{q}^{1/2}\text{d}\check{q}\text{d}s\text{d}\check{w}.\end{split}$
-It is easy to see that $G_{0}(\check{t}+r,\check{q})-G_{0}(\check{t},\check{q})$ satisfies the same estimates as $\Delta G$ in (7.5), so the integral in (7.17) is absolutely convergent.
+(7.19) $\begin{split}&\Delta F_{0}(\check{x})\\ &=-\frac{1}{\pi}\int_{\mathbb{R}}\int_{0}^{1}\big{[}G_{0}\big{(}\text{d}_{x}\Phi\cdot\check{x}+(Q_{\alpha\alpha}^{\prime\prime}/2)s^{2},\check{q}\big{)}-G_{0}\big{(}(Q_{\alpha\alpha}^{\prime\prime}/2)s^{2},\check{q}\big{)}\big{]}\text{d}\check{q}\text{d}s\\ &=-\frac{a_{0}(y_{0})}{\pi}\int_{\mathbb{R}}\int_{\mathbb{R}}\int_{0}^{\infty}\frac{1}{\check{w}}\big{[}h_{0}^{\prime}\big{(}\check{w}+\text{d}_{x}\Phi\cdot\check{x}+(Q_{\alpha\alpha}^{\prime\prime}/2)s^{2}-\check{q}\big{)}\\ &\hskip 99.58464pt-h_{0}^{\prime}\big{(}\check{w}+(Q_{\alpha\alpha}^{\prime\prime}/2)s^{2}-\check{q}\big{)}\big{]}\check{q}^{1/2}\text{d}\check{q}\text{d}s\text{d}\check{w}.\end{split}$
+Comparing (6.14) with (7.18), we see that the functions $H$ and $H_{0}$ are analogous and they have the same asymptotic behavior as $\check{t}\to\infty$.
+More precisely, $H_{0}$ satisfies Lemma D.1 with $\epsilon=0$.
+To apply this lemma to obtain the asymptotics of $H_{0}$, we set $\epsilon=0$ and drop all the restrictions of the kind $\check{t}<A/\epsilon$.
+The same modifications should be made in the proof of Lemma D.1 in Appendix D.2.
+Likewise, the functions $G$ of (6.17) and $G_{0}$ of (7.18) are analogous and they also have the same asymptotic behavior as $\check{t}\to\infty$.
+Repeating the argument in the proof of (7.7) (see Appendix E.1 from the beginning through the paragraph following (E.5)) and replacing $H$ with $H_{0}$, we see that $G_{0}(\check{t}+r,\check{q})-G_{0}(\check{t},\check{q})$ satisfies the same estimates as $\Delta G$ in (7.7), so the first integral in (7.19) is absolutely convergent.
 The following result is proven in Appendix E.3.
@@
 Under the assumptions of Theorem 3.6 one has
-(7.18) $\Delta F_{0}(\check{x})=C\int_{0}^{\Phi_{x}^{\prime}\check{x}}h_{0}(-t)\text{d}t,\ C:=\frac{a_{0}(y_{0})}{4(Q_{\alpha\alpha}^{\prime\prime}/2)^{1/2}}.$
+(7.20) $\Delta F_{0}(\check{x})=C\int_{0}^{\text{d}_{x}\Phi\cdot\check{x}}h_{0}(-t)\text{d}t,\ C:=\frac{a_{0}(y_{0})}{4(Q_{\alpha\alpha}^{\prime\prime}/2)^{1/2}}.$
 8. Computing the DTB function. End of proof of Theorem 3.6
-Pick any $A_{0}>0$. Throughout this section we assume $|\check{x}|\leq A_{0}$.
+Pick any $A_{0}>0$.
+Consistent with Theorem 3.6, throughout this section we assume $|\check{x}|<A_{0}$.
+As established in the preceding section, it suffices to consider only the segment $\Gamma_{l}$ (and the corresponding $a_{0}(y),P(\alpha)$) that is associated with $x_{0}$.
 8.1. Reduction of the DTB function to an integral
-Now we study $G_{\epsilon}$ introduced in (5.7).
+Now we study $\mathcal{G}_{\epsilon}$ introduced in (5.7).
 As usual, suppose $x_{0}=0_{2}$.
-Change variables $\hat{\xi}=\epsilon\xi$ and $\check{x}=x/\epsilon$, $\check{z}=z/\epsilon$ in (5.7) and express $G_{\epsilon}$ as follows:
-(8.1) $\begin{split}&G_{\epsilon}(\check{x}):=\frac{1}{(2\pi)^{2}}\int_{\mathcal{U}/\epsilon}\int_{\mathbb{R}^{2}}\big{(}\tilde{D}_{0}(\epsilon\check{x},\hat{\xi})+\kappa|\hat{\xi}|^{3}\big{)}^{-1}e^{-i\hat{\xi}\cdot(\check{x}-\check{z})}\text{d}\hat{\xi}\,F_{\epsilon}(\epsilon\check{z})\text{d}\check{z}.\end{split}$
-For clarity, $G_{\epsilon}$ in (LABEL:to_scl_PDO_2) is different from the functions $G$ and $G_{0}$ used in sections 6 and 7. Since
+Change variables $\hat{\xi}=\epsilon\xi$ and $\check{x}=x/\epsilon$, $\check{z}=z/\epsilon$ in (5.7) and express $\mathcal{G}_{\epsilon}$ as follows:
+(8.1) $\begin{split}&\mathcal{G}_{\epsilon}(\check{x}):=\frac{1}{(2\pi)^{2}}\int_{\mathcal{U}/\epsilon}\int_{\mathbb{R}^{2}}\big{(}\tilde{D}_{0}(\epsilon\check{x},\hat{\xi})+\kappa|\hat{\xi}|^{3}\big{)}^{-1}e^{-i\hat{\xi}\cdot(\check{x}-\check{z})}\text{d}\hat{\xi}\,F_{\epsilon}(\epsilon\check{z})\text{d}\check{z}.\end{split}$
+Since
 (8.2) $\tilde{D}_{0}(\epsilon\check{x},\hat{\xi})+\kappa|\hat{\xi}|^{3}\asymp|\hat{\xi}|^{3},\ |\hat{\xi}|\to\infty,$
 the integral with respect to $\hat{\xi}$ in (LABEL:to_scl_PDO_2) is absolutely convergent.
-Our aim is to reconstruct the jump of $f$ at $x_{0}=0_{2}$, so we compute $\Delta G_{\epsilon}(\check{x}):=G_{\epsilon}(\check{x})-G_{\epsilon}(0_{2})$.
-By (LABEL:to_scl_PDO_2), $\Delta G_{\epsilon}(\check{x})=\Delta G_{\epsilon}^{(1)}(\check{x})+\Delta G_{\epsilon}^{(2)}(\check{x})$, where
-(8.3) $\begin{split}\Delta G_{\epsilon}^{(1)}(\check{x})=\frac{1}{(2\pi)^{2}}\int_{\mathcal{U}/\epsilon}\int_{\mathbb{R}^{2}}&\big{[}(\tilde{D}_{0}(\epsilon\check{x},\hat{\xi})+\kappa|\hat{\xi}|^{3})^{-1}-(\tilde{D}_{0}(0_{2},\hat{\xi})+\kappa|\hat{\xi}|^{3})^{-1}\big{]}\\ &\times e^{-i\hat{\xi}\cdot(\check{x}-\check{z})}\text{d}\hat{\xi}\,F_{\epsilon}(\epsilon\check{z})\text{d}\check{z},\end{split}$
-(8.4) $\begin{split}&\Delta G_{\epsilon}^{(2)}(\check{x})=\frac{1}{(2\pi)^{2}}\int_{\mathcal{U}/\epsilon}\int_{\mathbb{R}^{2}}\frac{e^{-i\hat{\xi}\cdot\check{x}}-1}{\tilde{D}_{0}(0_{2},\hat{\xi})+\kappa|\hat{\xi}|^{3}}e^{i\hat{\xi}\cdot\check{z}}\text{d}\hat{\xi}\,F_{\epsilon}(\epsilon\check{z})\text{d}\check{z}.\end{split}$
+Our aim is to reconstruct the jump of $f$ at $x_{0}=0_{2}$, so we compute $\Delta\mathcal{G}_{\epsilon}(\check{x}):=\mathcal{G}_{\epsilon}(\check{x})-\mathcal{G}_{\epsilon}(0_{2})$.
+By (LABEL:to_scl_PDO_2), $\Delta\mathcal{G}_{\epsilon}(\check{x})=\Delta\mathcal{G}_{\epsilon}^{(1)}(\check{x})+\Delta\mathcal{G}_{\epsilon}^{(2)}(\check{x})$, where
+(8.3) $\begin{split}\Delta\mathcal{G}_{\epsilon}^{(1)}(\check{x})=\frac{1}{(2\pi)^{2}}\int_{\mathcal{U}/\epsilon}\int_{\mathbb{R}^{2}}&\big{[}(\tilde{D}_{0}(\epsilon\check{x},\hat{\xi})+\kappa|\hat{\xi}|^{3})^{-1}-(\tilde{D}_{0}(0_{2},\hat{\xi})+\kappa|\hat{\xi}|^{3})^{-1}\big{]}\\ &\times e^{-i\hat{\xi}\cdot(\check{x}-\check{z})}\text{d}\hat{\xi}\,F_{\epsilon}(\epsilon\check{z})\text{d}\check{z},\end{split}$
+and
+(8.4) $\begin{split}&\Delta\mathcal{G}_{\epsilon}^{(2)}(\check{x})=\frac{1}{(2\pi)^{2}}\int_{\mathcal{U}/\epsilon}\int_{\mathbb{R}^{2}}\frac{e^{-i\hat{\xi}\cdot\check{x}}-1}{\tilde{D}_{0}(0_{2},\hat{\xi})+\kappa|\hat{\xi}|^{3}}e^{i\hat{\xi}\cdot\check{z}}\text{d}\hat{\xi}\,F_{\epsilon}(\epsilon\check{z})\text{d}\check{z}.\end{split}$
 The following result is proven in Appendix F.1.
 Lemma 8.1.
-Let $I=(a,b)$ be an interval where $-\pi<a<0<b<\pi$. Let $D(\lambda,\theta,\epsilon)\in C^{\infty}\big{(}(0,\infty)\times I\times(0,\epsilon_{0})\big{)}$ be a function which satisfies
-(8.5) $\big{|}\partial_{\theta}^{l}\partial_{\lambda}^{k}D(\lambda,\theta,\epsilon)\big{|}\leq c_{k,l}(1+\lambda)^{-3},\ (\lambda,\theta,\epsilon)\in(0,\infty)\times I\times(0,\epsilon_{0}),$
-for any $k=0,1,2$, $l\in\mathbb{N}_{0}$, some $c_{k,l}$, which can be independent of $\epsilon$, and some $\epsilon_{0}>0$. Define
+Let $I\subset\mathcal{I}_{\alpha}$ be an interval. Let $D(\lambda,\theta,\epsilon)\in C^{\infty}\big{(}(0,\infty)\times I\times(0,\epsilon_{0})\big{)}$ be a function which satisfies
+(8.5) $\big{|}\partial_{\theta}^{l}\partial_{\lambda}^{k}D(\lambda,\theta,\epsilon)\big{|}\leq c_{k,l}\epsilon^{n}(1+\lambda)^{-3},\ (\lambda,\theta,\epsilon)\in(0,\infty)\times I\times(0,\epsilon_{0}),$
+for any $k=0,1,2$, $l\in\mathbb{N}_{0}$, and some $c_{k,l}$ independent of $\epsilon$, $\epsilon_{0}>0$, and $n\in\mathbb{N}_{0}$. Define
 (8.6) $J(r):=\int_{0}^{\infty}\int_{I}D(\lambda,\theta,\epsilon)e^{ir\lambda\sin\theta}\text{d}\theta\lambda\text{d}\lambda.$
 Then
-(8.7) $J(r)=\begin{cases}O(r^{-2})&\\ O(r^{-3}),&\text{if }D(\lambda=0^{+},\theta,\epsilon)\equiv 0,\end{cases}\ r\to\infty,$
-uniformly in $\epsilon\in(0,\epsilon_{0})$. Moreover, if $c_{k,l}=O(\epsilon)$, then
-(8.8) $J(r)=\begin{cases}O(\epsilon)O(r^{-2})\\ O(\epsilon)O(r^{-3}),&\text{if }D(\lambda=0^{+},\theta,\epsilon)\equiv 0,\end{cases}\ \epsilon\to 0,r\to\infty.$
-The expression in brackets in (8.3) satisfies (8.5) with $c_{k,l}=O(\epsilon)$ uniformly in $\check{x}$ within bounded sets (see (C.7)).
+(8.7) $J(r)=\epsilon^{n}\begin{cases}O(r^{-2})&\\ O(r^{-3}),&\text{if }D(\lambda=0^{+},\theta,\epsilon)\equiv 0,\end{cases}\ r\to\infty,$
+uniformly in $\epsilon\in(0,\epsilon_{0})$.
+The expression in brackets in (8.3) satisfies (8.5) with $n=1$ uniformly in $\check{x}$ within bounded sets (see (C.9)).
 Here $\hat{\xi}=\lambda(\cos\theta,\sin\theta)$, $\lambda=|\hat{\xi}|$.
-By the top case in (8.8), the integral with respect to $\hat{\xi}$ in (8.3) is $O(\epsilon)O(|\check{z}|^{-2})$, $|\check{z}|\to\infty$.
+By the top case in (8.7) with $n=1$, the integral with respect to $\hat{\xi}$ in (8.3) is $\epsilon O(|\check{z}|^{-2})$, $|\check{z}|\to\infty$.
 Recall that $\check{x}$ is confined to a bounded set.
@@
 Hence
-(8.9) $|\Delta G_{\epsilon}^{(1)}(\check{x})|\leq c\epsilon\int_{|\check{z}|\leq 1/\epsilon}(1+|\check{z}|)^{-2}\text{d}\check{z}=O(\epsilon\ln(1/\epsilon)).$
-The fraction in (LABEL:del_bfe_2) satisfies (8.5) uniformly in $\check{x}$ within bounded sets.
+(8.8) $|\Delta\mathcal{G}_{\epsilon}^{(1)}(\check{x})|\leq c\epsilon\int_{|\check{z}|\leq 1/\epsilon}(1+|\check{z}|)^{-2}\text{d}\check{z}=O(\epsilon\ln(1/\epsilon)).$
+The fraction in (LABEL:del_bfe_2) satisfies (8.5) with $n=0$ uniformly in $\check{x}$ within bounded sets.
 Moreover, its value at $\lambda=0^{+}$ equals zero.
-By the bottom case in (8.7), the integral with respect to $\hat{\xi}$ is $O(|\check{z}|^{-3})$, $|\check{z}|\to\infty$.
-This means that the integral with respect to $\check{z}$ in (LABEL:del_bfe_2) is absolutely convergent.
+By the bottom case in (8.7) with $n=0$, the integral with respect to $\hat{\xi}$ is $O(|\check{z}|^{-3})$, $|\check{z}|\to\infty$.
+This means that the integral with respect to $\check{z}$ in (LABEL:del_bfe_2) is absolutely convergent and uniformly bounded for $0<\epsilon\ll 1$.
 In both integrals we have used Lemma 6.1.
 In Section 7 we computed the limit
-(8.10) $\Delta F_{0}(\check{x})=\lim_{\epsilon\to 0}\Delta F_{\epsilon}(\epsilon\check{x}),\ |\check{x}|\leq A,$
-for any $A>0$ (see (7.1) and (7.18)). Pick any $A\gg A_{0}$ and use (8.9) to obtain from (LABEL:to_scl_PDO_2), (8.3), (LABEL:del_bfe_2):
-(8.11) $\begin{split}\Delta G_{\epsilon}(\check{x})=&\frac{1}{(2\pi)^{2}}\int_{|\check{z}|\leq A}\bigg{[}\int_{\mathbb{R}^{2}}\frac{e^{-i\hat{\xi}\cdot\check{x}}-1}{\tilde{D}_{0}(0_{2},\hat{\xi})+\kappa|\hat{\xi}|^{3}}e^{i\hat{\xi}\cdot\check{z}}\text{d}\hat{\xi}\bigg{]}\Delta F_{\epsilon}(\epsilon\check{z})\text{d}\check{z}\\ &+\frac{F_{\epsilon}(0)}{2\pi}\int_{\mathbb{R}^{2}}\frac{e^{-i\hat{\xi}\cdot\check{x}}-1}{\tilde{D}_{0}(0_{2},\hat{\xi})+\kappa|\hat{\xi}|^{3}}\frac{AJ_{1}(A|\hat{\xi}|)}{|\hat{\xi}|}\text{d}\hat{\xi}\\ &+O(1/A)+O(\epsilon\ln(1/\epsilon)),\end{split}$
-where $J_{1}$ is the Bessel function of the first kind.
-By Lemma 8.1, the integral in brackets above is $O(|\check{z}|^{-3})$, $|\check{z}|\to\infty$.
+(8.9) $\Delta F_{0}(\check{x})=\lim_{\epsilon\to 0}\Delta F_{\epsilon}(\epsilon\check{x}),\ |\check{x}|\leq A_{1},$
+for any $A_{1}>0$ (see (7.3) and (7.20)). Pick any $A_{1}\gg 1$ and use (8.8) to obtain from (LABEL:to_scl_PDO_2), (8.3), (LABEL:del_bfe_2):
+(8.10) $\begin{split}\Delta\mathcal{G}_{\epsilon}(\check{x})=&\frac{1}{(2\pi)^{2}}\int_{|\check{z}|\leq A_{1}}\bigg{[}\int_{\mathbb{R}^{2}}\frac{e^{-i\hat{\xi}\cdot\check{x}}-1}{\tilde{D}_{0}(0_{2},\hat{\xi})+\kappa|\hat{\xi}|^{3}}e^{i\hat{\xi}\cdot\check{z}}\text{d}\hat{\xi}\bigg{]}\Delta F_{\epsilon}(\epsilon\check{z})\text{d}\check{z}\\ &+\frac{F_{\epsilon}(0)}{(2\pi)^{2}}\int_{\mathbb{R}^{2}}\frac{e^{-i\hat{\xi}\cdot\check{x}}-1}{\tilde{D}_{0}(0_{2},\hat{\xi})+\kappa|\hat{\xi}|^{3}}\tilde{\vartheta}_{A_{1}}(\hat{\xi})\text{d}\hat{\xi}\\ &+O(1/A_{1})+O(\epsilon\ln(1/\epsilon)),\end{split}$
+where $\tilde{\vartheta}_{A_{1}}(\hat{\xi})$ is the Fourier transform of the characteristic function of the ball $\{|\check{x}|\leq A_{1}\}$.
+By Lemma 8.1 with $n=0$, the integral in brackets above is $O(|\check{z}|^{-3})$, $|\check{z}|\to\infty$.
 Use (8.9), and the Lebesgue dominated convergence theorem to take the limit as $\epsilon\to 0$ in the first integral in (8.10) to obtain:
-(8.12) $\begin{split}\frac{1}{(2\pi)^{2}}\int_{|\check{z}|\leq A}\int_{\mathbb{R}^{2}}\frac{e^{-i\hat{\xi}\cdot\check{x}}-1}{\tilde{D}_{0}(0_{2},\hat{\xi})+\kappa|\hat{\xi}|^{3}}e^{i\hat{\xi}\cdot\check{z}}\text{d}\hat{\xi}\,\Delta F_{0}(\check{z})\text{d}\check{z}.\end{split}$
-Since $A\gg A_{0}$ can be arbitrarily large and $AJ_{1}(A|\hat{\xi}|)/|\hat{\xi}|\to\delta(\hat{\xi})$ as $A\to\infty$, (8.11), (8.12), and Lemma 6.1 imply
-(8.13) $\begin{split}\lim_{\epsilon\to 0}\Delta G_{\epsilon}(\check{x})=&\frac{1}{(2\pi)^{2}}\int_{\mathbb{R}^{2}}\int_{\mathbb{R}^{2}}\frac{e^{-i\hat{\xi}\cdot\check{x}}-1}{\tilde{D}_{0}(0_{2},\hat{\xi})+\kappa|\hat{\xi}|^{3}}e^{i\hat{\xi}\cdot\check{z}}\text{d}\hat{\xi}\,\Delta F_{0}(\check{z})\text{d}\check{z}.\end{split}$
+(8.11) $\begin{split}\frac{1}{(2\pi)^{2}}\int_{|\check{z}|\leq A_{1}}\int_{\mathbb{R}^{2}}\frac{e^{-i\hat{\xi}\cdot\check{x}}-1}{\tilde{D}_{0}(0_{2},\hat{\xi})+\kappa|\hat{\xi}|^{3}}e^{i\hat{\xi}\cdot\check{z}}\text{d}\hat{\xi}\,\Delta F_{0}(\check{z})\text{d}\check{z}.\end{split}$
+Since $A_{1}\gg 1$ can be arbitrarily large and $\tilde{\vartheta}_{A_{1}}(\hat{\xi})\to\delta(\hat{\xi})$ as $A_{1}\to\infty$, (8.10), (8.11), and Lemma 6.1 imply
+(8.12) $\begin{split}\lim_{\epsilon\to 0}\Delta\mathcal{G}_{\epsilon}(\check{x})=&\frac{1}{(2\pi)^{2}}\int_{\mathbb{R}^{2}}\int_{\mathbb{R}^{2}}\frac{e^{-i\hat{\xi}\cdot\check{x}}-1}{\tilde{D}_{0}(0_{2},\hat{\xi})+\kappa|\hat{\xi}|^{3}}e^{i\hat{\xi}\cdot\check{z}}\text{d}\hat{\xi}\,\Delta F_{0}(\check{z})\text{d}\check{z}.\end{split}$
 8.2. Evaluation of the integral in (8.12)
-The last integral simplifies using that $\Delta F_{0}(\check{x})=\varphi(\Phi_{x}^{\prime}\check{x})$ for some $\varphi$ (see (7.18)). Simple transformations give
-(8.14) $\begin{split}\Delta G_{0}(\check{x}):=&\lim_{\epsilon\to 0}\Delta G_{\epsilon}(\check{x})\\ =&\frac{1}{2\pi}\int_{\mathbb{R}}\bigg{[}\int_{\mathbb{R}}\frac{e^{-i\lambda\vec{\Theta}\cdot\check{x}}-1}{\tilde{D}_{0}(0_{2},\vec{\Theta}_{0})+\kappa|\lambda|^{3}}e^{i\lambda t}\text{d}\lambda\bigg{]}\varphi(|\Phi_{x}^{\prime}|t)\text{d}t,\\ \varphi(t):=&C\int_{0}^{t}h_{0}(-s)\text{d}s,\ \vec{\Theta}_{0}:=\Phi_{x}^{\prime}/|\Phi_{x}^{\prime}|,\end{split}$
+The last integral simplifies using that $\Delta F_{0}(\check{x})=\varphi(\text{d}_{x}\Phi\cdot\check{x})$ for some $\varphi$ (see (7.20)). Simple transformations give
+(8.13) $\begin{split}\Delta\mathcal{G}_{0}(\check{x}):=&\lim_{\epsilon\to 0}\Delta\mathcal{G}_{\epsilon}(\check{x})\\ =&\frac{1}{2\pi}\int_{\mathbb{R}}\bigg{[}\int_{\mathbb{R}}\frac{e^{-i\lambda\vec{\Theta}\cdot\check{x}}-1}{\tilde{D}_{0}(0_{2},\vec{\Theta}_{0})+\kappa|\lambda|^{3}}e^{i\lambda t}\text{d}\lambda\bigg{]}\varphi(|\text{d}_{x}\Phi|t)\text{d}t,\\ \varphi(t):=&C\int_{0}^{t}h_{0}(-s)\text{d}s,\ \vec{\Theta}_{0}:=\text{d}_{x}\Phi/|\text{d}_{x}\Phi|,\end{split}$
 where $C$ is defined in (7.20). Denote
 (8.14) $\tilde{R}(\lambda):=\big{[}\tilde{D}_{0}(0_{2},\vec{\Theta}_{0})+\kappa|\lambda|^{3}\big{]}^{-1},\ R:=\mathcal{F}^{-1}\tilde{R}.$
-Since $\tilde{R}(\lambda)$ is $C^{2}$ near $\lambda=0$, $R(t)$ is absolutely integrable. Combine (7.18), (8.14), and (8.15) to obtain
-(8.16) $\begin{split}&\Delta G_{0}(\check{x})=C\int_{\mathbb{R}}(R(r-t)-R(-t))\int_{-\infty}^{|\Phi_{x}^{\prime}|t}h_{0}(-s)\text{d}s\text{d}t,\ r:=\vec{\Theta}_{0}\cdot\check{x}.\end{split}$
-We have extended the integral with respect to $s$ to $(-\infty,|\Phi_{x}^{\prime}|t]$ because $\int(R(r-t)-R(-t))\text{d}t\equiv 0$ for all $r$. Simple transformations give
-(8.17) $\begin{split}&\Delta G_{0}(\check{x})=(C/|\Phi_{x}^{\prime}|)\int_{\mathbb{R}}h_{0}(s)\int_{s}^{s+\Phi_{x}^{\prime}\check{x}}R(t/|\Phi_{x}^{\prime}|)\text{d}t\text{d}s.\end{split}$
+Since $\tilde{R}(\lambda)$ is $C^{2}$ near $\lambda=0$ (and smooth otherwise), $R(t)$ is absolutely integrable. Combine (7.20), (8.13), and (8.14) to obtain
+(8.15) $\begin{split}&\Delta\mathcal{G}_{0}(\check{x})=C\int_{\mathbb{R}}(R(r-t)-R(-t))\int_{-\infty}^{|\text{d}_{x}\Phi|t}h_{0}(-s)\text{d}s\text{d}t,\ r:=\vec{\Theta}_{0}\cdot\check{x}.\end{split}$
+We have extended the integral with respect to $s$ to $(-\infty,|\text{d}_{x}\Phi|t)$ because $\int(R(r-t)-R(-t))\text{d}t\equiv 0$ for all $r$. Transforming (LABEL:rhs_lim_4) further gives
+(8.16) $\begin{split}&\Delta\mathcal{G}_{0}(\check{x})=(C/|\text{d}_{x}\Phi|)\int_{\mathbb{R}}h_{0}(s)\int_{s}^{s+\text{d}_{x}\Phi\cdot\check{x}}R(t/|\text{d}_{x}\Phi|)\text{d}t\text{d}s.\end{split}$
 For convenience, we collect here all the constants that will be used in subsequent calculations.
-(8.18) $\tilde{D}(0_{2},\Phi_{x}^{\prime})=2\pi\frac{W^{2}|\Phi_{x}^{\prime}|}{|\Delta_{\Phi}|}\quad$ (by (C.5), (C.6))
-(8.19) $K(x_{0},\alpha_{0})=|\Phi_{x}^{\prime}|W\quad$ (by (6.4))
-(8.20) $Q_{\alpha\alpha}^{\prime\prime}(x_{0},\alpha_{0})=\frac{|\Delta_{\Phi}||x_{\alpha}^{\prime}|}{|\Phi_{x}^{\prime}|}\quad$ (by (D.21), (D.22))
-(8.21) $\psi(y_{0})=\frac{|H_{x}^{\prime}|}{|\Phi_{x}^{\prime}|}\quad$ (by (4.7))
+(8.17) $\tilde{D}(0_{2},\text{d}_{x}\Phi)=2\pi\frac{W^{2}|\text{d}_{x}\Phi|}{|\Delta_{\Phi}|}\quad$ (by (C.7), (C.8))
+(8.18) $K(x_{0},\alpha_{0})=|\text{d}_{x}\Phi|W\quad$ (by (6.3))
+(8.19) $Q_{\alpha\alpha}^{\prime\prime}(x_{0},\alpha_{0})=\frac{|\Delta_{\Phi}||x_{\alpha}^{\prime}|}{|\text{d}_{x}\Phi|}\quad$ (by (D.23), (D.24))
+(8.20) $\psi(y_{0})=\frac{|\text{d}_{x}H|}{|\text{d}_{x}\Phi|}\quad$ (by (4.7))
 […]
 For simplicity, we dropped all the arguments on the right.
-Let $\bar{h}_{0}$ be defined the same way as $h_{0}$ in (7.16), but without the factor $K(x_{0},\alpha_{0})$. Then, by the above formulas and (8.15):
-(8.23) $\begin{split}&\Delta G_{0}(\check{x})=\tilde{f}_{0}\Upsilon\big{(}\Phi_{x}^{\prime}\check{x}\big{)},\\ &\Upsilon(r)=2\pi\frac{W^{2}}{|\Delta_{\Phi}|}\int_{\mathbb{R}}\bar{h}_{0}(u)\int_{u}^{u+r}\mathcal{F}^{-1}\bigg{(}\frac{1}{2\pi\frac{W^{2}|\Phi_{x}^{\prime}|}{|\Delta_{\Phi}|}+\kappa|\lambda|^{3}}\bigg{)}(t/|\Phi_{x}^{\prime}|)\text{d}t\text{d}u.\end{split}$
+Let $\bar{h}_{0}$ be defined the same way as $h_{0}$ in (7.18), but without the factor $K(x_{0},\alpha_{0})$. By (8.14) and (LABEL:rhs_lim_4_v2)–(8.21):
+(8.22) $\begin{split}&\Delta\mathcal{G}_{0}(\check{x})=\Delta f\Upsilon\big{(}\text{d}_{x}\Phi\cdot\check{x}\big{)},\\ &\Upsilon(r)=2\pi\frac{W^{2}}{|\Delta_{\Phi}|}\int_{\mathbb{R}}\bar{h}_{0}(u)\int_{u}^{u+r}\mathcal{F}^{-1}\bigg{(}\frac{1}{2\pi\frac{W^{2}|\text{d}_{x}\Phi|}{|\Delta_{\Phi}|}+\kappa|\lambda|^{3}}\bigg{)}(t/|\text{d}_{x}\Phi|)\text{d}t\text{d}u.\end{split}$
 A simple transformation yields:
-(8.24) $\begin{split}\Upsilon(r)=&\frac{1}{|\Phi_{x}^{\prime}|}\int_{\mathbb{R}}\bar{h}_{0}(u)\int_{u}^{u+r}\mathcal{F}^{-1}\bigg{(}\frac{1}{1+\kappa C_{1}|\lambda|^{3}}\bigg{)}(t/|\Phi_{x}^{\prime}|)\text{d}t\text{d}u,\\ C_{1}=&\frac{|\Delta_{\Phi}|}{2\pi W^{2}|\Phi_{x}^{\prime}|}.\end{split}$
+(8.23) $\begin{split}\Upsilon(r)=&\frac{1}{|\text{d}_{x}\Phi|}\int_{\mathbb{R}}\bar{h}_{0}(u)\int_{u}^{u+r}\mathcal{F}^{-1}\bigg{(}\frac{1}{1+\kappa C_{1}|\lambda|^{3}}\bigg{)}(t/|\text{d}_{x}\Phi|)\text{d}t\text{d}u,\\ C_{1}=&\frac{|\Delta_{\Phi}|}{2\pi W^{2}|\text{d}_{x}\Phi|}.\end{split}$
 Further simple transformations prove (3.18), (3.19). The rest of the claims follow from the next lemma, which is proven in Appendix F.2.
@@
 One has $\Upsilon(0)=0$ and $\Upsilon(\pm\infty)=\pm 1/2$.
+Truncated-view artifacts are reasonably low. This will likely allow for successful image segmentation and motion estimation.
 9. Redundant data. Proof of Theorem 3.7
-If there are multiple critical points $y_{l}$ from which the singularity at $(x_{0},\xi_{0})$ is visible, the analog of (5.7) becomes
-(9.1) $\begin{split}&G_{\epsilon}=\text{Op}\big{(}\big{[}\textstyle\sum_{l}\tilde{D}_{l}(x,\xi/|\xi|)+\kappa|\epsilon\xi|^{3}\big{]}^{-1}\big{)}F_{\epsilon},\ x\in\mathcal{U}_{0},\\ &\tilde{D}_{l}(x_{0},\vec{\Theta}_{0}):=2\pi\frac{W^{2}(x_{0},y_{l})|\Phi_{x}^{\prime}(x_{0},\alpha_{l})|}{|\Delta_{\Phi}(x_{0},\alpha_{l})|}.\end{split}$
+Suppose the GRT data are redundant and there are multiple points $y_{l}$ from which the singularity at $(x_{0},\xi_{0})$ is visible.
+This means that for all $(x,\xi)$ in a conic neighborhood of $(x_{0},\xi_{0})$ there are multiple local solutions $\nu_{l}(x,\xi)$ and $\alpha_{l}(x,\xi)$ to the equation $\xi=-\nu\text{d}_{x}\Phi(x,\alpha)$.
+The local uniqueness and smoothness of the solutions follows from Assumption 3.2(1).
+By Assumption 3.3(2), their number is uniformly bounded.
+Set $p_{l}(x,\xi)=\Phi(x,\alpha_{l}(x,\xi))$ and $y_{l}=(\alpha_{l},p_{l})$.
+The principal symbol of $\mathcal{R}^{*}\mathcal{R}$ is given by (cf.
+(C.6))
+(9.1) $\begin{split}\textstyle\sum_{l}\frac{2\pi}{|\nu_{l}|}\frac{W^{2}(x,y_{l})}{|\Delta_{\Phi}(x,\alpha)|},\ y_{l}=y(x,\xi),\nu_{l}=\nu(x,\xi).\end{split}$
+Similarly to (C.8), define
+(9.2) $\begin{split}\tilde{D}_{0}(x,\xi):=&2\pi\textstyle\sum_{l}\frac{|\xi|}{|\nu_{l}|}\frac{W^{2}(x,y_{l})}{|\Delta_{\Phi}(x,\alpha_{l})|},\ y_{l}=y_{l}(x,\xi),\nu_{l}=\nu_{l}(x,\xi),\\ (x,\xi)\in&\,\mathcal{U}\times(\mathbb{R}^{2}\setminus 0_{2}).\end{split}$
+The analog of (5.7) becomes
+(9.3) $\begin{split}&\mathcal{G}_{\epsilon}=\text{Op}\big{(}\big{[}\textstyle\sum_{l}\tilde{D}_{l}(x,\xi/|\xi|)+\kappa|\epsilon\xi|^{3}\big{]}^{-1}\big{)}F_{\epsilon},\ x\in\mathcal{U}_{0},\\ &\tilde{D}_{l}(x_{0},\vec{\Theta}_{0}):=2\pi\frac{W^{2}(x_{0},y_{l})|\text{d}_{x}\Phi(x_{0},\alpha_{l})|}{|\Delta_{\Phi}(x_{0},\alpha_{l})|}.\end{split}$
 Here we have used that
-(9.2) $\vec{\Theta}_{0}=\xi_{0}/|\xi_{0}|=H_{x}^{\prime}(x_{0})/|H_{x}^{\prime}(x_{0})|=\Phi_{x}^{\prime}(x_{0},\alpha_{l})/|\Phi_{x}^{\prime}(x_{0},\alpha_{l})|$
-for all $l$, cf. the convention about the directions of $H_{x}^{\prime}$ and $\Phi_{x}^{\prime}$ stated above assumption 3.3. Therefore, we get from (8.15) that the analog of the function $R$ is
+(9.4) $\vec{\Theta}_{0}=\xi_{0}/|\xi_{0}|=\text{d}_{x}H(x_{0})/|\text{d}_{x}H(x_{0})|=\text{d}_{x}\Phi(x_{0},\alpha_{l})/|\text{d}_{x}\Phi(x_{0},\alpha_{l})|$
+for all $l$, cf. the convention about the directions of $\text{d}_{x}H$ and $\text{d}_{x}\Phi$ stated above Assumption 3.3. Therefore, we get from (8.14) that the analog of the function $R$ is
 (9.5) $\tilde{R}(\lambda):=\big{[}\textstyle\sum_{l}\tilde{D}_{l}(x_{0},\vec{\Theta}_{0})+\kappa|\lambda|^{3}\big{]}^{-1},\ R:=\mathcal{F}^{-1}\tilde{R}.$
-The analog (LABEL:rhs_lim_5) becomes
-(9.4) $\begin{split}&\Delta G_{0}(\check{x})=\tilde{f}_{0}\textstyle\sum_{l}\Upsilon_{l}\big{(}\Phi_{x}^{\prime}(x_{0},\alpha_{l})\check{x}\big{)},\\ &\Upsilon_{l}(r)=2\pi\frac{W^{2}(x_{0},y_{l})}{|\Delta_{\Phi}(x_{0},\alpha_{l})|}\int_{\mathbb{R}}\bar{h}_{l}(u)\\ &\hskip 28.45274pt\times\int_{u}^{u+r}\mathcal{F}^{-1}\left(\big{[}\textstyle\sum_{l}\tilde{D}_{l}(x_{0},\vec{\Theta}_{0})+\kappa|\lambda|^{3}\big{]}^{-1}\right)\bigg{(}\frac{t}{|\Phi_{x}^{\prime}(x_{0},\alpha_{l})|}\bigg{)}\text{d}t\text{d}u,\\ &\bar{h}_{l}(u):=\int_{\mathbb{R}}\varphi_{\alpha}(s)\varphi_{p}(\mu s\Phi_{\alpha}^{\prime}(x_{0},\alpha_{l})+u)\text{d}s.\end{split}$
+Using (4.13) and (6.8), we see that $\Delta\mathcal{G}_{0}(\check{x})$ equals to the sum of the terms analogous to (LABEL:rhs_lim_5):
+(9.6) $\begin{split}&\Delta\mathcal{G}_{0}(\check{x})=\Delta f\textstyle\sum_{l}\Upsilon_{l}\big{(}\text{d}_{x}\Phi(x_{0},\alpha_{l})\cdot\check{x}\big{)},\\ &\Upsilon_{l}(r)=2\pi\frac{W^{2}(x_{0},y_{l})}{|\Delta_{\Phi}(x_{0},\alpha_{l})|}\int_{\mathbb{R}}\bar{h}_{l}(u)\\ &\hskip 28.45274pt\times\int_{u}^{u+r}\mathcal{F}^{-1}\left(\big{[}\textstyle\sum_{l}\tilde{D}_{l}(x_{0},\vec{\Theta}_{0})+\kappa|\lambda|^{3}\big{]}^{-1}\right)\bigg{(}\frac{t}{|\text{d}_{x}\Phi(x_{0},\alpha_{l})|}\bigg{)}\text{d}t\text{d}u,\\ &\bar{h}_{l}(u):=\int_{\mathbb{R}}\varphi_{\alpha}(s)\varphi_{p}(\mu s\Phi_{\alpha}^{\prime}(x_{0},\alpha_{l})+u)\text{d}s.\end{split}$
 Simplifying we get similarly to (LABEL:rhs_lim_5), (8.23)
-(9.5) $\begin{split}&\Delta G_{0}(\check{x})=\tilde{f}_{0}\frac{\textstyle\sum_{l}\rho_{l}\Upsilon_{l}\big{(}\Phi_{x}^{\prime}(x_{0},\alpha_{l})\check{x}/|\Phi_{x}^{\prime}(x_{0},\alpha_{l})|\big{)}}{\sum_{l}\rho_{l}},\\ &\Upsilon_{l}(r)=\int_{\mathbb{R}}\bar{h}_{0}(u)\int_{u_{l}}^{u_{l}+r}\mathcal{F}^{-1}\bigg{(}\frac{1}{1+\kappa C_{1}|\lambda|^{3}}\bigg{)}(t)\text{d}t\text{d}u,\\ &\rho_{l}:=\frac{W^{2}(x_{0},y_{l})|\Phi_{x}^{\prime}(x_{0},\alpha_{l})|}{|\Delta_{\Phi}(x_{0},\alpha_{l})|},\ u_{l}:=\frac{u}{|\Phi_{x}^{\prime}(x_{0},\alpha_{l})|},\ C_{1}=\frac{1}{2\pi\textstyle\sum_{l}\rho_{l}}.\end{split}$
+(9.7) $\begin{split}&\Delta\mathcal{G}_{0}(\check{x})=\Delta f\frac{\textstyle\sum_{l}\nu_{l}\Upsilon_{l}\big{(}\text{d}_{x}\Phi(x_{0},\alpha_{l})\cdot\check{x}/|\text{d}_{x}\Phi(x_{0},\alpha_{l})|\big{)}}{\sum_{l}\nu_{l}},\\ &\Upsilon_{l}(r)=\int_{\mathbb{R}}\bar{h}_{0}(u)\int_{u_{l}}^{u_{l}+r}\mathcal{F}^{-1}\bigg{(}\frac{1}{1+\kappa C_{1}|\lambda|^{3}}\bigg{)}(t)\text{d}t\text{d}u,\\ &\nu_{l}:=\frac{W^{2}(x_{0},y_{l})|\text{d}_{x}\Phi(x_{0},\alpha_{l})|}{|\Delta_{\Phi}(x_{0},\alpha_{l})|},\ u_{l}:=\frac{u}{|\text{d}_{x}\Phi(x_{0},\alpha_{l})|},\ C_{1}=\frac{1}{2\pi\textstyle\sum_{l}\nu_{l}}.\end{split}$
 Using (9.4) finishes the proof.
@@
 We choose $f$ to be the characteristic function of the disk centered at $x_{c}=(1,1)$ with radius $r=2$.
-Thus, $\mathcal{S}=\{x\in\mathbb{R}^{2}:\|x-x_{c}\|=r\}$.
-The center of a local region of interest (ROI) is $x_{0}=x_{c}+r\vec{\beta}_{0}$, where $\beta_{0}=0.17\pi$.
+Thus, $\mathcal{S}=\{x\in\mathbb{R}^{2}:|x-x_{c}|=r\}$.
+The center of a local region of interest (ROI), i.e.
+the point on the boundary $x_{0}\in\mathcal{S}$, is $x_{0}=x_{c}+r\vec{\beta}_{0}$, where $\beta_{0}=-0.17\pi$.
 The GRT integrates over circles $\mathcal{S}_{y}$, $y=(\alpha,\rho)$, with various radii $\rho>0$ and centers $R\vec{\alpha}$, $\alpha\in[0,2\pi)$, where $R=10$.
-Thus, $\rho=\Phi(x,\alpha)=\|x-R\vec{\alpha}\|$ and
-(10.1) $\begin{split}&\Phi_{x}^{\prime}(x,\alpha)=\frac{x-R\vec{\alpha}}{\|x-R\vec{\alpha}\|}=:\vec{\Theta},\ \Phi_{\alpha}^{\prime}(x,\alpha)=R\vec{\Theta}\cdot\vec{\alpha}^{\perp},\\ &\Delta_{\Phi}(x,\alpha)=\det\begin{pmatrix}\vec{\Theta}&-\frac{R}{\|x-R\vec{\alpha}\|}\vec{\Theta}^{\perp}+c\vec{\Theta}\end{pmatrix}=-\frac{R}{\|x-R\vec{\alpha}\|}.\end{split}$
-Here $\alpha^{\perp}=(-\sin\alpha,\cos\alpha)$, and $\Theta^{\perp}$ is defined similarly.
-The continuous data corresponds to $\alpha\in[0,2\pi)$ and $\rho\in(0,2R)$.
-The global reconstruction region is the square $\mathcal{U}_{b}:=\{(x_{1},x_{2})\in\mathbb{R}^{2}:|x_{1}|<R_{rec},\ |x_{2}|<R_{rec}\}$, $R_{rec}=3.7$.
+The integration weight is $W\equiv 1$.
+Thus, $\rho=\Phi(x,\alpha)=|x-R\vec{\alpha}|$ and
+(10.1) $\begin{split}&\text{d}_{x}\Phi(x,\alpha)=\frac{x-R\vec{\alpha}}{|x-R\vec{\alpha}|}=:\vec{\Theta},\ \Phi_{\alpha}^{\prime}(x,\alpha)=R\vec{\Theta}\cdot\vec{\alpha}^{\perp},\\ &\text{d}_{x}\Phi_{\alpha}^{\prime}(x,\alpha)=-\frac{R}{|x-R\vec{\alpha}|}\vec{\alpha}^{\perp}+c\vec{\Theta},\\ &\Delta_{\Phi}(x,\alpha)=\det\begin{pmatrix}\vec{\Theta}\\ -\frac{R}{|x-R\vec{\alpha}|}\vec{\alpha}^{\perp}+c\vec{\Theta}\end{pmatrix}=\frac{R\vec{\alpha}\cdot\vec{\Theta}}{|x-R\vec{\alpha}|}\not=0,\ |x|<R.\end{split}$
+Here $\vec{\alpha}^{\perp}=(-\sin\alpha,\cos\alpha)$, and $\vec{\Theta}^{\perp}$ is defined similarly.
+The continuous data corresponds to $\alpha\in\mathcal{I}_{\alpha}:=[0,2\pi)\sim S^{1}$ and $\rho\in\mathcal{I}_{p}:=(0,2R)$.
+The image domain is $\mathcal{U}=\{|x|<R\}$.
+The support of the object is contained in the square reconstruction region $\mathcal{U}_{b}:=\{(x_{1},x_{2})\in\mathbb{R}^{2}:|x_{1}|<R_{rec},\ |x_{2}|<R_{rec}\}$, $R_{rec}=3.7$.
 This corresponds to the set $\mathcal{U}_{b}$ used in (3.17).
+Further (cf. (3.22)),
+(10.2) $\begin{split}&\nu_{l}=|\Delta_{\Phi}(x_{0},\alpha_{l})|^{-1},\ u_{l}=u,\ l=1,2.\end{split}$
+This implies that the DTB function given by the fraction in (3.21) becomes
+(10.3) $\frac{\textstyle\sum_{l}\nu_{l}\Upsilon_{l}(r)}{\textstyle\sum_{l}\nu_{l}}=\int_{\mathbb{R}}\frac{\textstyle\sum_{l}\nu_{l}h_{l}(u)}{\textstyle\sum_{l}\nu_{l}}\int_{u}^{u+r}R(t)\text{d}t\text{d}u.$
+Let us now make sure that the circular GRT used in this section satisfies Assumptions 3.1 and 3.2.
+Clearly, Assumption 3.1 is satisfied.
+We next check Assumption 3.2.
+By the first two lines in (10.1), Assumption 3.2(1) is satisfied since $\vec{\alpha}\cdot\vec{\Theta}\not=0$ when $|x|<R$.
+Next, suppose $|x_{1}-R\vec{\alpha}|=|x_{2}-R\vec{\alpha}|$ for some $\vec{\alpha}$ and $x_{1}\not=x_{2}$, $|x_{1}|,|x_{2}|<R$.
+The first line in (10.1) implies $\Phi_{\alpha}^{\prime}(x_{1},\alpha)\not=\Phi_{\alpha}^{\prime}(x_{2},\alpha)$, therefore Assumption 3.2(2) is satisfied.
+For any $(x_{0},\vec{\Theta}_{0})$, $|x_{0}|<R$, there are two circles $\mathcal{S}_{y}$ in the data, which contain $x_{0}$ and are orthogonal to $\vec{\Theta}_{0}$ at $x_{0}$.
+Thus, the conormal bundles of all the $\mathcal{S}_{y}$ cover $T^{*}\mathcal{U}_{b}$ and Assumption 3.2(3) is satisfied.
+Since $W\equiv 1$, Assumption 3.2(4) trivially holds.
+By [6, Corollary 5], $\mathcal{R}$ is injective.
+As stated at the end of Section 3.1, together with the visibility of every singularity and the fact that $\mathcal{R}$ and $\mathcal{R}^{*}$ are FIOs of order $-1/2$, this ensures that Assumption 3.2(5) is satisfied.
 The discrete data are given at the points
 […]
-Clearly, for each $x\in\mathcal{S}$ there are two points, $y_{l}=(\alpha_{l},\rho_{l})$, $l=1,2$, such that $\mathcal{S}_{y_{l}}$ is tangent to $\mathcal{S}$ at $x$.
+As mentioned above, for each $x\in\mathcal{S}$ there are two points, $y_{l}=(\alpha_{l},\rho_{l})$, $l=1,2$, such that $\mathcal{S}_{y_{l}}$ is tangent to $\mathcal{S}$ at $x$.
 To find $y_{1,2}$ for the selected $x_{0}$ we first solve $\|x_{c}+t\vec{\beta}_{0}\|=R$.
 This gives two values $t_{1},t_{2}$, $t_{2}<0<t_{1}$.
-Then $\alpha_{l}$ are determined from $x_{c}+t_{l}\vec{\beta}_{0}=R\vec{\alpha}_{l}$, and $\rho_{l}=\|x_{0}-R\vec{\alpha}_{l}\|$, $l=1,2$.
-In view of (10.1), in numerical computations we use $\|x_{0}-R\vec{\alpha}_{1}\|=t_{1}-r$ and $\|x_{0}-R\vec{\alpha}_{2}\|=-t_{2}+r$.
-The regularization parameter is set at $\kappa=0.1$.
+Then $\alpha_{l}$ are determined from $x_{c}+t_{l}\vec{\beta}_{0}=R\vec{\alpha}_{l}$, and $\rho_{l}=|x_{0}-R\vec{\alpha}_{l}|$, $l=1,2$.
+In view of (10.1), in numerical computations we use $|x_{0}-R\vec{\alpha}_{1}|=t_{1}-r$ and $|x_{0}-R\vec{\alpha}_{2}|=-t_{2}+r$.
+By a simple geometric argument (cf.
+(10.1)),
+(10.5) $-R\vec{\alpha}_{l}\cdot\vec{\Theta}_{l}=(t_{1}-t_{2})/2,\ l=1,2.$
+The regularization parameter is set at $\kappa=0.5$.
 In order to approximate the continuous GRT $\mathcal{R}$ and its adjoint $\mathcal{R}^{*}$, the reconstruction is performed on a dense $801\times 801$ grid covering the same square region $\mathcal{U}_{b}$, and the interpolated data $g_{\epsilon}$ are assumed to be given on a dense grid of the kind (10.4) with $N_{\alpha}^{\prime}=800$, $N_{\rho}^{\prime}=1201$.
@@
 The kernel is a piecewise-cubic polynomial with continuous $\varphi,\varphi^{\prime}$ and bounded $\varphi^{\prime\prime}$, so $\varphi\in C_{0}^{2}(\mathbb{R})$.
-Minimization of the functional (3.16) is performed using gradient descent.
-Results of reconstructions are shown in Figures 2 and 3.
-Reconstruction of the entire region $\mathcal{U}_{b}$ is shown in Figure 2.
+The minimization of the functional (3.17) is performed using gradient descent.
+The stopping criterion is met when the $L^{\infty}$-norm of the image update is less than $10^{-6}$ for three consecutive iterations.
+The algorithm converged in 214 iterations.
+The reconstruction results are shown in Figures 2 and 3.
+The reconstruction of the entire region $\mathcal{U}_{b}$ is shown in Figure 2.
 A profile of the reconstruction through the center of the ball, $x_{c}$, is shown on the right.
@@
 The right panel of Figure 3 shows the reconstructed (green) and predicted (blue) profiles.
-The latter is computed by using (10.1) in (3.20).
-The slight discrepancy in the interval $[1,2]$ is due to random numerical errors, which are clearly visible in the right panel of Figure 2.
+The latter is computed by using (10.1) in (3.21), (3.22).
 The plot in Figure 3 corresponds to the right jump in Figure 2.
@@
 Appendix A Proof of Lemma 3.5
-It is easy to see that $\Psi:H_{c}^{1}(\mathcal{U}_{b})\to\mathbb{R}$ is strongly convex [7, Definition 1.1.48] and [41, Section 2.3] for each $\epsilon>0$.
-Strong convexity means that there exists $c>0$ such that
+As the following shows, $\Psi:H_{0}^{1}(\mathcal{U}_{b})\to\mathbb{R}$ is strongly convex [9, Definition 1.1.48] and [45, Section 2.3] for each $\epsilon>0$.
+Indeed, strong convexity means that there exists $c>0$ such that
 (A.1) $\Psi(\lambda f_{1}+(1-\lambda)f_{2})\leq\lambda\Psi(f_{1})+(1-\lambda)\Psi(f_{2})-c\lambda(1-\lambda)\|f_{2}-f_{1}\|_{H^{1}(\mathcal{U}_{b})}^{2}$
-for any $f_{1,2}\in H_{c}^{1}(\mathcal{U}_{b})$ and $\lambda\in(0,1)$. By an easy calculation,
-(A.2) $\begin{split}\|\lambda\nabla f_{1}+(1-\lambda)\nabla f_{2}\|_{L^{2}(\mathcal{U})}^{2}\leq&\lambda\|\nabla f_{1}\|_{L^{2}(\mathcal{U})}^{2}+(1-\lambda)\|\nabla f_{2}\|_{L^{2}(\mathcal{U})}^{2}\\ &-2\lambda(1-\lambda)\|\nabla(f_{1}-f_{2})\|_{L^{2}(\mathcal{U})}^{2}\end{split}$
+for any $f_{1,2}\in H_{0}^{1}(\mathcal{U}_{b})$ and $\lambda\in(0,1)$. By an easy calculation,
+(A.2) $\begin{split}\|\lambda\partial_{x}f_{1}+(1-\lambda)\partial_{x}f_{2}\|_{L^{2}(\mathcal{U})}^{2}\leq&\lambda\|\partial_{x}f_{1}\|_{L^{2}(\mathcal{U})}^{2}+(1-\lambda)\|\partial_{x}f_{2}\|_{L^{2}(\mathcal{U})}^{2}\\ &-\lambda(1-\lambda)\|\partial_{x}(f_{1}-f_{2})\|_{L^{2}(\mathcal{U})}^{2}\end{split}$
 and
-(A.3) $\begin{split}\|\mathcal{R}(\lambda f_{1}&+(1-\lambda)f_{2})-g_{\epsilon}\|_{L^{2}(\mathcal{V})}^{2}\\ \leq&\lambda\|\mathcal{R}f_{1}-g_{\epsilon}\|_{L^{2}(\mathcal{V})}^{2}+(1-\lambda)\|\mathcal{R}f_{2}-g_{\epsilon}\|_{L^{2}(\mathcal{V})}^{2}\\ &-2\lambda(1-\lambda)\|\mathcal{R}(f_{1}-f_{2})\|_{L^{2}(\mathcal{V})}^{2}.\end{split}$
+(A.3) $\begin{split}\|\mathcal{R}(\lambda f_{1}&+(1-\lambda)f_{2})-g_{\epsilon}\|_{L^{2}(\mathcal{V})}^{2}\\ \leq&\lambda\|\mathcal{R}f_{1}-g_{\epsilon}\|_{L^{2}(\mathcal{V})}^{2}+(1-\lambda)\|\mathcal{R}f_{2}-g_{\epsilon}\|_{L^{2}(\mathcal{V})}^{2}\\ &-\lambda(1-\lambda)\|\mathcal{R}(f_{1}-f_{2})\|_{L^{2}(\mathcal{V})}^{2}.\end{split}$
 Therefore
-(A.4) $\begin{split}\Psi(\lambda f_{1}&+(1-\lambda)f_{2})\leq\lambda\Psi(f_{1})+(1-\lambda)\Psi(f_{2})\\ &-c\lambda(1-\lambda)\big{(}\|\mathcal{R}(f_{1}-f_{2})\|_{L^{2}(\mathcal{V})}^{2}+\|\nabla(f_{2}-f_{1})\|_{L^{2}(\mathcal{U})}\big{)}.\end{split}$
+(A.4) $\begin{split}\Psi(\lambda f_{1}&+(1-\lambda)f_{2})\leq\lambda\Psi(f_{1})+(1-\lambda)\Psi(f_{2})\\ &-\lambda(1-\lambda)\big{(}\|\mathcal{R}(f_{1}-f_{2})\|_{L^{2}(\mathcal{V})}^{2}+\|\partial_{x}(f_{2}-f_{1})\|_{L^{2}(\mathcal{U})}\big{)}.\end{split}$
 By (3.6),
@@
 Also
-(A.6) $\|f\|_{H^{1}(\mathcal{U})}\leq c\big{(}\|f\|_{H^{-1/2}(\mathcal{U})}+\|\nabla f\|_{L^{2}(\mathcal{U})}\big{)},\ f\in C_{0}^{\infty}(\mathcal{U}_{b}).$
+(A.6) $\|f\|_{H^{1}(\mathcal{U})}\leq c\big{(}\|f\|_{H^{-1/2}(\mathcal{U})}+\|\partial_{x}f\|_{L^{2}(\mathcal{U})}\big{)},\ f\in C_{0}^{\infty}(\mathcal{U}_{b}).$
 Applying (A.5) and (A.6) to (A.4) proves (A.1).
-Moreover, the functional is proper (its domain is not empty) and continuous $H_{c}^{1}(\mathcal{U}_{b})\to\mathbb{R}$.
-Thus the solution to (3.16) exists and is unique for each $\epsilon>0$ [7, Theorem 1.3.1], [41, Corollary 2.20].
+Moreover, the functional is proper (its domain is not empty) and continuous $H_{0}^{1}(\mathcal{U}_{b})\to\mathbb{R}$.
+Thus the solution to (3.17) exists and is unique for each $\epsilon>0$ [9, Theorem 1.3.1], [45, Corollary 2.20].
 Appendix B Proof of Lemmas 4.1 and 4.2
@@
 Consider the function $H(x_{*}(y))$. Differentiating the first equation in (4.4) with respect to $p$ and then using the second equation gives
-(B.1) $\begin{split}\partial_{p}H(x_{*}(y))=&\nu_{1*}(y)=\frac{|H_{x}^{\prime}(x_{*}(y))|}{|\Phi_{x}^{\prime}(x_{*}(y),\alpha)|}>0.\end{split}$
-Another easy calculation, combined with (B.1) and (3.11), shows
-(B.2) $\begin{split}\det M(y)&=-\big{(}[H_{xx}^{\prime\prime}(x_{*}(y))-\nu_{1*}(y)\Phi_{xx}^{\prime\prime}(x_{*}(y),\alpha)]e,e\big{)}|\Phi_{x}^{\prime}(x,\alpha)|^{2}\\ &=(\varkappa_{\mathcal{S}}(x)-\varkappa_{\mathcal{S}_{y}}(x)){|\Phi_{x}^{\prime}(x,\alpha)|^{2}}{|H_{x}^{\prime}(x)|},\end{split}$
-where $e$ is a unit vector tangent to $\mathcal{S}$ at $x_{*}(y)$. Using assumption 3.3(2) proves the first statement.
-Now we prove the second assertion in (4.5). Denote
-(B.3) $e_{0}:=-\Phi_{x}^{\prime}(x_{*}(y),\alpha)/|\Phi_{x}^{\prime}(x_{*}(y),\alpha)|,\ V:=\begin{pmatrix}e_{0}&e\end{pmatrix}.$
-Clearly, $V$ is a $2\times 2$ orthogonal matrix. By construction,
-(B.4) $M_{1}:=\begin{pmatrix}1&0_{2}^{T}\\ 0_{2}&V^{T}\end{pmatrix}M\begin{pmatrix}1&0_{2}^{T}\\ 0_{2}&V\end{pmatrix}=\begin{pmatrix}0&|\Phi_{x}^{\prime}|&0\\ |\Phi_{x}^{\prime}|&*&*\\ 0&*&([H_{xx}^{\prime\prime}-\nu_{1*}\Phi_{xx}^{\prime\prime}]e,e)\end{pmatrix},$
-$\text{sgn}M=\text{sgn}M_{1}$, and * denote various quantities, whose values are irrelevant. Next we use Haynsworth’s inertia additivity formula [18],
-(B.5) $\begin{split}\text{sgn}M_{1}=&\text{sgn}\bigg{(}\big{(}\big{[}H_{xx}^{\prime\prime}-\nu_{1*}\Phi_{xx}^{\prime\prime}\big{]}e,e\big{)}-\begin{pmatrix}|\Phi_{x}^{\prime}|&0\end{pmatrix}\tilde{M}_{1}^{-1}\begin{pmatrix}|\Phi_{x}^{\prime}|\\ 0\end{pmatrix}\bigg{)}\\ &+\text{sgn}\tilde{M}_{1},\quad\tilde{M}_{1}:=\begin{pmatrix}0&|\Phi_{x}^{\prime}|\\ |\Phi_{x}^{\prime}|&*\end{pmatrix}.\end{split}$
-An easy calculation shows that $\text{sgn}\tilde{M}_{1}=0$ and
-(B.6) $\begin{split}\text{sgn}M_{1}=&\text{sgn}\big{(}\big{[}H_{xx}^{\prime\prime}-\nu_{1*}\Phi_{xx}^{\prime\prime}\big{]}e,e\big{)}=-1.\end{split}$
+(B.1) $\begin{split}\partial_{p}H(x_{*}(y))=&\nu_{1*}(y)=\frac{|\text{d}_{x}H(x_{*}(y))|}{|\text{d}_{x}\Phi(x_{*}(y),\alpha)|}>0.\end{split}$
+The last equality follows from (3.12) and the second equation in (4.4).
+Let $e$ be a unit vector tangent to $\mathcal{S}$ at $x_{*}(y)$. Denote
+(B.2) $e_{0}:=-\partial_{x}\Phi(x_{*}(y),\alpha)/|\partial_{x}\Phi(x_{*}(y),\alpha)|,\ V:=\begin{pmatrix}e_{0}&e\end{pmatrix}.$
+Clearly, $V$ is a $2\times 2$ orthogonal matrix. Recall that $M$ is defined in (4.3). As is easily checked,
+(B.3) $M_{1}:=\begin{pmatrix}1&0_{2}^{T}\\ 0_{2}&V^{T}\end{pmatrix}M\begin{pmatrix}1&0_{2}^{T}\\ 0_{2}&V\end{pmatrix}=\begin{pmatrix}0&|\text{d}_{x}\Phi|&0\\ |\text{d}_{x}\Phi|&a&b\\ 0&b&\text{d}_{x}^{2}[H-\nu_{1*}\Phi](e,e)\end{pmatrix},$
+where $a,b$ are some quantities, whose values are irrelevant. Therefore, $\det M_{1}=\det M$ and $\text{sgn}M=\text{sgn}M_{1}$.
+Using (B.1) and (3.11) gives
+(B.4) $\begin{split}\det M(y)&=-\text{d}_{x}^{2}[H(x_{*}(y))-\nu_{1*}(y)\Phi(x_{*}(y),\alpha)](e,e)|\text{d}_{x}\Phi(x,\alpha)|^{2}\\ &=(\varkappa_{\mathcal{S}}(x)-\varkappa_{\mathcal{S}_{y}}(x)){|\text{d}_{x}\Phi(x,\alpha)|^{2}}{|\text{d}_{x}H(x)|}.\end{split}$
+Assumption 3.3(2) completes the proof of the first statement.
+To prove the second assertion in (4.5) we use Haynsworth’s inertia additivity formula [21],
+(B.5) $\begin{split}\text{sgn}M_{1}=&\text{sgn}\bigg{(}\text{d}_{x}^{2}[H-\nu_{1*}\Phi](e,e)-{\color[rgb]{0,0,0}\begin{pmatrix}0&b\end{pmatrix}}\tilde{M}_{1}^{-1}{\color[rgb]{0,0,0}\begin{pmatrix}0\\ b\end{pmatrix}}\bigg{)}\\ &+\text{sgn}\tilde{M}_{1},\quad\tilde{M}_{1}:=\begin{pmatrix}0&|\text{d}_{x}\Phi|\\ |\text{d}_{x}\Phi|&a\end{pmatrix}.\end{split}$
+An easy calculation shows that
+(B.6) $\text{sgn}\tilde{M}_{1}=0,\quad\begin{pmatrix}0&b\end{pmatrix}\tilde{M}_{1}^{-1}\begin{pmatrix}0\\ b\end{pmatrix}=b^{2}\big{(}\tilde{M}_{1}^{-1}\big{)}_{22}=0.$
+Therefore, arguing as in (B.4) and using Assumption 3.3(2), we conclude that
+(B.7) $\begin{split}\text{sgn}M_{1}=&\text{sgn}\big{(}\text{d}_{x}^{2}[H-\nu_{1*}\Phi](e,e)\big{)}=-1.\end{split}$
 B.2. Proof of Lemma 4.2
-Eq.
-(4.9) implies that $\mathcal{R}f\in I^{-3/2}(\mathcal{V},\Gamma)$ is a conormal (or, more generally, Lagrangian) distribution, see [22, Section 25.1].
-By [22, Proposition 25.1.5], $\tilde{g}(\eta)=O(|\eta|^{-3/2})$, $|\eta|\to\infty$.
-Substituting this into (2.4) and using the equivalence of norms (2.3) and (2.4) gives $g\in C^{1/2}(\mathcal{V})$.
-As mentioned above, assumptions 3.1(1,3) imply that $\text{supp}\,g$ is compact.
-The leading order term in (4.13) is obtained using [14, eq. 26, p. 360]:
-(B.7) $\mathcal{F}((t-i0)^{a})=\frac{2\pi e(-a)}{\Gamma(-a)}\mu_{+}^{-(a+1)}.$
-Using the decay of the lower order terms in $\tilde{\upsilon}$ as $\lambda\to\infty$ (see $\tilde{R}$ in (4.9) and [22, eq. (25.1.4)]), gives $\Delta g\in C_{0}^{3/2}(\mathcal{V})$.
+Eqs.
+(4.9), (4.10) imply that $\hat{f}=\mathcal{R}f\in I^{-3/2}(\mathcal{V},\Gamma)$ is a conormal (or, more generally, Lagrangian) distribution, see [26, Section 25.1].
+By [16, Lemma 2.1 and its proof] with $\mu=-3/2$ and $k=1$, $g=\chi_{V}\hat{f}\in C_{0}^{1/2}(V)$.
+The leading order term in (4.13) is obtained using [17, eq.
+26, p.
+360]:
+(B.8) $\mathcal{F}((t-i0)^{-\sigma})=\frac{2\pi e(\sigma)}{\Gamma(\sigma)}\mu_{+}^{\sigma-1},\ \sigma\in\mathbb{R},$
+where $\mu$ is the Fourier transform variable.
+Using the decay of the lower order terms in $\tilde{\upsilon}$ as $\lambda\to\infty$ (namely, $\tilde{R}$ in (4.10)), gives $\Delta g\in C_{0}^{3/2}(V)$.
+Here we apply [16, Lemma 2.1 and its proof] with $\mu=-3/2$ and $k=1$ to $\partial_{p}\Delta g$ and $\partial_{\alpha}\Delta g$ and then use the first line in (2.3) with $s=3/2$ and $k=1$.
 It remains to prove the properties of the interpolated functions $g_{\epsilon}$ and $\Delta g_{\epsilon}$.
-Let $V\subset\mathbb{R}^{n}$ be any domain.
-Pick any $g\in C_{0}^{s}(V)$ and suppose first $10<s<$ 1.
-Let $\varphi\in C_{0}^{\lceil s\rceil}(\mathbb{R}^{n})$ be an interpolation kernel exact to degree $1$, and $g_{\epsilon}$ be the corresponding interpolated function.
-Denote $y_{h}:=y+h$, $\Delta g(y,h):=g(y_{h})-g(y)$ and similarly for $\Delta g_{\epsilon}(y,h)$.
-We have two equivalent representations
-(B.8) $\begin{split}&\Delta g_{\epsilon}(y,h)-\Delta g(y,h)\\ &=\sum_{j}\varphi\bigg{(}\frac{y_{h}}{\epsilon}-j\bigg{)}\big{(}g(\epsilon j)-g(y_{h})\big{)}-\sum_{j}\varphi\bigg{(}\frac{y}{\epsilon}-j\bigg{)}\big{(}g(\epsilon j)-g(y)\big{)}\\ &=\sum_{j}\bigg{[}\varphi\bigg{(}\frac{y_{h}}{\epsilon}-j\bigg{)}-\varphi\bigg{(}\frac{y}{\epsilon}-j\bigg{)}\bigg{]}\big{(}g(\epsilon j)-g(y_{h})\big{)}\\ &\quad+\sum_{j}\varphi\bigg{(}\frac{y}{\epsilon}-j\bigg{)}\big{(}g(y_{h})-g(y)\big{)}.\end{split}$
+Pick any $g\in C_{0}^{s}(V)$.
+Let $\varphi\in C_{0}^{\lceil s\rceil}(\mathbb{R}^{2})$ be an interpolation kernel exact to degree $1$, and $g_{\epsilon}$ be the corresponding interpolated function.
+Denote $y_{h}:=y+h$.
 We assume here for simplicity that $y_{j}=\epsilon j$, $j\in\mathbb{Z}^{2}$.
-When $h\geq\epsilon$, we use the second line above.
-Since $\varphi$ is compactly supported, the number of terms in each sum is finite.
-Moreover, $|\epsilon j-y_{h}|\leq c\epsilon$ in the first sum and $|\epsilon j-y|\leq c\epsilon$ in the second sum.
-Using that $g\in C_{0}^{s}(V)$, implies
-(B.9) $|\Delta g_{\epsilon}(y,h)-\Delta g(y,h)|\leq ch^{s},\ h\geq\epsilon.$
-When $0<h<\epsilon$, we use the second representation in (B.8). The kernel $\varphi$ is at least as smooth as $g$, so
-(B.10) $\begin{split}\bigg{|}\frac{\varphi\big{(}(y_{h}/\epsilon)-j\big{)}-\varphi\big{(}(y/\epsilon)-j\big{)}}{(h/\epsilon)^{s}}\frac{g(\epsilon j)-g(y_{h})}{\epsilon^{s}}\bigg{|}<c.\end{split}$
-The term on the last line in (B.8) is bounded by $ch^{s}$, and the assertion follows. The fact that $\|g_{\epsilon}\|_{L^{\infty}(V)}<c$ is obvious.
-Suppose now $g\in C_{0}^{s}(V)$ and $21<s<$ 2. As above, $\|g_{\epsilon}\|_{L^{\infty}(V)}<c$. By the exactness of $\varphi$ up to order one, differentiating (3.13) gives
-(B.11) $\begin{split}&g_{\epsilon}^{\prime}(y)=g^{\prime}(y)+\sum_{j}\varphi^{\prime}\bigg{(}\frac{y}{\epsilon}-j\bigg{)}\frac{g(\epsilon j)-(g(y)+(\epsilon j-y)\nabla g(y))}{\epsilon},\end{split}$
-where all the primes stand for the same derivative $\partial_{y_{l}}$, $1\leq l\leq n$, and $\nabla g=(\partial_{y_{1}}g,\dots,\partial_{y_{n}}g)^{T}$.
+Suppose first $10<s<$ 1.
+If $|h|>\epsilon$, it suffices to prove that $|g_{\epsilon}(y)-g(y)|\leq c\epsilon^{s}$.
+Then $|g_{\epsilon}(y_{h})-g_{\epsilon}(y)|\leq c|h|^{s}$ follows from the triangle inequality and the assumption $g\in C_{0}^{s}(V)$.
+By the exactness of $\varphi$,
+(B.9) $\begin{split}g_{\epsilon}(y)-g(y)=\sum_{j}\varphi\bigg{(}\frac{y-\epsilon j}{\epsilon}\bigg{)}\big{(}g(\epsilon j)-g(y)\big{)}.\end{split}$
+Since $\varphi$ is compactly supported, $|\epsilon j-y|\leq c\epsilon$ and the number of terms in the sum is finite, and the desired assertion follows.
+If $|h|<\epsilon$, we have
+(B.10) $\begin{split}&g_{\epsilon}(y_{h})-g_{\epsilon}(y)=\sum_{j}\bigg{[}\varphi\bigg{(}\frac{y_{h}-\epsilon j}{\epsilon}\bigg{)}-\varphi\bigg{(}\frac{y-\epsilon j}{\epsilon}\bigg{)}\bigg{]}\big{(}g(\epsilon j)-g(y)\big{)}.\end{split}$
+By assumption, $\varphi$ is at least as smooth as $g$, so
+(B.11) $\begin{split}\bigg{|}\frac{\varphi\big{(}(y_{h}-\epsilon j)/\epsilon\big{)}-\varphi\big{(}(y-\epsilon j)/\epsilon\big{)}}{(|h|/\epsilon)^{s}}\frac{g(\epsilon j)-g(y)}{\epsilon^{s}}\bigg{|}<c.\end{split}$
+Hence $|g_{\epsilon}(y_{h})-g_{\epsilon}(y)|\leq c|h|^{s}$. The fact that $\|g_{\epsilon}\|_{L^{\infty}(V)}<c$ is obvious.
+Suppose now $g\in C_{0}^{s}(V)$, $21<s<$ 2, i.e., we study $\Delta g$ of Lemma 4.2.
+As above, $\|g_{\epsilon}\|_{L^{\infty}(V)}<c$.
+By the exactness of $\varphi$ up to order one, differentiation of (3.14) gives
+(B.12) $\begin{split}&g_{\epsilon}^{\prime}(y)=g^{\prime}(y)+\sum_{j}\varphi^{\prime}\bigg{(}\frac{y-\epsilon j}{\epsilon}\bigg{)}\frac{g(\epsilon j)-\big{[}g(y)+\text{d}_{y}g(y)\cdot(\epsilon j-y)\big{]}}{\epsilon},\end{split}$
+where all the primes stand for the same derivative $\partial_{y_{l}}$ for some $l=1,2,\dots n$.
 The fraction on the right is $O(\epsilon^{s-1})$, therefore $\|g_{\epsilon}\|_{C^{1}(V)}<c$, $0<\epsilon\ll 1$.
-To prove $\|g_{\epsilon}^{\prime}\|_{C^{s-1}(V)}<c$, we argue similarly to (B.8). The case $h\geq\epsilon$ is completely analogous. If $0<h<\epsilon$, from (B.11) and the exactness of $\varphi$:
-(B.12) $\begin{split}&\Delta g_{\epsilon}^{\prime}(y,h)-\Delta g^{\prime}(y,h)\\ &=\sum_{j}\bigg{[}\varphi^{\prime}\bigg{(}\frac{y_{h}}{\epsilon}-j\bigg{)}-\varphi^{\prime}\bigg{(}\frac{y}{\epsilon}-j\bigg{)}\bigg{]}\frac{g(\epsilon j)-[g(y_{h})+(\epsilon j-y_{h})\nabla g(y_{h})]}{\epsilon}\\ &\quad+\sum_{j}\varphi^{\prime}\bigg{(}\frac{y}{\epsilon}-j\bigg{)}\frac{\epsilon j-y}{\epsilon}[\nabla g(y_{h})-\nabla g(y)].\end{split}$
-Dividing the first sum by $h^{s-1}=(h/\epsilon)^{s-1}\epsilon^{s-1}$ we get similarly to (B.10) that the ratio is bounded.
-The second sum divided by $h^{s-1}$ is bounded, because $g^{\prime}\in C^{s-1}(V)$.
-The proof is complete.
+To prove $\|g_{\epsilon}^{\prime}\|_{C^{s-1}(V)}<c$, we argue similarly to the case $s<1$.
+As before, if $|h|\geq\epsilon$, we only need to show that $|g_{\epsilon}^{\prime}(y)-g^{\prime}(y)|\leq c\epsilon^{s-1}$.
+The latter inequality immediately follows from (B.12).
+If $|h|<\epsilon$, from (B.12) and the exactness of $\varphi$:
+(B.13) $\begin{split}&g_{\epsilon}^{\prime}(y_{h})-g_{\epsilon}^{\prime}(y)\\ &=\sum_{j}\bigg{[}\varphi^{\prime}\bigg{(}\frac{y_{h}-\epsilon j}{\epsilon}\bigg{)}-\varphi^{\prime}\bigg{(}\frac{y-\epsilon j}{\epsilon}\bigg{)}\bigg{]}\frac{g(\epsilon j)-[g(y)+\text{d}_{y}g(y)\cdot(\epsilon j-y)]}{\epsilon}.\end{split}$
+Dividing by $|h|^{s-1}=(|h|/\epsilon)^{s-1}\epsilon^{s-1}$ we get similarly to (B.11) that the ratio is bounded. The proof is complete.
 Appendix C Proof of Lemma 5.2
-The idea of the proof is to reduce $f_{\epsilon}$ to $G_{\epsilon}$ by first neglecting smooth terms in $f_{\epsilon}$ and then by neglecting lower order terms of finite smoothness, which still do not contribute to the DTB.
+The idea of the proof is to reduce $f_{\epsilon}$ to $\mathcal{G}_{\epsilon}$ by first neglecting smooth terms in $f_{\epsilon}$ and then by neglecting lower order terms of finite smoothness, which still do not contribute to the DTB.
 This is done in Sections C.1 and C.2, respectively.
@@
 (C.1) $\mathcal{T}_{\epsilon}=\text{Op}(\chi(x)|\xi|)\chi(x)\text{Op}\big{(}\tilde{D}(x,\xi)+\kappa\epsilon^{3}|\xi|^{2}\big{)}\in L^{3}(\mathcal{U}).$
-Let $\mathcal{B}_{\epsilon}\in L^{-3}(\mathcal{U}_{b})$ be a local parametrix for $\mathcal{T}_{\epsilon}$ on the open set $\{x\in\mathcal{U}_{b}:\chi(x)>1/2\}\supset\mathcal{U}_{0}$, see Figure 4.
+Recall that $\chi\in C_{0}^{\infty}(\mathcal{U}_{b})$.
+Let $\mathcal{B}_{\epsilon}\in L^{-3}(\mathcal{U}_{b})$ be a local parametrix for $\mathcal{T}_{\epsilon}$ on the open set $\mathcal{U}_{1/2}:=\{x\in\mathcal{U}_{b}:\chi(x)>1/2\}\Supset\mathcal{U}_{0}$, see Figure 4.
 Below, we omit the part “$x\in\mathcal{U}_{b}$” in similar sets.
 Then
-(C.2) $\begin{split}&f_{\epsilon}(x)=(\mathcal{B}_{\epsilon}F_{\epsilon})(x)+(\mathcal{H}_{\epsilon}f_{\epsilon})(x),\\ &\tilde{B}_{\epsilon}(x,\xi)-\big{(}|\xi|\tilde{D}(x,\xi)+\kappa|\epsilon\xi|^{3}\big{)}^{-1}\in S^{-1}(\mathcal{U}_{0}),\end{split}$
-where $\tilde{B}_{\epsilon}\in S^{-3}(\mathcal{U}_{b})$ is the complete symbol of $\mathcal{B}_{\epsilon}$, and $\mathcal{H}_{\epsilon}:\mathcal{E}^{\prime}(\mathcal{U}_{b})\to C^{\infty}(\mathcal{U}_{0})$ is an operator with a smooth kernel, $H_{\epsilon}(x,z)$.
-Figure 4. Illustration of the sets $\mathcal{U}_{0}$, $\mathcal{U}_{b}$, $\mathcal{U}$, $\{\chi(x)=1\}$, and the level set $\{\chi(x)=c\}$, $10<c<$ 1.
-To prove the latter claim, use a partition of unity to write $f=f_{1}+f_{2}$, where
-$\text{supp}\,f_{1}\subset\{\chi(x)>1/2\},\quad\text{supp}\,f_{2}\subset\{\chi(x)<3/4\}.$
-Then $\mathcal{B}_{\epsilon}\mathcal{T}_{\epsilon}(f_{1}+f_{2})=f_{1}+\mathcal{H}_{\epsilon}^{\prime}f_{1}+\mathcal{B}_{\epsilon}\mathcal{T}_{\epsilon}f_{2}$, where $\mathcal{H}_{\epsilon}^{\prime}\in L^{-\infty}(\{\chi(x)>1/2\})$.
-The first two terms on the right follow by the properties of a parametrix.
-The fact that
-$\mathcal{B}_{\epsilon}\mathcal{T}_{\epsilon}:\mathcal{E}^{\prime}(\{\chi(x)<3/4\})\to C^{\infty}(\mathcal{U}_{0})$
+(C.2) $\begin{split}&f_{\epsilon}(x)=(\mathcal{B}_{\epsilon}F_{\epsilon})(x)+(\mathcal{H}_{\epsilon}f_{\epsilon})(x),\\ &\tilde{B}_{\epsilon}(x,\xi)-\chi_{1}(\xi)\big{(}|\xi|\tilde{D}(x,\xi)+\kappa|\epsilon\xi|^{3}\big{)}^{-1}\in S^{-1}(\mathcal{U}_{1/2}),\end{split}$
+where
+  1. (1)
+    $\tilde{B}_{\epsilon}\in S^{-3}(\mathcal{U}_{b})$ is the complete symbol of $\mathcal{B}_{\epsilon}$, i.e. $\mathcal{B}_{\epsilon}=\text{Op}(\tilde{B}_{\epsilon})$,
+  2. (2)
+    $\mathcal{H}_{\epsilon}:\mathcal{E}^{\prime}(\mathcal{U}_{b})\to C^{\infty}(\mathcal{U}_{0})$ is an operator with a smooth kernel, $H_{\epsilon}(x,z)\in C^{\infty}(\mathcal{U}_{0}\times\mathcal{U}_{b})$, and
+  3. (3)
+    $\chi_{1}\in C^{\infty}(\mathbb{R}^{2})$ satisfies $\chi_{1}(\xi)\equiv 0$, $|\xi|\leq R$, and $\chi_{1}(\xi)\equiv 1$, $|\xi|\geq R+1$, for some $R\gg 1$.
+Figure 4. The sets $\mathcal{U}_{0}$, $\mathcal{U}_{b}$, $\mathcal{U}$, $\{\chi(x)=1\}$, and the level set $\{\chi(x)=c\}$, $10<c<$ 1.
+To prove (C.2), use a partition of unity to write $f_{\epsilon}=f_{\epsilon}^{(1)}+f_{\epsilon}^{(2)}$, where
+(C.3) $\text{supp}\,f_{\epsilon}^{(1)}\subset\mathcal{U}_{1/2}=\{\chi(x)>1/2\},\quad\text{supp}\,f_{\epsilon}^{(2)}\subset\{\chi(x)<3/4\}.$
+Then $\mathcal{B}_{\epsilon}\mathcal{T}_{\epsilon}f_{\epsilon}=f_{\epsilon}^{(1)}+\mathcal{H}_{\epsilon}^{\prime}f_{\epsilon}^{(1)}+\mathcal{B}_{\epsilon}\mathcal{T}_{\epsilon}f_{\epsilon}^{(2)}$, where $\mathcal{H}_{\epsilon}^{\prime}\in L^{-\infty}(\mathcal{U}_{1/2})$.
+The first two terms on the right follow by the properties of a parametrix, [55, Section I.4, Appendix].
+By construction, $\text{dist}(\{\chi(x)<3/4\},\mathcal{U}_{0})>0$, and the property
+(C.4) $\mathcal{B}_{\epsilon}\mathcal{T}_{\epsilon}:\mathcal{E}^{\prime}(\{\chi(x)<3/4\})\to C^{\infty}(\mathcal{U}_{0})$
 follows by the pseudolocality of $\Psi$DOs.
-Derivatives of $H_{\epsilon}(x,z)$ with respect to $x$ and $z$ are uniformly bounded with respect to $x\in\mathcal{U}_{0}$, $z\in\mathcal{U}_{b}$, and $0<\epsilon\ll 1$.
-This follows from the way a symbol of $\mathcal{B}_{\epsilon}$ is constructed (see [52, Ch.
+The derivatives $\partial_{x}^{m_{1}}\partial_{z}^{m_{2}}H_{\epsilon}(x,z)$, $m_{1},m_{2}\in\mathbb{N}_{0}^{2}$, are uniformly bounded with respect to $x\in\mathcal{U}_{0}$, $z\in\mathcal{U}_{b}$, and $0<\epsilon\ll 1$.
+This follows from the above discussion, the way a symbol of $\mathcal{B}_{\epsilon}$ is constructed (see [55, Ch.
 I, eqs.
-(4.39), (4.40)]) and observing that all the seminorms of $\big{(}|\xi|\tilde{D}(x,\xi)+\kappa|\epsilon\xi|^{3}\big{)}^{-1}$ as member of $S^{0}(\mathcal{U}_{b})$ (i.e., the minimal constants $c_{m}$ and $c_{m_{1},m_{2}}$ in (2.7), where $r=0$) are uniformly bounded for $0<\epsilon\ll 1$.
+(4.39), (4.40)]), and the observation that all the seminorms of $\chi_{1}(\xi)\big{(}|\xi|\tilde{D}(x,\xi)+\kappa|\epsilon\xi|^{3}\big{)}^{-1}$ as member of $S^{0}(\mathcal{U}_{1/2})$ (i.e., the minimal constants $c_{m}$ and $c_{m_{1},m_{2}}$ in (2.7), where $r=0$) are uniformly bounded for $0<\epsilon\ll 1$.
+Likewise, the essence of the second line in (C.2) is that all the seminorms of the left-hand side as member of $S^{-1}(\mathcal{U}_{1/2})$ are uniformly bounded for $0<\epsilon\ll 1$ as well.
 Therefore, by Lemma 5.1,
@@
 C.2. Extracting the leading local singularity of $f_{\epsilon}$
-In the preceding section we converted $f_{\epsilon}$ to $\mathcal{B}_{\epsilon}F_{\epsilon}$ modulo a smooth function.
+In the preceding section we converted $f_{\epsilon}$ to $\mathcal{B}_{\epsilon}F_{\epsilon}$ modulo a $C^{\infty}(\mathcal{U}_{0})$ function.
 We now represent $\mathcal{B}_{\epsilon}=\mathcal{B}_{\epsilon}^{(0)}+\Delta\mathcal{B}_{\epsilon}$ with some convenient $\mathcal{B}_{\epsilon}^{(0)}\in L^{0}(\mathcal{U}_{b})$, $\Delta\mathcal{B}_{\epsilon}\in L^{-1}(\mathcal{U}_{b})$.
 Then we use the mapping properties of $\Delta\mathcal{B}_{\epsilon}$ to show that $\Delta\mathcal{B}_{\epsilon}F_{\epsilon}$ does not contribute to the DTB.
-Using the results in [53, Section 6.2, Chapter VIII] and [11, Proposition 4.2.4, Section 4.2] and (3.5), the principal symbol of $\mathcal{R}^{*}\mathcal{R}$ is given by
-(C.4) $\begin{split}\frac{2\pi}{|\nu|}\frac{W^{2}(x,y)}{|\Delta_{\Phi}(x,\alpha)|},\ y=y(x,\xi),\nu=\nu(x,\xi),\end{split}$
-where the functions $y(x,\xi)\in\mathcal{V}$ and $\nu(x,\xi)\in\mathbb{R}$ are obtained by solving the equations:
-(C.5) $p=\Phi(x,\alpha),\ \xi=-\nu\text{d}_{x}\Phi(x,\alpha).$
-The existence of a solution follows from assumption 3.2(3), and its local uniqueness follows from assumption 3.2(1).
+Using the results in [56, Chapter VIII, Section 6.2] and (3.5), the principal symbol of $\mathcal{R}^{*}\mathcal{R}$ is given by
+(C.6) $\begin{split}\frac{2\pi}{|\nu|}\frac{W^{2}(x,y)}{|\Delta_{\Phi}(x,\alpha)|},\ y=y(x,\xi),\nu=\nu(x,\xi),\ (x,\xi)\in\mathcal{U}\times(\mathbb{R}^{2}\setminus 0_{2}),\end{split}$
+where the functions $y(x,\xi)\in V$ and $\nu(x,\xi)\in\mathbb{R}$ are obtained by solving the equations:
+(C.7) $p=\Phi(x,\alpha),\ \xi=-\nu\text{d}_{x}\Phi(x,\alpha),\ (\alpha,p)\in V.$
+The existence of a local solution follows from Assumption 3.2(3).
+Its local uniqueness and smoothness follows from Assumption 3.2(1).
+In Theorem 3.6 we assume that the GRT data are not redundant, so only one solution exists.
+Note that the symbol in (C.6) is integrable at the origin $\xi=0$.
+More precisely, it satisfies the first inequality in (2.7) with $n=2$ and $\delta=1$.
 Define
@@
 (C.10) $|\xi|\tilde{D}(x,\xi)-\tilde{D}_{0}(x,\xi/|\xi|)\in S^{-1}(\mathcal{U}_{0}).$
-Thus we can find $\Delta\tilde{B}_{\epsilon}\in S^{-1}(\mathcal{U}_{b})$ (with all seminorms uniformly bounded when $0<\epsilon\ll 1$) such that
+Thus we can find $\Delta\tilde{B}_{\epsilon}\in S^{-1}(\mathcal{U}_{b})$ (with all the seminorms uniformly bounded for $0<\epsilon\ll 1$) such that (see (C.2))
 (C.11) $\begin{split}&\tilde{B}_{\epsilon}(x,\xi)=\big{(}\tilde{D}_{0}(x,\xi/|\xi|)+\kappa|\epsilon\xi|^{3}\big{)}^{-1}+\Delta\tilde{B}_{\epsilon}(x,\xi),\ x\in\mathcal{U}_{0}.\end{split}$
-Note that (LABEL:tilde_B) is required to hold only for $x\in\mathcal{U}_{0}$.
-Since $\mathcal{B}_{\epsilon}$ is a parametrix for $T_{\epsilon}$ on $\{\chi(x)>1/2\}$, we can select $\tilde{B}_{\epsilon}$ in (C.2) and $\Delta\tilde{B}_{\epsilon}$ so that
-(C.10) $\begin{split}\tilde{B}_{\epsilon}(x,\xi)\equiv 0,\ \Delta\tilde{B}_{\epsilon}(x,\xi)\equiv 0,\ x\in\{\chi(x)<1/4\}.\end{split}$
 From (5.6),
-(C.11) $\text{Op}\big{(}\Delta\tilde{B}_{\epsilon}(x,\xi)\big{)}F_{\epsilon}=\mathcal{W}_{\epsilon}g_{\epsilon},\ \mathcal{W}_{\epsilon}:=\text{Op}\big{(}\Delta\tilde{B}_{\epsilon}(x,\xi)\big{)}\text{Op}(\chi(x)|\xi|)\chi(x)\mathcal{R}^{*}.$
-By construction, $\mathcal{W}_{\epsilon}$ is an FIO of order $-1/2$, $\mathcal{W}_{\epsilon}\in I^{-1/2}(\mathcal{U}\times\mathcal{V}\times\mathcal{C}^{*})$.
-The canonical relation $\mathcal{C}^{*}$ is the image of $\mathcal{C}$ (see (3.3)) under the exchange of $T^{*}\mathcal{U}$ and $T^{*}\mathcal{V}$.
-By (LABEL:tilde_B), all the seminorms of a symbol of $\mathcal{W}_{\epsilon}$, $W_{\epsilon}\in S^{0}(\mathcal{U}\times\mathcal{V}\times\mathbb{R})$, $0<\epsilon\ll 1$, are bounded.
-Recall that $B_{\infty,\infty}^{s}(\mathbb{R}^{n})=C^{s}(\mathbb{R}^{n})$, $s>0$, $s\not\in\mathbb{N}$ [1, Remark 6.4(2)]. We use [23, Proposition 5.5] with $p,q=\infty$, which asserts that
-(C.12) $\mathcal{W}_{\epsilon}:C^{s-(1/2)-m_{c}(\infty)}(\mathbb{R}^{n})\to C^{s}(\mathbb{R}^{n}),\ m_{c}(p)=-(n-1)\big{|}(1/p)-(1/2)\big{|},$
-is continuous.
-Here $s-(1/2)-m_{c}(\infty)>0$ and $s>0$.
-See also [45, Theorem, p.
-173].
-For the validity of (C.12) it is required that the phase function of $\mathcal{W}_{\epsilon}$ be non-degenerate (see the paragraph following (3.2)) and the amplitude be compactly supported in $x$ (see (C.10)).
-Clearly, $m_{c}(\infty)=-1/2$.
-Using (C.12) with $s=1/2$ and that $g_{\epsilon}\in C_{0}^{1/2}(\mathcal{V})$, $\|g_{\epsilon}\|_{C^{1/2}(\mathcal{U}_{b})}<c$, $0<\epsilon\ll 1$ (see Lemma 4.2), we conclude that $\mathcal{W}_{\epsilon}g_{\epsilon}\in C_{0}^{1/2}(\mathcal{U}_{b})$ and $\|\mathcal{W}_{\epsilon}g_{\epsilon}\|_{C^{1/2}(\mathcal{U}_{b})}<c$, $0<\epsilon\ll 1$.
-Therefore, by (2.3),
-(C.13) $\|\mathcal{W}_{\epsilon}g_{\epsilon}\|_{L^{\infty}(\mathcal{U}_{b})}<c,0<\epsilon\ll 1;\ |\mathcal{W}_{\epsilon}g_{\epsilon}(x_{0}+\epsilon\check{x})-\mathcal{W}_{\epsilon}g_{\epsilon}(x_{0})|=O(\epsilon^{1/2})$
-when $\check{x}$ is confined to any bounded set.
-Combining (C.3) and (C.13) finishes the proof.
+(C.12) $\begin{split}&\text{Op}\big{(}\Delta\tilde{B}_{\epsilon}(x,\xi)\big{)}F_{\epsilon}=\mathcal{B}_{1}\mathcal{R}^{*}g_{\epsilon},\\ &\mathcal{B}_{1}:=\text{Op}\big{(}\Delta\tilde{B}_{\epsilon}(x,\xi)\big{)}\text{Op}(\chi(x)|\xi|)\chi(z)\in L^{0}(\mathcal{U}).\end{split}$
+Arguing similarly to Section 6.1 (from the beginning through (6.7)), we get
+(C.13) $\begin{split}&\|\mathcal{B}_{1}\mathcal{R}^{*}g_{\epsilon}\|_{L^{\infty}(\mathcal{U}_{b})}<c,0<\epsilon\ll 1;\\ &|(\mathcal{B}_{1}\mathcal{R}^{*}g_{\epsilon})(x_{0}+\epsilon\check{x})-(\mathcal{B}_{1}\mathcal{R}^{*}g_{\epsilon})(x_{0})|=O(\epsilon^{1/2}),\ \check{x}\in\mathcal{K},\end{split}$
+where $\mathcal{K}$ is any compact set.
+Combining (5.7), (C.2), (C.5), (LABEL:tilde_B) and (C.13) finishes the proof.
 Appendix D Proof of Lemmas 6.2, 6.3
@@
 Suppose Assumptions 3.13.4 are satisfied. Pick any $A>0$. One has
-(D.1) $H(\check{t},\check{q};\cdot)\equiv 0,\ \check{t}<-c.$
-Also, for some smooth and bounded $\varkappa(t,q;\cdot)$,
-(D.2) $\begin{split}&\partial_{\check{t}}^{k}\big{[}H(\check{t},\check{q};\cdot)-\varkappa(\epsilon(\check{t}+\check{q});\cdot)\check{t}^{-1/2}\big{]}=O(\check{t}^{-3/2}),\\ &k=0,1,\ \check{t}\to+\infty,\ |\check{t}|<A/\epsilon,\end{split}$
+(D.1) $H(\check{t},\check{q};*)\equiv 0,\ \check{t}<-c.$
+Also, for some smooth and bounded $\varkappa(t,q;\theta)$,
+(D.2) $\begin{split}&\partial_{\check{t}}^{k}\big{[}H(\check{t},\check{q};*)-\varkappa(\epsilon\check{t},\epsilon\check{q};\theta)\check{t}^{-1/2}\big{]}=O(\check{t}^{-3/2}),\\ &k=0,1,\ \check{t}\to+\infty,\ \check{t}<A/\epsilon,\end{split}$
 and
-(D.3) $\begin{split}&\partial_{\check{t}}\big{[}\varkappa(\epsilon(\check{t}+\check{q});\cdot)\check{t}^{-(1/2)}\big{]}=O(\check{t}^{-3/2}),\ \check{t}\to+\infty,\ |\check{t}|<A/\epsilon.\end{split}$
-The big-$O$ terms are uniform with respect to $\check{q}$ and the second group of variables collectively denoted $\cdot$’, namely $\theta,\epsilon$.
-These variables are restricted to the sets indicated in Lemma 6.2.
-For simplicity, we drop the arguments $\check{q}$, $\theta$ and $\epsilon$ from all the functions.
-The limit as $\check{t}\to-\infty$ in (6.16) follows immediately from (D.1).
-Next, suppose $\check{t}\to\infty$.
-Then
+(D.3) $\begin{split}&\partial_{\check{t}}\big{[}\varkappa(\epsilon\check{t},\epsilon\check{q};\theta)\check{t}^{-1/2}\big{]}=O(\check{t}^{-3/2}),\ \check{t}\to+\infty,\ \check{t}<A/\epsilon.\end{split}$
+The big-$O$ terms are uniform with respect to $\check{q}$, $\theta$, and $\epsilon$.
+These variables are restricted to the sets $|\check{q}|\leq A/\epsilon$, $\theta\in I_{\alpha}$, and $0<\epsilon\ll 1$.
+For simplicity, we drop the arguments $\check{q}$, $\theta$ and $*=(x,\theta,\epsilon)$ from all the functions.
+From (D.1) and (LABEL:H_props_2_a) with $k=0$ we get $H(\check{t})=O(\check{t}^{-1/2})$, $s\to\infty$.
+From Assumption 3.4(1), (6.13), and (6.14), it follows that $H(\check{t},\check{q};*)$ and $\partial_{\check{t}}H(\check{t},\check{q};*)$ are uniformly bounded.
+The limit as $\check{t}\to-\infty$ in (6.19) easily follows from (6.17).
+Next, suppose $\check{t}\to\infty$. Then
 (D.4) $\begin{split}&G(\check{t})=\int_{\mathbb{R}}\chi_{1\text{D}}(\epsilon(s-\check{t}))\frac{H_{1}(s)}{s-\check{t}}\text{d}s+\int_{0}^{\infty}\chi_{1\text{D}}(\epsilon(s-\check{t}))\frac{\varkappa(\epsilon(s+\check{q}))}{(s-\check{t})s^{1/2}}\text{d}s,\\ &H_{1}(s):=H(s)-\varkappa(\epsilon(s+\check{q}))s_{+}^{-1/2},\end{split}$
-where $\varkappa(\cdot)$ is the same as in (LABEL:H_props_2_a), (LABEL:H_props_2_b). In view of (D.1), the first of the integrals is split into five
-(D.5) $\begin{split}&\bigg{(}\int_{-c}^{0}+\int_{0}^{1}+\int_{1}^{\check{t}-1}+\int_{\check{t}-1}^{\check{t}+1}+\int_{\check{t}+1}^{\infty}\bigg{)}\chi_{1\text{D}}(\epsilon(s-\check{t}))\frac{H_{1}(s)}{s-\check{t}}\text{d}s\\ &=J_{1}+\dots+J_{5}.\end{split}$
+where $\varkappa(\cdot)$ is the same as in (LABEL:H_props_2_a), (LABEL:H_props_2_b). Using (D.1), split the first of the integrals into four
+(D.5) $\begin{split}&\bigg{(}\int_{-c}^{t/2}+\int_{t/2}^{\check{t}-1}+\int_{\check{t}-1}^{\check{t}+1}+\int_{\check{t}+1}^{\infty}\bigg{)}\chi_{1\text{D}}(\epsilon(s-\check{t}))\frac{H_{1}(s)}{s-\check{t}}\text{d}s\\ &=J_{1}+\dots+J_{4}.\end{split}$
 By (LABEL:H_props_2_a) with $k=0$, $H_{1}(s)=O(s^{-3/2})$, $s\to\infty$.
-It is trivial to see that $J_{1},J_{2}=O(1/\check{t})$.
-Breaking up the integral over $[1,\check{t}-1]$ into the integrals over $[1,\check{t}/2]$ and $[\check{t}/2,\check{t}-1]$ we get $J_{3}=O(1/\check{t})$.
-Breaking up the integral over $[\check{t}+1,\infty)$ into the integrals over $[\check{t}+1,2\check{t}]$ and $[2\check{t},\infty)$ gives $J_{5}=O(1/\check{t})$.
+It is trivial to see that $J_{1}=O(1/\check{t})$ and $J_{2}=O(\check{t}^{-3/2}\ln\check{t})$.
+In the first integral we use $s-\check{t}\asymp\check{t}$, and in the second one we use that $s\asymp\check{t}$ to bound above $H_{1}(s)$ by $c\check{t}^{-3/2}$.
+Breaking up the integral over $[\check{t}+1,\infty)$ into the integrals over $[\check{t}+1,2\check{t}]$ and $[2\check{t},\infty)$ gives $J_{4}=O(\check{t}^{-3/2}\ln\check{t})$.
 Finally, using (LABEL:H_props_2_a) with $k=1$ and that $|s-\check{t}|\leq 1$, $\epsilon\check{t}\leq c$, we get
 (D.6) $\partial_{s}\big{[}\chi_{1\text{D}}(\epsilon(s-\check{t}))H_{1}(s)\big{]}=O(\check{t}^{-3/2}),$
-which implies $J_{4}=O(\check{t}^{-3/2})$. Combining the five estimates proves that the first integral in (D.4) is $O(1/\check{t})$.
-Consider the second integral in (D.4). Recall that $\check{t}\to\infty$. Write the integral in the form
-(D.7) $\begin{split}J(\check{t},\epsilon):=&\int_{0}^{\infty}\frac{\chi_{1\text{D}}(\epsilon(\check{s}-\check{t}))\varkappa(\epsilon(\check{s}+\check{q}))}{(\check{s}-\check{t})\check{s}^{1/2}}\text{d}\check{s}=\epsilon^{1/2}\int_{0}^{\infty}\frac{\chi_{1\text{D}}(s-t)\varkappa(s+q)}{(s-t)s^{1/2}}\text{d}s\\ =&\epsilon^{1/2}\int_{0}^{\infty}\frac{(\chi_{1\text{D}}(s-t)-\chi_{1\text{D}}(0))\varkappa(s+q)}{(s-t)s^{1/2}}\text{d}s\\ &+\epsilon^{1/2}\chi_{1\text{D}}(0)\int_{0}^{\infty}\frac{\varkappa(s+q)}{(s-t)s^{1/2}}\text{d}s,\ q=\epsilon\check{q}.\end{split}$
-The integral on the second line is clearly uniformly bounded. The identity [43, Eq. 2.2.5.26] with $\alpha=1/2$:
+which implies $J_{3}=O(\check{t}^{-3/2})$. Combining the four estimates proves that the first integral in (D.4) is $O(1/\check{t})$.
+Consider the second integral in (D.4), which we denote $J(\check{t},\epsilon)$. Recall that $\check{t}\to\infty$. Write $J(\check{t},\epsilon)$ in the form
+(D.7) $\begin{split}J(\check{t},\epsilon):=&\epsilon^{1/2}\int_{0}^{\infty}\frac{\chi_{1\text{D}}(s-t)\varkappa(s+q)}{(s-t)s^{1/2}}\text{d}s\\ =&\epsilon^{1/2}\int_{0}^{\infty}\frac{(\chi_{1\text{D}}(s-t)-\chi_{1\text{D}}(0))\varkappa(s+q)}{(s-t)s^{1/2}}\text{d}s\\ &+\epsilon^{1/2}\chi_{1\text{D}}(0)\int_{0}^{\infty}\frac{\varkappa(s+q)}{(s-t)s^{1/2}}\text{d}s,\ q=\epsilon\check{q},\ t=\epsilon\check{t}.\end{split}$
+It is not hard to show that the integral on the second line is uniformly bounded. Recall that $\chi_{1\text{D}}\in C_{0}^{\infty}(\mathbb{R})$. The identity [47, Eq. 2.2.5.26] with $\alpha=1/2$:
 (D.8) $\int_{0}^{\infty}\frac{\text{d}s}{(s-t)s^{1/2}}=0,\ t>0,$
@@
 (D.9) $\int_{0}^{\infty}\frac{\varkappa(s+q)}{(s-t)s^{1/2}}\text{d}s=\int_{0}^{\infty}\frac{\varkappa(s+q)-\varkappa(t+q)}{(s-t)s^{1/2}}\text{d}s=O(1),$
-Here the term $O(1)$ is uniform with respect to $t,q\in\mathbb{R}$. Therefore $J(\check{t},\epsilon)=O(\epsilon^{1/2})$. Combining the results proves (6.16).
+Here the term $O(1)$ is uniform with respect to $0<t<A,|q|<A$. Therefore $J(\check{t},\epsilon)=O(\epsilon^{1/2})$.
+Given that $\partial_{\check{t}}H$ is bounded, it is clear that $G$ is bounded with $\check{t}$ in compact sets. Combining all the results proves (6.19).
 D.2. Proof of Lemma D.1
-Assertion (D.1) is obvious from (6.11). To prove (LABEL:H_props_2_a), (LABEL:H_props_2_b), begin by proving the following auxiliary lemma.
+Assertion (D.1) is obvious from (6.14) because $h$ is compactly supported. To prove (LABEL:H_props_2_a), (LABEL:H_props_2_b), begin by proving the following auxiliary lemma.
 Lemma D.2.
 Pick any $A>0$ and a function $f\in C^{2}(\mathbb{R})$. Let $\varphi\in C_{0}^{2}(\mathbb{R})$ be an interpolating kernel, which is exact up to degree one. One has
-(D.10) $\begin{split}\partial_{t}^{k}\big{[}&\textstyle\sum_{j}\varphi(t+b-j)f(\epsilon j)(j-b)_{+}^{1/2}-f(\epsilon(t+b))t^{1/2}\big{]}=O\big{(}t^{-3/2}\big{)},\\ &t\to+\infty,\ \epsilon\to 0,\ \epsilon|t|<A,\ k=0,1,2.\end{split}$
-The big-$O$ term is uniform in $b\in\mathbb{R}$.
+(D.10) $\begin{split}\partial_{\check{t}}^{k}\big{[}&\textstyle\sum_{j}\varphi(\check{t}+\check{q}-j)f(\epsilon j)(j-\check{q})_{+}^{1/2}-f(\epsilon(\check{t}+\check{q}))\check{t}^{1/2}\big{]}=O\big{(}\check{t}^{-3/2}\big{)},\\ &\check{t}\to+\infty,\ \epsilon\to 0,\ \epsilon\check{t}<A,\ k=0,1,2.\end{split}$
+The big-$O$ term is uniform in $|\check{q}|\leq A/\epsilon$.
 Proof.
-Denote $g(t):=f(\epsilon t)(t-b)^{1/2}$ and $u:=t+b$. Taylor expanding $g$ gives
-(D.11) $\begin{split}&g(j)=g(u)+g^{\prime}(u)(j-u)+R_{1}(j,u),\\ &R_{1}(j,u)=\int_{u}^{j}g^{\prime\prime}(s)(j-s)\text{d}s.\end{split}$
+Denote $g(\check{t}):=f(\epsilon\check{t})(\check{t}-\check{q})^{1/2}$ and $u:=\check{t}+\check{q}$.
+The difference in brackets in (D.10) becomes $g_{\epsilon}(u)-g(u)$, where $g_{\epsilon}$ stands for the interpolated $g$ similarly to (3.14) (see (D.12)).
+Taylor expanding $g$ gives
+(D.11) $\begin{split}&g(j)=g(u)+g^{\prime}(u)(j-u)+R_{1}(j,u),\ j-u\in\text{supp}\,\varphi,\\ &R_{1}(j,u)=\int_{u}^{j}g^{\prime\prime}(s)(j-s)\text{d}s.\end{split}$
 The exactness of $\varphi$ up to degree one implies:
-(D.12) $\sum_{j}\varphi(u-j)g(j)=g(u)+\textstyle\sum_{j}\varphi(u-j)R_{1}(j,u).$
-Clearly
-(D.13) $|g^{\prime\prime}(u)|=O\big{(}t^{-3/2}\big{)},\ \epsilon|t|<A,\ t\to\infty.$
-Hence, by (D.11)
-(D.14) $\begin{split}&R_{1}(j,u)=O(t^{-3/2}),\ |j-u|<A,\ \epsilon|t|<A,\ t\to\infty.\end{split}$
+(D.12) $g_{\epsilon}(u):=\textstyle\sum_{j}\varphi(u-j)g(j)=g(u)+\textstyle\sum_{j}\varphi(u-j)R_{1}(j,u).$
+By assumption, $0<\epsilon t<A$ and $\epsilon|\check{q}|\leq A$, so $|\epsilon\check{t}+\epsilon\check{q}|<2A$ and
+(D.13) $|g^{\prime\prime}(u)|=\left|\partial_{\check{t}}^{2}\left(f(\epsilon\check{t}+\epsilon\check{q})\check{t}^{1/2}\right)\right|=O\big{(}\check{t}^{-3/2}\big{)},\ \epsilon\check{t}<A,\ \check{t}\to+\infty.$
+The values of $j,s,u$ used in (D.11) and (D.12) satisfy $|j-s|\leq|j-u|\leq\text{diam}(\text{supp}\,\varphi)$. Hence, by (D.11) and (D.13)
+(D.14) $\begin{split}&R_{1}(j,u)=O(\check{t}^{-3/2}),\ j-u\in\text{supp}\,\varphi,\ \epsilon\check{t}<A,\ \check{t}\to+\infty.\end{split}$
 Furthermore, using that $\varphi$ is exact up to order one,
-(D.15) $\begin{split}\partial_{u}R_{1}(j,u)=&-g^{\prime\prime}(u)(j-u),\\ \partial_{u}\textstyle\sum_{j}\varphi(u-j)R_{1}(j,u)=&\textstyle\sum_{j}\varphi^{\prime}(u-j)R_{1}(j,u)-g^{\prime\prime}(u)\textstyle\sum_{j}\varphi(u-j)(j-u)\\ \hskip 56.9055pt=&\textstyle\sum_{j}\varphi^{\prime}(u-j)R_{1}(j,u),\end{split}$
+(D.15) $\begin{split}\partial_{u}R_{1}(j,u)=&-g^{\prime\prime}(u)(j-u),\\ g_{\epsilon}^{\prime}(u)-g^{\prime}(u)=&\partial_{u}\textstyle\sum_{j}\varphi(u-j)R_{1}(j,u)\\ =&\textstyle\sum_{j}\varphi^{\prime}(u-j)R_{1}(j,u)-g^{\prime\prime}(u)\textstyle\sum_{j}\varphi(u-j)(j-u)\\ \hskip 56.9055pt=&\textstyle\sum_{j}\varphi^{\prime}(u-j)R_{1}(j,u),\end{split}$
 and
-(D.16) $\begin{split}\partial_{u}^{2}\textstyle\sum_{j}\varphi(u-j)&R_{1}(j,u)\\ =&\textstyle\sum_{j}\varphi^{\prime\prime}(u-j)R_{1}(j,u)-g^{\prime\prime}(u)\textstyle\sum_{j}\varphi^{\prime}(u-j)(j-u)\\ \hskip 56.9055pt=&\textstyle\sum_{j}\varphi^{\prime\prime}(u-j)R_{1}(j,u)-g^{\prime\prime}(u).\end{split}$
+(D.16) $\begin{split}g_{\epsilon}^{\prime\prime}(u)-g^{\prime\prime}(u)=&\textstyle\sum_{j}\varphi^{\prime\prime}(u-j)R_{1}(j,u)-g^{\prime\prime}(u)\textstyle\sum_{j}\varphi^{\prime}(u-j)(j-u)\\ \hskip 56.9055pt=&\textstyle\sum_{j}\varphi^{\prime\prime}(u-j)R_{1}(j,u)-g^{\prime\prime}(u).\end{split}$
 Combining (D.12)–(D.16) and using that $\varphi^{\prime\prime}\in L^{\infty}(\mathbb{R})$ finishes the proof. ∎
-To prove (LABEL:H_props_2_a), by (6.10) and (6.11) we consider
+To prove (LABEL:H_props_2_a), in view of (6.13) and (6.14), we consider
 (D.17) $H_{aux}(\check{t}):=\textstyle\sum_{j_{2}}\varphi_{p}(\check{t}+\check{q}-j_{2})a_{0}(\theta,\epsilon j_{2})(j_{2}-\check{q})_{+}^{1/2}.$
@@
 We moved one derivative inside the parentheses because $H$ is defined in terms of $h^{\prime}$ in (6.14).
-Replacing $\check{t}$ with $[\Phi(\theta+\epsilon\mu s)-\Phi(\theta)]/\epsilon+\check{t}+\check{q}$, multiplying both sides by $K(\theta+\epsilon\mu s)\varphi_{\alpha}(s)$, and integrating with respect to $s$ (cf.
+Replacing $\check{t}$ with $[\Phi(\theta+\epsilon\mu s)-\Phi(\theta)]/\epsilon+\check{t}$, multiplying both sides by $K(\theta+\epsilon\mu s)\varphi_{\alpha}(s)$, and integrating with respect to $s$ (cf.
 (6.13), (6.14)) finishes the proof.
-This also gives a formula for $\varkappa$.
-Our derivation shows that $\varkappa$ is smooth and bounded with all derivatives, but the formula for $\varkappa$ is not needed.
-The claim (LABEL:H_props_2_b) immediately follows because $\varkappa$ is smooth and $\epsilon|\check{t}|<A$.
+This also gives a formula for $\varkappa$:
+(D.19) $\varkappa(t,q;\theta)=t\partial_{t}a_{0}(\theta,t+q)+(1/2)a_{0}(\theta,t+q).$
+This shows that $\varkappa$ is smooth and bounded with all derivatives on any compact set. The claim (LABEL:H_props_2_b) immediately follows as well.
 D.3. Proof of Lemma 6.3
-For simplicity, in what follows we omit the dependence of $G(\check{t},\check{q};\theta,\epsilon)$ on $\check{q},\theta$, and $\epsilon$, and denote $Q^{\prime}:=Q_{\alpha}^{\prime}$ and $Q^{\prime\prime}:=Q_{\alpha\alpha}^{\prime\prime}$.
-By (6.15) we study the convergence of the sum
-(D.19) $J_{\epsilon}:=\epsilon^{1/2}\sum_{\alpha_{j_{1}}\in\text{dom}\,P}G(Q(\alpha_{j_{1}})/\epsilon).$
-Recall that $Q$ depends on $x$, which is confined to a sufficiently small open set $U^{\prime}$ (see the paragraph following (6.7)), and on the selected $P(\alpha)$.
-If $|Q(\alpha)|\geq c$ on $\text{dom}\,P$, using (6.16) in (D.19) gives for each of the partial sums, over $Q(\alpha_{j_{1}})>0$ and $Q(\alpha_{j_{1}})<0$, $|J_{\epsilon}|\leq c\epsilon^{1/2}(1/\epsilon)\epsilon^{1/2}=O(1)$.
-Thus, in what follows we assume $Q(a_{0})=0$ for some $a_{0}\in\text{dom}\,P$ and replace $\text{dom}\,P$ in (D.19) with $I_{\alpha}^{\prime}$, which is a sufficiently small neighborhood of $a_{0}$.
-If $Q^{\prime}(\alpha)\not=0$ on $I_{\alpha}^{\prime}$, then $|Q^{\prime}(\alpha)|\asymp|\alpha-a_{0}|$ on $I_{\alpha}^{\prime}$ and
+D.3.1. Estimation away from the stationary points of $Q$.
+It follows from (6.19) that the magnitude of $G$ can be controlled only by its first argument, $\check{t}$.
+Therefore, in this section, we omit the dependence of $G(\check{t},\check{q};*)$ on $\check{q}$ and $*=(x,\theta,\epsilon)$.
+In view of (6.18), we study the convergence of the sum
+(D.20) $J_{\epsilon}:=\epsilon^{1/2}\textstyle\sum_{j_{1}\in J_{\alpha}}G(Q(x,\alpha_{j_{1}})/\epsilon),\ x\in U^{\prime}.$
+Recall that $Q$ depends on the selected $P(\alpha)$.
+Fix any $0<\delta\ll 1$ and $x\in U^{\prime}$.
+Since $x$ is fixed, we can drop it from notation.
+If $|Q(\alpha)|\geq\delta$ on $I_{\alpha}$, using (6.19) in (D.20) gives the same bound $|J_{\epsilon}|\leq c\epsilon^{1/2}(1/\epsilon)\epsilon^{1/2}=O(1)$ for each of the partial sums: over $Q(\alpha_{j_{1}})>0$ and over $Q(\alpha_{j_{1}})<0$.
+Thus, in what follows we may assume $|Q(\alpha)|\leq\delta$ for some $\alpha\in I_{\alpha}$.
+If $|Q_{\alpha}^{\prime}(\alpha)|\geq\delta$ on $I_{\alpha}$, then $|Q_{\alpha}^{\prime}(\alpha)|\asymp|\alpha-a_{0}|$ on $I_{\alpha}$ and
 (D.21) $|J_{\epsilon}|\leq c\epsilon^{1/2}\sum_{j_{1}=1}^{1/\epsilon}\big{(}j_{1}^{-1/2}+\epsilon^{1/2}+j_{1}^{-1}\big{)}<c,\ \epsilon\to 0.$
-Hence, in what follows, we assume that $Q^{\prime}(\tilde{\alpha})=0$ for some $\tilde{\alpha}\in I_{\alpha}^{\prime}$ and $\tilde{x}\in U^{\prime}$.
-As is easily seen, $Q(\tilde{\alpha})=0$ and $Q^{\prime}(\tilde{\alpha})=0$ for some $\tilde{\alpha}\in\text{dom}\,P$ and $\tilde{x}\in U^{\prime}$ if and only if $\tilde{x}\in\mathcal{S}$ and $P(\alpha)$ is associated with $\tilde{x}$.
-The necessity is established in Lemma D.3 below, and the sufficiency follows from the Bolker condition, assumption 3.2(2).
-Recall that due to possible data redundancy, there can be several $P(\alpha)$ associated with $\tilde{x}$.
-In the rest of this subsection we select one of them.
-We need the following lemma, which is proven in appendix D.4.
+D.3.2. Estimation near a stationary point of $Q$.
+Based on the results of the preceding subsection, it remains to consider the case when $|Q(a_{1})|$, $|Q_{\alpha}^{\prime}(a_{2})|\leq\delta$ for some $a_{1},a_{2}\in I_{\alpha}$.
+Since $I_{\alpha}$ can be made as small as we like (but finitely small), it is instructive to consider the case when $Q(z,\alpha)=Q_{\alpha}^{\prime}(z,\alpha)=0$ for some $(z,\alpha)\in\mathcal{U}\times\mathcal{I}_{\alpha}$.
+This happens if and only if $z\in\mathcal{S}$ and $P(\alpha)$ is associated with $z$.
+To prove sufficiency, suppose there is a point $(z,\alpha)\in\mathcal{U}\times\mathcal{I}_{\alpha}$ such that
+(D.22) $\begin{split}&Q(z,\alpha)=\Phi(z,\alpha)-\Phi(x(\alpha),\alpha)=0,\\ &Q_{\alpha}^{\prime}(z,\alpha)=\Phi_{\alpha}^{\prime}(z,\alpha)-\Phi_{\alpha}^{\prime}(x(\alpha),\alpha)=0.\end{split}$
+The Bolker condition, Assumption 3.2(2), gives $z=x(\alpha)$.
+The necessity is established in the following Lemma D.3, which is proven in Appendix D.4.
 Lemma D.3.
@@
 One has
-(D.21) $Q_{\alpha}^{\prime}(\tilde{x},\tilde{\alpha})=0,\ Q_{\alpha\alpha}^{\prime\prime}(\tilde{x},\tilde{\alpha})=\det M(\tilde{y})\frac{|x_{\alpha}^{\prime}(\tilde{\alpha})|^{2}}{|\Phi_{x}^{\prime}(\tilde{x},\tilde{\alpha})||H_{x}^{\prime}(\tilde{x})|}>0$
+(D.23) $\begin{split}&Q(\tilde{x},\tilde{\alpha})=Q_{\alpha}^{\prime}(\tilde{x},\tilde{\alpha})=0,\\ &Q_{\alpha\alpha}^{\prime\prime}(\tilde{x},\tilde{\alpha})=\det M(\tilde{y})\frac{|x_{\alpha}^{\prime}(\tilde{\alpha})|^{2}}{|\text{d}_{x}\Phi(\tilde{x},\tilde{\alpha})||\text{d}_{x}H(\tilde{x})|}>0\end{split}$
 and
-(D.22) $|\Delta_{\Phi}(\tilde{x},\tilde{\alpha})|=|\det M(\tilde{y})|\frac{|x_{\alpha}^{\prime}(\tilde{\alpha})|}{|H_{x}^{\prime}(\tilde{x})|}.$
-By continuity, $Q_{\alpha\alpha}^{\prime\prime}(x,\alpha)>0$ for all $(x,\alpha)\in U^{\prime}\times I_{\alpha}^{\prime}$.
-D.3.1. Critical point in $I_{\alpha}^{\prime}$
-Fix any $x\in U^{\prime}$. Suppose $Q^{\prime}(a_{1})=0$ for some $a_{1}\in I_{\alpha}^{\prime}$. By shifting $\alpha$, we can assume without loss of generality that $a_{1}=0$ and
-(D.23) $Q(\theta)=-h+\int_{0}^{\theta}(\theta-s)Q^{\prime\prime}(s)\text{d}s$
-for some $h$.
-The value of $h$ is determined from $Q(a_{0})=0$.
-Since $Q^{\prime\prime}(\alpha)\asymp 1$ on $I_{\alpha}^{\prime}$ (see the sentence following Lemma D.3) and $Q(a)=0$, we have $h\geq 0$.
-For any $c>0$, the number of $\alpha_{j_{1}}$ inside the set $\{\alpha\in I_{\alpha}^{\prime}:|Q(\alpha)|\leq c\epsilon\}$ is $O(\epsilon^{-1/2})$ and is uniformly bounded with respect to $a_{0}\in I_{\alpha}^{\prime}$.
+(D.24) $|\Delta_{\Phi}(\tilde{x},\tilde{\alpha})|=|\det M(\tilde{y})|\frac{|x_{\alpha}^{\prime}(\tilde{\alpha})|}{|\text{d}_{x}H(\tilde{x})|}.$
+By the above argument, we can suppose that the point $x$ selected following (D.20) is located sufficiently close to the curve $x(\alpha),\alpha\in I_{\alpha}$.
+Otherwise, by shrinking $I_{\alpha}$ if necessary, we can ensure that either $|Q(\alpha)|\geq\delta$ or $|Q_{\alpha}^{\prime}(\alpha)|\geq\delta$ on $I_{\alpha}$.
+By Lemma D.3, $Q_{\alpha\alpha}^{\prime\prime}(x(\alpha),\alpha)\geq c>0$ for all $\alpha\in I_{\alpha}$.
+By continuity, $Q_{\alpha\alpha}^{\prime\prime}(\alpha)\geq c>0$ for all $\alpha\in I_{\alpha}$.
+Due to the lower bound on $Q_{\alpha\alpha}^{\prime\prime}$, the number of $\alpha_{j_{1}}$ in the set $\{\alpha\in I_{\alpha}:|Q(\alpha)|\leq c\epsilon\}$ is at most $A\epsilon^{-1/2}$, where $A$ is independent of $x$.
 Since $G$ is bounded, the total contribution of such $\alpha_{j_{1}}$ to $J_{\epsilon}$ is bounded.
-Hence it remains to estimate the following two sums
-(D.24) $\begin{split}J_{\epsilon}^{+}:=&\epsilon^{1/2}\sum_{\alpha_{j_{1}}\in I_{\alpha}^{\prime},Q(\alpha_{j_{1}})>\epsilon}\bigg{[}\epsilon^{1/2}+\frac{\epsilon}{Q(\alpha_{j_{1}})}\bigg{]},\\ J_{\epsilon}^{-}:=&\epsilon^{1/2}\sum_{\alpha_{j_{1}}\in I_{\alpha}^{\prime},Q(\alpha_{j_{1}})<-\epsilon}\bigg{(}-\frac{\epsilon}{Q(\alpha_{j_{1}})}\bigg{)}^{1/2}.\end{split}$
+Hence, by (6.19), it remains to estimate the following two sums
+(D.25) $\begin{split}J_{\epsilon}^{+}:=&\epsilon^{1/2}\sum_{\alpha_{j_{1}}\in I_{\alpha}^{+}}\bigg{[}\epsilon^{1/2}+\frac{\epsilon}{Q(\alpha_{j_{1}})}\bigg{]},\ I_{\alpha}^{+}:=\{\alpha\in I_{\alpha}:\,Q(\alpha)\geq\epsilon\},\\ J_{\epsilon}^{-}:=&\epsilon^{1/2}\sum_{\alpha_{j_{1}}\in I_{\alpha}^{-}}\bigg{(}-\frac{\epsilon}{Q(\alpha_{j_{1}})}\bigg{)}^{1/2},\ I_{\alpha}^{-}:=\{\alpha\in I_{\alpha}:\,Q(\alpha)\leq-\epsilon\}.\end{split}$
+For estimation purposes, we can assume $Q^{\prime\prime}(\alpha)\geq 1$, $\alpha\in I_{\alpha}$ (rather than the more cumbersome $Q^{\prime\prime}(\alpha)\geq c$).
+When the argument $x$ of $Q$ is omitted and we write $Q(\alpha)$, then a prime denotes the derivative with respect to $\alpha$.
+Since $Q^{\prime\prime}(\alpha)\geq 1$, $I_{\alpha}^{+}$ is the union of no more than two intervals, and $I_{\alpha}^{-}$ is a single interval.
+Also, the equation $Q^{\prime}(\alpha)=0$, $\alpha\in I_{\alpha}$, has at most one solution.
 Begin with $J_{\epsilon}^{+}$.
-Clearly the sums over $\alpha_{j_{1}}>0$ and $\alpha_{j_{1}}<0$ lead to the same upper bound, so we consider only $\alpha_{j_{1}}>0$.
-Also, for estimation purposes, we can assume $Q^{\prime\prime}(s)\geq 1$, $s\in I_{\alpha}^{\prime}$ (rather than the more cumbersome $Q^{\prime\prime}(s)\geq c$).
-Let $\theta_{1}>0$ satisfy
-(D.25) $Q(\theta_{1})=-h+\int_{0}^{\theta_{1}}(\theta_{1}-s)Q^{\prime\prime}(s)\text{d}s=\epsilon.$
-For any $\theta\geq\theta_{1}$ one has
-(D.26) $Q(\theta)\geq-h+\int_{0}^{\theta_{1}}(\theta_{1}-s)Q^{\prime\prime}(s)\text{d}s+\int_{\theta_{1}}^{\theta}(\theta-s)\text{d}s=\epsilon+(\theta-\theta_{1})^{2}/2.$
-Since the shift $\theta\to\theta-\theta_{1}$ does not affect the upper bound we get
-(D.27) $\begin{split}|J_{\epsilon}^{+}|\leq&c\epsilon^{1/2}\sum_{j_{1}=0}^{1/\epsilon}\bigg{[}\epsilon^{1/2}+\frac{\epsilon}{\epsilon+(\epsilon j_{1})^{2}}\bigg{]}\leq c+c\epsilon^{1/2}\int_{0}^{1}\frac{\text{d}x}{\epsilon+x^{2}}=O(1).\end{split}$
+Let $[a,b]$ be one of at most three intervals that make up $I_{\alpha}^{+}$, subject to the condition $Q^{\prime}(\alpha)\not=0$ on $(a,b)$.
+Suppose $Q^{\prime}(\alpha)>0$ on $(a,b)$.
+The proof in the case $Q^{\prime}(\alpha)<0$ is analogous.
+By construction, $Q(a)\geq\epsilon$.
+Consider the function $Q_{1}(\alpha)=\epsilon+(\alpha-a)^{2}/2$, $\alpha\in[a,b]$, which satisfies $Q_{1}(a)=\epsilon$, $Q_{1}^{\prime}(a)=0$, and $Q_{1}^{\prime\prime}(\alpha)=1$ on $[a,b]$.
+Clearly, $Q_{1}(\alpha)\leq Q(\alpha)$, $\alpha\in[a,b]$.
+Therefore, the value of $J_{\epsilon}^{+}$ is increased if we replace $Q$ with $Q_{1}$.
+There are $O(1/\epsilon)$ terms in the first sum in (D.25), so the term $O(\epsilon^{1/2})$ in the brackets in $J_{\epsilon}^{+}$ contributes an $O(1)$ term.
+Then
+(D.26) $\begin{split}J_{\epsilon}^{+}\leq&c+c\epsilon^{1/2}\sum_{j}\frac{\epsilon}{\epsilon+(\epsilon j)^{2}/2}\leq c.\end{split}$
 Next, consider $J_{\epsilon}^{-}$.
-If $h<\epsilon$, then $Q(\theta)\geq-\epsilon$ for all $\alpha\in I_{\alpha}^{\prime}$ and $J_{\epsilon}^{-}=0$.
-Hence we can assume $h>\epsilon$.
-In this case we initially assume for simplicity that $Q(0)=Q(a)=-\epsilon$ for some $a>0$.
-Therefore
-(D.28) $\begin{split}-(Q(\theta)+\epsilon)=&-\int_{0}^{\theta}(\theta-s)Q^{\prime\prime}(s)\text{d}s+\theta\int_{0}^{a}(1-(s/a))Q^{\prime\prime}(s)\text{d}s\\ =&\int_{0}^{\theta}s(1-(\theta/a))Q^{\prime\prime}(s)\text{d}s+\theta\int_{\theta}^{a}(1-(s/a))Q^{\prime\prime}(s)\text{d}s.\end{split}$
-Both integrands on the second line are positive, so the minimum on the right is achieved when $Q^{\prime\prime}(s)\equiv 1$.
-Hence, shifting the argument of $Q$ again we can assume $Q(\theta)=-h+\theta^{2}/2$.
-Similarly to (D.27),
-(D.29) $\begin{split}|J_{\epsilon}^{-}|\leq&c\epsilon^{1/2}\sum_{h-(\epsilon j_{1})^{2}>\epsilon}\bigg{[}\frac{\epsilon}{h-(\epsilon j_{1})^{2}}\bigg{]}^{1/2}\leq c\int_{0}^{h^{1/2}}\frac{\text{d}x}{(h-x^{2})^{1/2}}<\infty.\end{split}$
-Note that the value of the last integral is independent of $h>0$.
-It is clear that in the above arguments all the bounds can be made uniform with respect to $x\in U^{\prime}$.
+Let $[a,b]$ be one of at most two intervals that make up $I_{\alpha}^{-}$, subject to the condition $Q^{\prime}(\alpha)\not=0$ on $(a,b)$.
+Again, we can assume $Q^{\prime}(\alpha)>0$ on $(a,b)$.
+Consider the function $Q_{1}(\alpha)=Q(a)+(\alpha-a)^{2}/2$, $\alpha\geq a$, which satisfies $Q_{1}(a)=Q(a)$, $Q_{1}^{\prime}(a)=0$, and $Q_{1}^{\prime\prime}(\alpha)=1$ for $\alpha\geq a$ (see the green curve in Fig.
+5).
+By construction, $Q_{a}^{\prime}(\alpha)\leq Q^{\prime}(\alpha)$, $\alpha\in[a,b]$.
+Pick any $t\in[a,b]$, and find $t_{1}$ by solving $Q_{1}(t_{1})=Q(t)$, i.e.
+$t_{1}=a+[2(Q(t)-Q(a))]^{1/2}$ (see Fig.
+5).
+We claim that $Q_{1}^{\prime}(t_{1})\leq Q^{\prime}(t)$.
+Figure 5. Construction used to estimate $J_{\epsilon}^{-}$.
+To prove the claim, we need to show that $t_{1}-a=[2(Q(t)-Q(a))]^{1/2}<Q^{\prime}(t)$. This is equivalent to
+(D.27) $2\bigg{[}Q^{\prime}(a)t+\int_{a}^{t}(t-s)Q^{\prime\prime}(s)\text{d}s\bigg{]}\leq\bigg{[}Q^{\prime}(a)+\int_{a}^{t}Q^{\prime\prime}(s)\text{d}s\bigg{]}^{2},\ a\leq t\leq b.$
+Squaring the right-hand side and using that $Q^{\prime}(a)\geq 0$ and $Q^{\prime\prime}(s)\geq 1$, we see that the desired result follows if we prove
+(D.28) $2\int_{a}^{t}(t-s)Q^{\prime\prime}(s)\text{d}s\leq\bigg{[}\int_{a}^{t}Q^{\prime\prime}(s)\text{d}s\bigg{]}^{2}.$
+Writing $Q^{\prime\prime}(s)=1+\phi(s)$, where $\phi(s)\geq 0$, and simplifying proves the claim.
+Let $b_{1}$ be obtained by solving $Q_{1}(b_{1})=Q(b)$.
+We have $Q_{1}^{\prime}(\alpha)\leq Q^{\prime}(\alpha)$, so $b_{1}\geq b$.
+The preceding argument implies that if the graph of $Q_{1}$ is shifted by $b_{1}-b$ to the left (to obtain the blue curve in Fig.
+5 as indicated by the black arrow), then it will lie above the graph of $Q$ (shown in red).
+This means that $Q_{1}(\alpha+(b_{1}-b))\geq Q(\alpha)$, $\alpha\in[a,b]$.
+Hence, if we replace $Q(\alpha)$ by $Q_{1}(\alpha+(b_{1}-b))$ and $[a,b]$ by $[a-(b_{1}-b),b]$, then the value of $J_{\epsilon}^{-}$ is increased.
+The shift, of course, is irrelevant for estimation purposes, so we instead consider $Q_{1}(\alpha)$ and $[a,b_{1}]$.
+This leads to the following bound
+(D.29) $\begin{split}J_{\epsilon}^{-}&\leq c\epsilon^{1/2}\sum_{h+(\epsilon j)^{2}/2<-\epsilon}\bigg{[}\frac{\epsilon}{-h-(\epsilon j)^{2}/2}\bigg{]}^{1/2}\leq c\int_{0}^{|2h|^{1/2}}\frac{\text{d}x}{(|2h|-x^{2})^{1/2}}<\infty.\end{split}$
+Note that the value of the last integral is independent of $h$.
+It is clear that in the above arguments all the bounds are uniform with respect to $x\in U^{\prime}$.
 D.4. Proof of Lemma D.3
@@
 (D.30) $Q(\alpha)=\Phi(\tilde{x},\alpha)-P(\alpha)=\Phi(\tilde{x},\alpha)-\Phi(x(\alpha),\alpha).$
-The first claim in (D.21) follows by differentiating (4.8) and (D.30) with respect to $\alpha$ and using that $x^{\prime}(\alpha)$ is tangent to $\mathcal{S}$, $x(\tilde{\alpha})=\tilde{x}$, and $\text{d}_{x}\Phi(\tilde{x},\tilde{\alpha})$ is conormal to $\mathcal{S}$ at $\tilde{x}$.
-Differentiating (D.30) with respect to $\alpha$ twice and using (6.7) gives:
-(D.31) $Q_{\alpha\alpha}^{\prime\prime}=-\big{[}(\Phi_{xx}^{\prime\prime}x_{\alpha}^{\prime},x_{\alpha}^{\prime})+2\Phi_{x\alpha}^{\prime\prime}x_{\alpha}^{\prime}+\Phi_{x}^{\prime}x_{\alpha\alpha}^{\prime\prime}\big{]}.$
-Differentiate the second equation in (4.8) with respect to $\alpha$:
-(D.32) $H_{xx}^{\prime\prime}x_{\alpha}^{\prime}=\nu_{1}^{\prime}\Phi_{x}^{\prime}+\nu_{1}\big{[}\Phi_{xx}^{\prime\prime}x_{\alpha}^{\prime}+\Phi_{x\alpha}^{\prime\prime}\big{]}.$
-Compute the dot product of (D.32) with $x_{\alpha}^{\prime}$ and use that $\Phi_{x}^{\prime}x_{\alpha}^{\prime}\equiv H_{x}^{\prime}x_{\alpha}^{\prime}\equiv 0$ (which follows from (4.8) to get
-(D.33) $(H_{xx}^{\prime\prime}x_{\alpha}^{\prime},x_{\alpha}^{\prime})=\nu_{1}\big{[}(\Phi_{xx}^{\prime\prime}x_{\alpha}^{\prime},x_{\alpha}^{\prime})+\Phi_{x\alpha}^{\prime\prime}x_{\alpha}^{\prime}\big{]}.$
-Differentiate the identity $\Phi_{x}^{\prime}x_{\alpha}^{\prime}\equiv 0$ to get
-(D.34) $(\Phi_{xx}^{\prime\prime}x_{\alpha}^{\prime},x_{\alpha}^{\prime})+\Phi_{x\alpha}^{\prime\prime}x_{\alpha}^{\prime}+\Phi_{x}^{\prime}x_{\alpha\alpha}^{\prime\prime}\equiv 0.$
-Using (D.33) and (D.34) in (D.31) and applying Lemma 4.1 finally gives:
-(D.35) $\begin{split}Q_{\alpha\alpha}^{\prime\prime}=&-\big{[}(1/\nu_{1})(H_{xx}^{\prime\prime}x_{\alpha}^{\prime},x_{\alpha}^{\prime})-(\Phi_{xx}^{\prime\prime}x_{\alpha}^{\prime},x_{\alpha}^{\prime})\big{]}\\ =&-\big{(}\big{[}H_{xx}^{\prime\prime}-\nu_{1}\Phi_{xx}^{\prime\prime}\big{]}e,e\big{)}\frac{|x_{\alpha}^{\prime}|^{2}}{\nu_{1}},\ e:=x_{\alpha}^{\prime}/|x_{\alpha}^{\prime}|.\end{split}$
-Combining this with (B.1) and (B.2) proves the second claim in (D.21).
-Next we prove (D.22). Similarly to (B.2),
-(D.36) $|\Delta_{\Phi}(x,\alpha)|=|\Phi_{x}^{\prime}||\Phi_{x\alpha}^{\prime\prime}e|.$
-Using (D.33) and then (B.2) gives
-(D.37) $\begin{split}|\Phi_{x}^{\prime}||\Phi_{x\alpha}^{\prime\prime}e|&=\big{|}\big{(}[H_{xx}^{\prime\prime}-\nu_{1}\Phi_{xx}^{\prime\prime}]e,e\big{)}\big{|}\frac{|x_{\alpha}^{\prime}||\Phi_{x}^{\prime}|}{|\nu_{1}|}=\frac{|\text{det}M||x_{\alpha}^{\prime}|}{|\nu_{1}||\Phi_{x}^{\prime}|},\end{split}$
+By construction, $x^{\prime}(\alpha)$ is tangent to $\mathcal{S}$ and $\text{d}_{x}\Phi(x(\alpha),\alpha)$ is conormal to $\mathcal{S}$ at $x(\alpha)$ (see (4.8)).
+Differentiation of (D.30) with respect to $\alpha$ gives
+(D.31) $\begin{split}Q^{\prime}(\alpha)=&\Phi_{\alpha}^{\prime}(\tilde{x},\alpha)-\big{[}\text{d}_{x}\Phi(x(\alpha),\alpha)\cdot x_{\alpha}^{\prime}+\Phi_{\alpha}^{\prime}(x(\alpha),\alpha)\big{]}\\ =&\Phi_{\alpha}^{\prime}(\tilde{x},\alpha)-\Phi_{\alpha}^{\prime}(x(\alpha),\alpha).\end{split}$
+Using that $x(\tilde{\alpha})=\tilde{x}$ proves the first line in (D.23).
+Next, differentiate (D.31) with respect to $\alpha$ and set $\alpha=\tilde{\alpha}$ to obtain
+(D.32) $Q_{\alpha\alpha}^{\prime\prime}=-\text{d}_{x}\Phi_{\alpha}^{\prime}\cdot x_{\alpha}^{\prime}.$
+Differentiate the second equation in (4.8) with respect to $\alpha$ and set $\alpha=\tilde{\alpha}$:
+(D.33) $\text{d}_{x}^{2}H\,x_{\alpha}^{\prime}=\nu_{1}^{\prime}\text{d}_{x}\Phi+\nu_{1}\big{[}\text{d}_{x}^{2}\Phi\,x_{\alpha}^{\prime}+\text{d}_{x}\Phi_{\alpha}^{\prime}\big{]}.$
+Apply (D.33) to $x_{\alpha}^{\prime}$ and use that $\text{d}_{x}\Phi\cdot x_{\alpha}^{\prime}\equiv 0$ a second time to get
+(D.34) $\text{d}_{x}^{2}H(x_{\alpha}^{\prime},x_{\alpha}^{\prime})=\nu_{1}\big{[}\text{d}_{x}^{2}\Phi(x_{\alpha}^{\prime},x_{\alpha}^{\prime})+\text{d}_{x}\Phi_{\alpha}^{\prime}\cdot x_{\alpha}^{\prime}\big{]}.$
+Using (D.34) in (D.32) and applying Lemma 4.1 finally gives:
+(D.35) $\begin{split}Q_{\alpha\alpha}^{\prime\prime}=&-\big{[}(1/\nu_{1})\text{d}_{x}^{2}H(x_{\alpha}^{\prime},x_{\alpha}^{\prime})-\text{d}_{x}^{2}\Phi(x_{\alpha}^{\prime},x_{\alpha}^{\prime})\big{]}\\ =&-\text{d}_{x}^{2}[H-\nu_{1}\Phi](e,e)\frac{|x_{\alpha}^{\prime}|^{2}}{\nu_{1}},\ e:=x_{\alpha}^{\prime}/|x_{\alpha}^{\prime}|.\end{split}$
+Combining this with (B.1) and (B.4) proves the second line in (D.23).
+Next we prove (D.24).
+By (3.4), (3.5), $\Delta_{\Phi}(x,\alpha)=\det\begin{pmatrix}\text{d}_{x}\Phi\\ \text{d}_{x}\Phi_{\alpha}^{\prime}\end{pmatrix}$.
+Considering the matrix product $\Delta_{\Phi}(x,\alpha)V$, where $V=\begin{pmatrix}e_{0}&e\end{pmatrix}$ (see (B.2)), gives
+(D.36) $|\Delta_{\Phi}(x,\alpha)|=|\text{d}_{x}\Phi||\text{d}_{x}\Phi_{\alpha}^{\prime}\cdot e|.$
+Using (D.32), (D.35), and then (B.4) gives
+(D.37) $\begin{split}|\text{d}_{x}\Phi||\text{d}_{x}\Phi_{\alpha}^{\prime}\cdot e|&=\big{|}\text{d}_{x}^{2}[H-\nu_{1}\Phi](e,e)\big{|}\frac{|\text{d}_{x}\Phi||x_{\alpha}^{\prime}|}{|\nu_{1}|}=\frac{|\text{det}M||x_{\alpha}^{\prime}|}{|\nu_{1}||\text{d}_{x}\Phi|},\end{split}$
 and (D.24) follows from (B.1).
@@
 E.1. Proof of Lemma 7.1
-For convenience, we drop the arguments $\check{q}$, $\theta$ and $\epsilon$ from all the functions.
-By (LABEL:H_props_2_a) with $k=1$ and (LABEL:H_props_2_b), $\Delta H(r,\check{t})=O(\check{t}^{-3/2})$, $\check{t}\to+\infty$.
-The statement $\Delta G=O(1/|\check{t}|)$, $\check{t}\to-\infty$, follows immediately from (D.1).
-Next, suppose $\check{t}\to+\infty$.
-Then
-(E.1) $\begin{split}&\Delta G(r,\check{t})=\int_{\mathbb{R}}\chi_{1\text{D}}(\epsilon(s-\check{t}))\frac{\Delta H(r,s)}{s-\check{t}}\text{d}s.\end{split}$
-In view of (D.1), the integral in (E.1) is split into three
-(E.2) $\begin{split}&\bigg{(}\int_{-c}^{\check{t}-1}+\int_{\check{t}-1}^{\check{t}+1}+\int_{\check{t}+1}^{\infty}\bigg{)}\chi_{1\text{D}}(\epsilon(s-\check{t}))\frac{\Delta H(r,s)}{s-\check{t}}\text{d}s\\ &=J_{1}+J_{2}+J_{3}.\end{split}$
-Breaking up the integral over $[-c,\check{t}-1]$ into the integrals over $[-c,\check{t}/2]$ and $[\check{t}/2,\check{t}-1]$ we get $J_{1}=O(1/\check{t})$.
-Breaking up the integral over $[\check{t}+1,\infty)$ into the integrals over $[\check{t}+1,2\check{t}]$ and $[2\check{t},\infty)$ gives $J_{3}=O(1/\check{t})$.
-Also, from (LABEL:H_props_2_a) with $k=1$ and (LABEL:H_props_2_b) we get
-(E.3) $|H_{\check{t}}^{\prime}(\check{t}+r)-H_{\check{t}}^{\prime}(\check{t})|\leq O(\check{t}^{-3/2})+c\big{|}\partial_{\check{t}}\big{[}\chi(\epsilon(\check{t}+\check{q}))\check{t}^{-1/2}\big{]}\big{|}=O(\check{t}^{-3/2}).$
-Therefore, using $|s-\check{t}|\leq 1$ gives
-(E.4) $\partial_{s}\big{[}\chi_{1\text{D}}(\epsilon(s-\check{t}))\Delta H(r,s)\big{]}=\epsilon O(\check{t}^{-3/2})+O(\check{t}^{-3/2}),\ \check{t}\to+\infty.$
-Combining the three estimates proves that the integral in (E.1) is $O(1/\check{t})$. Combining the results finishes the proof of the estimate of $\Delta G$.
-The derivative $\partial_{r}\Delta G(r,\check{t})$ in (6.16) is estimated in exactly the same fashion.
-In particular, we use $\partial_{r}\Delta H(r,\check{t})=\partial_{r}H(r+\check{t})=O(\check{t}^{-3/2})$ as $\check{t}\to+\infty$.
-To prove (7.6), use (E.1) to write
-(E.5) $\begin{split}&\Delta G(r,\check{t}+h)-\Delta G(r,\check{t})\\ &=\bigg{(}\int_{-1}^{1}+\int_{|s|>1}\bigg{)}\bigg{[}\frac{\chi_{1\text{D}}(\epsilon(s-h))}{s-h}-\frac{\chi_{1\text{D}}(\epsilon s)}{s}\bigg{]}\Delta H(r,s+\check{t})\text{d}s\\ &=:J_{1}(\check{t},h)+J_{2}(\check{t},h).\end{split}$
-When $0<\epsilon\ll 1$ and $|s|<1$, we have $\chi_{1\text{D}}(\epsilon(s-h))\equiv\chi_{1\text{D}}(\epsilon s)\equiv 1$. Then
-(E.6) $\begin{split}J_{1}(\check{t},h)&=\bigg{[}\ln\bigg{|}\frac{s-h}{s}\bigg{|}\Delta H(r,s+\check{t})\bigg{]}\bigg{|}_{s=-1}^{1}-\int_{-1}^{1}\Delta H_{\check{t}}^{\prime}(r,s+\check{t})\ln\bigg{|}\frac{s-h}{s}\bigg{|}\text{d}s\\ &=:J_{11}(\check{t},h)+J_{12}(\check{t},h).\end{split}$
-By (LABEL:H_props_2_a) with $k=0$ and (LABEL:H_props_2_b),
-(E.7) $\begin{split}\bigg{|}\ln\bigg{|}\frac{s-h}{s}\bigg{|}\Delta H(r,s+\check{t})\bigg{|}\leq c\frac{h\ln(1/h)}{(1+|\check{t}|)^{3/2}},\ s=\pm 1.\end{split}$
-To bound all the remaining terms, we can look at them individually instead of estimating differences at $\check{t}+r$ and $\check{t}$ as above. Indeed,
-(E.8) $\begin{split}|J_{12}(\check{t},h)|&\leq\int_{-1}^{1}\big{|}H^{\prime}(s+\check{t})\ln|1-(h/s)|\big{|}\text{d}s\\ &\leq c\int_{-1}^{1}(1+|s+\check{t}|)^{-3/2}\big{|}\ln|1-(h/s)|\big{|}\text{d}s\\ &\leq c\frac{h}{(1+|\check{t}|)^{3/2}}\int_{-1/h}^{1/h}\big{|}\ln|1-(1/u)|\big{|}\text{d}u\leq c\frac{h\ln(1/h)}{(1+|\check{t}|)^{3/2}}\end{split}$
-and
-(E.9) $\begin{split}&|J_{2}(\check{t},h)|\leq ch\int_{|s|>1}\bigg{(}\frac{1}{s^{2}}+\frac{\epsilon}{|s|}\bigg{)}\frac{1}{(1+|s+\check{t}|)^{3/2}}\text{d}s\\ &\leq ch(|\check{t}|^{-3/2}+\epsilon|\check{t}|^{-1}).\end{split}$
-Combining the results finishes the proof.
+For convenience, we drop the arguments $\check{q}$, $\theta$ and $\epsilon$ from all the functions. By (6.17) and (7.5),
+(E.1) $\begin{split}&\Delta G(r,\check{t})=G(r+\check{t})-G(\check{t})=\int_{\mathbb{R}}\chi_{1\text{D}}(\epsilon(s-\check{t}))\frac{\Delta H(r,s)}{s-\check{t}}\text{d}s.\end{split}$
+Begin with (7.7).
+Adding (LABEL:H_props_2_a) with $k=1$ and (LABEL:H_props_2_b) yields $|\Delta H(r,\check{t})|\leq c|r|(1+|\check{t}|)^{-3/2}$.
+The bound $|\Delta G|\leq c|r|/|\check{t}|$, $\check{t}\to-\infty$, follows from (D.1).
+Suppose $\check{t}\to+\infty$. Using (D.1), the integral in (E.1) is split into three
+(E.2) $\begin{split}&\bigg{(}\int_{-c}^{\check{t}-1}+\int_{\check{t}-1}^{\check{t}+1}+\int_{\check{t}+1}^{\infty}\bigg{)}\chi_{1\text{D}}(\epsilon(s-\check{t}))\frac{\Delta H(r,s)}{s-\check{t}}\text{d}s\\ &=:J_{1}+J_{2}+J_{3}.\end{split}$
+Breaking up the integral over $[-c,\check{t}-1]$ into the integrals over $[-c,\check{t}/2]$ and $[\check{t}/2,\check{t}-1]$ we get $|J_{1}|\leq c|r|/\check{t}$.
+Breaking up the integral over $[\check{t}+1,\infty)$ into the integrals over $[\check{t}+1,2\check{t}]$ and $[2\check{t},\infty)$ gives $|J_{3}|\leq c|r|\check{t}^{-3/2}\ln\check{t}$.
+Using the function $(\chi_{1\text{D}}(\epsilon\check{w})-1)/\check{w}$, we obtain
+(E.3) $\begin{split}&J_{2}=J_{2}^{\prime}+O(\epsilon|r|\check{t}^{-3/2}),\\ &J_{2}^{\prime}=\int_{-1}^{1}\frac{1}{\check{w}}\Delta H(r,\check{w}+\check{t})\text{d}\check{w}=\int_{-1}^{1}\frac{1}{\check{w}}\int_{w}^{r+w}\partial_{\check{t}}H(s+\check{t})\text{d}s\text{d}\check{w}.\end{split}$
+To estimate $J_{2}^{\prime}$, we may assume without loss of generality that $10<r<$ 1 and rewrite $J_{2}^{\prime}$ as follows
+(E.4) $\begin{split}J_{2}^{\prime}=&\int_{-1}^{-1+r}\partial_{\check{t}}H(s+\check{t})\int_{-1}^{s}\frac{\text{d}\check{w}}{\check{w}}\text{d}s+\int_{-1+r}^{1}\partial_{\check{t}}H(s+\check{t})\int_{s-r}^{s}\frac{\text{d}\check{w}}{\check{w}}\text{d}s\\ &+\int_{1}^{1+r}\partial_{\check{t}}H(s+\check{t})\int_{s-r}^{1}\frac{\text{d}\check{w}}{\check{w}}\text{d}s.\end{split}$
+Adding (LABEL:H_props_2_a) with $k=1$ and (LABEL:H_props_2_b) yields $\partial_{\check{t}}H(\check{t})=O(\check{t}^{-3/2})$. After simple but tedious calculations, (E.4) implies:
+(E.5) $\begin{split}|J_{2}^{\prime}|\leq&c\check{t}^{-3/2}\bigg{[}\int_{-1}^{-1+r}|\ln|s||\text{d}s+\int_{-1+r}^{1}\bigg{|}\ln\frac{|s|}{|s-r|}\bigg{|}\text{d}s+\int_{1}^{1+r}|\ln|s-r||\text{d}s\bigg{]}\\ =&O(r\ln(1/r)\check{t}^{-3/2}).\end{split}$
+Therefore, $J_{2}=O(r\ln(1/r)\check{t}^{-3/2})$.
+The above argument also proves that $\Delta G(r,\check{t})$ is bounded on compact sets. Combining the estimates of $J_{1},J_{2},J_{3}$ finishes the proof of (7.7).
+It is easy to check that
+(E.6) $\begin{split}&\Delta G(r+h,\check{t})-\Delta G(r,\check{t})=\Delta G(h,r+\check{t}),\\ &\Delta G(r,\check{t}+h)-\Delta G(r,\check{t})=\Delta G(h,r+\check{t})-\Delta G(h,\check{t}),\end{split}$
+Since $|r|\leq A_{2}$, (7.8) and (7.9) follow from (7.7).
 E.2. Proof of Lemma 7.2
+Recall that the parameter $A\gg 1$ used in Lemma 7.2 was selected above the equation (7.13).
 First we investigate the dependence of $G$ (which is implicitly used in (7.16) as part of $\Delta G$) on $\epsilon$ in its last argument.
@@
 Lemma E.1.
-Pick any $A_{1}>0$. Suppose $|\check{t}|<A_{1}$, $|\epsilon\check{q}-P(0)|<\epsilon^{1/2}A_{1}$, and $|\theta|<\epsilon^{1/2}A_{1}$. Under the assumptions of Theorem 3.6 one has
-(E.10) $\begin{split}G(\check{t},\check{q};\theta,\epsilon)=\mu G_{0}(\check{t},\check{q})+O(\epsilon^{1/2}).\end{split}$
+Pick any $A_{1}\gg 1$. Suppose $|\check{t}|<A_{1}$, $|\epsilon\check{q}-P(0)|<\epsilon^{1/2}A_{1}$, and $|\theta|<\epsilon^{1/2}A_{1}$. Under the assumptions of Theorem 3.6 one has
+(E.7) $\begin{split}G(\check{t},\check{q};\theta,\epsilon)=\mu G_{0}(\check{t},\check{q})+O(\epsilon^{1/2})\end{split}$
+uniformly in $\check{t}$, $\check{q}$, and $\theta$ restricted to the indicated sets.
 Partition the interval $[-A,A]$ into the union of $K\gg 1$ non-overlapping intervals of length $2A/K$.
 Let these intervals be denoted $B_{k}$, and $s_{k}$ be their centers, $k=1,2,\dots,K$.
-Recall that $P(\alpha)$ is associated with $x_{0}\in\mathcal{S}$.
+Recall that $P(\alpha)$ is associated with $x_{0}=0_{2}\in\mathcal{S}$.
 Clearly,
 (E.8) $\begin{split}\frac{P(\epsilon^{1/2}s)}{\epsilon}=&\frac{P(\epsilon^{1/2}s_{k})}{\epsilon}+P^{\prime}(\epsilon^{1/2}s_{k})\frac{s-s_{k}}{\epsilon^{1/2}}+O(1/K^{2})\\ =&\frac{P(\epsilon^{1/2}s_{k})}{\epsilon}+\big{[}P^{\prime}(0)+O(\epsilon^{1/2}s_{k})\big{]}\frac{s-s_{k}}{\epsilon^{1/2}}+O(1/K^{2})\\ =&\frac{P(\epsilon^{1/2}s_{k})}{\epsilon}+P^{\prime}(0)\frac{s-s_{k}}{\epsilon^{1/2}}+O(1/K),\ s\in B_{k}.\end{split}$
-Here we have used that $A>0$ is fixed and $|s_{k}|<A$.
-Otherwise, the term $O(1/K)$ may grow with $A$.
-The same applies to many other bounds and big-$O$ terms in this section.
-Using that $H_{0}$ is compactly supported and $|\check{t}|<A$, we get from (7.16)
-(E.12) $\begin{split}&G_{0}(\check{t}+a,\check{q})-G_{0}(\check{t},\check{q})=-\int_{-c}^{c}\ln\bigg{|}\frac{w-a}{w}\bigg{|}\partial_{\check{t}}H_{0}(w+\check{t},\check{q})\text{d}w,\ |\check{t}|<A.\end{split}$
-Hence, by assumption 3.4(1) and (7.16), $|\partial_{\check{t}}H_{0}(w+\check{t},\check{q})|<c$ and
-(E.13) $\begin{split}&|G_{0}(\check{t}+a,\check{q})-G_{0}(\check{t},\check{q})|<c\int_{-c}^{c}\bigg{|}\ln\bigg{|}\frac{w-a}{w}\bigg{|}\bigg{|}\text{d}w<ca\ln(1/a),\\ &0<a<1,\ |\check{t}|<A.\end{split}$
-Furthermore, assumption 3.4(1) implies
+Here we have used that $A>0$ is fixed and $|s_{k}|<A$. Otherwise, the term $O(1/K)$ may grow with $A$. The same applies to many other bounds and big-$O$ terms in this section.
+Observe that $G_{0}(\check{t},\check{q})$ in (7.18) is obtained from $G(\check{t},\check{q};x,\theta,\epsilon)$ by setting $x=0_{2}$, $\theta=0$, $\epsilon=0$ and $a_{0}(\theta,\epsilon j_{2})\equiv a_{0}(y_{0})$ in (6.13), (6.14) and (6.17).
+Therefore, $G_{0}$ satisfies (7.7):
 […]
-Hence
+To study the smoothness of $G_{0}$ with respect to the second argument, denote similarly to (7.18):
+(E.10) $\begin{split}&\check{G}_{0}(\check{t},\check{q}_{1},\check{q}_{2}):=\int_{\mathbb{R}}\frac{H_{0}(\check{w}+\check{t},\check{q}_{1},\check{q}_{2})}{\check{w}}\text{d}\check{w},\\ &\check{H}_{0}(\check{t},\check{q}_{1},\check{q}_{2}):=a_{0}(y_{0})\sum_{j_{2}\in\mathbb{Z}}h_{0}^{\prime}(\check{t}+\check{q}_{1}-j_{2})(j_{2}-\check{q}_{2})_{+}^{1/2}.\end{split}$
+Then, $G_{0}(\check{t},\check{q})=\check{G}_{0}(\check{t},\check{q},\check{q})$. Furthermore,
+(E.11) $\begin{split}G_{0}(\check{t},\check{q}+r)-G_{0}(\check{t},\check{q})=&\big{[}G_{0}(\check{t},\check{q}+r)-G_{0}(\check{t}-r,\check{q}+r)\big{]}\\ &+\big{[}\check{G}_{0}(\check{t},\check{q},\check{q}+r)-\check{G}_{0}(\check{t},\check{q},\check{q})\big{]}.\end{split}$
+The difference in the first brackets is bounded as in (E.9).
+Clearly, $G_{0}(\check{t},\check{q}+1)\equiv G_{0}(\check{t},\check{q})$. Hence, we can assume $10\leq\check{q}<$ 1. Using that $h_{0}$ is compactly supported and $|\check{t}|<A$, we obtain
+(E.12) $\begin{split}\int_{\mathbb{R}}\frac{h_{0}^{\prime}(\check{w}+\check{t}+\check{q}-j_{2})}{\check{w}}\text{d}\check{w}=O(j_{2}^{-2}),\ j_{2}\to\infty.\end{split}$
+Also,
+[…]
+The estimates (E.12), (E.13) and (E.10) imply that the difference in the second brackets in (E.11) is $O(r^{1/2})$. Together with (E.9) this gives
 […]
@@
 (E.15) $\begin{split}G\bigg{(}\check{t},\frac{P(\epsilon^{1/2}s)}{\epsilon};\theta,\epsilon\bigg{)}=&\mu G_{0}\bigg{(}\check{t},\frac{P(\epsilon^{1/2}s)}{\epsilon}\bigg{)}+O(\epsilon^{1/2})\\ =&\mu G_{0}\bigg{(}\check{t},\frac{P(\epsilon^{1/2}s_{k})}{\epsilon}+P^{\prime}(0)\frac{s-s_{k}}{\epsilon^{1/2}}\bigg{)}\\ &+O(1/K^{1/2})+O(\epsilon^{1/2}),\end{split}$
-where $s=\epsilon^{1/2}\mu j_{1}\in B_{k}$ and $|\theta|<\epsilon^{1/2}A$. The equation (E.10) applies because by the assumptions of the lemma,
+where $s=\epsilon^{1/2}\mu j_{1}\in B_{k}$ and $|\theta|<\epsilon^{1/2}A$. The equation (E.7) applies because the assumptions of Lemma 7.2 imply
 (E.16) $\begin{split}&\check{t}=\frac{Q_{\alpha\alpha}^{\prime\prime}}{2}\epsilon(\mu j_{1})^{2}=O(1),\\ &\theta=\alpha_{j_{1}}=\epsilon\mu j_{1}=O(\epsilon^{1/2}),\\ &\epsilon\check{q}=P(\epsilon\mu j_{1})=P(0)+O(\epsilon^{1/2}),\ \epsilon^{1/2}\mu|j_{1}|\leq A,\end{split}$
@@
 (E.17) $\begin{split}G\bigg{(}\frac{Q_{\alpha\alpha}^{\prime\prime}}{2}s^{2},\frac{P(\epsilon^{1/2}s)}{\epsilon};\theta,\epsilon\bigg{)}=&\mu G_{0}\bigg{(}\frac{Q_{\alpha\alpha}^{\prime\prime}}{2}s_{k}^{2},\frac{P(\epsilon^{1/2}s_{k})}{\epsilon}+P^{\prime}(0)\frac{s-s_{k}}{\epsilon^{1/2}}\bigg{)}\\ &+O(1/K^{1/2})+O(\epsilon^{1/2}),\ s\in B_{k},|\theta|<\epsilon^{1/2}A.\end{split}$
-There are $|B_{k}|/(\epsilon^{1/2}\mu)$ values of $j_{1}$ such that $\epsilon^{1/2}\mu j_{1}\in B_{k}$. Sum (E.18) with $s=\epsilon^{1/2}\mu j_{1}$ to obtain
-(E.19) $\begin{split}\epsilon^{1/2}&\sum_{\epsilon^{1/2}\mu j_{1}\in B_{k}}G\bigg{(}\frac{Q_{\alpha\alpha}^{\prime\prime}}{2}\epsilon(\mu j_{1})^{2},\frac{P(\epsilon\mu j_{1})}{\epsilon};\theta,\epsilon\bigg{)}\\ =&\epsilon^{1/2}\sum_{\epsilon^{1/2}\mu j_{1}\in B_{k}}G_{0}\bigg{(}\frac{Q_{\alpha\alpha}^{\prime\prime}}{2}s_{k}^{2},u_{\epsilon}+P^{\prime}(0)\mu j_{1}\bigg{)}\\ &+|B_{k}|\big{[}O(1/K^{1/2})+O(\epsilon^{1/2})\big{]},\ u_{\epsilon}:=\frac{P(\epsilon^{1/2}s_{k})-P^{\prime}(0)\epsilon^{1/2}s_{k}}{\epsilon}.\end{split}$
+There are $|B_{k}|/(\epsilon^{1/2}\mu)$ values of $j_{1}$ such that $\epsilon^{1/2}\mu j_{1}\in B_{k}$. Substitute $s=\epsilon^{1/2}\mu j_{1}$ into (E.17) and sum with respect to $j_{1}$ to obtain
+(E.18) $\begin{split}\epsilon^{1/2}&\sum_{\epsilon^{1/2}\mu j_{1}\in B_{k}}G\bigg{(}\frac{Q_{\alpha\alpha}^{\prime\prime}}{2}\epsilon(\mu j_{1})^{2},\frac{P(\epsilon\mu j_{1})}{\epsilon};\theta,\epsilon\bigg{)}\\ =&\epsilon^{1/2}\mu\sum_{\epsilon^{1/2}\mu j_{1}\in B_{k}}G_{0}\bigg{(}\frac{Q_{\alpha\alpha}^{\prime\prime}}{2}s_{k}^{2},u_{\epsilon}+P^{\prime}(0)\mu j_{1}\bigg{)}\\ &+|B_{k}|\big{[}O(1/K^{1/2})+O(\epsilon^{1/2})\big{]},\ u_{\epsilon}:=\frac{P(\epsilon^{1/2}s_{k})-P^{\prime}(0)\epsilon^{1/2}s_{k}}{\epsilon}.\end{split}$
+The term $O(1/K^{1/2})$ is uniform with respect to $1\leq k\leq K$ and $0<\epsilon\ll 1$.
+The term $O(\epsilon^{1/2})$ is uniform with respect to $1\leq k\leq K$ and $K\geq 1$.
 It follows from (7.18) and (E.14) that $G_{0}$ is 1-periodic in its last argument and Hölder continuous.
-From (4.8) and (D.30),
-$P^{\prime}(\alpha)=[\Phi(x(\alpha),\alpha)]_{\alpha}^{\prime}=\Phi_{x}^{\prime}(x(\alpha),\alpha)x_{\alpha}^{\prime}(\alpha)+\Phi_{\alpha}^{\prime}(x(\alpha),\alpha)=\Phi_{\alpha}^{\prime}(x(\alpha),\alpha).$
-By the assumption of the theorem, $P^{\prime}(0)\mu$ is irrational, so the points $P^{\prime}(0)\mu j_{1}$ are uniformly distributed mod 1 [35, Exercise 5.6].
-Take the limit $\epsilon\to 0$ in (E.19) and use [35, Corollary 1.2] to find
+From (4.8) and the definition $P(\alpha)=\Phi(x(\alpha),\alpha)$,
+(E.19) $\begin{split}P^{\prime}(\alpha)&=[\Phi(x(\alpha),\alpha)]_{\alpha}^{\prime}=\text{d}_{x}\Phi(x(\alpha),\alpha)\cdot x_{\alpha}^{\prime}(\alpha)+\Phi_{\alpha}^{\prime}(x(\alpha),\alpha)\\ &=\Phi_{\alpha}^{\prime}(x(\alpha),\alpha).\end{split}$
+By the assumption of the theorem, $P^{\prime}(0)\mu$ is irrational, so the points $P^{\prime}(0)\mu j_{1}$ are well distributed mod 1 [38, Definition 5.1 and Example 5.2].
+Take the limit $\epsilon\to 0$ in (E.18) and use [38, Theorem 5.1] to find
 (E.20) $\begin{split}\lim_{\epsilon\to 0}\epsilon^{1/2}&\sum_{\epsilon^{1/2}\mu j_{1}\in B_{k}}G\bigg{(}\frac{Q_{\alpha\alpha}^{\prime\prime}}{2}\epsilon(\mu j_{1})^{2},\frac{P(\epsilon\mu j_{1})}{\epsilon};\epsilon\mu j_{1},\epsilon\bigg{)}\\ =&\frac{|B_{k}|}{\mu}\bigg{[}\mu\int_{0}^{1}G_{0}\bigg{(}\frac{Q_{\alpha\alpha}^{\prime\prime}}{2}s_{k}^{2},\check{q}\bigg{)}\text{d}\check{q}+O(1/K^{1/2})\bigg{]}.\end{split}$
@@
 (E.21) $\begin{split}\lim_{\epsilon\to 0}\epsilon^{1/2}&\sum_{\epsilon^{1/2}\mu|j_{1}|\leq A}G\bigg{(}\frac{Q_{\alpha\alpha}^{\prime\prime}}{2}\epsilon(\mu j_{1})^{2},\frac{P(\epsilon\mu j_{1})}{\epsilon};\epsilon\mu j_{1},\epsilon\bigg{)}\\ =&\int_{|s|\leq A}\int_{0}^{1}G_{0}\bigg{(}\frac{Q_{\alpha\alpha}^{\prime\prime}}{2}s^{2},\check{q}\bigg{)}\text{d}\check{q}\text{d}s.\end{split}$
-Clearly the same arguments apply when we add $\Phi_{x}^{\prime}\check{x}$ to the first argument of $G$. Therefore, combining (E.21), (7.14) and (7.2), proves (7.15).
+Clearly the same arguments apply when we add $\text{d}_{x}\Phi\cdot\check{x}$ to the first argument of $G$, cf. (7.16). Therefore, combining (E.21), (7.16) and (7.4), proves (7.17).
 E.2.1. Proof of Lemma E.1
-Fix any $A_{1}$. Throughout the proof we assume $|\check{t}|<A_{1}$. Fix any $\bar{h}\in C_{0}^{2}(\mathbb{R})$ and denote
+Let $A_{1}$, $\check{t}$, $\check{q}$ and $\theta$ be selected as in the statement of Lemma E.1. Fix any $\bar{h}\in C_{0}^{2}(\mathbb{R})$ and denote
 (E.22) $\begin{split}\mathcal{Q}(u):=\int_{\mathbb{R}}\frac{\chi_{1\text{D}}(\epsilon\check{w})}{\check{w}}\bar{h}^{\prime}(\check{w}+u)\text{d}\check{w}.\end{split}$
@@
 for some $c_{1}>0$, which is the same in both places.
-In view of (6.11), (6.14), and (E.22), we estimate
-(E.24) $\begin{split}&\bigg{|}\sum_{j_{2}}\mathcal{Q}(\check{t}+\check{q}-j_{2};\theta,\epsilon)\big{[}a_{0}(\theta,\epsilon j_{2})-a_{0}(\theta,\epsilon\check{q})\big{]}(j_{2}-\check{q})_{+}^{1/2}\bigg{|}\\ &\leq c\epsilon\sum_{j_{2}=1}^{1/\epsilon}\frac{1}{j_{2}^{2}}j_{2}^{3/2}=O(\epsilon^{1/2}).\end{split}$
+Substitute $\bar{h}(u)=h(u;\theta,\epsilon)$ of (6.13) into (E.22) to obtain $\mathcal{Q}(u;\theta,\epsilon)$. In view of (6.14), (6.17), and (E.22), we estimate
+(E.24) $\begin{split}&\bigg{|}\sum_{j_{2}\in J_{p}}\mathcal{Q}(\check{t}+\check{q}-j_{2};\theta,\epsilon)\big{[}a_{0}(\theta,\epsilon j_{2})-a_{0}(\theta,\epsilon\check{q})\big{]}(j_{2}-\check{q})_{+}^{1/2}\bigg{|}\\ &\leq c\epsilon\sum_{j_{2}=1}^{1/\epsilon}\frac{1}{j_{2}^{2}}j_{2}^{3/2}=O(\epsilon^{1/2}).\end{split}$
 This implies
-(E.25) $\begin{split}G(\check{t},\check{q};\theta,\epsilon)=&a_{0}(\theta,\epsilon\check{q})\sum_{j_{2}}\mathcal{Q}(\check{t}+\check{q}-j_{2};\theta,\epsilon)(j_{2}-\check{q})_{+}^{1/2}+O(\epsilon^{1/2}).\end{split}$
-Using the assumptions of Lemma E.1, we find
-(E.26) $\begin{split}G(\check{t},\check{q};\theta,\epsilon)=&\big{[}a_{0}(0,P(0))+O(\epsilon^{1/2})\big{]}\\ &\times\sum_{j_{2}}\mathcal{Q}(\check{t}+\check{q}-j_{2};\theta,\epsilon)(j_{2}-\check{q})_{+}^{1/2}+O(\epsilon^{1/2})\\ =&a_{0}(0,P(0))\sum_{j_{2}}\mathcal{Q}(\check{t}+\check{q}-j_{2};\theta,\epsilon)(j_{2}-\check{q})_{+}^{1/2}+O(\epsilon^{1/2}).\end{split}$
+(E.25) $\begin{split}G(\check{t},\check{q};\theta,\epsilon)=&a_{0}(\theta,\epsilon\check{q})\sum_{j_{2}\in J_{p}}\mathcal{Q}(\check{t}+\check{q}-j_{2};\theta,\epsilon)(j_{2}-\check{q})_{+}^{1/2}+O(\epsilon^{1/2}).\end{split}$
+Since $|\check{t}|<A_{1}$, (E.23) implies that the sum in (E.25) is absolutely convergent and its value is uniformly bounded. Using the assumptions of Lemma E.1, we find
+(E.26) $\begin{split}G(\check{t},\check{q};\theta,\epsilon)=&\big{[}a_{0}(0,P(0))+O(\epsilon^{1/2})\big{]}\\ &\times\sum_{j_{2}\in J_{p}}\mathcal{Q}(\check{t}+\check{q}-j_{2};\theta,\epsilon)(j_{2}-\check{q})_{+}^{1/2}+O(\epsilon^{1/2})\\ =&a_{0}(0,P(0))\sum_{j_{2}\in J_{p}}\mathcal{Q}(\check{t}+\check{q}-j_{2};\theta,\epsilon)(j_{2}-\check{q})_{+}^{1/2}+O(\epsilon^{1/2}).\end{split}$
 To show that $\chi$ can be dropped from the integral in (E.26) (which is inside the definition of $\mathcal{Q}$), we consider similarly to (E.22):
@@
 for some $c_{1}>0$, which is the same in both places. Given that $|\check{t}|<A_{1}$, this implies
-(E.29) $\begin{split}\sum_{j_{2}}\mathcal{Q}_{c}(\check{t}+\check{q}-j_{2};\theta,\epsilon)(j_{2}-\check{q})_{+}^{1/2}=O(\epsilon^{1/2}).\end{split}$
-Therefore, by (E.26) and (E.29),
-(E.30) $\begin{split}G(\check{t},\check{q};\theta,\epsilon)=&a_{0}(0,P(0))\int_{\mathbb{R}}\frac{1}{\check{w}}\sum_{j_{2}}h^{\prime}(\check{w}+\check{t}+\check{q}-j_{2};\theta,\epsilon)(j_{2}-\check{q})_{+}^{1/2}\text{d}\check{w}\\ &+O(\epsilon^{1/2}).\end{split}$
-Further,
+(E.29) $\begin{split}\sum_{j_{2}\in J_{p}}\mathcal{Q}_{c}(\check{t}+\check{q}-j_{2};\theta,\epsilon)(j_{2}-\check{q})_{+}^{1/2}=O(\epsilon^{1/2}).\end{split}$
+Due to the term $\epsilon/|u|$ in the bound in (E.28), it is essential that the sum in (E.29) is over a subset of the range $0<j_{2}-\check{q}<c/\epsilon$. Therefore, by (E.26) and (E.29),
+(E.30) $\begin{split}G(\check{t},\check{q};\theta,\epsilon)=&a_{0}(0,P(0))\int_{\mathbb{R}}\frac{1}{\check{w}}\sum_{j_{2}\in J_{p}}h^{\prime}(\check{w}+\check{t}+\check{q}-j_{2};\theta,\epsilon)(j_{2}-\check{q})_{+}^{1/2}\text{d}\check{w}\\ &+O(\epsilon^{1/2}).\end{split}$
+It is obvious that if $\epsilon=0$ in (E.22) (i.e., if $\chi_{1\text{D}}$ is replaced by the constant function 1), then $\mathcal{Q}(u)=O(u^{-2})$, $u\in\mathbb{R}$. This implies that
+(E.31) $\begin{split}\sum_{j_{2}-\check{q}>c/\epsilon}\mathcal{Q}(\check{t}+\check{q}-j_{2};\theta,\epsilon)(j_{2}-\check{q})_{+}^{1/2}=O(\epsilon^{1/2}).\end{split}$
+Hence, in (E.30), we can extend the summation to all of $j_{2}\in\mathbb{Z}$. Further,
 (E.32) $\begin{split}K(\theta+\epsilon\mu s)=&K(\theta)+O(\epsilon),\ \frac{\Phi(\theta+\epsilon\mu s)-\Phi(\theta)}{\epsilon}=\mu s\Phi^{\prime}(\theta)+O(\epsilon),\end{split}$
-where $s\in\text{supp}\,\varphi_{\alpha}$, which implies (see (6.10))
-(E.32) $\begin{split}&\partial_{u}^{k}\big{(}h(u;\theta,\epsilon)-\mu h_{0}(u)\big{)}=O(\epsilon),\ k=0,1,\\ &|\partial_{u}^{2}\big{(}h(u;\theta,\epsilon)-\mu h_{0}(u)\big{)}|<c,\ u\in\mathbb{R},\theta=O(\epsilon^{1/2}),\end{split}$
-Similarly to (E.22) and (E.23), pick any $\bar{h}\in C_{0}^{2}(\mathbb{R})$ and define
-(E.33) $\begin{split}\mathcal{Q}_{\epsilon}(u):=\int_{\mathbb{R}}\frac{1}{\check{w}}\Delta\bar{h}_{\epsilon}^{\prime}(\check{w}+u)\text{d}\check{w},\ \Delta\bar{h}_{\epsilon}(u):=\bar{h}(u+\epsilon)-\bar{h}(u).\end{split}$
+where $s\in\text{supp}\,\varphi_{\alpha}$.
+Here and in what follows, we drop the first argument of $K(x_{0}=0_{2},\alpha)$ from notation (see the paragraph following (7.3)).
+Using (6.13), introduce the variable
+(E.33) $s_{1}=s_{1}(s;\theta,\epsilon)=\frac{\Phi(\theta+\epsilon\mu s)-\Phi(\theta)}{\epsilon\mu\Phi^{\prime}(0)}.$
+By the assumption in Theorem 3.6, $\mu\Phi_{\alpha}^{\prime}(0_{2},\alpha_{0})$ is irrational.
+Hence $\Phi^{\prime}(0)=\Phi_{\alpha}^{\prime}(0_{2},\alpha_{0})\not=0$.
+The assumption $|\theta|=O(\epsilon^{1/2})$ implies
+(E.34) $s_{1}(s;\theta,\epsilon)-s=O(\epsilon^{1/2}),\ \partial s_{1}(s;\theta,\epsilon)/\partial s=1+O(\epsilon^{1/2}).$
+Changing variable $s\to s_{1}$ in (6.13) gives
+(E.35) $\begin{split}h(u;\theta,\epsilon)&=\mu\int_{\mathbb{R}}K(\theta+\epsilon\mu\zeta)\varphi_{\alpha}(\zeta)|\partial\zeta/\partial s|\varphi_{p}(\mu s\Phi^{\prime}(0)+u)\text{d}s,\ \zeta=\zeta(s;\theta,\epsilon).\end{split}$
+To simplify the next step, in (E.35), the new variable $s_{1}$ is denoted $s$, and the old variable $s$ is denoted $\zeta$. Together with (7.18) this implies
+(E.36) $\begin{split}&h(u;\theta,\epsilon)-\mu h_{0}(u)\\ &=\mu\int_{\mathbb{R}}\big{[}K(\theta+\epsilon\mu\zeta)\varphi_{\alpha}(\zeta)|\partial\zeta/\partial s|-K(0)\varphi_{\alpha}(s)\big{]}\varphi_{p}(\mu s\Phi^{\prime}(0)+u)\text{d}s.\end{split}$
+Using that $\varphi_{\alpha},\varphi_{p}\in C_{0}^{2}(\mathbb{R})$ (see Assumption 3.4(1)) and (E.34), we conclude
+(E.37) $\partial_{u}^{k}(h(u;\theta,\epsilon)-\mu h_{0}(u))=O(\epsilon^{1/2}),\ k=0,1,2.$
+Similarly to (E.22) and (E.23), pick any $\bar{h}_{\epsilon}\in C_{0}^{2}(\mathbb{R})$ such that $\bar{h}_{\epsilon}^{(k)}=O(\epsilon^{1/2})$, $k=0,2$, and define
+(E.38) $\begin{split}\mathcal{Q}_{\epsilon}(u):=\int_{\mathbb{R}}\frac{1}{\check{w}}\bar{h}_{\epsilon}^{\prime}(\check{w}+u)\text{d}\check{w}.\end{split}$
 Then
-(E.34) $\begin{split}&|\mathcal{Q}_{\epsilon}(u)|\leq c\epsilon\ln(1/\epsilon)(1+u^{2})^{-1},\ u\in\mathbb{R}.\end{split}$
-For large $|u|$, (E.34) is proven integrating by parts in (E.33). For $|u|$ in a bounded set, we argue similarly to (E.12), (E.13).
-Using (E.32) and (E.33), (E.34) with $h_{\epsilon}(u)=h(u;\theta,\epsilon)-\mu h_{0}(u)$ we get from (E.30)
-(E.35) $\begin{split}&|G(\check{t},\check{q};\theta,\epsilon)-\mu G_{0}(\check{t},\check{q})|\leq c|J_{\epsilon}|+O(\epsilon^{1/2}),\\ &J_{\epsilon}:=\sum_{j_{2}}\mathcal{Q}_{\epsilon}(\check{t}+\check{q}-j_{2})(j_{2}-\check{q})_{+}^{1/2}=O(\epsilon\ln(1/\epsilon)).\end{split}$
+(E.39) $\begin{split}&|\mathcal{Q}_{\epsilon}(u)|\leq c\epsilon^{1/2}(1+u^{2})^{-1},\ u\in\mathbb{R}.\end{split}$
+For large $|u|$, (E.39) is proven by integrating by parts in (E.38) and using the bound on $\bar{h}_{\epsilon}$. For $|u|$ in a bounded set, we use the bound on $\bar{h}_{\epsilon}^{\prime\prime}$.
+Set $\bar{h}_{\epsilon}(u)=h(u;\theta,\epsilon)-\mu h_{0}(u)$. Using (E.36) – (E.39), we get from (E.30)
+(E.40) $\begin{split}&|G(\check{t},\check{q};\theta,\epsilon)-\mu G_{0}(\check{t},\check{q})|\leq c|J_{\epsilon}|+O(\epsilon^{1/2}),\\ &J_{\epsilon}:=\textstyle\sum_{j_{2}\in\mathbb{Z}}\mathcal{Q}_{\epsilon}(\check{t}+\check{q}-j_{2})(j_{2}-\check{q})_{+}^{1/2}=O(\epsilon^{1/2}).\end{split}$
 This proves the lemma.
 E.3. Proof of Lemma 7.3
-We evaluate the integrals in (7.17) similarly to [25, eqs. (3.16)–(3.20)]. Begin by considering
+We evaluate the integrals in (7.19) similarly to [28, eqs. (3.16)–(3.20)]. Begin by considering
 (E.41) $\begin{split}J(r):=&\int_{\mathbb{R}}\int_{\mathbb{R}}\int_{0}^{\infty}\frac{1}{\check{w}}\big{[}h_{0}^{\prime}\big{(}\check{w}+r+(Q_{\alpha\alpha}^{\prime\prime}/2)s^{2}-\check{q}\big{)}\\ &\hskip 85.35826pt-h_{0}^{\prime}\big{(}\check{w}+(Q_{\alpha\alpha}^{\prime\prime}/2)s^{2}-\check{q}\big{)}\big{]}\check{q}^{1/2}\text{d}\check{q}\text{d}s\text{d}\check{w}\\ =&(Q_{\alpha\alpha}^{\prime\prime}/2)^{-1/2}\int_{\mathbb{R}}\int_{\mathbb{R}}\int_{\mathbb{R}}\frac{1}{\check{w}}\big{[}h_{0}^{\prime}\big{(}\check{w}+r+u-\check{q}\big{)}\\ &\hskip 85.35826pt-h_{0}^{\prime}\big{(}\check{w}+u-\check{q}\big{)}\big{]}u_{+}^{-1/2}\check{q}_{+}^{1/2}\text{d}\check{q}\text{d}u\text{d}\check{w}.\\ \end{split}$
-Substitute $h_{0}$ expressed in terms of its Fourier transform, $\tilde{h}_{0}=\mathcal{F}h$, and use the identity [14, p. 360, eq. 20]:
-(E.37) $\mathcal{F}(x_{+}^{a-1})=e^{ai\pi/2}\Gamma(a)(\lambda+i0)^{-a},$
+Substitute $h_{0}$ expressed in terms of its Fourier transform, $\tilde{h}_{0}=\mathcal{F}h$, and use the identities [17, p. 360, eqs. 19 and 21]:
+(E.42) $\begin{split}\mathcal{F}(x_{+}^{a-1})&=e^{ai\pi/2}\Gamma(a)(\lambda+i0)^{-a},\ a\not=0,-1,-2,\dots;\\ \mathcal{F}(1/x)&=i\pi\text{sgn}\,\lambda,\end{split}$
 to obtain
@@
 (E.44) $\begin{split}J(r)=&-\frac{i\Gamma(1/2)\Gamma(3/2)}{2}\int_{\mathbb{R}}\frac{e^{-i\lambda r}-1}{\lambda}\tilde{h}_{0}(\lambda)\text{d}\lambda\\ =&\frac{\pi}{8}\int_{\mathbb{R}}\big{[}\text{sgn}(-r-t)-\text{sgn}(-t)\big{]}h_{0}(t)\text{d}t=-\frac{\pi}{4}\int_{0}^{r}h_{0}(-t)\text{d}t.\end{split}$
-Substitution of (E.39) back into (7.17) gives (7.18).
+Substitution of (E.44) back into (7.19) yields (7.20).
 Appendix F Proof of Lemmas 8.1 and 8.2
 F.1. Proof of Lemma 8.1
-Integrating by parts in (8.6) with respect to $\lambda$ twice and using a partition of unity we can assume that $I=(-\delta,\delta)$ for some small $0<\delta\ll 1$, and $D(\lambda,\theta,\epsilon)\in C_{0}^{\infty}(I)$ in $\theta$.
-On $I$ we can change variable $\theta\to t$ to get
+We may assume $n=0$ in (8.5). The case $n\geq 1$ is completely analogous.
+If 00 is not in the closure of $I$, integration by parts in (8.6) twice with respect to $\lambda$ yields the top line in (8.7).
+Therefore, using a partition of unity, we can assume that $I=(-\delta,\delta)$ for some small $0<\delta\ll 1$, and $D(\lambda,\theta,\epsilon)\in C_{0}^{\infty}(I)$ as a function of $\theta$.
+Change variable $\theta\to t=\sin\theta$ on $I$ to get
 (F.1) $J(r)=\int_{0}^{\infty}\int_{I_{1}}D_{1}(\lambda,t,\epsilon)e^{ir\lambda t}\text{d}t\lambda\text{d}\lambda=\int_{0}^{\infty}\tilde{D}_{1}(\lambda,r\lambda,\epsilon)\lambda\text{d}\lambda.$
@@
 (F.2) $I_{1}=(-\sin\delta,\sin\delta),\ D_{1}(\lambda,t,\epsilon)=D(\lambda,\sin^{-1}t,\epsilon)/(1-t^{2})^{1/2},$
-and $\tilde{D}_{1}(\lambda,\mu,\epsilon)$ is the Fourier transform of $D(\lambda,\theta,\epsilon)$ with respect to $\theta$.
+and $\tilde{D}_{1}(\lambda,\mu,\epsilon)$ is the Fourier transform of $D_{1}(\lambda,t,\epsilon)$ with respect to $t$.
 Thus $\tilde{D}_{1}(\lambda,\mu,\epsilon)$ decays rapidly as $\mu\to\infty$, i.e.
@@
 (F.4) $|J(r)|\leq c\int_{0}^{\infty}\frac{\lambda}{1+|r\lambda|^{3}}\text{d}\lambda.$
-Changing variable $\lambda\to r\lambda$ proves the top cases in (8.7).
-All the other assertions of the lemma can be proven in a similar fashion. For example, if $c_{k,l}=O(\epsilon)$ in (8.5), then $c_{N}$ in (F.3) and $c$ in (F.4) are $O(\epsilon)$.
-If $D(\lambda=0,\theta,\epsilon)\equiv 0$, then we integrate by parts in (8.6) three times instead of two. The analog of (F.1) becomes
+A change of variable $\lambda\to r\lambda$ yields the top case in (8.7).
+If $D(\lambda=0,\theta,\epsilon)\equiv 0$, then, away from $\theta=0$, we integrate by parts in (8.6) three times instead of two to prove the bottom line in (8.7).
+In a neighborhood of $\theta=0$, the analog of (F.1) becomes
 (F.5) $J(r)=\int_{0}^{\infty}(\tilde{D}_{1}(\lambda,r\lambda,\epsilon)/\lambda)\lambda^{2}\text{d}\lambda.$
 Selecting $N=4$ in (F.3) then gives
-(F.6) $|J(r)|\leq c\int_{0}^{\infty}\frac{\lambda^{2}}{1+|r\lambda|^{4}}\text{d}\lambda=O(r^{-3}),\ r\to\infty.$
+(F.6) $|J(r)|\leq c\int_{0}^{\infty}\frac{\lambda^{2}}{1+|r\lambda|^{4}}\text{d}\lambda=O(r^{-3}),\ r\to\infty,$
+and the proof is complete.
 F.2. Proof of Lemma 8.2
@@
 The last equality follows from (3.16) and (7.18) (recall that $\bar{h}_{0}$ is defined without the prefactor in (7.18)). Similarly,
-(F.8) $\begin{split}&\Upsilon(+\infty)+\Upsilon(-\infty)\\ &=\frac{1}{|\Phi_{x}^{\prime}|}\int_{\mathbb{R}}\bar{h}_{0}(u)\int_{\mathbb{R}}\text{sgn}(t-u)\mathcal{F}^{-1}\bigg{(}\frac{1}{1+\kappa C_{1}|\lambda|^{3}}\bigg{)}(t/|\Phi_{x}^{\prime}|)\text{d}t\text{d}u=0.\end{split}$
+(F.8) $\begin{split}&\Upsilon(+\infty)+\Upsilon(-\infty)\\ &=\frac{1}{|\text{d}_{x}\Phi|}\int_{\mathbb{R}}\bar{h}_{0}(u)\int_{\mathbb{R}}\text{sgn}(t-u)\mathcal{F}^{-1}\bigg{(}\frac{1}{1+\kappa C_{1}|\lambda|^{3}}\bigg{)}\left(\frac{t}{|\text{d}_{x}\Phi|}\right)\text{d}t\text{d}u=0.\end{split}$
 The last equality follows because the signum function is odd, while the functions $\bar{h}_{0}$ and $\mathcal{F}^{-1}\big{(}(1+\kappa C_{1}|\lambda|^{3})^{-1}\big{)}$ are even.